A Complete Bibliography of Publications in Communications in Statistics: Theory and

Methods: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

30 August 2019 Version 1.03

Title word cross-reference

(1, 3) [RC07]. (2, 3) [RC07]. (2^{m−p_{III}}) [YL06a]. (2^{m−p_{IV}}) [YL06a]. (m + 1) [AH04b]. (n, f, k(i, j)) [GCLG06]. (p, q) [CWG06]. +(1) [Sal06]. 1 [Blo07, WS00]. 2^{k} [BS03, Hua08, Sal03, SBM01]. 2^{M} [BS06a]. 2^{m1+m2} [KLHT06]. 2^{n−p} [LI00, LH07]. 2 ≤ m_{1}, m_{2} ≤ 4 [KLHT06]. 2 × 2 [AFM08, Smi00b, WW07, CL09, CY09]. 2 × c [RC03]. 2 × k [PB00]. 4 [Fly05]. −X [LC05b]. _{3}F_{2} [AJSC08]. _{pk} [PL00]. A [CS09a, LFRLFG07]. α [BLZ08, Cha08, LN09]. AR(1) [BFZ03, LW07a]. AR(p) [IWC05, KM02]. ARCH(0, 1) [SW03]. C_{5} [HCL00]. C_{p} [Sca02, SS05b]. W [AGHMP05]. χ^{2} [CC02, De 08, Gra03, CQ05]. D [FW04, HAS00, TH06]. D_{s} [HH09]. DBL(p, 0, 1) [LWW03]. E [HAS00]. E(s^{2}) [CCG09]. ϵ [CHW08]. F [GEGMMI09, Kab05, NO01, Wil09]. Γ [JP08]. garch(1, 1) [KCL00]. K [SB06, ADP05, Ayd05, Bie07c, CP04, DLb05, Dem07, DD08, iHA00, JS09, JP08, KOF02, Kli07, MPP07, MS07c, RR02, TPP02, TPP03]. L [AGH08, Gor09, KO05a, PC02, SV08b, MPP07]. L^{1} [JC08, Lou08]. l^{p} [Blo07]. L_{1} [DW00, EAN00, LT04, Wan05b]. L_{2} [WQC07]. ⟨n, f, k(i, j)⟩ [GCLG06]. Logodds(p) [Par09]. M [CSV04, DM00a, Han08a, Hwa02, QL09, SVC07, Wan07b, XLZL07]. MA(1) [AZ02]. MA(∞) [SYCZ02]. μ [DFP04]. n [Ili03]. p [Blo07, CL09, CY09, De 05, Par09, Pum07, Wan06a, WW03, XW08b, Zha07c, dlHRB00, dlHRB03]. P (X > Y) [SM08a]. P (Y < X) [RMK08]. ϕ [JS09, MMPV01, Rao03]. ±k [Bor03]. Pr(X < Y) [Bak08b]. q [Cha09, JN09, Kem01]. R [AFK04, JS06b, MW09, Cha07]. R({00, 10, 01, 20}|Ω) [KLHT06]. r(X, Y) [AL05]. s [WS00]. S^{2} [ZBLG05]. Σ [PGF03]. t [BCV02, DGLS03, Dut07, FDR08, FM06, GTB07, JPB09, Kha00, KS08, Kim02, LP05, LCL01, NC02, SN06, TD07, Tsi08, VC00, WGP08]. t(GT) [AG03]. T^{2} [And09, AdL09, JR05, Kak07, MS00, SBJ^{+}07, WSB09, FM06]. U [AS07a, CC00, HA01, PC08, YTN07]. U(0, θ) [Wan05a]. X [AL05]. Y [AL05].

-Admissible [CHW08]. -Bayesian [Han08a]. -beta-geometric [Kem01]. -Block [AH04b]. -Discrepancy [WQC07]. -disparity [MMPV01]. -Distribution [CQ05, FDR08]. -Distributions [LP05]. -disturbances [KM02]. -Divergences [JS09]. -Estimation [AFK04, JS06b]. -Estimators [CSV04, QL09, SVC07, SV08b, Wan05b, Wan07b, DM00a]. -Exploring [Sho02]. -Factorially [Cha09]. -Fold [Ayd05]. -Limit [JC08]. -Mass [DFP04]. -Minimax [JP08, HAS00]. -Mixing [LN09, Rao03]. -Modified [Cha08]. -Moments [AGH08]. -Norm [Blo07]. -Optimal [FW04, HH09, CS09a, TH06]. -Optimality [LFRLFG07]. -Overlapping [MPP07]. -PA [LCL01]. -Parameter [Fly05]. -Ratio [GEGMMI09]. -Record [ADP05, Cha07]. -Records [Cha07, DD08, JP08]. -Sample [SB06]. -Series [BLZ08]. -Shaped [AS07a, CC00]. -Smoother [Hwa02]. -Spacings [JS09, XLZL07]. -Statistic [PC08]. -Statistics [Gor09, KO05a, YTN07, PC02]. -Symmetry [MW09]. -Tabulated [Han08b]. -TDT [WS00]. -Test [FM06, FM06, HA01, Kab05, NO01]. -Type [Gra03, AGHMP05]. -Value [Wan06a, WW03, XW08b, dlHRB03, dlHRB00]. -Values [CL09, CY09, De 05, Zha07c]. -Weibull [JN09].

1 [AM07]. 1273 [Shi08a].

2007 [CNSV08].

35 [MBS17]. 36 [CNSV08].

Aalen [ENL05, Ros09]. Aberration [Jac09, DLT00]. Abrupt [GKB06]. absence [MD00]. Absolute [CHL01, GT03c]. Absolutely [DJ08, EGB06]. Absolutes [TB07]. Abundance [Mac02]. Accelerated [AGAAG02, CN09, FP01, Lia07b, LHQ07, LNP02, YH06, YCW^{+}08, YKC06, WL00]. Acceptance [Ami03, GR05]. Acceptance-Sampling [Ami03].

Accessibility [EGS05]. Accounting [Pul02]. Accuracy [Bro07]. Accurate [Sho05]. Across [Leu02]. Across-Location [Leu02]. Activity [AF09, BB04]. Adaptation [SC09a]. Adaptive [AFK04, BBB09, BC08a, CZR02, GSYB03, GF06, KKP07, Sal06, Wan06b, YW07]. Addendum [Mon00]. Addition [ABN07]. Additive [BFZ03, CS09b, DM08a, Deb08, SAT05, SSF04, WC07, Che00]. Adjacent [MPG08]. Adjusted [AWJ05, CS04b, DW07, FFCN05, Fro03, KM05, KS04b, YGKD09]. Adjustment [Nai02, Vas08, WL08, WLQ09, DM00b]. Adjustments [CCC06, Tu01]. Admissibility [NM07, WL01a]. Admissible [CHW08, LWZ07, Xu09]. Advances [Edi09b]. advice [GT00]. After [FM06]. Against [AO09, BD03, FW07, Mur01, SB03, Tak01, SS00a]. Agatston [HK06]. Age [Han08b, Fu,00]. age-period-cohort [Fu,00]. Ageing [HMN03]. Aggregate [TWH08]. Aggregation [MW01]. Aging [FV09, MGD08, NS08]. Agreement [HLS09, SQ04, ZS02]. AIC [YNI05, NWTI03, YYK07]. Aided [SPG07]. AIDS [LD00]. Aircraft [CM05]. Akaike [HM04, YNI05]. al [Muk06, Che06]. Al-Sultan [Che06]. Alarm [And09, MPS08]. Algebraic [AL05]. Algorithm [Aga09, BDH08, BA06, GT03a, HW06, Kan07, Led07, RK01, SL08a, WW05]. Algorithm-Based [HW06]. Algorithms [Ma06, MT00, PR08b]. alias [Joh00]. Aligned [MPS04]. Allocation [BW03, CZ03, PNJ07, ZW00]. Allow [BL07]. Almost [YLX09]. Alternative [AFdBP05, JW05, LTHK01]. Alternatives [AS07a, Gur06, Mur01, SB03, SJ04, Wil09, SS00a, WH00]. AMOC [Gur06]. Among [BC05b, BC08b, LJ06, SQ04, YWL08]. analogs [MS00]. Analyses [HML04, SC06, PK00a]. Analysis [Abo05, Akd01, AF09, Ano04c, AND04, Art08, BHN05, BBL04, Bar02, BC05a, BGS07, CS07a, CCK02, Che01, CS04b, CT04, CB07, CS09b, Coo09, DLN08, DMGFRQG07, Dra02, Dut01, Dut07, ES09, EGMPMRPM05, FS04, GBSW09, GM02, Ger09, Gia03, GVP05, GVMS07, GVW05, GO05, HZSX06, HK06, HS09, HC02, JC08, KA04, KKSA07, KC04, KS02, LMM01, Li01, Li09b, Liu01b, LR09, LBZ09, MPS04, MPP01, MLB09, MA01, MB07c, Nag02, NS09, NM05, NM09, OS05, Pat01, Pav03, Pul02, Pul06, RDI01, RT09, RCA02, Rod03, SHRA06, SM06a, SA04, SL08b, ST01a, SAT05, SSF04, TZY02, TXG03, TZ01, Tsi02, Tsi03, TC04, Tso09a, Tu01, VL06, WHZ03, WCL04, Wan07a, XSD^{+}08, XD09, YNI05, Yu07, YKC06, ZF03, ZYL06, Zha09, Zum07, Bar00, Fu,00, GT00]. analysis [Gön00, LZ00, Man00b, Mar00, Nor00, PGPG00, SHA00, Tsi00]. Analytical [RCX06, WSPP08, CNZ00]. Analyze [BN09]. Analyzing [Han08b, JM09, SBJ^{+}07, ZM09b]. Anant [Dir06]. Ancillary [WH08]. and/or [LW04, LW07a]. Anderson [TH08]. Angular [Fal03]. Animal [CS07a]. Ankylosing [CSP^{+}09]. ANOVA [KSG03, Nor03, XW08b, Yan00a]. Antonio [Nad09]. Appearances [Ant03]. Application [ASS05, BDH08, CM05, CHYL08, CWG06, CSP^{+}09, GAAA07, Gia03, GM08, GB01, JLBB07, LAR07, LRG08, LR09, LCL09, MN08, PL09, RK01, RCBM09, SLB08, SHRA06, SM08a, SP09, SSM06, SC09b, SM01, TC09, WQC07, WH08, WL01b, XZ08, ZC09, Fu,00, Hür00, LD00]. Applications [Agg01, Ani04, ACS01a, Bak08a, BM01, Cha08, CQ05, CKL08, CA08, Coo09, CL08b, CLY08, DGLSG02, DMGFRQG07, Edi09b, Efr09, ELF02, GSMGVM08, Gru08, Han08a, Har04, IPS08, IM04, JC08, JN09, JP08, KP09a, Lin04, Pat07, PGTM06, PJH09, ST05a, SG05a, Sha01b, Shi01, Sul07, TPS08, ZZ09, PG00, TY00]. Applied [CS07a, Dra02, Nor00, Vex08]. Applying [CDL01]. Appreciation [Dir06]. Approach [Ahm04, AA09, AND04, Ant03, CCK02, CS04b, CS09b, CKK04, DATM05, DH06, DGRQ01, DGR07, GZ03, GLW05, GVMS02, HKR02, HLS09, HRVW07, HRB01, HL09, HC02, Hug04, Jaś07, KG08, KR03, Kov09, LLX09, LQ05, Li01, Li02b, LH01, LTHK01, Mic06, Muk05b, NN09, NA07, OZMC02, RSS03, SY08, SM08b, TBM03, Wan09b, WTZL01, XSD^{+}08, YNI05, ZT09, ZF03, dlCMM03, dlHRB06, BL00a, CP00, iHA00, LZ00, QTR00]. Approaches [Car09, KG00]. Appropriate [CCGB03]. Approximate [GK04a, SW09b, CYHL00]. Approximated [AAG03, LBSM01]. Approximately [Zho01]. Approximating [Ham06, Raq04]. Approximation [CL01, DD04, KL04, Liu01a, OSS05, SL08a, SO06, WJ06, XK08, ZK08, ZZ09, CC02, WDRB00]. Approximations [CCD02, FJJZ05, FJSZ07, Kol07, KSG03, LCL09, MMPV01, PR04, PJH09, SB06, Sho05, TS08a]. Arbitrary [Ram08]. Arch [LL02, Hin09, SP08b, ZW08]. ARCH/GARCH [Hin09]. Archimedean [Rod08]. Arcsin [VH01]. Arcsine [KK04b]. Area [BRG07, CL08a, LSW05]. Arfima [HL02b, Rav01]. Argument [LX09]. ARIMA [SP03a]. Arising [GH06, Hui05, KO07, OS09b]. Arithmetic [Kim07, SL07]. ARMA [BA05b, ITV06, Lee05a, Pou07, SDZ01]. Aroc [App02]. Arrays [EK04b, Has08, OS05]. Arrival [Led07]. Arthritis [Kov09, She09, TC09, ZC09]. article [Ano07]. Artificial [Tsi09]. Artificially [Kim09, MD00]. Ascertainment [KE01]. Aspect [MS06]. Aspects [SDM08]. Assays [HN05, ZS02]. Assembly [MS07a]. Assess [LL01]. Assessing [DG00, FKB09, LC08, MW01, Mei08, SJ00, Woo02, WC08, YGKD09]. Assessment [HLS09, Kab05, SHRA06, Sil06, XG09, pLM00]. Asset [FHL03]. Associated [BL09, Jan05b]. Association [BL07, CD07, Gup04a, LLCNM01, MPP01, NQMRLÚF01, NÚF05, SG08, XG09, QTR00, Shu00]. Associative [Rao04]. Assumption [CL06a]. Assumptions [LN09]. Asymetric [FZ06]. Asymmetric [AT01b, AT01a, FZ05, FJC07, PX03, QC09, SC05, SvDM07, Sol08, VL06, Vis09, WL01b, Wil05b, Wil07, WC08, YZ05, PL00, Sai00]. Asymmetry [Özt01, SM01]. Asymptotic [AO07, ALOS09, BGAH09, Bar09, BCMRPAEG07, BHH07, BL04, CCBB09, Cho01, Col01, Col03, CSV04, CF05, CL08b, DGR07, DM04, EM01, EOS08, FDR08, FR08, Gau06, Ger09, Gil07, GCB00, GT03b, Hod09, HC02, ITV06, Jeo01, JO05, JS07, Kak07, KH03, Kol07, KN01b, LOS07, LN09, Lin05, LB05, Liu01b, Myn06, ON03, Oga08, Pak09, PPL00, Pum07, QL09, SL02, Sco07, SLZ08, Sim08, SV08b, TF04, Usa02, VT04, Wan06b, WWY08, WYZ05, XKWW08, Yan00a, Yi01, ZM09a, Cho00, Coe00, KM02, MM02, PC02]. Asymptotically [DFP04, MLY02, Rah08]. Asymptotics [CS06b, LM04, Wan05b, BCV02]. Attainable [SL04]. Attaining [Tan03, Tan06]. Attainment [AO07]. Attracted [LR01]. Attraction [Fal03, Has08]. Attribute [CK07, NA07, NA09, GR05]. Audit [RvdGM05]. Auditing [MS07b]. Auditors [RMvdG06]. Augmented [Pre04]. Augmenting [Tsi09]. Author [Ano03a]. Autocorrelated [Dut07, MI08, NAS08, Sca02]. Autocorrelation [CS04a, LW07a, Cho00]. autocovariance [Cho00]. Automatic [Yua05]. Automobile [GDSPSVP08]. Autoregression [Ter02]. Autoregressions [AFK04, CY07, Myn06]. Autoregressive [AT01a, AT05, BR09, BA06, CC01, Gau06, HN01, Lag04, LL02, LLN09, Li09a, LL01, Nie05, PS05, SA03, SA08, Sha09, Shi08a, Shi08b, Ter01, Wei09, YC02, Guo00, Had00, MM02, Ter00]. Auxiliary [GSA08, KK09, SG07, SV08a, SK09b]. Availability [BM09, GS04b, KSW09]. Average [AH04a, BM09, BR07, BL04, CHC07, GK05, Jaś07, Lag04, Lue08, LVVA04, SSA07, sS03b, US03, VG09a, Wan05b, MM02]. Average-Power [CHC07]. Averages [SAAAS07]. Averaging [Ani04]. axial [HAS00]. Axiomatic [HHL07].

B [Nad09, ADS02, LM04]. B-Spline [LM04]. B-Splines [ADS02]. Back [KP07]. Backward [SL09b]. Bags [Arn01a]. Bahadur [KS09, MM02]. Balance [BKB02]. Balanced [Agg06, AJ06, AWA05, Ano07, BA07, BCRW02, GKK07, Gru04, Gup06, IT09, KRB05, KLHT06, PH02]. Band [Luc01]. Bandit [Yue04]. Bands [SSD01, WYY00]. Bandwidths [BP09a]. Barroso [MBS17]. Bartlett [AC00, CM01, CCC06, FUOC02, Tu01, WW03]. Bartlett-corrected [AC00]. Bartlett-Type [FUOC02]. Based [ANN08, ADP05, AA03, AB05a, AG03, BT07, Bak06, BA05a, BK05a, BBB09, BAA07, Bie07a, BS06b, BH05, CAB02, CZL08, CGT02, Cra01, DGZ01, DV05, Dem07, DF08, DGR07, Dix03, DCN07, DD04, Dub06, Eks02, EBF09, Fie07, Fig07, FMAA08, GHMGMPP06, Ghi05, GY05, GR09, Gra03, GGS04, HZSX06, HA01, HM02, HN03, Hin09, HW06, HL09, JH08, Jah04, Jah05, JSV01, JS09, KK03, Kaw04, KH02, LBAVM07, Led06, LNP01, LL07, LWB03, LWB04, LWB06, LHP06, LCC08, Lun02, MR09, McE08, Men02, Men07, MvdM01a, MPV06, NS09, NA09, NGYZ09, PC08, PS07b, PRVHVA09, Pew04, Pre04, PJH09, QS06, RS04, RAK08, RC07, RYS09, SL08a, SB06, SP02, SHRA06, SP04, Sho05, SC05, SG05b, Sol06, TA08, TWH08, TLW08, TS07, Tza09, UP03, VLLS06]. Based [VH07, Wan07b, WSB09, WL03, YWL08, YZ07, YC05, DM00a, ENS00, GA00, KM02, Mei05, Neu00, WJ00, dlHRB00]. baseline [KK00a]. Basing [CDB06]. Basis [AKKM09, GF08, JN03]. Baumgartner [Neu00]. Bayes [Ano07, BH01, Agg06, AH01, Arn01a, BA07, BMW08, BR05, Bha00, CMNP07, sC07, CDL01, FSZ07, GKK07, HC01, HL09, Jah04, KKL06, Koi06, Koy06, KS04a, Li02b, LY05, Lia08a, Lia00, Lia03, Lia05, Lia06, Lia07a, Lia08b, Lia08c, LH09, MG09, Ohy02, PN00, PK00b, Rau09, SP07, SP09, Sol00, SP05, TTX01, TBM03, Wan05a, WZ07, WL01a]. Bayesian [Ami03, ANMSL06, AE03b, Bar00, Bar01, Bar02, BRM00, Bro07, CS07a, CP00, Che01, CL09, CDKK05, DRB07, De 07, De 08, DH06, Dem09, ED08, FCL06, FW04, GM02, GDSPSVP08, GSMGVM08, GVMS02, GVP05, GVGP06, GVMS07, Gön00, Han08a, Hsi01, HRB01, HL02a, KA04, KKP07, KDCW09, KLK00, Kov09, LLX09, LM06, LH01, Lu00, MR09, Man00b, MM00, MV00, MS07b, MD07, MI08, MS05, MAS05, MB07c, Nan09, NE05, NN09, OJ08, PJ06, QAT05, RT09, RCA02, Rod03, Rod06, SDZ01, SA03, SSA07, SA08, She09, SB08, SB09, SAT05, SS05b, SC06, SP08b, TTX01, TD07, Tsi01a, Tsi01b, Tsi02, Tsi03, Tsi08, Tsi09, TS08b, UP03, VPS08, WL01b, Wil09, WGP08, XD09, Zha06, ZC09, Zhe00, Zum07, dlCMM03, dlHRB03, dlHRB06]. be [Bie08]. Beads [Arn01a]. Beal [RT06]. Bechhofer [Pan05]. behaved [LCL01]. Behavior [LOS07, Pak09, Sim08, VT04, WS05, ZM05, ZBN09]. Behrens [Kab05]. Beijing [CHYL08]. Benchmarking [ZT09]. Bernoulli [BBQ02, BC08b, BC09, SCL09]. Bernstein [AM07]. Bessel [AMP05]. Best [App02, CP09, GD05, HL09, Kli07, Lin04, NN09, RLÚF05, Spu05b, Tak09, Wul05, WL00]. Best-Location [CP09]. Best-Possible [RLÚF05]. Beta [BR09, ELF02, GN05, NG04, Nad05, Nad09, dSFM07, CC00, Kem01, PG00]. Beta-Gamma [BR09]. Beta-Normal [ELF02]. Better [HA01, LZJ00, SS00a]. Betting [LB07]. Between [Bar09, BL07, Blo07, GT03c, Hud07, Jon07, LLCNM01, LZ01, LNRS07, LBSM01, NR04, PS07a, PS07b, Tan03, UP03, Wil03, Yu07, YCW^{+}08, PPL00, PLL00]. Bhattacharyya [Koi06, Tan03, Tan06]. Bi [MS06]. Bi-Aspect [MS06]. Bias [CGR09, CU08, CCC09, ES09, Iba03, KR08, Mac02, Sai01, VC00]. Bias-Corrected [CCC09, Mac02]. Biased [Akd01, AE03a, GH06, SKA01, SA01, YZ07, dUÁRC06]. Biases [CSV04, GBSW09]. Biasing [ÖK07b, Pak04]. Bibliography [Dra02, Nor00]. Bilinear [BH06, MF03]. Bimodality [AGH08]. Binary [App02, Bru08, CD07, Che08, FJSZ07, HM04, Kan07, Kim02, LC05a, MXZ08, MM05, Ora06, Rao04, SS01, SS02b, TLF04, YW07, ZM09b, FC00, QTR00, ZW00]. Binning [MS07a]. Binomial [Arn01b, BC05b, Cha08, CZA08, Dre06, Ili03, JM09, Kat01, KK04a, LBSM01, NN09, OS09b, PS07a, RT06, SZ06, SDM08, SP05, SP03b, Wei09, Woo02, XP04, XK08, FL00, iHA00, Lu00, SSG00]. Binomial-Type [KK04a, iHA00]. Bioavailability [XSD^{+}08]. Bioequivalence [CM09, XSD^{+}08]. Biography [Ano04a]. Bioinformatic [Efr09]. Bioinformatics [CSP^{+}09, MN08]. Biomedical [RCPO04]. Biostatistics [CSP^{+}09]. Bipolar [Fig07]. Birnbaum [FL06, Rie03, SLB08, Tsi01a]. Birth [Jan05b]. Birthday [FS05, PO06]. Bisexual [GMM01]. Bispectral [Hin09, JH08]. Bispectral-Based [JH08]. Bivariate [ASS05, BCPK05, BR09, BK05a, BLK09, Bar06, DATM05, DK02, DB06, DW07, DSPS09, EB07, FV02, FV07, FV09, HN09, HdMS01, IM04, IM08, JB08, JBS09, JHL07, KK01, KP08, KP09a, KP07, MY07, NQMRLÚF01, Nie05, OJ06, Pul06, Quo06, Rao04, SAS04, SASAS06, SG08, She05b, SL04, SS07b, SS01, SC09b, Wil05a, XD04, Yam07, YC09, YH06, AKP00, DC00, FFvR00, JLR00, SM00, WH00]. Blended [GKB06]. BLINEX [WL01a, WL01b]. Block [AH04b, BS08, BD03, GKFM^{+}08, GV05, Hua08, PM02, SPG07, Smi05, Tin01, Tin02, JS00]. Block-Savits [BS08]. Blocked [GV05, GT00]. Blocking [Guo03]. blocks [Pre00]. BMP’s [GDCOCO06]. Board [Ano00h, Ano00e, Ano00a, Ano00b, Ano00c, Ano00d, Ano00f, Ano00g, Ano00i, Ano00j]. Bobecka [Bou09]. BOLD [KRB^{+}09]. Bootstrap [BRG07, DWB01, GBSW09, Gat06, Gui00, HCL00, HT07, JW06, Pat01, SP01, WJ06, CNZ00, WJ00]. Borel [LVVA04]. Both [Mon01, NM05, NM09, LZ00]. Bound [AO07, GIT08, Koi06, NS06, Spu02, Tan03, Tan06, WC07, XW08b]. Boundaries [GM08, HC01]. Boundary [DW07, Hua06, JK04b, DJ00]. Bounded [Kli07, Myn06, WL01b]. Bounding [CKM07]. Bounds [Bie07a, Bie07b, Bie07c, Bie08, CZL08, Dan05, EEPL08, Fro07, GT03c, Gor09, GR09, KO05a, KO05b, Kli07, Lee03, NQMRLÚF01, NÚF05, QN08, QC09, RCX06, Raq04, RLÚF05, Sil04, Spu05a, VJ07, WQC07, Wil05a, YL06a, RR02]. Box [FM06]. Branching [Rah08, VKT00]. Breaking [Joh00]. Breast [Han08b]. Breiman [PD07]. Brief [NS08, Pan05]. Bringing [Mak08]. Brownian [BHH07, JK04b, Ree07]. Buehler [KL05]. Burn [CN09]. Burn-In [CN09]. Burr [Bra03, Jah05, JW09, Mok05, PS01b]. BVBGAR [BR09].

c [PL00]. Calculating [Kat01]. Calculation [BA05b]. Calculations [ACS01a, RCL08]. Calibration [AA02, CKB00, GBG06, GAB01, KKM07, Kit02, MR05, NP01, NP08, RMAM09]. Can [Bie08]. Cancer [Han08b, LHQ07, sS08a, BRM00, WCT00]. Candidate [DLN08]. Canonical [AM06, CK05, Rud02, TZY02, Tri07b]. Canonization [Stȩ07]. Capability [HL09, Lin05, PX03, SC05, SS05b, PL00]. Capture [AG08, DLZ09, Hug04, SC06, ZLY05, GA00]. Capture-Recapture [DLZ09, Hug04, SC06, ZLY05]. Carcinogenicity [HN05]. Cards [OS09a]. Career [Edi09a]. Carlo [FP01, GS04b, Kim07]. Case [AS07a, BGM08, CSM08, CHB09, GK04a, Hod09, KC08, Lia06, Nav08b, PD07, SA04, SS07b, SSF04, Tin09, ZN03, dlHRB06, PK00b, Yan00a, ZFY00]. Case-Cohort [SSF04]. case-control [ZFY00]. Case-Crossover [Nav08b]. Case-Study [SA04]. Cases [AH04a, CP09, LMM01, Zew08]. Castillo [NK05]. Catalogue [DLT00]. CATANOVA [DBA05]. Catastrophe [NS06]. Catchability [GA05]. Categorical [BL07, Cha04a, FS04, KML01, McG07, Nan09, TTB08]. Category [BL07]. Cauchy [DSC01, JB08]. Causal [Cho01, Zha08]. Causality [CY07]. CDF [Sho05]. cDNA [Gup06]. Cell [ACP03, DH06, Nan09]. Cells [CG04, DFP04, HZSX06]. Censored [AGAAG02, AE08, AT02, BA05a, BK05a, BS06b, BD05, BC08a, CAB02, CCGPO03, Cra02, DF08, EM01, GZ03, GAAA07, HJ06, JIP08, JE01, JC08, Kan02, LOS07, Li02a, LWB04, LWB06, LNC09, LSD07, Lue08, MP09, MR09, MAS05, OSS05, PH03, PP07, Pon07, QS06, SQ04, SB06, She05b, She07, Sol06, Sol08, SSD01, SC09b, TC09, WZ01, WJ06, WL03, YSZ09, YYL09, ZZW07, ZP04, ZL05, ZYSZ08, Zum07, AKP00, SMS00, WJ00]. Censoring [Agg01, AO01, Asg06, BZX09, BV05b, BDF06, BCK07, HL02a, Hwa05, JN03, KP09b, KK07b, Lia08c, LH09, Mar02a, PS07b, SY08, WL05, Wan08, ZYSZ08, AKP00, AO00, KLK00]. Censorship [Dau03, Jeo03, MLY02, Ros02, Wan05a]. Censorship/Truncation [MLY02]. Center [Bon08, FJJZ05, JS06b]. Central [CMLB05, CQ05, CS08b, Fab08, Fab09, GH04, MYS07, Nad09, OS09b, PR08b, Wil03, ZZ01, dSFM07, DGLS03, Nad05]. Certain [ASA02, CK04a, KK04a, MS03, Ram06, Sha07]. Cessation [QWSL09]. Chain [CS07b, CG07b, XL06]. Chains [AE03b, CKL08, GL04, Led06, SL02, Joh00]. Chance [AFM08]. Chance-Corrected [AFM08]. Change [BDH08, Bar02, BMS01, DI08, FCL06, GH09, GW01, HSAN05, HT06, Jac09, KA04, KWW04, KR02, Las08, LL07, LZZ08a, PC08, PJ06, RG04, Ram05, SL08b, VG06, Vex08, XM00, YWK01, ZLH^{+}09]. Change-Point [BDH08, HT06, LZZ08a]. Change-points [XM00]. Changepoint [Gur06, LQ05, Lia08a]. Changes [Bro09, CW01, HT07, LLN09, MPT09, XGZ09, ZC02, KCL00]. Changing [BBQ02, CC01, Kan07, PP07]. Characteristic [Fab09, LSW05, Mei05, SM01, TS07]. Characteristics [BR05, YX08, VKT00]. Characterization [Ahs04, AA03, BÖ07, BS08, Bou09, DLb05, DB06, Dod04, Gov01, GK04b, GGS04, HV03, HVW07, HM02, Mar02a, NS06, Nas06, Pak04, Roy04, TA08, VH07, WSPP08, ZM05]. Characterizations [ABN04, Asa04, BAA07, CKK04, Dem07, GS04a, KK04a, KG04, NG04, NR04, Nav08a, RS04, SG04, SW04, SSM09, XD04]. Characterized [EH04, TA07]. Characterizing [Ham06]. Chart [AdL09, CK07, CMY01, GS08, KK03, LC05b, LXG06, MS09, MBPP02, Men02, ZBLG05, Alw00]. Charts [Bak06, GD06, Men07, MOPS08, NWA02, Rog06, SC03, ZC02]. Checking [RMvdG06, lTTNlT08]. checks [SMS00]. Chi [BM01, CMLB05]. Chi-Square [BM01, CMLB05]. Chinese [XL06]. Choice [BS03, CDKK05, KLBW07, MA01, UP03]. Choices [BL07]. Choosing [AKS06, KS05a, ÖK07b]. circle [Pew00]. Circular [Art08, Gat06, HN01, IT09, JK04a, Lun02, MPP07]. Claim [GDSPSVP08]. Class [AVGQ04, CCGB03, CNG03, EOOPGB09, HSAN05, Hud02, IM08, Jon07, Mei05, OB05, PSZ07, PK02, Pre04, QWSL09, RT09, SG05a, SP02, Sha07, She07, SB03, SP04, SK09b, Sun02, WWY08, Wei09, XMX08, YH08, Fou00, Kim02, KCV00, LD00]. Classes [FV09, Kem04, Kim09, Rei06]. Classic [CF05]. Classical [De 05, KKM07, LVNF07, Min02, Pan05]. Classification [CV08, JLBB07, JWSP04, KRB^{+}09, Kim09, LY05, Lin01, MK05, SDZ01, SS01]. Classifications [MPP01]. Clear [LH07, YL06a]. Clinical [CL08b, FJJZ05, FJSZ07, HML04, HC01, Li09b, LW07b, LHQ07, LR09, MT05, Ruk09, Tin09, YNI05, LD00, LC00, WCT00, ZW00]. Clock [KP07]. Closed [CK05, DMGFRQG07, RL09, CAC00]. Closeness [GK03, HI09, HS05, Rei06, Wen02, Yan07, Fou00]. Closure [Mak08]. Cluster [NMS08, Sal06, ZM09b, Man00b, PJ00]. Clustered [WCL04, WH08]. Clustering [DD04, EH04, HM04, Por08]. Clusters [GEGMMI09, VTKB04, YZ01]. Cochran [CDB06]. Cochrane [ST01b]. Coefficient [ASS05, Bur08, CS08a, CK02, Gir04, HKR02, HK05, HWL09, Jeo01, KM05, KS04b, LFRLFG07, LM04, LZZ08b, NC05, PS05, QL09, SL07, SW03, TN08, Tso05, WP02, YC02, ZL06, ZC09, Bha00, Guo00, Sep00, WYY00, BH01]. Coefficients [AENT00, For09, Guo03, LW02, Sak02, SGC09, Tso09b, VJ07, BA00, Mon00]. Cohen [YC09]. Coherent [NRS07a]. Cohort [SSF04, TF09, Fu,00]. Cointegrated [CY07]. Cointegration [CKL07, MM00]. Collinearity [Fu05, Liu03b]. Colored [CH09]. Column [PGP03, Sal06]. Combat [Liu03b]. Combination [CM05, YTN07]. Combinations [BK09, GdT05, Tan03, dBSSvV01]. Combinatorial [RSS03]. Combined [EK04b, HN05, LH07, TH06]. Combining [BÖG09, KS01, Shi01, Wu06]. Comment [Mar02b, Özk09]. Comments [Muk06, Nad09, dSFM07]. Common [For09, GZ05, HK05, HRB01, JE01, YW08]. Communications [CNSV08, MBS17, Shi08a, Ano07]. Community [JLBB07]. Companies [XL06]. Comparative [GBG06, GAB01, NS04]. Comparing [AJ09, Bur08, HS05, MM02, Tso03, dUÁRC06]. Comparison [BK05b, CZR02, HI09, Hus05, IR07, JZ01, KKM07, KM05, LSRD04, MPT09, NÚF05, Ohy02, Rie03, SP01, Sho07, Sol00, SBJ^{+}07, Swi09, Vas08, XK08, Yan07, YWL08, BH01, SBM01]. Comparisons [AE03a, Aos01, Fro07, HZ00, HWZ06, KAHS09, KS04b, MPS04, OZMC02, ÖK07a, PS07b, PM02, SKA01, Shi01, Spu05b, VH01, Yam07, YLX09, PLL00]. Compartmental [Gan04, VKT00]. Compatibility [lTTNlT08]. Competing [Bas09, ENL05, EB08, Nag02, SSF04]. Complementary [RYS09]. Complete [QS06, LC00]. Completeness [DZK04]. Complex [AJSC08, Ric07, RACSSCJ05]. Component [BÖG09, EHT04, EGMPMRPM05, KS01, MS07c, RSS03, Tak09, Ton02, YC03, Zha07b, WYY00]. component-wise [WYY00]. Components [Aos01, Dut01, Güv05, Güv06, JW05, KSS03, KN01a, Kna02, LL05a, LM07, Li08c, LW02, Mar07, Mar08b, MXX08, NL07, NRS07a, Sad08a, ST01b, SSJ02, VL06]. Composite [Coo09, HHL07, SSG00]. Compositional [RDI01, SBJ^{+}07]. Compound [PP07, BRM00, JS00]. Compromise [PNJ07]. Computation [GKFM^{+}08]. Computer [JJ08, SPG07]. Computer-Aided [SPG07]. Computing [BES04, CS09a, DRB07]. Concentration [KS04b]. Concentrations [LMG06]. Concerning [GBS05, Mon00]. Concomitant [BES04]. Concomitants [Ery05, HN09]. Concordance [QS06]. concordant [LZ00]. Concurrent [ZYZ06]. Condition [GVGP06, Zha01]. Conditional [ACS01b, BÖ07, BS06a, BB04, CD03, CC01, DGZ01, EOS08, GSYB03, HKR02, HN05, LL02, LL01, Mar08a, Nas06, NR04, OGM06, Par09, SN00, SSM09, lTTNlT08, WW07, BP00, SJ00, WC00]. Conditionally [ACS01a, AG08, GA05, MS05, SS07b, GA00]. conditionally-unbiased [GA00]. Conditionals [ACS07]. Conditioning [LR01]. Conditions [Dut07, FFVFVF03, Tza08, PK00a]. Cones [Bou09]. conference [Ano04c]. Confidence [AR09, BBK05, Blo08, BH05, Bur08, BK01, BK05b, CZA08, CGT02, Dub06, EEPL08, Fab09, FK06, Fra04, Fro03, GYX06, KL05, LS02, LL05a, LM07, Li08c, LX08, LS09, Mac02, PH02, PB04, Par09, QT02, RAK08, RYS09, RW02, RT06, SB08, SB09, SCL09, SSD01, Swi09, TN08, VJ07, WF05, Wil05b, Wil07, DM00a, FL00, FTV00, Gui00, HCL00, KOF02, Lu00, WYY00]. Configurations [MN08]. Confounding [Zha08]. Conjoint [Gia03]. Connected [KP08]. Conscious [Sta09]. Consensus [JS06b]. Conservative [FTV00]. Consideration [KKP07]. Considerations [Swi09, Tin09]. considering [PJ00]. Consistency [AGHMP05, Bel02, FFVFVF03, FP09, OST09, XWW07, YA09, YZZ08]. Consistent [Bar05, CNG03, LH02, LT04, SL07]. Constant [ASN03, Bha01, Hwa02, Laj06]. Constrained [Ano07, CS07a, GKK07, RK01, TM07]. Constraint [BL04]. Constraints [HK06, LWZ07, Ma06, ST05b, XW08a, Xu09]. Construct [Hua08]. Constructed [KKMS07]. Constructing [CZA08, FY00, SS05c]. Construction [Ahm00, EGB06, Sri02, SSD01, MT00]. Constructions [AJ09, KRB05]. Constructive [Spi07]. Contact [Pum07]. contagious [PJ00]. Contains [BK03]. Contaminated [Iba03, TC04]. Contamination [MI08]. Context [FHK07, SM06b, YA09, ZS06]. Contiguous [BCRW02]. Contingency [CL09, CY09, DBA05, FW07, GVP05, RC03, PB00]. Continuity [XK08]. Continuous [BBB09, DM08a, DJ08, EGB06, Fab08, Har04, Hug04, KBC02, KH02, Las08, LLCNM01, MPV06, NS06, PvdL07, VH07, dlCMM03, CBsK00, GA00]. Contoured [BZ08, Ng00]. contours [FFvR00]. Contraction [BK08]. Contractions [Nav08a]. Contrast [HH09]. Contrasts [Jan05a, KM08b, LR01]. Contributing [MCY08]. Control [AJ09, AdL09, Bak06, BGM08, BC09, CKF09, CK07, CMY01, GS08, GD06, HL03, HRVW07, KK03, LC05b, Luc01, LVVA04, MBPP02, Men02, Men07, NWA02, OZMC02, Rog06, SP09, SC03, Spu05b, ZBLG05, CAC00, PLL00, ZFY00]. Controlled [SS04]. Controller [JT09]. Controls [RvdGM05]. Controversy [Mak08]. Conventional [PD07]. Convergence [Aga09, AD02, GBR01, Ham08, Lia09a, SL09a, SYCZ02, Tri07a, Wan05a, XWW07, PR08b]. Conversion [Das07]. Convex [ANMSL06, BK09, MY07, WD04, YTN07, LZJ00]. Convolution [Ayd05, Sch02, vDK03]. Cook [LL01]. Coordinate [EB07]. Copula [ANN08, BMN06, DV05, DW07, Sun05b]. Copula-Based [ANN08, DV05]. Copulas [AB05a, BM08a, DJ08, DSPS09, EGB06, IM08, Jav09a, Jav09b, KKMS07, Lie05, Rod08, SS08b, Sun02, SN05]. Cordeiro [MBS17]. Cores [Fab01]. Corrected [AFM08, CM01, CD08, CCC09, Mac02, AC00, VC00]. Correction [CU08, Gau06, Hua06, Wan09a]. Corrections [FUOC02, XK08, CNZ00]. Correlated [AT02, AND04, DLN08, FFVF01, Gen06, GZ05, IM04, JBS09, Kan02, LW04, LSRD04, MA01, ST01b, SSJ02, WW07, WYZ05, Zho01, Zum07, RJ00, SSG00]. Correlation [ASS05, BFZ03, BH05, Bur08, Dut07, Guo03, HN09, HN07, HWL09, LW04, LWW03, LCC08, Men01, NC05, PR04, Sak02, SL04, SM08b, TZY02, Tso05, BA00, Bha00, BH01, Shu00]. Correlations [AM06, CK05, HN01, SC09a, Gön00]. correspondence [FC00]. Corrigendum [AT01b, MBS17, Shi08a]. cosine [Wu00]. Cost [AH04a, Car09, Jaś07, LVVA04, RCBM09, Sta09]. Cost-Conscious [Sta09]. Costs [JZ01]. Count [BB04, GO05, KM08a, MPS08, PSZ07, Quo06, XL09, XD09]. Countable [CG07b, GL04]. Counting [BL09, CHYL08, Fie07, Hug04, SK06]. Counts [DH06, Wei09]. Covariable [WZ01, ZL05]. Covariance [Art08, CB07, CBB06, CS08c, CNG03, GH09, HN01, Hod09, MBS17, Pan02, Vas08, WSB09, YSH05, AC00, JS00]. Covariate [CB07, DE09, DLZ09, DM04, LT02, PD07, SK01, XY07b]. Covariates [CU08, JH01, KDCW09, Kra08, LLCNM01, MVB01, Ora06, sS05a, YH06, Zha09, WH00]. Covariation [GdT05]. Covariogram [FRMAA08]. Coverage [AF08, CZA08, RW02]. Cox [BV05b, Bro02, Bru08, CL06a, DE09, DM04, FM06, JO05, Kra08, PD07, Sam03, YCW^{+}08]. Cox-Regression [Sam03]. Cramér [Sco00]. Credible [Han08a, Zha06]. Cressie [GHMGMPP06]. Criteria [BCK07, YIS04, Fou00]. Criterion [Bar09, CCG09, DG08, HI09, HM04, KK02a, MP08, ÖK07b, PC08, YNI05, ZCL09]. Criterion-Robust [ZCL09]. Cross [AAGM05, APdR01, KP02, MPP01, Spi07, Wan06b, WTZL01]. Cross-Classifications [MPP01]. Cross-Sectional [AAGM05]. Cross-Validation [APdR01, Spi07, Wan06b, WTZL01]. crosscovariance [Cho00]. Crosses [BD03, GD05]. Crossing [BHH07, JK04b]. Crossover [JW05, Nav08b]. Csiszar [Bha01]. Cuboidal [WSPP09]. Cumulants [Wil03]. Cumulative [EHT04, GKFM^{+}08, Han08b, ZQ07]. Cure [LT02]. Current [CS09b, Raq07, SP08a]. curtailed [BLB00]. Curtailment [BC08b]. Curvature [DYKG09]. Curve [BRG07, BR07, Blo08, DATM05, Pan02, SO06, Vas08, BTH00, KOF02]. Curves [BBB08, LZ01, LSW05, MK05, SP02]. Customers [Sta09]. cusum [Alw00, LXG06, MPT09, Rog06, KCL00]. Cycle [BP09b].

D [MBS17]. Damage [LMR09, SC09b]. Dams [BP09b]. Darling [TH08]. Data [AGAAG02, AT02, AAGM05, AF09, Art08, BHN05, BN09, BK05a, BS04, BD03, Bha01, BS06b, BB04, BC08a, CCD02, Cha04a, Che01, CT04, Che08, CS09b, Chi09, CCGPO03, CA08, Coo09, CNSV07, CNSV08, CLY08, DATM05, Dat03, Dem09, DF08, DW07, DLZ09, Dut07, Eks02, EB08, EM01, FHK07, Gal02, Gat06, GAAA07, GY05, GG05b, Gil07, GBS05, GDSPSVP08, GMPG04, GVW05, GN03a, GO05, Gur07, Han08b, HMM03, HC02, JIP08, JK04a, JM09, JC08, JP08, JR07, KC08, Kan07, KD08, KDCW09, KN01a, KH02, KR02, KR03, KML01, LS02, LOS07, Li02a, Lia07b, LWB04, LC05a, LWB06, LNC09, LB05, LXG06, LR09, LSRD04, LN05, LSD07, Lue08, LTHK01, MP09, MR09, MFPP07, MB07a, Mon01, MAS05, MB07c, Nag02, Nie05, Par01, PH03, PSZ07, PD07]. Data [Por08, Pul06, QWSL09, QL09, Que09, Quo06, RYS09, RACSSCOJ07, SB06, SS07a, SG08, SBT02, SL08b, ST05b, SS03a, She05b, She07, SC06, SL05, Sol06, Sol08, ST01b, SBJ^{+}07, SSD01, SGC09, SC09b, Tar05, TC09, TF09, TC04, VTKB04, WZ01, WJ06, WV06, Wan07a, WH08, Wan09a, WLQ09, XW08a, XL09, XD09, YNI05, YC09, YA09, YH06, YYL09, YKC06, ZF03, Zha07c, ZZW07, ZM09b, ZP04, ZL05, ZYSZ08, Zum07, dL07, dUÁRC06, AKP00, CR00, CC88, DC00, Gil00, KLK00, LCL01, Man00b, MD00, MV00, QTR00, Sca02, Shu00, WJ00]. Data-Based [GY05]. Data-Driven [GG05b]. Data-Transformation [MB07a]. Database [PD07]. DD [Bie07b]. DDA [Bie07b]. Dead [Luc01]. Dead-Band [Luc01]. Dealing [Don00]. Death [AAGM05, CJ04]. Deaths [Han08b]. Debugging [LNP01]. Decision [ANMSL06, BL04, DD04, GVGP06, Jaś07, LY05]. Decks [OS09a]. Decomposing [Mon01]. Decomposition [AAGM05, Fal03, GF08, IPS08, SN05]. Decompositions [TT07, Kim00]. Deconvolution [LH02, LT04, LT08]. Decreasing [LCF06]. Default [CDL01]. Defects [Hsu05]. Deficiency [GN04, MP08]. Defined [EE09, Fig07, Jon07, SW04]. Deformation [GVW05]. Deformed [MS03]. Degradation [BHN05, BN09, CM08]. Degree [BQ05, YWL08]. Degree-Based [YWL08]. Delay [And09]. Deletion [WCS03, Zew08, BP00, WC00]. Deliberately [Ber05]. Dementia [CJ04]. Demétrio [KM08a]. Demonstration [AL05]. Deng [PTY09]. Deng-Yuan [PTY09]. Densities [Tak01]. Density [BF02, CDL01, DGLSG02, Eid09, GY05, HN07, IT03, Jon07, KKP03, LH02, LT04, LT08, LN09, Lie05, MT06, ON03, Pak09, PGTM06, Rod08, SA09, SW09a, SLZ08, WS05, WL08, dlHRB06]. departures [Bar00]. Dependence [BLK09, Bar09, CKF09, Dol08, DSPS09, ES09, FFVFVF03, FMAA08, Gal02, GdT05, HWZ06, Jav09b, KC04, Li08b, Mud03, Mud08, PvdL07, SG04, SL04, Sun02, Sun05a, YH08, YBCH08, DR00, DW00]. Dependencies [Grz03]. Dependent [APdR01, Bar05, BZ08, BM08b, CCK02, DM04, HT06, HRB01, Jeo03, KO05a, Li09c, Lim07, Nag02, NL07, NRS07a, RCA02, Rod05, Sak02, SPG07, sS05a, WLQ09, XP04, MV00, WYY00]. Depth [CLY08, PH03]. Derivation [Kaw04, SAVG07, Sty09, TPS08]. Derivative [Bar02, Dar09, ON03]. Derivatives [GG05b, MT06]. Derived [Bie08, JS06a, Min03, NÚF05, SSG00]. Deriving [FJSZ07]. Design [AdL09, BBB09, BM08b, Chi07, CKM07, DG08, Fri03, Güv05, HML04, Hua08, Hud07, Kit02, LH07, LHQ07, MP08, Nav08b, NWA02, RS03, RCBM09, Rog06, Sal06, ST01a, SB08, SB09, Smi05, Sty09, SSF04, Tin01, Tin02, ZBLG05, ZYZ06, Fu,00, KOF02]. Designed [BW03, DG00]. Designing [Alw00, Tin09]. Designs [Abo05, AB01, AJ06, AJ09, Arn01b, BS06a, BD03, BKB02, Bor03, BH05, BS03, But05, But06, CCG09, CS09a, CW01, CJ02, FW04, GD05, Gup06, HMA05, HLLL08, HH09, IT09, Jac09, Jan02, KRB05, KMS09, KS02, KLHT06, Li01, LL05a, LM07, MVB01, MM05, Muk03, PGP03, Pre04, Pre08, Sal03, SPG07, Sri02, SS05c, TTB08, TH06, TLF04, WCT00, WSPP08, WSPP09, Wul05, YL06a, YLZ09, YW07, ZCL09, Zha07a, Zho01, Cha00, CDL00, DLT00, GT00, HAS00, KCV00, LI00, LC00, Pre00, SBM01]. Desirability [FKB09]. Detect [GEGMMI09]. Detecting [Bro09, CP04, DLN08, GLW05, GKB06, LMM01, MPT09, ZN03, PLL00]. Detection [AAG03, Bro07, CC08, CDL01, Dut01, Eid09, GG05b, GH09, HSAN05, VG06, Vex08, WCL08]. Deteriorating [AH04a]. Determination [Aos01, CCGB03, CZ07, GW01, KM05, KS04b, VH01, YNI05, VHS00]. determinations [AENT00]. Determine [Muk05b, NN09]. Determining [Che06, LMR09, VTKB04, WW03]. Deterministic [Myn06]. Developing [AS08, IM08]. Development [HHL07, PL09]. deviance [DM00b]. Deviation [Dau03, MJS09]. Deviations [CM05, CHL01, HPR07, Mac08, MP09, PR08a]. Diabetes [Gan04]. Diagnosis [KK03, She09]. Diagnostic [McE08, RS01]. Diagnostics [CHL01, DGRLS03, Hat09, LAR07, LW07b, SP03b, WCL04, Zew08, BP00, BL00b]. Diagonal [DJ08, SM01]. Diallel [BD03, GD05]. Dichotomized [Kim09]. dichotomizinlg [CBsK00]. Dickey [CDL01]. Difference [GT03c, Hud07, HL02b, OS09b, PS07a, RT06, SB08, TLW08, XK08, FL00, pLM00]. Difference-Based [TLW08]. Differences [Bie07a, Dan05, Nav08b, WW07, WSB09, PLL00]. Different [JO05, KK07a, ZL06]. Differential [Rav01, SA04, SL09b, FY00]. differentiation [LC00, WCT00]. Diffusion [DI08, RT09, YYW05]. Digestibility [CS07a]. Dilation [BK08]. Dilution [Blo02]. Dimension [BQ05, MCY09, TF04, YA09, Coo00]. Dimensional [BGS07, BCRW02, DFP04, KN01b, Pav03, Pon07, Ryn07]. Dimensionality [YIS04]. Dimensionless [PvdL07]. Direction [DR00, WF05, DW00]. Directional [Mud03, Mud08, SL08b, Sun05a]. Dirichlet [FS05, Hui05, SC02]. Dirichlet-Related [SC02]. Disability [AAGM05]. Disaggregation [ZT09]. Disassembly [PS01a]. Disassembly-Reassembly [PS01a]. disasters [VKT00]. Discontinuity [GG05b]. discordant [LZ00]. Discount [Tri07a]. Discounted [AH04a]. Discovery [De 07, GMKFH08]. Discrepancy [De 08, WQC07, ZF03]. Discrete [AJSC08, BBL04, Bro09, CCD02, CKL08, Dem07, DD08, Ger09, GIT08, HC01, HMN03, JS06a, Kem04, Lia08c, Luc01, LTHK01, MBB08, MPT04, Min02, MA01, OJ06, OJ08, OS09b, Por08, RACSSCOJ07, Roy03, Sim08, lTTNlT08, Ton05]. Discrete-Time [BBL04, Ton05]. Discriminant [BC05a, BGS07, Leu02, Pav03]. Discriminate [AFdBP05]. Discrimination [AND04, AKKM09, CK02, CLY08, ZCL09]. Discussion [BS06a]. Disease [AF09, LMR09, YZ01, Fu,00]. Diseases [FKB09, TF09]. disequilibrium [WS00]. disparity [MMPV01]. Dispersion [CS08b, CBB06, CU08, FP01, TWW01, Vin07, LI00, MBS17]. dissolution [pLM00]. Distance [EBF09, Fan09, Gru08, JR07, LL01, CC00, Kim00]. Distance-Based [EBF09]. Distinguish [Muk05a]. Distortion [DFP04, SL05]. Distributed [BZX09, BZ08, Che01, Ery05, LS02, MK03, PR08a, ST05a]. Distribution [AS07a, AS07b, Adc07, AAS01, Ahm08, ADP05, Ahs04, ASS05, AA03, ALT04, AMP05, AG03, Asg06, ALOS09, AJSC08, Ayd05, Bak06, BA05a, BK05a, BES04, BR05, BD05, Bol07, Bou09, BM01, BOR02, CMLB05, CB06, Cha09, CQ05, CGT03, sC07, CK07, Cho01, CMY01, CGA06, CA08, DSC01, DP06, De 08, DLb05, DF08, DK02, DB06, DGLSG02, Dix03, DMGFRQG07, Dre06, DLZ09, EHT04, ELF02, FDR08, Fly05, For05, FV09, FL06, Gen06, Gen07, GT03a, GAAA07, GAMG09, GKFM^{+}08, GT03b, GTB07, GMM01, Gov01, GMO05, Grz03, GN03a, GB01, GK03, Gup04a, GN05, Güv03, HKR02, HN05, HM02, Hod09, Jah04, JW09, JB08, JBS09, Jan05b, JS07, JJ08, JRM09, Jon07, JN09, KSW09, Kha05, KO07, KF07, KK01, KK04b, KS05b, KP08]. Distribution [Kra05, KP09a, KP09b, KK07b, Laj06, LL03, Li02b, LG03, Li09c, Lia03, LOS05, LWB03, LWB04, LWB06, Liu03a, MLY02, Mac08, MR09, MM06, MFPP07, MFPG07, Mei05, Mei08, Men07, Min03, MB07b, NG04, NG05, NC05, NE03, NQMRLÚF01, NÚF05, NN09, NW09, OGM06, OR07, OS09b, OB05, PK06, PS07b, PRVHVA09, PK02, Pew02, Pew04, Pol05, Pon07, Pul06, RC02, Ram06, Ram08, RMK08, RJ04, RCL08, RC03, Rie03, RLÚF05, Roy03, RF02, SAVG07, SAS04, SLB08, SB06, SK09a, SZ06, SBT02, Sco06, Sco07, SM09a, SB03, SP04, SC03, Sho02, Sho05, Sho07, SvDM07, SPS07, SSM06, ST08, Sul07, Tar05, TA08, TH08, Tom03, TF04, TS08b, Tza08, Tza09, Vau02, VJ07, VL06, VPS08, Vin07, Vis09, hWlF03, Wan05a, Wan06a, Wan08, WF05, Wil05a]. Distribution [Wil05b, Wil07, WL03, XP04, YZ05, ZK08, ZBN09, Aos00, BRM00, CFB00, CL00, HA00, HZ00, JLR00, Kem01, NC02, PK00b, Pew00, SS00b, Sol00, TPP02, Wu00, Yan00a]. Distribution-Free [Bak06, SB03, SP04]. Distributional [LLN09]. Distributions [ASA02, AO01, AVGQ04, ACS01a, ACS07, Ars09, Asa04, BÖ07, BT07, Bak08b, BLK09, BGAH09, BCMRPAEG07, BK09, Bie07a, Bie08, Bra03, BL09, CM05, Cha08, CAB02, CS06a, CG07a, Cha07, CF01, CCGPO03, CF05, CHB09, Dem07, DD08, DK02, DGLS03, EOOPGB09, Fab08, Fab09, Fan06, FJ09, Fig07, FHK07, For09, GG05a, GS04a, Gav08, Ger09, GSMGVM08, GT03c, GN04, GK04b, GIT08, HSAN05, HK05, HMK05, HA01, Hud02, IA09, ITY01, JBS09, JPB09, JK04a, JS06a, JW06, JC08, Jon07, Kak07, KKC06, Kan02, Kaw04, Kem04, KK04a, Kim06a, KM08a, KH02, LG03, Li08b, Lia09a, Lia06, Lia07a, LH09, Lin01, Lin05, LP05, LNRS07, LBSM09, LB07, Ma06, MP09, MPP07, Mar02a, MY07, Men02, MF03, Min02, MvdM01b, MGD08, MRT09, Mok05]. Distributions [MPT09, MT01, MT02, Muk06, MS03, NS06, Nas06, Ohy02, PS01b, PvdL07, PG00, RJ04, Rod03, Rod06, RACSSCJ03, RACSSCJ05, RACSSCOJ07, SMGVMD02, SG05a, ST01a, Sha07, Sha01b, Sha02a, SLZ08, SI02, SL04, SDM08, SK05, Sty09, SM06b, SSM09, Sür06, Tan03, Tan06, TS08a, TA07, lTTNlT08, TIS01, TK01, TPP03, Tsi01c, Tsi02, Tso09b, VLLS06, Vin04, VH07, WHZ03, Wil03, WGP08, XD04, YH08, Yeh04, Yeh09, Yu07, vDK03, Cho00, Coe00, CM00b, CC00, FFvR00, GO00, iHA00, MT00, Ng00, PJ00, Sco00, Tsi00]. Disturbance [TOS09]. Disturbances [CWS01, KM02]. Divergence [Bha01, GHMGMPP06, GVMS07, HMM03, MPP08]. Divergences [JS09]. Diversity [Fro03]. Divisibility [Yeh09]. Divisible [Sri02, SS05c, Vin07]. Do [Muk05a, HS09, SN06]. Domain [Fal03, Has08]. Dominated [ANMSL06]. Dose [Che08, HR03, HR08, Jan05a, Par01, CAC00]. Dose-Response [HR03, HR08, Par01]. Double [GT03b, Gov01, HK02, JT09, Lia07a, PN08, RMvdG06, RJ04, ST07, BLB00]. Doubly [BS06b, DGLS03, SB06, SSM09, CM00a, NC02]. doubly-truncated [CM00a]. Down [HMA05, Sty09]. Drifted [JT09]. Driven [GG05b]. Dropouts [LR09, QWSL09]. Dual [Ahs04, ZY08, ZW00]. dual-objective [ZW00]. Duality [Li08b]. Due [CM08]. Dummy [KK03]. Duration [SY08]. Durbin [LR00]. Dynamic [ANN08, HRVW07, MOS04, NN09, SP08b, Tri07a, Tri07b, Tri08].

Early [Tin09]. Easy [Abo05, IM08]. Easy-to-Generate [IM08]. Econometric [GM02]. Economic [Ami03, NWA02, Pap04, ZT09]. Economical [HL03]. Edgeworth [WJ06]. Editor [AT05, Ano04b, Ano05b, Ano05c, Cho05, GGVM01, GTB08, GN03b, Jan01]. Editorial [Bal06, Bal07, Edi09b, Smi00a, Ano00h, Ano00e, Ano00a, Ano00b, Ano00c, Ano00d, Ano00f, Ano00g, Ano00i, Ano00j]. Eduardo [Nad09]. Effect [CL06a, CBsK00, DE09, DW07, Dup08, LW07a, SK01, AKP00, PJ00]. Effective [Che08, Jan05a, Alw00, LY00]. Effectiveness [Car09]. Effects [AJ06, AJ09, BH05, CZ03, CJ02, CDL01, EH04, FJSZ07, GBSW09, GV05, HMA05, HR03, JO05, LMR09, Li01, LT02, Li08c, LH07, OJ06, OJ08, Rao01b, RCBM09, SJS09, SHLT09, ST01c, Tin01, Tin02, XM08, XG09, XD09, YYAZ01, YC09, LD00, LI00, Mon00, SJ00, YD00]. Efficacy [Guo03]. Efficiency [EK04b, FM06, Gru04, Guo03, JSV01, Jeo01, Jeo03, JO05, KL05, KS07, KRB05, Liu01b, PM02, ST01b, YWL08, GCB00, MM02]. Efficient [AS07b, Ahm04, Bon08, CZ03, CG04, GMW04, Hua08, KMS09, KV05, Mar00, MPV06, OS09a, OJ08, RJ00, SPG07, SZ06, TTB08, Set04]. Eig [SBT02]. Eigenvectors [KS02]. Elimination [Sai01, Sal06]. Elliptic [DGLS03]. Elliptical [DGLSG02, DGRLS03, Has08, LAR07, Liu03a, VLLS06, YIS04, FFvR00]. Elliptically [BZ08, Hod09, Ng00]. EM-Algorithm [Led07]. EM-Based [SL08a]. Embedded [Sha02b]. Empirical [Ano07, BMW08, sC07, CC09, FSZ07, FHL03, GKK07, Gra03, HN03, HL09, Jah04, JW06, JRM09, KKC06, KZ04, Koy06, KS04a, Li02b, LY05, LX08, Lia08a, Lia00, Lia03, Lia05, Lia06, Lia07a, Lia08b, Lia08c, LH09, LOS05, LZZ08a, LCC08, Mac08, Mei05, NK08, QT02, Rao03, Rau09, TBM03, TS07, Wan05a, WV06, YSZ09, ZQ06, ZQ07, ZL05, PPL00, PN00, PK00b]. Endpoints [AF08]. Engin [Mud08]. Enhanced [JT09]. Enrollment [FJJZ05]. Entrance [CZL08]. Entrance-Based [CZL08]. Entropies [ZS06]. Entropy [BAA07, BGNR01, CG07b, Fro03, GL04, IT03, JP08, Kaw04, Lim07, Liu01b, MT01, MT02, NMS09, Ric07, Rod05, Rod08, ST07, CYHL00, SSW00]. Entry [KWW04]. Environment [Han08b, Nag02]. Environmental [Hsu05]. Epidemic [CHYL08, RG04, Sha01a, TY00]. Epsilon [GTB07]. Epsilon-Skew- [GTB07]. Equal [Bur08, CSM08, FK06, GA05, Kha08, LM07]. Equal-Catchability [GA05]. Equal-Tail [Bur08]. Equality [HK05, JW06, Kan02, KM08b, Li09a, Men01, Sak02, TM07, BA00]. Equation [Jaś07, LM06, OS09b]. Equations [FGM03, FR08, KDCW09, LQ05, Lee05b, SL09b, Tza08, Wan07b, FY00]. equicorrelated [GK00]. Equicorrelation [GBS05]. Equipment [LHP06, Pul01, Pul02]. Equivalence [CM09, EE09, EM01, Kol07, LZ01, PPL00]. Equivalent [YSH05]. Equivariant [CAB02, RC07, WD04]. Erlang [YYW05]. Errata [CNSV08, Vin05]. Erratum [Ano03b, Ano03c, Ano05a, Ano06, Ano07, Lee07, Yu06]. Error [AE03a, BR07, CM05, CL08a, CKM07, DG08, DRB07, FZ05, FZ06, Fu05, GHL07, Güv06, HL03, Iba03, KS08, Kna02, LL02, LN09, Ma06, MS07a, MA01, NW09, OS05, PH02, PB04, QN08, Raq04, SKA01, ST01a, Sho02, ST01b, SSJ02, ST01c, SV08b, TD07, Tsi08, TS08b, Usa02, WZ01, XW08b, YWK01, YZ07, YLX09, ZL05, ZBN09, AC00, CFB00, GCB00, KG00, SYCZ02, YD00]. Error-In-Covariable [WZ01, ZL05]. Errors [APdR01, BZ08, BM08b, CC09, FFVF01, LR01, LX08, LW04, LW07a, LWW03, LCC08, MK03, MI08, She09, SK09b, SVC07, TLW08, WYZ05, XY07b, YC02, Zho01, ZW08, dlCMM03, Don00, MM02, Sca02]. Errors-in-Variables [CC09, LX08, LCC08, XY07b]. Erythematosus [AF09]. Estimability [HS08, WWJ09]. Estimate [CL06a, DM08b, GSYB03, MD04, SC09a, sS02a, sS03b, Ter01, Ton02, hWlF03, WWY08, Ter00]. Estimated [Hud07, MFPP07, UT01]. Estimates [AGAAG02, BW03, CSV04, JO05, Myn06, SP08a, TB07, WL01a, Syk00, VC00]. Estimating [AT01a, AZ02, BD05, CSWM09, CGT02, Dau03, DCN07, EHT04, FGM03, GK05, GG05a, GV05, Hsu05, HC02, JW05, Kim07, KP09b, LQ05, LB07, Mak02, Muk02, Muk07, PL00, PC02, RT04, SM06a, SC05, SK09b, SV09, SSJ02, TBM03, Wan07b, WH08, WTZL01, WFZ08, ZGS03, ZLY05, Zha08, ZC09, GZ03, HAS00, KH03, LD00, QTR00]. Estimation [AS07b, ADP05, ASS05, AFK04, AMP05, Ano07, App02, ACS01a, AG03, Asg06, AG08, BMN06, BP09a, Bak08b, BJ03, BM09, BK05a, BMW08, BR05, Bel02, BGNR01, BR07, BA06, BB01, BM08a, BH06, BS06b, Bla09, Bon08, BM01, BDF06, BC08a, CS04a, CAB02, CS06a, CD03, CS04b, CL08a, CHYL08, CC88, Che05, CKL07, CG04, CG07b, Col01, Col03, CZ07, CC00, DSC01, DG08, Dar09, DI08, DH06, Dix03, DJ09, DLZ09, DM04, Eid09, Eks02, EK04b, Fab01, FCL06, Fan06, FSZ07, FZ05, FZ06, FL04, FFVF01, FBH03, FRMAA08, FMAA08, FL06, GHL07, Gau06, Ghi05, GY05, GKK07, GM08, GPM05, GMPG04, GMM01, GMW04, Grz03, GSA08, Güv06, HA00, Had09, HI09, HT06, HN01, HH05, HMK05, HJ06, HN09, HN07, HRVW07, HL02b]. Estimation [HK02, IT03, JIP08, Jah05, JW09, JP08, JS06b, KC07, KRB^{+}09, KV05, KS08, KKP03, KZ04, KN01a, Kna02, KS05b, KH02, KML01, KG02, KK07b, Las08, LT08, Li02b, Li02a, Lia08b, LH09, LS09, LM06, Lim07, LWB03, LWB04, LSW05, LSD07, LZZ08b, Lue08, LVNF07, MLY02, Mac02, MBB08, Mar07, Mar05, MM06, MG09, MvdM01b, MI08, MPV06, Muk03, MS07c, NK05, NS04, NE05, NMS09, OJ08, OS05, ÖK07a, ÖK07b, Özk09, PJ06, PY08, PP07, Pol05, Pon07, PR08b, QAT05, RTW06, RMK08, RC07, RACSSCJ03, RACSSCJ05, Ros02, RMAM09, Ryn07, SA09, ST05a, SAS04, Sam09, SA01, SZ06, Sca02, SM08a, SM09a, SG07, SP07, SP09, ST05b, SL07, Sha02a, Sha09, sS05a, She05b, SY08, SM09b, SS00b, SS05b, SST^{+}09, SSM06, SDM08, SM08b, ST08, ST01c, SW02]. Estimation [SL09b, TOS09, Tar05, TD07, Tin02, TC09, Tza09, UNN05, Vas08, Vis09, WS05, WZ01, Wan05a, Wan08, WL08, Wan09a, WZ07, WL01b, WL03, Wul05, XM08, XY07a, XW08a, YYAZ01, YC02, YC03, YSH05, ZM09a, ZCL09, ZL06, ZY07, ZW08, AY00, AKP00, Bha00, BH01, ENS00, Guo00, Had00, HZ00, Hür00, KK00a, Kha00, KK00b, Man00a, Mar00, MD00, PN00, PK00b, Sai00, SSW00, TY00, Tan00, WDRB00, Zhe00, Muk06]. Estimation-Equivalent [YSH05]. Estimation/Testing [KZ04]. Estimations [DF08, ÖK07a, SW03, Sol08]. Estimator [Akd02, ASW06, ALOS09, Ayd05, BCPK05, Bar05, Ber05, Blo07, CWS01, CDB06, CK02, DP06, ENL05, EOS08, Fan09, For09, Gir04, GA05, Hod09, HKW09, Hwa05, JN03, KS01, KC05, KM05, KKT05, KN01b, LH02, LT04, LL07, LOS07, LN09, LOS05, LW02, Liu03b, Liu04, MYS07, MJS09, NWTI03, ON03, OST09, PS07b, RZ02, SAS04, SG06, SS07b, ST01b, SG05c, TU06, TS08b, WJ06, WSB09, XWW07, XKWW08, YX08, YCL09, YZZ08, CFB00, CH00, Fu,00, GA00, Tan00]. Estimators [Akd01, AE03a, AWA05, AT09, AT05, AT02, AWJ05, AD02, AGHMP05, BF02, BCMRPAEG07, BÖG09, CGR09, CHW08, Cho01, CSV04, CD08, CCC09, CNG03, DGR07, DM04, Eid05, GBSW09, GM08, Gru04, Gru08, GN04, HS05, Ili03, JSV01, KC08, KKM07, Kha08, KBC02, Koy06, KK09, KS04a, KR08, Li09c, Lie05, LP05, LB05, Liu03a, LM04, LWZ07, Mac08, NA06, NK08, Nor03, Ohy02, ÖK07c, Pak09, PK02, Pum07, QL09, Rah08, Rao01b, RT04, Rei06, Ros09, RMAM09, SL02, Sai01, SKA01, SS06, SW09a, Set04, SL07, She05a, SLZ08, SV08a, SK09b, Sol06, SVC07, SV08b, Ter02, TT07, Tom03, TLW08, Usa02, Wan05b, WL05, Wan07b, WC07, Yan07, YZ07, YLX09, YSH05, Zie06, Zie09, dUÁRC06, Ahm00, Cho00, DM00a, FY00, Sol00, SYCZ02, ZFY00]. Euler [CWG06]. EV [MYS07]. Evaluate [Wen02]. Evaluating [Jan02, SA04]. Evaluation [GF08, JH01, LHP06, MT06, Men02, PL09, XP04]. Evaluations [Car09, FBH03]. Event [KMS09, LN05, Pul06, SK05, SC09b]. Event-Related [KMS09]. Evidence [HAA06, HHL07, MD07, SHRA06]. Evidence-Based [SHRA06]. EWMA [MOPS08, ZC02]. Exact [AS07a, BRG07, BMS01, BK01, Cha07, Fri03, GR05, HN05, LBSM01, Sty09, Tsi01c, Yam07, YZ01, ZM09a, Zha07c, vdM08, Neu00]. Examination [Akd01, UT01]. Examples [GVMS02, Tu01]. Exceedances [BT07]. Excess [Pak04]. exchange [Yan00b]. Exchangeable [LNRS07, ZM09b]. Exchangeably [MK03]. Excluding [BCRW02, MPG08]. Exclusion [Li08a]. Existence [DM08b, Tza08]. Expansion [CS09a, Oga08, TF04, WJ06, Yan00a]. Expansions [AKKM09, CHB09, EGMPMRPM05, Kak07, CL00]. Expectation [Nas06, SS00a]. Expectations [BÖ07, Bie07a, Bie07b, Bie07c, Bie08, Dan05, SSM09]. Expected [BES04, DH06, FJ09, GM01, GMKFH08, Liu01a, LVVA04, Raq07]. experience [LCL01]. Experiment [GV05, RJ00]. Experimental [Jan02, Kit02, KS02, MVB01, MM05, RS03, ST01a, Tin01, Wan07a]. Experiments [BS03, But06, Gup06, HAA06, HH09, Hug04, HR03, Mon01, MA01, OS05, PS01a, PM02, Pre04, Pre08, WP02, Cha00, DG00, KCV00, Mon00, PLL00, Pre00, SBM01]. Explanatory [BN09]. Explicit [CZA08, RW02, TTX01]. exploration [Don00]. Exploratory [SL08b]. Exploring [Sho02]. Exponent [Fan09]. Exponential [AS07a, Agg06, Ahm08, ADP05, Bak08b, BS08, BBB08, BD05, CS07a, CCBB09, CAB02, CQ05, CDKK05, CF05, CS08c, Cra01, Dod04, DM08b, FJ09, FW04, FUOC02, FFCN05, FV02, FV07, FV09, GK05, GAMG09, GSMGVM08, Gov01, GN04, GN03a, GK03, HKR02, HK05, HS05, HM02, HWZ06, HLW08, HL02a, Jah04, KBC02, KH02, KP08, KP09b, LLX09, Lee05b, Li02b, Lia06, Lia07a, Lia08b, LXG06, LR08, MP09, MR09, MKvdMT05, MS05, OR07, PK06, PK02, Pet09, RG04, RMK08, SAVG07, SMGVMD02, SB06, SP02, Sch02, SM08a, SM09a, Sil04, SSM06, ST07, Sul07, Tan03, TS08b, VPS08, Vin07, WC07, WTZL01, XWW07, ZN03, Alw00, Hür00, Lia00, Sai00, WH00, WL00]. Exponential-Bound [WC07]. Exponentiality [Dau03, Mur01, ZZ01, SS00a]. Exponentials [FV07]. Exponentiated [NG05, NE03, NE05]. Expressions [ASW06, HS08]. Extended [CD08, CLY08, Dix03, GAAA07, HM04, PN00, SAVG07, KOF02]. Extending [ZXM07]. Extension [LHP06, Min03, OS09b, RK01, SS07b, TXG03, WW05, YZ05]. Extensions [Adc07, Car02, SK00]. Externality [ED08]. Extinction [Fro07]. Extraction [McE08]. Extreme [CG07a, CS06b, DGR07, GG05a, GM08, Wu00, WL03, Yu07]. extremely [LZ00]. Extremes [HH05].

F [Nad09]. Factor [CMNP07, CSP^{+}09, HRB01, Mon01, NM05, SL08a, YL06a, pLM00, Mon00, PLL00, RJ00]. Factor-Gene [CSP^{+}09]. Factorial [Abo05, ADM08, BS06a, CW01, Gup06, HMA05, Jac09, KK04a, KLHT06, Sal03, WP02, YLZ09, LI00]. Factorially [Cha09]. Factorials [QC09, Joh00]. Factorization [GGS04]. Factors [AH04b, CDL01, Jac09, KKL06, Mon01, NM09, Tin01, YLZ09, Mon00]. Failure [ASN03, BHN05, BN09, Bha01, BL04, CCK02, Coo09, Dix03, FE06, GW01, GK04b, Li02b, MLY02, NR04, PP07, Pul06, SL02, SS07a, SK09a, hWlF03, YH06, ZYSZ08, SS00b]. Failures [Pul01]. Fall [DSW07]. Fallible [RMvdG06]. False [De 07, GMKFH08]. Falsifying [Hin09]. Familial [CD07, HN01, HMM03, MW01]. Families [AB05a, AGHMP05, Bie07b, Bie07c, Bie08, BYP07, EGB06, EM01, Gra03, HMN03, JK04a, KH03, KBC02, LLX09, LR08, Muk05a, TK01, VH07]. Family [Agg06, CCBB09, CG07a, sC07, Coo09, CF05, CS08c, FUOC02, FFCN05, FV09, GN04, KM08a, KK09, Lia08b, MBB08, MS05, Ohy02, Pet09, RG04, SC09a, SMGVMD02, SL04, SV08a, Tan03, Tan06, VLLS06, Vin04, XWW07, YZ01, BA00, FFvR00, GO00, Lia00]. Farlie [Güv03]. Farlie-Gumbel-Morgenstern [Güv03]. Fast [Ani04, GKFM^{+}08]. Fatigue [BN09, VLLS06]. Feasibility [CDB06]. Feature [Pav03]. Features [JWSP04]. fecundability [Kem01]. Feedback [Luc01]. Feeding [BB04]. Felix [Nad09]. Feller [Jaś07]. Fertilization [MD03]. Fiducial [OR07]. Field [AN07, LC05a, TN08, Cho00]. Fields [FP09, Li09c]. Fieller [Syk00]. Filter [DATM05]. Filtering [Led07]. Finance [TPS08]. Financial [FHK07, Quo06, Ree07]. Finding [CV08]. Finite [ADM08, ASA02, AGHMP05, Car02, CY07, Dar09, Dub06, EHT04, GZ03, Lia09b, LRG08, Mac08, NGYZ09, RL09, Sam09, SK09b, SST^{+}09, Sol06, lTTNlT08, Tsi02, Wan09b, WD04, WCL08, Yan00b]. Finite-Population [SST^{+}09]. Finitely [MPV06]. First [AH04b, Dar09, EB07, Gau06, Lag04, PS05, VT04, Guo00, Had00, HZ00]. First-Order [Gau06, PS05]. Fish [BB04]. Fisher [AE08, AO00, BA05b, BB01, EM01, Kab05, PF05, PK06, Smi07, WF05, ZP04, vdM08]. FISSAR [Shi08a, Shi08b]. Fit [AO01, BHD04, BBB08, Byo08, CK04b, Dam05, EH04, GMO05, GF06, GF08, Gra03, GN03a, HM02, JH08, KKC06, Mar02a, MM06, MFPP07, MB07a, Mei08, OGM06, PRVHVA09, SN06, Sür06, TH08, TS07, Woo02, XL09, MMPV01]. Fitting [Blo08, EB08, GDSPSVP08, RACSSCOJ07, Sho07, YX08, Mon00]. Fixed [BM08b, Fra04, Jaś07, NM05, NM09, Tin02, YC09]. Fixed-Effects [YC09]. Flexible [HZSX06]. Flight [CM05]. Flowgraph [Huz04]. focus [FS04]. Fold [Ayd05, LM07, Li08c]. Folded [CGA06, Kim06a]. Folding [HLLL08]. Foldover [CJ02]. Following [Chi07, MvdM01b, Sty09, TS08b]. Footrule [QS06, SQ01, SQ04]. Forecasting [SP08b, XL06]. Forecasts [Wen02]. foreign [Yan00b]. Forest [PD07]. Forewarning [Pul06]. Form [RL09, ST00]. Forms [For05, GS04a, PF05]. Formula [CZA08, CBB06, CS08c, OG08, SK06, TSSI02, MBS17]. Formulae [Hui05, TTX01, Wil03]. Formulas [Rog06, TPS08]. Formulations [Pan05]. Forward [SL09b]. Forward-Backward [SL09b]. Four [Guo03, Tsi01c].

Four-Parameter [Tsi01c]. Fractional [ADM08, BK09, IT03, Jac09, KLHT06, QC09, YLZ09, Joh00, LI00]. Fractionally [Shi08b, Shi08a]. Frailty [KD08, LL03, Rod05, SAT05, LD00]. Frame [NMS08]. Framework [CGR09, GBSW09, HW06]. Free [AS07b, Bak06, SB03, SP04]. Frequencies [BQ05, CWG06, RTW06]. Frequency [GDSPSVP08, KN01b]. Frequential [Dut01]. Frequentist [CMNP07, GVGP06, MD07]. Fresh [Pan05]. Friday [FV07, FV09]. Frontier [DMGFRQG07, TD07]. Full [BS06a, RT09, Rod06]. Fully [SP08b]. Function [AAS01, Agg06, ADP05, AWA05, AWJ05, Ami03, Aos01, AB05b, AJSC08, Ayd05, BK02, BA07, BB01, BS06b, DY08, DF08, DCN07, Eid09, FZ06, FP09, GS04b, GG05b, Gru04, GGS04, HN07, HKW09, HC02, HK02, JN03, JRM09, JP08, Jun02, KKC06, Las08, Leu02, Li08c, Mei05, MRT09, Nas06, NA06, NMS09, OSS05, Pon07, RCX06, RT04, Ros02, SG04, SM09a, Sha01b, sS05a, SM09b, SPS07, SSM06, ST01c, ST07, Ton05, TS07, Tza09, UNN05, WL05, WL01b, Wil05a, WFC01, ZQ06, ZQ07, CFB00, Sai00]. Functional [AKKM09, ALOS09, CS08a, Col01, Col03, EOS08, GBG06, GH04, KMS09, LM06, Lim07, ZL06, ZQ07]. Functional-Coefficient [ZL06]. Functioning [SA04]. Functions [ADS02, AB05a, Ano07, ANMSL06, BF02, CQ05, CHYL08, CGT02, EHT04, Gen06, GKK07, GM01, GKFM^{+}08, GK04b, JBS09, JK04b, Jon07, Li09c, Li09a, LX09, MW09, Muk02, Muk07, NR04, NQMRLÚF01, NÚF05, Pol05, PJH09, Raq04, RCBM09, RLÚF05, Sad08a, SP02, SM06a, Sun05b, WS05, WH08, WD04, WCL08, YC03, SN00, SSQ00]. Fundamental [Arn01a]. Further [Gil07, Swi09]. Future [AH01, Pul01, Rao01a]. Fuzzy [HZSX06, HW06, TB07]. FWER [CKF09].

G [MBS17]. GA [KLHT06]. GA-Optimal [KLHT06]. Galton [Fro07, GMM01]. Gambler [TT03]. Gamma [AAS01, AGH08, AO01, BR09, BOR02, CM05, CL01, FZ05, FV07, GK03, Hsi01, KKC06, LS02, Lia03, LWB04, LWB06, Liu01a, Mur01, NMS09, Ram05, Tsi01c, Tza08, Tza09, Wil03, CM00a, JLR00]. GARCH [Hin09, Yan00b]. Gauss [IP06, TU06]. Gaussian [Ami03, AO01, BLK09, CV08, CSM08, CK02, CK04a, FJC07, GAMG09, GVMS07, GMO05, Had09, Iba03, Kaw04, LC05a, Mar08a, Min03, MT01, MT02, MW09, RACSSCJ03, RACSSCOJ07, Sch02, Tan00, TM03, TM07, Ton05, WFC01, YA09]. Gaussian-Type [BLK09]. Gaussianity [JH08]. GEE [FGM03, OZMC02]. Gene [CSP^{+}09, DLN08]. General [ASW06, Asp06, BHN05, Bar06, Bar09, BR07, Bie07a, CCGB03, HSAN05, HS08, Hür00, IP06, KK02a, KSW09, KH03, NM07, OB05, Pew02, Pew04, RvdGM05, Rei06, SG05a, She07, Shi00, Sho07, SV09, Sol08, ST07, WCL08, Yeh09, Zie09, ZBN09, BP00, NO01, PGPG00]. Generalised [HL02b, Sha02a]. Generalization [CA08, JB08, JR05, KO07, LVNF07, Min02, Roy04, SMGVMD02, VG09b, BRM00]. Generalizations [vDK03]. Generalized [AAS01, Ahm08, Ahs04, AWA05, ACS07, AG03, Asa04, Asg06, BT07, BK08, Bie07a, Bie07b, BDF06, BL09, CM05, CM09, CCG09, CB06, Che00, dHCC01, CCC06, CBB06, CU08, CL08b, Dre06, EA05, FZ06, FGM03, FJC07, FV07, GK03, GGT05, GP03, HdMS01, Jah04, Jah05, KO05b, Kaw04, KKP03, KF07, KP09b, LD08, LQ05, LR08, Ma06, MR09, Mar02a, Mar05, MB07a, Men07, Nas06, NK08, PS01b, RC02, RMK08, SLB08, SK09a, Sco06, Sho07, ST08, Sri02, SS05c, Sul07, Tar05, TA07, TA08, Ter02, Ton02, TPP03, Tsi01c, Tsi08, Vau02, WC07, XZ08, XGZ09, XMX08, XW08b, YW08, Yeh04, YZZ08, YYL09, Yue04, Zha01, Zha07c, Coe00, DLT00, GO00, HA00, MBS17, PC02, QTR00, TPP02]. Generate [IM08]. Generated [EOOPGB09, HHL03, LC05a]. Generating [AJSC08, KKC06, Sha01b]. Generation [JJ08]. Generators [Rie03]. Genes [DLN08]. Genetic [HW06]. Genome [XG09]. Genome-Wide [XG09]. Genotypes [LBZ09]. Geometric [BLZ08, DLb05, DB06, Dod04, KP08, KP09a, LBSM09, NK08, PP07, SP07, YYY06, Kem01]. Geometry [AM06, ADM08, BOR02, Rav01]. Gibbs [Kan07, Zhe00]. Gini [Fro03]. Given [DJ08, EGB06, NQMRLÚF01, SAAAS07]. GLD [Sho07]. Glimpse [Edi09a]. global [FC00]. Gls [HL02b]. Gluing [SS08b]. GMANOVA [Yam07]. GMM [QAT05]. Goal [Hsu05]. Gompertz [JJ08]. Good [SP08a, VH04, YX08]. Goodness [BHD04, BBB08, Byo08, Dam05, EH04, GF06, GF08, Gra03, GN03a, HM02, JH08, Mar02a, MM06, MFPP07, OGM06, PRVHVA09, SN06, Sür06, TH08, TS07, Woo02, XL09, MMPV01]. Goodness-Of-Fit [HM02, Mar02a, BHD04, BBB08, Byo08, Dam05, GF06, GF08, JH08, MFPP07, OGM06, PRVHVA09, SN06, Sür06, TH08, TS07, XL09, MMPV01]. GOS [Bie08]. Govindarajulu [Ano04a]. Grade [Das07]. graded [WCT00]. Granger [CY07]. Graphical [KC04, LMM01, Mal09, SL08b]. graphics [Coo00]. Graphs [EGS05]. Green [BDF06]. Grey [GD06]. Gross [SVC07]. Group [CZR02, CZ03, EA05, FSZ07, Hsu05, JZ01, MT05, PP07, Sri02, SS05c, TBM03]. Group-Testing [TBM03]. Grouped [Bha01, GBS05, GN03a, KH02, Tar05, Wan09a, ZGS03, CC88]. Groups [Sri02, SS05c, CAC00]. Growing [Myn06]. Growth [BR07, BBB08, DATM05, MK05, Pan02, SO06, Vas08, BTH00, KOF02]. Guarantee [MKvdMT05, Ruk09]. Guaranteed [CZA08, RW02, VG06]. Guesstimation [Aga09]. Gumbel [Güv03, AA03, Che01, PRVHVA09]. Gupta [Pan05].

Haar [TH06]. Haar-Wavelet [TH06]. Hadamard [DLT00, EK04a]. Hadi [NK05]. Haenszel [ZFY00]. Half [BA05a, CA08, Eid09, Pew02, Pew04, Wan08, WGP08]. Half- [WGP08]. Half-Logistic [BA05a, Wan08]. Half-Normal [CA08, Eid09, Pew02, Pew04, WGP08]. Handling [Car09, Cha04b]. Haplotype [LBZ09]. Hard [Jac09]. Hartley [CDB06]. Hassairi [Muk06]. Hastings [GT03a]. Having [MKvdMT05, NM05, NM09]. Hazard [AAS01, BP09a, BC08a, GN01, GGS04, JO05, KWW04, KG04, LOS07, Li02a, LL08, LWB03, Raq04, ZQ07]. Hazards [CS09b, DE09, DWB01, OJ06, OJ08, SS07a, sS08a, SSF04, ZZW07, DM00b]. Health [SA04]. Health-Related [SA04]. Healthcare [Car09]. Heavy [Gal02, HT06, HT07, WWY08]. Heavy-Tailed [Gal02, HT06, HT07]. Hellinger [CC00]. Heterogeneity [Ayi09, LL03, LD00]. Heterogeneous [Fok08, HLW08, Hug04]. Heteroscedastic [BA07, CD03, CCC09, DYKG09, Dut07, NW09, Pou07, SM09b, Yan00a]. Heteroscedasticity [Leu02, LW03, LW04, LW07a, TZ01, XMX08]. Heteroskedastic [MOS04, Usa02, BCV02]. Heteroskedasticity [BB04, CNG03, CNSV07, CNSV08, HHL03]. Heteroskedasticity-Consistent [CNG03]. Hidden [BV05a, FHL03, MV00]. Hierarchical [BMW08, GDSPSVP08, MG09, KOF02]. High [BGS07, Pat01, Pav03, YA09]. High-Dimensional [BGS07, Pav03]. Higher [IA09]. Higher-Order [IA09]. Highly [DLN08, TTB08, WDRB00]. Hill [Gir04]. Hinde [KM08a]. Hitting [Sil04]. HIV [TY00]. Hodges [DP06]. Hoeffding [LL05b]. Hölder [Ham08]. Hollander [Spu05a]. Holm [Mal09]. Homogeneity [CY09, Fie07, GVGP06, Gur07, LWW03, NGYZ09, SB03, TS08a, TM03, WW07, WW03, JS00]. Homogeneous [BJM04, Bou09, MPT04, Pap04, VT04]. Homogenetic [sS02a]. Homoscedasticity [LL02]. Horizontal [KKMS07]. Horvitz [CDB06]. Hotelling [And09, FM06, JR05, Kak07, MS00, SBJ^{+}07, WSB09]. Housing [ED08]. Huang [PTY09]. Hybrid [GM02, JM09, KP09b, LNC09, LN05]. Hyperbolic [KF07, Kra05, Vau02]. Hypercube [But05]. Hypercubes [CS08b]. Hypercubic [Hud07, HAS00]. Hypergeometric [CQ05, LX09, LBSM01, LR08, RC03, RACSSCJ03, RACSSCOJ07]. Hypersphere [Fig07]. Hypotheses [DRB07, HHL07, HMM03, HS08, MPP08, NM07]. Hypothesis [Arn01b, Art08, CMNP07, EE09, GBG06, GVMS02, Jac08, Kim06b, LSW05, MB07a, OST09, TF04, Ton05, VJ07, WC08, DC00]. Hypothesized [Muk05a].

I.I.D. [Ram06, Ram08]. Identical [BK09]. Identically [PR08a]. Identification [BQ05, CCGPO03, HRVW07, IH08, KLBW07, Pan02, SP03a, SA03, SSA07, SA08, US03, HR00]. Identify [CSP^{+}09]. Identifying [Jan05a, Kra08, MBPP02, MCY08]. Identity [ACS01a]. Ignorability [Nan09]. II [AE08, Asg06, BA05a, BK05a, CAB02, Cra02, DRB07, EM01, JC08, KLK00, LC00, LWB06, LNC09, MAS05, SB06, TIS01, Tin09, VTN09, Wan08, WL03]. III [Has08, Mok05]. Ill [Lou08, LCL01]. ill-behaved [LCL01]. Ill-Posed [Lou08]. Illustration [CDL01, JPB09, Li01, SS03a, BL00b]. Images [Pat01]. Imaging [KMS09]. Imbalance [Nor03]. Immigration [Rah08]. Immune [MD03]. Impact [Ayi09, HS09, GK00]. Imperfect [AO09, CSWM09, Hsu05]. Implications [NS06, CM00b]. Implicit [HMK05, Muk06]. Implied [Sho02]. Importance [DWB01, RSS03]. Imprecise [AV05]. Improbability [Blo02]. Improve [SP08a]. Improved [BCV01, CY09, CNG03, FSZ07, GHL07, HL02a, Kha00, KBC02, KS04a, Lu00, PK02, Pew04, SW09a, Spu02, Spu05a, UOFC08, Ahm00, Aos00, ST00]. Improvement [KC07, NWTI03, PS01a, RZ02, SG07, YCL09, Zhe03]. Improvements [MT06]. Improving [Eid05, CC02]. impulse [Cho00]. Imputation [CS09b, Dem09, LR09]. Incapability [PL09]. Include [JLBB07]. Inclusion [OG08]. Incompatibility [ACS01b]. Incomplete [CQ05, DD04, PGP03, PM02, TTX01, WW07]. Inconsistent [SL07]. Incorporating [ST01c]. Increasing [Cha09, MLY02]. incubation [TY00]. Independence [CL09, DV05, Deh07, GMKFH08, Güv03, JHL07, Que09, SJ00]. Independent [ABN07, Bol07, Ery05, HPR07, Kak07, KK07a, Las08, MXX08, PR08a, ST05a, PG00]. Index [CGR09, CC09, GPM05, HL09, Mar05, PL09, PX03, SO05, SS03a, SS05b, Gui00, PL00]. Indicator [PGF03]. Indices [Lin05, SC05]. Individual [CMY01, TF09, ZS02]. Individual-Level [TF09]. individuals [KOF02]. Induce [MF03]. Induced [Fab01]. Induct [PD07]. Inductive [dC02]. Industry [JZ01]. Inefficiency [Gau06]. Inequalities [GO00, HMN03, LJ06, MJS09, Rao03, Raq07]. Inequality [AO07, Fio08, iK02, LL05b, Lin04, LWZ07, She05a, XW08a, Xu09]. infection [TY00]. Inference [Agg01, AO09, AND04, Ayi09, BA05a, BZX09, Bas09, BV05a, BMS01, CB06, CSM08, Che08, CKL07, CCGPO03, Cra01, CNSV07, DW07, GZ05, GTB07, Güv05, Jah04, KSW09, LBAVM07, Lia08a, LWB06, LNC09, LH01, MR09, MS05, MXX08, Nan09, NN09, Pew02, Pew04, PS05, RACSSCJ05, SB06, Sch05, Smi00b, SP08b, SS02b, Tsi01a, Tsi01c, Tsi08, Tsi09, Tso05, TS08b, WGP08, XMX08, YW07, YZ01, ZQ06, ZQ07, dC02, KM02, MM00, Ng00, WJ00, CNSV08]. Inferences [HS09, LH01, OR07, YW08, Zha07b]. Inferential [GK04a]. Infinite [Ham06, UT01, Yeh09]. Infinitely [Vin07]. Infinity [Blo07]. Inflated [GGT05, Rod03, Rod06, XD09]. Inflexions [AGH08]. Infliximab [RCBM09]. Influence [ABN07, Bar03, DGRLS03, EAN00, EGMPMRPM05, GHMGMPP06, Jun02, LAR07, LL01, MB07b, MLB09, NA06, OJ06, PGPG00, RS01, Sil06, TZY02, TWW01, BL00b, FC00, WC00]. Influential [KH00, LMM01, Pan02, SK01]. Information [AE08, AF08, AO07, ACS01b, AV05, BA05b, BYP07, Bra03, BOR02, CK02, Das07, DD04, Dup08, EM01, Gav08, HM04, KS08, Mar08a, MOS04, NK08, NWTI03, PC08, PF05, PK06, PRVHVA09, SG07, SV08a, SK09b, Smi07, WSPP08, YNI05, ZF03, ZP04, AO00, DW00, DG00, Lia00]. Informative [BV05b, QWSL09]. Informatively [LBZ09]. Informativeness [Pav03]. INID [CHB09]. Innovation [SLZ08]. Innovations [AT01b, AT01a, HT07]. Inspection [Lia09b, Rau09]. Institutionalization [CJ04]. Insurance [KN01a]. Integer [Zho01]. Integer-Valued [Zho01]. Integral [Jan05b, Koi06, Lee05b, Muk03, NC05, NC02]. Integrals [FS05]. Integrated [HN03, JW06, SL08b, Shi08a, Shi08b, ZL06]. Integration [GM02, She05a]. Intensity [CHYL08, LN05]. Inter [Cha07, LWB03, SM08b]. Inter- [Cha07]. Inter-record [LWB03]. Inter-Variable [SM08b]. Interaction [CB07, EBF09, KKSA07, Mon01]. Interactions [YL06a]. Intercept [GV05, Kha08]. interclass [BA00]. Interim [MT05]. Interlitter [HR08]. Internal [CKM07, CM00b]. interpolation [PRW00]. Interval [Agg01, AR09, Asg06, BK05a, Blo08, BM01, CZA08, DF08, GS04b, GH06, Kan02, LS09, LWB04, MS09, QT02, SAS04, SB08, SB09, SSM09, SC09b, TN08, TC09, Wan08, Wil05b, Wil07, WCL08, XM08, ZYSZ08, Zum07]. Interval-Censored [DF08, SC09b, TC09, ZYSZ08, Zum07]. Intervals [BBK05, Bro09, BH05, Bur08, BK01, BK05b, CS08b, CGT02, Dub06, EE09, Fab09, FK06, Fra04, Fro03, GYX06, KS07, LS02, LL05a, LM07, Li08c, Mac02, NP08, PH02, PB04, RAK08, RYS09, RW02, RT06, SHC02, SCL09, Swi09, Zha06, DM00a, FTV00, Gui00, Lu00, QHB00]. Intervened [Sco06]. Intraclass [BH05, Bur08, Men01, Bha00, BH01, Gön00]. Intrinsic [KK02b, KKL06]. Invariance [Jav09a, Smi07]. Invariant [LRG08, RL09, Sun02, NO01]. Invariants [Muk05a]. Inverse [Ami03, AO01, CSM08, CK02, CK04a, GAMG09, GMO05, GN07, KKM07, Kaw04, KO07, LD08, Lou08, MXZ08, MPP07, Min03, MT01, MT02, MW09, NA09, PJ06, Raq04, Sch02, sS03b, SY08, TTX01, TM03, TM07, Wan07b, WFC01]. Inverse-Gaussian [GMO05]. Inverse-Probability-Weighted [sS03b, SY08]. Inversion [Ma06]. Invertible [SW09a]. investigation [HR00]. Involving [HLS09, JH01, PJH09, TC09]. IPPS [MPG08]. Irregularly [Hat09]. ISE [WYZ05]. Issue [Shi08a]. Issues [HML04, Rya00]. Item [SA04, Zhe00]. Items [Cha04b]. Iterated [GP03]. IV [Bra03].

Jackknife [GPM05, RT04, YC03]. Jackknifing [KKP03, YTN07]. January [Shi08a]. Jeffrey [AE03b]. Jeffreys [MM00]. Jockeying [Sta09]. Johnson [DZK04]. Joint [Cho01, MOS04, NÚF05, RCBM09, SC09b, QTR00]. Jonckheere [LZ07]. Judge [NK03]. Jump [RT09]. Jump-Diffusion [RT09]. Jumps [Bar01]. Juvenile [LMR09].

Kalman [DATM05]. Kaplan [BCPK05, RZ02, Ros09, sS02a]. Kappa [YC09, ZC09]. Karhunen [GF08]. Kelly [WW02]. Kendall [LZ07]. Kernel [BBB09, BF02, BGNR01, BS06b, DGZ01, DY08, Eid05, Eid09, KKP03, LOS07, Li09c, OST09, PS07b, Sai01, WCL08, DJ00]. Kernel-Based [BBB09, DGZ01]. Kidney [PD07]. Kiefer [AO07]. Kiefer-Type [AO07]. Kind [CF01, HZ00]. Kinds [YZ07]. Knowing [Tso09a]. Known [BYP07, CK02, HKR02, HK05, Tak01, WL05, YL06b, ZLY05]. Kolmogorov [KL02]. Kotz [Ars09, HN07]. Kotz-Type [Ars09, HN07]. Koziol [BDF06]. Ks [App02]. Kshirsagar [iK02, Dir06, QN08]. Kshirsagar-Type [QN08]. Kullback [NWTI03, PRVHVA09]. Kulldorff [FWB08]. Kurtosis [DZK04, Fio08, KF07, Mar08b, TPS08].

L [MBS17, Nad09]. Lack [Rao04, Roy04]. LAD [TS08b]. lag [HR00]. Lagrange [Byo08]. Lagrangian [CF01]. lambda [Coe00, Sho07, Spu02, Spu05a, Tar05]. Laplace [KS05b, Mei05, Ree07, Vis09, YZ05]. Large [BBQ02, GTB07, HPR07, Mac08, MP09, MCY09, Pat01, PR04, PR08a, Pew02, TF04, WYY00, Zie06]. Large-Sample [GTB07, Pew02]. Largest [FJ09]. Latent [CBKP04, CK04b, CCGPO03, EA05, Fok08, LQ05, MM06, QWSL09, SL08a]. Latin [Bor03, But05]. Lattices [But06]. Law [GYX06]. Laws [GP03, LR01, Pak04, RS04, Sim08]. Layout [Spu05b]. LDV [MXZ08]. Leaks [GKB06]. Least [AA09, CHL01, DI08, DM08b, HdMS01, Hod09, HL02b, KG08, SC09a, TT07, TB07, Usa02, YSH05, Mil00]. Least-Absolutes [TB07]. Least-Squares [TT07]. Left [Fio08, Hwa05, JIP08, OST09, RCL08, sS05a, She05b, She07, SY08]. Left-Truncated [JIP08, RCL08, She05b, She07]. Left-Truncation [Hwa05, SY08]. Legitimate [Tso09a]. Lehmann [DP06, Spu02, Spu05a, YCW^{+}08]. Leibler [NWTI03, PRVHVA09]. Lemma [Adc07]. Length [Fra04, KO05a, Pak04, SA01, SC03, dUÁRC06]. Length-Biased [SA01, dUÁRC06]. Length-Biasing [Pak04]. Leonardo [Nad09]. Lepage [Neu00]. less [Bar00]. Letter [AT05, Ano04b, Ano05b, Ano05c, Cho05, GGVM01, GTB08, GN03b, Jan01]. Level [Asp06, CW01, Jac08, Jac09, OS05, Pat01, TF09, YLZ09]. Levels [Hsu05, NM05, NM09, SAAAS07, Shu00]. Leveraged [CNSV07, CNSV08]. Life [AGAAG02, Ahm04, AB05b, Bie07a, Cha04b, FBH03, GK04b, HA01, HML04, Kan02, KN01a, LCF06, Lia08c, LB05, NBA06, PTY09, PN08, RSG07, Sad08a, SA04, ST07, SSM09, VLLS06, YKC06, ZS06, ZQ06, Zhe03, Chi00, KLK00, PPL00, SN00, SS00b]. Life-Times [Kan02]. Lifetime [AWJ05, CA08, Kem04, Lia06, LNP02, MKvdMT05, MB07c, NSMP03, SG08, SBT02, YCW^{+}08]. Lifetimes [BZX09, LNC09]. Lightly [MS07c]. Like [Bak06, WS05]. Likelihood [AGAAG02, AT05, AGHMP05, Blu08, CCD02, CB06, CC09, CM01, CD08, CCC09, DSC01, DF08, DM04, FW07, FFCN05, Fig07, FWB08, HMM03, HLW08, JW09, Jun02, KZ04, KR03, LBAVM07, Lag04, LX08, LZZ08a, LCC08, Ma06, Men01, NK08, PP07, Pew04, QT02, RK01, RYS09, SL02, Sco07, SL08b, Sha02a, SW03, SS02b, Tak01, TM03, Tza08, Tza09, UY04, VG06, hWlF03, Wan06b, WV06, XWW07, XKWW08, XSD^{+}08, XLZL07, XW08a, YZZ08, YSZ09, YYL09, Zha01, ZK08, ZQ06, ZQ07, ZM09b, ZL05, dL07, BH01, Hür00, JS00, Man00a, MD00, QHB00]. Likelihood-Based [LBAVM07, RYS09]. Likelihoods [HHL07, Shi01]. Limit [ALT04, Deb08, Dre06, GH04, Hwa05, JS09, JC08, KL05, Kim07, Led06, LC05b, MLY02, MYS07, PR08b, sS03b, WJ06, ZZ01]. Limited [JPB09]. Limiting [Jav09b, LR00]. Limits [Han08a, Men02, RAK08, YIS04, FL00, HCL00]. Line [And09, BA06, Eid05, Eid06, Eid09, Mud03, Zha01, DR00]. Linear [ADS02, Ahm08, Akd01, AE03a, Aos01, AFdBP05, Bar05, BZ08, CV08, CS08a, CZ03, CWS01, CC09, dHCC01, CCC06, CBB06, CS06b, CKK04, DH06, DGRLS03, EA05, FCL06, Gal02, GHMGMPP06, GdT05, GS04a, Gen06, GD05, Güv06, Hod09, HS08, HWL09, HJ01, IWC05, IT05, IP06, IPS08, JBS09, JSV01, JK04b, KK02a, KM05, KS08, Kit02, KR02, Koy06, KS04b, LD08, LL07, Leu02, Li08c, LX08, LWB03, Liu04, LZZ08a, LCC08, LNP02, LWZ07, Mar07, MPP08, MYS07, MJS09, MLB09, NP01, NP08, NM05, NM09, NWTI03, NM07, NAS08, OS09b, Özk09, PH02, PS07b, PK00b, Rud02, SP02, SK09a, SO06, SW09a, ST05b, SLZ08, SM09b, SSJ02, Spi07, Stȩ07, Tan03, TH06, TT07, TF04, TB07, Tri07a, Tri07b, Tri08, UOFC08, WZ01, WJB01, WWJ09, WZ07, WD04]. Linear [WFZ08, XGZ09, XMX08, XY07b, XY07a, XW08a, Xu09, YYK07, Yan07, YYAZ01, YX08, YZZ08, YC02, YC03, YSH05, Yu07, YYL09, Zew08, Zha07b, ZQ07, ZZW07, ZL05, ZY07, Zho01, dBSSvV01, AC00, BP00, CH00, CM00b, CBsK00, Gön00, KM02, MBS17, Mil00, NO01, PN00, PGPG00, QTR00, Shi00, SK00, SJ00, SYCZ02, WDRB00, YD00]. Linearity [JH08]. Lines [Sta09]. LINEX [ADP05, HKW09, SPS07, Sol00, ST01c, ZM09a]. link [Kim02]. Listed [XL06]. Liu [AWA05, ASW06, KS01, Liu03b, Liu04, ÖK07b, TU06, YX08, YCL09]. Liu-Type [Liu03b, Liu04, YX08]. Lives [HJ06]. Loaded [MS07c]. Local [DFP04, FC00, FFVF01, FFVFVF03, Hwa02, KC04, LAR07, Lag04, Led08, LM04, MLB09, PS07b, RS03, RS01, SG04, SO06, TZY02, TWW01, VG06, ZZW07, ZW08, DJ00, Mar00]. Localization [Bro07]. Localized [ED08]. Locally [Mar02b, WH08]. Location [AS07a, AG03, CL01, CP09, CSV04, CM01, De 05, EM01, GHL07, Gra03, HS05, JC08, KZ04, Kra05, LJ06, Leu02, Lia03, LSD07, MS06, PK02, SASAS06, Tan06, TS09, Ter01, TK01, Wu06, AO00, CR00, CC02, DC00, DM00b, Hür00, LI00, Sai00, SM00, Ter00]. location-dispersion [LI00]. Location-Scale [CM01, EM01, JC08, LSD07, TK01, CC02]. Loci [WCL08, LZ00]. Loève [GF08]. Log [DH06, FZ05, FZ06, GAMG09, HH09, Jeo03, KR02, LD08, LWB04, LWB06, LNP02, Rud02, SBT02, Shi01, SBJ^{+}07, Tza08, Tza09, WHZ03, CH00, SJ00]. Log-Eig [SBT02]. Log-Exponential [GAMG09]. Log-Gamma [LWB04, LWB06]. Log-Likelihood [Tza08, Tza09]. Log-Likelihoods [Shi01]. Log-Linear [DH06, KR02, LD08, Rud02, CH00, SJ00]. Log-Non-Linear [LNP02]. Log-Odds [WHZ03]. Log-Rank [Jeo03]. Log-Ratio [SBJ^{+}07]. Logarithm [GP03]. Logarithmic [KO05b, KO07]. Logistic [AKKM09, AAG03, Asg06, Ayi09, BA05a, CGA06, DP06, EH04, FGM03, IH08, Jeo01, LC08, MPP08, MM05, Sha02a, ST08, TTB08, VG06, Wan08, ZLY05, CL00, Rya00]. Loglinear [CBKP04, KC04]. Lognormal [Fly05, Kim07, LS02, Mei08, RJ04, TH08, Tom03, UP03]. Lognormally [BZX09]. Logrank [Dup08]. Lomax [GAAA07]. Long [AH04a, BM08b, CP00, FMAA08, GO05, HRB01, HL02b, KV05, SHRA06, TOS09, Tau09]. Long-Memory [TOS09]. Long-Range [FMAA08, HRB01, KV05]. Long-Range-Dependent [BM08b]. Long-Run [AH04a]. Long-Tailed [GO05]. Longest [MKvdMT05]. Longitudinal [AF09, Dem09, Kim06b, KR02, LW07a, QL09, SS02b, XW08a, XL09, YWK01, Shi00, WYY00]. Look [EK04a, Pan05, Mil00, PB00]. Lorenz [SP02]. Loss [Agg06, ADP05, AWA05, Ami03, Ano07, ANMSL06, BA07, Das07, FZ05, FZ06, GHL07, GKK07, Gru04, HKW09, JP08, MS07a, MD04, Mic06, NMS09, PM02, SP02, SPS07, Sol08, ST01c, ST07, WL01a, WL01b, ZM09a, Sai00]. Low [Woo02, YA09]. Lower [Gor09, GR09, QN08, QC09, Spu02, WQC07, YL06a]. LP [PR08b]. LR [TF04]. LS [MYS07]. LSS [Sam09]. Lupus [AF09].

M [AM07, Dir06, MBS17, Nad09, AM07]. Machine [SHC02]. Magnetic [KMS09]. Mahalanobis [Kim00]. Main [CJ02, Mon00]. Maintenance [AH04a, CN09, KKP07]. Management [XL06]. Manifolds [CV08]. Mann [Ozt02, TS09]. MANOVA [YW08]. Mantel [ZFY00]. Manufacturing [HZSX06, PL09]. Many [DSW07, VH01, VHS00]. Many-One [VH01, VHS00]. MAP [CHC07]. Mapping [LZ00, WCL08]. Marginal [ACS01b, BL09, CY09, DK02, FMAA08, SB09, SGC09, WW07, Zha08]. Marginals [BM08a, KLBW07]. Margins [NQMRLÚF01]. Mark [HC02, SM06a]. Marked [SM06a]. Markers [SHRA06]. Market [CGT03]. Marketing [Gia03]. Markov [ANN08, AH04b, Ani04, AE03b, BBL04, BJM04, BV05a, BL04, CT04, CS07b, CKL08, CG07b, DD04, FHL03, GL04, GH04, GMW04, Har04, Huz04, IA09, IP06, Jaś07, KL04, Led06, Lee05a, LOS05, LL05b, LVVA04, Mac08, MPT04, MV00, Pap04, Pon07, RCPO04, SL02, SS04, Sil04, SP08b, TU06, VT04, XL06, ZY08]. Markov-Modulated [ZY08]. Markovian [Led07, LNP01]. Marshall [BGAH09, GAAA07, Li08b, Yeh04]. Martingale [Vex08, WJ00]. Martingale-based [WJ00]. Martingales [LL05b, SW04]. Masked [MB07c, Nag02]. Masking [Bas09, WCS03]. Mass [DFP04, PJH09]. Matched [Fri03, SB08, SB09]. Matched-Pair [SB09]. Matched-Paired [SB08]. Matched-Pairs [Fri03]. Matching [MN08]. Mathematical [Agg01]. Matrices [EM01, EK04a, GBS05, DLT00, KOF02]. Matrix [AE03a, BA05b, Bra03, CBB06, CS08c, CNG03, DMGFRQG07, Hod09, IPS08, LX09, LRG08, SMGVMD02, Tri08, WSPP08, WSB09, YZ07, YLX09, AO00, AC00, JS00, MBS17]. Matrix-Variate [Tri08]. Mauricio [Nad09]. Maverick [NK03]. Max [Has08]. Max-Domain [Has08]. Maxima [Dol08, IR07, KP08, RAK08]. Maximal [CSV04, GEGMMI09, Muk05a]. Maximization [Kaw04]. Maximum [AGAAG02, AT05, AGHMP05, BP09b, BS05, Car02, CB06, CHC07, CCC09, DSC01, DF08, DM04, EK04b, FV07, FV09, GL04, IT03, JW09, Kak07, KR03, Ma06, Man00a, MD04, PP07, RC03, Rod08, SL02, Sha02a, SW03, SL04, UY04, VG06, hWlF03, WSPP09, WSB09, XWW07, XKWW08, XW08a, YL06a, YZZ08, BH01, Hür00, MD00, SSW00]. Maximum-Likelihood [UY04]. Maxwell [BR05]. May [GS05]. MDL [DY08]. Mean [Agg06, AE03a, Aos01, Ars09, AB05b, BK03, Bie07c, BW03, BK01, CL08a, CK02, CZ07, DG08, DY08, FJ09, Gau06, GS08, GMPG04, GMM01, GK04b, HT06, Iba03, KC08, Kim07, KMPS09, KK09, Laj06, LCF06, Lia06, MKvdMT05, MKS01, NBA06, PH02, PN08, QT02, QN08, Rah08, RMAM09, Sad08a, SKA01, SG06, SL07, SV08a, SV09, Swi09, SV08b, Tin01, Ton05, VPS08, WF05, WL01a, Wil05b, Wil07, Woo02, YZ07, YLX09, YW08, ZM09a, ZC02, ZK08, ZQ06, Cho00, FY00, Kha00, PPL00, RR02, SN00, SS00b]. Mean-Correction [Gau06]. Mean-Squared-Error [SV08b]. Mean-Variance [Bie07c]. Means [CS06a, CHW08, CSM08, Eks02, KKL06, KM08b, KSG03, KKT05, Lin04, MW09, Pan05, Raq04, SST^{+}09, TM03, TM07, Tso03, ZS02]. Measure [DYKG09, Fro03, GVMS07, HS05, KL05, Mar08b, PF05, SM01, VTKB04, Wen02, Yan07, YC05, EAN00]. measured [CFB00]. Measurement [HL03, Li01, MS07a, She09, SK09b, YWK01, GCB00, KG00, Sca02, YD00]. Measurements [CS04a, Hat09, IT09, SSL08, dBSSvV01]. Measures [AFM08, AND04, BYP07, CD07, CS08b, Fab08, GHMGMPP06, Gup04a, Jan02, KLBW07, LJ06, LR08, NQMRLÚF01, NÚF05, Par01, PvdL07, SG08, SL09a, SL04, WCS03, BTH00, QTR00, WC00]. Measuring [Blo02, De 08, Fio08, JR07, MD04, Sch05]. Mechanical [Pul02]. Mechanisms [Tso09a]. Median [AH01, Dub06, GSYB03, GSA08, KG08, OG08, SW03, Ter01, YSZ09, Ter00]. Median-Ranked [OG08]. Medians [SP04]. Meier [BCPK05, RZ02, Ros09, sS02a]. Mellin [NC02, NC05]. Memory [HL02b, Rao04, Roy04, TOS09, Tau09, CP00]. Mendes [Nad09]. Meta [ES09, Jac08, SHRA06, LZ00]. Meta-Analysis [ES09, SHRA06, LZ00]. Meta-Regression [Jac08]. Method [AZ02, And09, CCGB03, DCN07, Eid09, Ger09, GDCOCO06, Han08a, Hin05, JW05, JR07, KKSA07, Kim07, LHP06, Mei08, SL08b, Sha02b, Smi05, lTTNlT08, Tso05, UY04, Vas08, VH04, WLQ09, WW02, YYK07, YSZ09, ZC09, Zhe03, BRM00, Coo00, Lu00]. Methodology [DGRQ01, GPM05, SLB08, Sho02, Sho07, BL00a]. Methods [CHL01, CKM07, FWB08, GL04, GK04a, HI09, HdMS01, Hua08, JZ01, Li02a, MS07b, Pat01, SP05, SS05c, SW02, TF09, YH06, CC00, HZ00, CNSV08, MBS17, Shi08a, Ano07]. Methotrexate [RCBM09]. Metric [JWSP04]. Metropolis [GT03a, WW05]. MEWMA [JT09]. Microarray [Dat03, DLN08, Gup06]. Microbial [JLBB07]. Midrange [ABN04]. Might [GS05]. Mimicking [Sam03]. MIMO [JT09]. Mini [Ano04c]. Mini-conference [Ano04c]. Minification [BJ03]. Minima [Dol08, RAK08, Yeh09]. Minimal [BQ05]. Minimally [FW04]. Minimax [Bel02, CHL01, Dar09, Hud07, JRM09, JP08, Koy06, MLY02, SV08b, TLF04, Zho01, HAS00]. Minimization [SL05]. Minimum [Blo07, Car02, CW01, Che08, CK02, Din01, EK04b, Fan09, Güv06, Jac09, Jan05a, KN01a, SVC07, WZ07, WSPP09, WD04, CC00, DLT00]. Minimum-Distance [Fan09]. Mining [PD07]. MINQUE [YSH05]. Miranda [Nad09]. Mis [Yu07, YD00]. Mis-Specification [Yu07]. mis-specifications [YD00]. Misalignment [YGKD09]. Misclassified [CS07b, RYS09]. Mises [WF05, AM07, BCV01, Sco00]. Misleading [Blu08, KMPS09]. Mismeasured [YH06]. Missing [BZ08, BD03, BDF06, Cha04b, DM04, Dup08, GMPG04, HI09, HS09, KC08, KDCW09, LBZ09, LTHK01, Nie05, RTW06, SK01, ST05b, WV06, Wan07b, Zha09]. Misspecification [Lia09a]. Mixed [CZ03, Dix03, EH04, FGM03, Fok08, Güv05, Koy06, KP08, LW07a, MLY02, Mal09, MM06, MLB09, RMvdG06, RCBM09, SL08a, SS02b, WFZ08, XD09, YYAZ01, YCL09, Zew08, dL07, Che00, DC00, Gön00, SS00b]. Mixed-Effects [XD09, YYAZ01]. Mixed-State [SL08a]. Mixing [DAN08, LN09, OSS05, OST09, Rao03, WLQ09]. Mixture [AGH08, ASA02, Arn01b, Ars09, AGHMP05, BC05a, EHT04, EB08, FE06, LL08, MP08, Nas06, NGYZ09, Por08, Pre04, Pre08, Rod08, Sha07, Sol06, Tak01, Wu02, WCL08, JLR00, KCV00, Pre00]. Mixtures [ACS07, BS08, CV08, CCGPO03, FJC07, FV07, GSMGVM08, GO05, Ham06, HS05, HH09, Pak09, Pol05, Tsi02, Cha00]. MLE [QAT05, YYY06]. MLEs [MP09]. Mode [Gav08, OST09]. Model [AT09, AAGM05, AS08, AN07, Ami03, AE03b, Ayi09, BHN05, BZX09, BA07, BBB03, BBL04, Bas09, BZ08, Bha01, BV05b, BDF06, Bru08, CS07a, CZ03, CWS01, Che06, CL06a, CS07b, CB07, CZL08, CC09, CS09b, CKL07, CDKK05, CJ04, DM08a, Deb08, De 08, Dem09, DE09, DM08b, DM04, Eid06, ENL05, EB08, FW04, FJSZ07, FV07, FBH03, GHMGMPP06, Gau06, GDSPSVP08, GR09, GA05, GK04a, GB01, GGT05, Güv05, Han08b, HM04, Hat09, Hod09, HH09, Hud07, HR03, HR08, Iba03, IT05, IP06, Jah05, Jeo01, JO05, Kan07, KH03, Kha05, KS08, KW05, Kna02, Kra08, KG04, LAR07, LL02, LNP01, LL03, LQ05, LMG06, Li08c, LX08, LL08, Lia08c, LN09, Lia07b, LS09, LB05, LZZ08a, LN05, LM04, MXZ08, MLY02, Mal09, Mar07, MM06, McE08, Men01, MK03]. Model [MI08, Mok05, MAS05, MOS04, Nag02, NK05, NE05, NP01, NWTI03, NM07, Nor03, OJ08, Özk09, Pan02, PH02, PB04, PP07, Por08, Pul02, QAT05, QWSL09, RvdGM05, RMvdG06, RG04, Rao01b, RJ04, RCBM09, Rod05, RCPO04, SASAS06, SS07a, SJS09, SO05, SO06, SBT02, SL07, She09, Shi08a, Shi08b, SAT05, Sol06, Sol08, ST01b, SSJ02, ST01c, ST07, SSF04, SGC09, SS02b, SC09b, Tak01, TS08a, TXG03, TZ01, TH06, TT07, TAS01, Tin02, Ton02, TU06, TS08b, UT01, VLLS06, Vin07, hWlF03, WCL04, WV06, Wan09b, WJB01, WWJ09, WZ07, WWY08, Wu02, XP04, XL06, Xu09, YWK01, YH06, YC02, YL06b, YSH05, Yu07, YYL09, ZCL09, ZLH^{+}09, ZZW07, Zha09, ZL05, ZY07, ZY08, Zie06, dlHRB06, CM00b, CBsK00, Gön00, KOF02, Kem01, Kha00, KLK00, KCV00, LI00, Man00a, MM00]. model [MV00, NO01, QTR00, Sep00, SSG00, SMS00, WH00, WC00, WDRB00, YD00, Yan00b, Zhe00]. Model-Based [McE08]. Modeling [Asp06, BL07, CM08, DATM05, DR01, ED08, EA05, JK04a, KD08, KDCW09, LD08, LBZ09, Mak02, NS06, Nie05, PD07, Quo06, Sho02, Sho07, Shu00, YC09, MV00]. Modelling [KR02, KR03, Ree07, WP02, Syk00]. Models [AS07b, AT01b, AT01a, AT05, ANV09, AZ02, APdR01, Ani04, Ano04c, AFdBP05, Art08, ASN03, AGHMP05, ALOS09, BMN06, BN09, BHD04, Bar05, BCV01, BA05b, BA06, BH06, BJM04, Bla09, BV05a, CS08a, CCBB09, CCD02, CD03, CT04, Che05, CBKP04, CK04b, CDL01, CDKK05, CCGB03, CM01, dHCC01, CF05, CCC06, CBB06, CS06b, CU08, CD08, CS08c, CCC09, Cra01, DG08, DY08, DV05, DH06, DGRLS03, DW07, EHT04, EH04, FCL06, FGM03, FL04, FUOC02, FFCN05, Fie07, Fok08, FV02, FRMAA08, FHL03, GBG06, GM02, Gan04, GO06, HK06, Hin09, HRB01, HWL09, Huz04, ITV06, IWC05, JHL07, KM05, Kha08, KCL00, KD08, Koy06, KC04, KS04b, LBAVM07, Lag04, Led06, LLN09, LT02, LL01, LW03, LW04, LW07a, LC08, LWW03, LW07b, LCC08, LSRD04, LSD07, LZZ08b]. Models [Mar05, MPT04, MPP08, MYS07, MG09, MVB01, MM05, MLB09, MPV06, MT01, Muk05a, MOS04, NP08, NW09, NGYZ09, NM05, NM09, OJ06, OZMC02, Par01, Pat07, Pou07, QL09, RSG07, RTT01, RCBM09, RT09, Rud02, Ryn07, SL08a, SP03a, SSL08, SA03, SSA07, ST05b, Sha02b, SM09b, Shi01, SW02, Stȩ07, SP03b, TWW01, TOS09, TD07, TTB08, TPS08, TAS01, TLF04, TB07, Tri07a, Tri07b, Tri08, Tsi08, US03, UP03, UOFC08, Vas08, WZ01, Wan05b, WC07, Wei09, WD04, WFC01, Woo02, WFZ08, WCL08, XWW07, XKWW08, XGZ09, XMX08, XM08, XY07b, XY07a, XLZL07, XW08a, XD09, YYAZ01, YW08, YZZ08, YC03, YCW^{+}08, YSZ09, ZCL09, Zew08, ZGS03, ZLY05, ZL06, Zha07b, Zho01, ZW08, dlCMM03, AC00, BP00, BCV02, Bar00, BTH00, Che00, CH00, CNZ00]. models [CBsK00, CC02, FC00, Gil00, GCB00, Kim02, KM02, LD00, MBS17, MM02, PJ00, PGPG00, Shi00, SJ00, SYCZ02, TY00, VC00, VKT00, YD00]. Models* [MT02]. Modes [BHN05]. Modification [WW02, Zha06, Neu00]. Modifications [AKS06, CY07]. Modified [AT05, APdR01, Cha08, Che06, Dix03, LHQ07, PC08, PK00a, Sha01b, UY04, Wan06a, ZK08]. Modulated [ZY08]. Modulus [BS05]. Molecular [SS03a]. Moment [ACS01a, Gav08, HMN03, KKC06, KK04a, NGYZ09, PJH09, Rao03, Sha01b]. Moment-Based [NGYZ09, PJH09]. Moments [Ahm08, AGH08, ASW06, Bar06, Cra02, DD08, Fly05, Gov01, GN05, Hsi01, Hwa05, Imo03, IT03, JJ08, KO05b, MS03, NM03, NG05, NR04, PS01b, Sil04, Sim08, Wil03, WFZ08, Raq00, ST00]. Monitored [Bro09]. Monitoring [DMH08, GS08, KNA08, LLN09, MCY09, MT05, NA07, NA09, NAS08, PN08, SHLT09, TN08, VTN09, XGZ09, ZC02]. Monotone [BZ08, RvdGM05, WHZ03]. Monotonic [AWJ05, GW01]. Monte [FP01, GS04b, Kim07]. Monte-Carlo [GS04b]. Morgan [MWM03]. Morgenstern [Güv03]. Mortality [RDI01, TF09]. Most [ACP03, Mar02b, VG09a]. Motion [BHH07, Ree07]. Mountains [BP09b]. Mover [SC09b]. Mover-Stayer [SC09b]. Moving [Lag04, SSA07, SAAAS07, US03, Wan05b, MM02]. Moving-Average [US03]. MPL [Pum07]. MSAE [AENT00]. MSE [PS07b]. Mu [Spu05a]. Much [GS05]. Multi [Asp06, CW01, CK04a, CJ04, CL08b, DGZ01, Dut01, EA05, FJJZ05, Fig07, FS04, GR05, JZ01, JS06b, NA07, NA09, Ryn07, Ton05, YLZ09, Man00b, Mon00]. Multi-Attribute [NA07, NA09, GR05]. Multi-Center [FJJZ05, JS06b]. Multi-Dimensional [Ryn07]. Multi-focus [FS04]. Multi-Frequential [Dut01]. Multi-Group [EA05]. Multi-Hypothesis [Ton05]. Multi-Level [Asp06, CW01, YLZ09]. Multi-Sample [Fig07]. Multi-Stage [DGZ01, JZ01, Man00b]. Multi-state [CJ04]. Multi-Updating [CL08b]. Multicenter [FJSZ07]. Multicollinearity [KS04a, MB07b, GK00]. Multidimensional [Cho01, Hin05, Cho00, Zhe00]. Multidimensionality [KC04]. Multilevel [CCG09, Dem09, KD08, KR03, MVB01, MM05]. Multinomial [ACP03, CG04, DJ09, KW05, Por08, TS08a, YC09, Ahm00]. Multiparameter [KBC02]. Multiple [AAGM05, Aos01, BP00, BHN05, BK03, BL07, BW03, CC08, CKF09, CZR02, CP04, CS09b, Chi09, Dem09, FKB09, HLS09, HJ01, IH08, IT05, KMS09, KRB^{+}09, KK02a, Kha05, KM08b, KG02, LMR09, LR09, MPS04, MN08, McG07, NC05, OZMC02, RTW06, Sam09, SP03a, SO05, Shi01, Spu05b, VH01, WC00, WCS03, XLZL07, Zew08, CBsK00, PLL00, Shu00, SK00]. Multiple-Outlier [XLZL07]. Multiplicative [CHYL08, Ma06, Nai02, TZ01, WL08, WLQ09]. Multiplier [Byo08]. Multiply [ZP04]. Multiscale [ON03]. Multistage [RT04]. Multistate [CS04b, CT04, Huz04, TC09, ZLH^{+}09]. Multivariate [And09, AdL09, AFdBP05, ACS01a, BBQ07, Bar03, BZ08, CM09, CK07, CMY01, Cra01, CZ07, DGZ06, Deh07, DF08, DGRLS03, DMH08, Fal03, FJC07, GLW05, GHMGMPP06, GS08, GSMGVM08, Gup04a, HHL03, HK02, JW06, Jun02, KK03, Kim00, KS05b, Koy06, LBAVM07, Li01, Li08b, MBPP02, MS06, Mar08b, MCY09, Min03, MB07b, MOPS08, NM03, NWA02, OJ08, Pak09, PvdL07, Que09, RC03, RLÚF05, SS07a, SA08, SJ04, Sür06, Tak09, TD07, Tom03, TF04, Tri07b, TPP02, TPP03, Tsi01b, VTKB04, VL06, VG09b, Vis09, WWJ09, Wu06, Yeh04, Yeh09, YZZ08, Aos00, Man00b, Ng00, PGPG00, Shi00, VC00]. Multiwavelets [Efr09]. My [Chi00]. Myositis [LMR09].

Nadarajah [dSFM07]. Natural [KBC02, LR08, Pat01]. Near [HH05, ZK08]. Near-Extremes [HH05]. Nearest [JWSP04]. Nearly [Pre08, Man00a]. Nearness [KM05, KS04b]. Necessary [Ruk09]. Negative [EOOPGB09, LBSM01, Ma06, OS09b, SZ06, SP03b, Woo02]. Neighbor [JWSP04, SJS09]. Neighborhood [DSW07]. Nelson [ENL05, Ros09]. Nested [AFdBP05, BKB02, Güv06, KRB05, Li01, LM07, Li08c, PH02, PB04, CH00, ZFY00]. Nesting [CJ02]. Network [CSP^{+}09, BL00a]. Networks [BS04, GVMS07, SP08b]. Neural [BS04, SP08b, BL00a]. Neutral [JLBB07]. Next [Hsi01]. Neyman [RC02]. NINDS [LCL01]. No [Dup08, Shi08a]. Noise [CH09]. Noises [SP08b]. Noisy [Dar09]. nomographs [Alw00]. Non [AT01a, Ano04c, AFdBP05, ANMSL06, BHH07, BJM04, CQ05, CY07, CWG06, CMY01, DYKG09, DGLS03, EOOPGB09, GM08, GW01, Jan02, Kak07, LMR09, LC05b, LNP02, Ma06, Nad05, Nad09, NC05, Nan09, OS09b, Pap04, PR08a, SL09a, SST^{+}09, UY04, VT04, Yan07, YA09, dSFM07, CH00, FFvR00, Gil00, Tan00, WDRB00]. Non-Central [Nad09, dSFM07, DGLS03, Nad05]. non-elliptical [FFvR00]. Non-Gaussian [YA09, Tan00]. Non-homogeneous [BJM04, Pap04, VT04]. non-linear [WDRB00]. Non-Monotonic [GW01]. non-nested [CH00]. Non-Normal [AT01a, CMY01, UY04]. Non-Orthogonality [Jan02]. Non-singular [DGLS03]. Non-Uniform [GM08]. Non-Weighted [SL09a]. non-zero [Gil00]. Non/Semi [Ano04c]. Non/Semi-parametric [Ano04c]. Noncentral [For05, LG03, NC02]. Nonconservative [vdM08]. Nonconstant [JH01]. Nonexponential [sC07]. Nonhomogeneous [RCA02]. Nonidentical [Sad08a]. Nonidentically [Ery05, ST05a]. Nonignorable [KML01]. Nonlinear [CCBB09, CM01, CU08, CS08c, CCC09, DG08, DYKG09, KR08, Lee05a, LW03, LW04, LW07a, LWW03, NW09, Pou07, RTT01, RCBM09, SSL08, SHLT09, SP03b, TWW01, Tan00, VTN09, XWW07, XKWW08, dlCMM03, CC02, HR00]. Nonlinearity [KL02]. Nonnegative [AZ02, Rao01b]. Nonnested [Shi01, WJB01]. Nonnormal [ITY01, ST01a, TIS01, CR00]. Nonnormality [IT05, Oga08, Yan00a]. Nonparametric [AT02, AND04, ALOS09, BBK05, BP09a, BM09, Bar09, Bel02, BR07, CC01, CB07, CHYL08, Chi07, Col03, DF08, DAN08, Eks02, EOS08, FS04, FP09, GMPG04, Gur06, HRVW07, HK02, KG02, LLCNM01, Li08a, Lin01, LC08, LSW05, MD00, MR05, Özt01, Pak09, PSZ07, Pol05, RAK08, RW02, Sak02, SO06, Sha09, She05b, SHLT09, SMS00, TLW08, VG06, VH04, Wan09b, WYZ05, XM08, Zha07b, Zha09, ZW08, ZYSZ08, Zie06, Zie09, dUÁRC06, VHS00, Yan00b]. Nonproportional [Kra08]. Nonrandomized [vdM08]. Nonregular [MvdM01b, DLT00]. Nonresponse [KML01]. nonscalar [AC00]. Nonspherical [CWS01]. Nonstandard [Pol05, PK00a]. Nonstationary [BR07, Eks02, GVW05]. Norm [Blo07, KN01a]. Normal [Adc07, AT01a, ASS05, AN07, AVGQ04, BK05a, BMW08, CHW08, CMY01, CA08, DYKG09, DMGFRQG07, Eid09, EEPL08, ELF02, For05, For09, Gen06, GZ05, GBS05, GSMGVM08, GK04a, GB01, Gup04a, GN05, HN09, JBS09, JPB09, Jun02, Kha05, KK02b, KKL06, Kim06a, Kim09, KKT05, LBAVM07, Lin04, LP05, LNRS07, MFPP07, MFPG07, MLB09, Ohy02, Pan05, Pew02, Pew04, PGTM06, Pou07, Rah08, RCL08, Roy03, SP09, Sho05, SS07b, Tak09, UY04, VJ07, VL06, Wan06a, WWJ09, WL01a, Wil05a, WGP08, Yu07, ZM09a, ZC02, AY00, Aos00, AC00, Coe00, KK00b, Pew00]. Normality [DGR07, Imo03, IWC05, Kak07, LC05b, Oga08, QL09, Sak02, SN06, VG09b]. Normalized [Mur01]. Normalizing [TSSI02]. Normals [Ham06]. NORTA [NA09]. Note [Ano04f, AB05b, CS04a, CL06a, Cra02, CZ07, DP06, Du08, Dut07, FWB08, Fu05, Gen06, GT03a, Jeo01, JJ08, KF07, KKT05, LT08, LZ07, LP05, MRT09, Mud08, Pak09, PK06, PR04, PD07, Sad08b, SO05, SK06, sS02a, SP04, Sil06, SC06, SSJ02, Sun05a, Tri07b, WTZL01, WW03, YNI05, Yan07, YL06a, Yue04, ZCL09, ZC02, BH01, Mon00, Ng00, SSW00]. Notes [QHB00]. Notion [BBQ07]. Novel [AA09, SvDM07]. Nuclear [MG09, SC09a]. Nuisance [ACP03, Mar02b, Muk02, Muk07]. Null [Chi09, GVMS02, MT01, MT02, NC05, TF04, Yan00a]. Number [Ant03, BBK05, Bol07, Cha07, CH09, CG04, Dar09, EK04b, FCL06, HZSX06, HH05, Hsu05, PR04, RS04, RTW06, Rie03, VTKB04, WTZL01, YL06a]. Numbers [CW01]. Numerical [BJM04, Fro07, GS04b, MT06, MPT09, RCX06].

Objective [CS07a, WGP08, ZW00]. Observable [JN03]. Observation [Pan02, SS01]. Observations [Bar05, CS04b, CT04, CMY01, Dar09, DSW07, Fok08, HT06, IP06, Kim09, Las08, MPV06, PM02, RTW06, SPG07, Sun05b]. Observed [BS06a, BW03, DI08, AO00]. Observing [LB07]. Obtain [GVGP06]. Obtained [DGLS03]. Occasion [SP08a]. Occasional [Bar01]. Occupancy [TC09]. Occurrences [CKL08]. Odds [LW07b, WHZ03, Zha08, dlHRB00]. Offspring [GMM01, Rah08]. Often [Blu08]. Ohio [PD07]. old [SS00a]. Olkin [BGAH09, GAAA07, Li08b, Yeh04]. OLS [KM02, Myn06]. OLS-based [KM02]. omitted [LY00]. Omnibus [Mei05]. On-Line [And09, BA06]. On-Target [SAAAS07]. One [Agg06, Bak08b, BH05, CR00, CF05, Dam05, DRB07, DFP04, GN04, Kim09, KSG03, MD07, Muk05a, NM05, Nor03, Pon07, Rao01b, SASAS06, SN06, SG05b, Spu05b, VH01, XM08, YTN07, ZGS03, dL07, DC00, VHS00, Yan00a]. One- [Pon07]. One-Parameter [Agg06, CF05, GN04, Muk05a]. One-Sample [Dam05, SASAS06, SN06, YTN07, dL07, DC00]. One-Sided [DRB07, MD07, ZGS03]. One-Way [BH05, KSG03, Nor03, Rao01b, Spu05b, XM08, Yan00a]. Onefold [Güv06, PB04]. Open [EE09]. Operating [LSW05]. Operation [BGAH09, Rao04]. Operations [Pak04]. Ophthalmology [Kim06b]. Optimal [AH04a, ADM08, Arn01b, Bie08, BH05, BS03, BCK07, Cha00, CL01, CJ02, DFP04, DM08b, FW04, GD05, HH09, HC02, JWSP04, Kat01, Kit02, Kra05, KLHT06, LNP01, LI00, LC00, LRG08, MS07a, MVB01, MM05, MPS08, Muk07, PR08b, RL09, SBM01, WL03, Yue04, ZCL09, Zha07a, Zie09, CS09a, HAS00, Tan00, TH06]. Optimality [BB01, BCK07, Jaś07, LLCNM01, LFRLFG07, LVVA04, SW09b, dlHRB00]. Optimization [DGRQ01, HW06, RK01, RCBM09]. optimizing [MT00]. Optimum [BW03, Che06, DG08, MP08, ZW00]. Option [TPS08]. Orcutt [ST01b]. Order [ADS02, Ahm08, Ahs04, AS07c, BÖ07, BT07, BBQ02, BES04, Bar06, BGAH09, Bar09, BAA07, BK08, BK09, Bie07a, Bie07b, BL09, CGR09, CCGB03, CF05, CS06b, Cra02, CHB09, Dan05, DSW07, FW07, GS04a, Gau06, Gen07, GKFM^{+}08, Gov01, HN09, IA09, Jah05, JBS09, JC08, KO05b, KH03, KK07a, KM08b, KR08, Lag04, LNRS07, MPP07, Mar05, NMPS05, NBA06, NS08, Nas06, NRS07b, Nav08a, PS01b, PS05, PJH09, Pum07, RC07, Ric07, Sad08b, ST05a, SJS09, SM08a, SM09a, SDZ01, She05a, SSM06, ST08, Sul07, TA07, TA08, TM03, TM07, TPP03, UT01, XLZL07, YBCH08, Aos00, BCV02, GO00, Guo00, HA00, Had00, iHA00, HAS00, Hür00, ST00, TPP02]. Order-preserving [She05a]. Ordered [Ber05, CS06a, Cha04a, CKK04, Gen06, Gur06, KKT05, NRS07b, SB03, SJ04, PB00]. ordered- [PB00]. Ordering [BS08, FE06, GN01, KF07, MY07, MOPS08, NS08, NL07, LZJ00, SSQ00, XM00]. Orderings [HLW08, KK07a]. Orders [AA01, Fio08, MGD08, RJ00]. Ordinal [DBA05, Jeo01, KDCW09, LC08, MW01, Que09]. Ordinary [SS06].

Oriented [ÖK07b]. Origin [KG08]. Orthant [MKS01]. Orthogonal [DBA05, OS05, SB06, WH08, CL00, FY00, Pre00]. Orthogonality [Jan02]. Orthogonally [GV05, Mon01]. OSCV [Yi01]. Osteoarthritis [Tin09, ZLH^{+}09]. Other [GBSW09, ÖK07a, SL04, LCL01]. Outcome [App02]. Outcomes [Blo02, BC05b, FKB09, ZW00]. Outdoor [CM08]. Outfluence [HS09]. Outlier [AAG03, BFZ03, CDL01, SB03, XLZL07]. Outliers [BK03, GLW05, IH08, MCY08, SP03a, SP05, ZN03, Kim00]. Output [MOPS08]. Over-estimating [KH03]. Overdispersed [CCC06]. Overdispersion [XL09]. Overhauls [Pul01]. Overlap [CSM08]. Overlapping [MPP07, MvdM01a, Smi00b]. Overshoot [AH04b]. Overview [BCMRPAEG07, Dat03].

P [MBS17]. PA [LCL01]. Pages [Shi08a]. Pain [ZLH^{+}09]. Pair [PM02, SB09]. Pair-Wise [PM02]. Paired [SB08, QTR00, SBM01]. Pairs [Fri03, Pre04, LZ00]. Pairwise [KRB05, SL05, VH01, ZM09b]. Panel [AS07b, CKL07, PSZ07, ST01b, SGC09, Zha07c]. Paper [Mud08]. Parallel [AB05b, FV02, Sad08a, Sta09]. Parallelism [EGS05]. Parameter [Agg06, ADP05, AZ02, Bak08b, BJ03, Bla09, CD03, CP09, CF05, DG08, De 05, Fan06, FZ05, FL04, FUOC02, FFCN05, Fly05, FRMAA08, GK05, GHL07, GYX06, Gau06, GN04, GK04a, HKR02, HJ06, Hsi01, HL02b, Ili03, JW09, Jeo01, JE01, KS05a, KP02, KF06, KH02, LL07, LX08, Lia03, LM06, LWB03, MBB08, MJS09, MD07, Muk02, Muk05a, NK05, NMS09, ÖK07b, ÖK07c, PK02, Ram05, Ram08, RMK08, SM09a, SPS07, SCL09, ST07, SV08b, Tan06, Ter01, Tin02, TH08, Tsi01c, Tza08, Tza09, Vis09, Wan05a, Wan06a, WDRB00, WL03, YYY06, YX08, Yi01, YC02, YZ05, ZCL09, FTV00, KCL00, Lu00, Sai00, Ter00, Tsi00]. Parameters [AGAAG02, AS07a, ASW06, AT09, AT01a, AR09, AKS06, AS08, AJSC08, BD05, CS07a, CAB02, CL01, Che06, dHCC01, DSC01, Dix03, FL04, FL06, Han08a, HK05, KA04, KK07a, Kha08, KBC02, KS05b, KP09b, KK07b, LWB04, Mar02b, MFPP07, MRT09, MS05, Muk07, PP07, RC07, RYS09, RACSSCJ03, RF02, SZ06, SDM08, ST08, SG05c, Tar05, Ter02, Tom03, Ton02, Wan09a, AO00, CC88, DM00a, FL00, HA00, SS00b]. Parametric [BHD04, DM08b, EB08, Fab01, Gur07, LTHK01, RJ04, SL09b, Tso03, TC04, Tso05, Tso09a, UT01, WV06, WL08, YC03, Ano04c, Mar00]. Parametrization [BGAH09, TLF04]. Pareto [Asa04, Bra03, CB06, Coo09, GHL07, Gui00, HA00, HZ00, Hür00, JE01, KK07b, Mar02a, MB07a, PS01b, RC02, Ram06, Ram08, RJ04, SPS07, Sol08, SG05c, TA07, TA08, Tsi03, XD04, Yeh04, Yeh09]. Pareto-Lognormal [RJ04]. Part [HK06]. Partial [BKB02, CCGPO03, Dat03, LSW05, LCC08, NS08, ZL05, Lia00, LC00, SYCZ02]. Partially [AGAAG02, AJ06, BV05a, BV05b, CC09, DI08, HWL09, KLHT06, LX08, YC03, Zha07b]. Particle [FP01]. particles [VKT00]. Particular [dlHRB06]. Partition [SW09b]. Partitioned [TT07]. Partitioning [BGM08, BC09, GH06]. Partly [WZ01, YC02]. Passage [AH04b, VT04]. Past [PN08, Rao01a, Sad08a].

Patchwork [DSPS09]. Path [BHN05, Blo07, CM05, SK06]. Patients [MD03]. Patil [FV07, FV09]. Pattern [QC09]. Patterns [SM06a, SP08a]. PCA [YA09]. Peaks [DGR07, McG07]. Peaks-Over-Threshold [DGR07]. Pearson [RACSSCJ05, SI02]. Penalized [LZZ08b, Ma06]. Penalty [Lou08]. Perfect [AAG03, LNP01, MB07b, SC06]. Performance [AdL09, FHL03, HKW09, Leu02, WC08]. Performances [Wan09b]. Performing [OZMC02, Tso09a]. Period [Han08b, Fu,00]. Periodic [AB05a, BK02, BA05b, BA06, BH06, CS04b, CT04, Dut01, Gau06, Sha09, US03]. Periodogram [Dut01]. Permanent [CHB09]. Permanents [GKFM^{+}08]. Permutation [BS06a, HJ01, LRG08, LCL01, Sal03, SBJ^{+}07]. Persistence [HT07]. Perspective [Bro07]. perturbation [WDRB00]. Perturbations [WCL04]. Perturbed [BS04, VT04, YYW05]. Peto [HN05]. Phantom [PD07]. Pharmaceutical [JZ01]. Pharmacokinetics [RCBM09]. Phase [GSA08, Tin09, VTN09, KNA08, LC00, LHQ07, MCY08, WCT00]. PHD [ZXM07]. Phi [MPP08]. Phi-Divergence [MPP08]. Phylogenetic [Shi01]. Physiological [Ric07]. Picking [GH06]. Piece [Las08, Zha09]. Piece-Wise [Las08, Zha09]. Piecewise [JK04b, NP01]. Pilot [CKM07, Muk05b, CM00b]. Pitman [HI09, HS05, KM05, KS04b, KKT05, MWM03, Rei06, SG05c, Wen02, Yan07]. Pitman-Closeness [HI09, Wen02]. Plan [Ami03, GR05, HL02a]. Plans [ADM08, BCRW02, CJ02, Hua08, MPG08]. Plant [AG08, GA00]. Plant-Capture [AG08, GA00]. Plants [MG09]. Platforms [Sho07]. Play [Cha04a]. Play-the-Winner [Cha04a]. Plot [Güv05, Jac09, WSPP08, WSPP09, Wul05, YLZ09]. Plotting [Mak08]. Point [Asg06, BK05a, BDH08, Das07, DI08, Din01, Ger09, GM08, GVMS02, GVGP06, HSAN05, HT06, Jaś07, Las08, LZZ08a, PC08, PJ06, Ram05, SM06a, SL08b, SP07, SP09, VG06]. Points [Bar02, FCL06, GW01, MN08, RC03, Vex08, XM00]. Pointwise [AD02, Ayd05]. Poisson [BK02, BMS01, Bro09, BK01, BK05b, Cha08, CS06a, CHW08, CK07, DK02, FWB08, GO05, GGT05, HdMS01, JHL07, KK01, KL04, Led06, Nav08b, RYS09, RCA02, Sco06, SJ00, Swi09, Tsi01b, Vin07, Wan07a, Woo02, Yua05]. Poisson-Exponential [Vin07]. Policies [AH04a, CN09]. Policy [KKP07, LNP01]. Pollutant [LMG06]. Pólya [CL08b, MPP07, TPP02]. Polynomial [BBB08, CS09a, FFVF01, FFVFVF03, GS05, KNA08, Led08, SB06, She05a, Usa02, ZCL09, Bar00, DJ00, GT00, Mon00]. Polynomials [DBA05, LX09, Mon00]. Pooled [Kan07]. Popular [TPS08]. Population [Agg06, AG08, Ber05, Cha09, CSWM09, CP09, CZ07, Dub06, Fie07, GZ03, HL02a, KC08, KK09, MKvdMT05, NS04, NMS09, Sam09, SM09a, SG06, SS07b, SV08a, SK09b, SST^{+}09, SV09, SSM06, Tak09, Tso03, VH04, WD04, Wil09, XG09, ZLY05, AY00, GA00, KK00b, WL00]. Populations [BMW08, BC08b, BC09, CSM08, CK04a, GK05, GKFM^{+}08, Hui05, JE01, Kat01, KK02b, Lia09b, LRG08, RL09, SM08a, YIS04, BA00]. Portmanteau [ALT04]. Posed [Lou08]. Position [Mak08]. Positive [De 07, EEPL08, GY05, GPM05, Gor09, Lia08b, Lia00]. Possible [HK06, RLÚF05]. Post [TTX01]. Posterior [FJ09, JP08, Lia09a, MP09, TZ01, Tsi00, dlHRB03, dlHRB00]. Posteriori [MD04]. Poverty [SL09a]. Power [CHC07, Chi09, DV05, EGMPMRPM05, GYX06, GMKFH08, GSMGVM08, JH01, JS06a, Kaw04, KSG03, Lag04, MG09, MW09, Nas06, Nav08a, OB05, RCL08, SAVG07, SMGVMD02, Sha01b, SvDM07, TS08b, VPS08, Vin04, WFC01, vDK03, CNZ00, LR00]. Power-Law [GYX06]. Power-Series [EGMPMRPM05]. Power-Variance [Vin04]. Powerful [Abo05, Mar02b, Mei08, VG09a, Set04]. Powers [Yam07, MMPV01]. Practical [CKM07, Tin09, GT00]. Practices [LB07]. Pragmatic [AF09]. Pre [Akd02, DLN08, Rei06, TTX01]. Pre-Post [TTX01]. Pre-Specified [DLN08]. Pre-Test [Akd02, Rei06]. Precise [CMNP07]. Precision [SP08a]. Predicting [Sha01a]. Prediction [ADP05, AH01, AAG03, BA07, CL08a, DGZ01, Fu05, GD05, IP06, JRM09, JP08, KS07, Kim07, LRG08, MAS05, ÖK07b, PGF03, Pul01, QN08, SHC02, Spi07, Tin01, WSPP09, WD04, Tsi00]. Prediction-Oriented [ÖK07b]. Predictions [KV05, CKB00]. Predictive [AH04a, De 05, GD06, Hsi01, Kha05, Men02, Men07, MB07b, dlHRB03, dlHRB06, Ng00, dlHRB00]. Predictor [Agg06, DGZ06]. Predictors [MM05, Pou07, RL09, Xu09]. Preface [Ano04d, Ano04e, Bai07, CHP05, CL06b, Gup04b, GO08, LP07, NR09, PTE09, PGF07, Won09]. Preferences [Kum01]. Preliminary [CSV04, Had09, HKW09, Kab05, SN06, TOS09]. Presence [BFZ03, CH09, HHL03, KP09b, Mar02b, MB07b, Muk02, Muk07, NK03, Nie05, Pul06, SK09b, SST^{+}09, WWY08]. Present [ACP03]. Preservation [MGD08]. preserving [She05a]. Presmoothed [ENL05, JIP08]. Presmoothing [JSV01]. Prevention [Sha01a]. Preventive [KKP07]. Price [Kum01]. Prices [ED08]. Pricing [FHL03, SS04, TPS08]. Primer [Pul06]. Principal [BÖG09, EGMPMRPM05, KS01, KSS03, LW02, Mar08b, VL06]. Principle [Byo08, Kaw04]. Prior [AV05, AE03b, De 05, KS08, Mic06, MRT09, dlHRB06, DG00, MM00]. Priors [FJ09, Hsi01, KK02b, KKL06, MS05, Tri08, Lia00]. Probabilistic [Bak08a, Lee05a]. Probabilities [AAGM05, Bas09, BHH07, CS07b, CZL08, DRB07, GN07, Hug04, JK04b, Jaś07, KH03, LBSM01, Nan09, OG08, SB09, TC09, VT04, WW07, ZY08]. Probability [ASN03, BBQ02, BF02, Bol07, Bur08, CZA08, DJ09, Fro07, Gav08, GT03c, JS06a, Kol07, KS04b, Lee03, Lin04, LB07, MF03, MS03, Pak04, PR04, PJH09, RC03, RACSSCOJ07, SA09, SBT02, sS03b, SY08, SC02, Vin04, Wan07b, WWY08, Gil00]. Probable [ACP03]. probit [Sep00]. Problem [AA02, Arn01a, BDH08, De 05, Gav08, GVGP06, Kab05, KKM07, KE01, Kit02, LH02, LT04, LT08, MS06, MD07, PR08b, SL08b, SP04, SW09b, TS09, TT03, WL01a, Wu06, Yue04, SM00]. Problems [ANMSL06, CL08a, FS05, HVW07, HW06, KS05a, Kim06b, Lia08a, Lou08, LTHK01, MS07b, PC08, PO06, SC02, TTX01, YGKD09, ZGS03]. Procedure [BC08b, CK05, CH09, DLN08, DD04, FSZ07, Gia03, JM09, LCF06, Mal09, MKvdMT05, SSL08, SV09, ST01c, TS09, WC08, YYAZ01, ZM09b, Aos00, PLL00, WL00]. Procedures [ACP03, AFdBP05, CZR02, CN09, DRB07, KML01, MPT09, NM07, Pan05, Rau09, She07, NO01]. Process [BR09, BR07, Bro09, BL09, CCK02, Che06, CHYL08, Chi07, Cho01, DMH08, DAN08, Fro07, GYX06, GS08, GT03b, GMM01, GD06, Ham08, HN03, HL09, Hug04, Jan05b, KMPS09, Led07, Lee05a, Lin05, Mac08, MCY09, Muk03, NA07, PL09, PX03, Pum07, Rah08, SP09, SC05, SS04, SS05b, SK05, Ton05, WC08, XZ08, YYY06, YYW05, ZYZ06, Guo00, KCV00, RJ00, SYCZ02]. Processes [BJ03, BFZ03, BL04, BLZ08, CWG06, DI08, DM08a, DD04, Fie07, GM08, GH04, Gra03, GMW04, Had09, Har04, HL02b, Huz04, Jaś07, JW06, LL07, LOS05, LL05b, LVVA04, MJS09, MPS08, NA09, Nie05, OSS05, Pon07, Rao03, RCA02, SW09a, SA08, SLZ08, Sil04, SSD01, ZC02, Had00, PPL00, PL00, Rav01]. Product [Ahm08, Hwa05, KLBW07, LCL09, NM03, PS01b, Sha01b, sS03b, TS08a, WJ06, ST00]. Product-Limit [Hwa05, sS03b, WJ06]. Production [Kum01]. Products [Nad05, Nad09, dSFM07]. Professor [Ano04a, Edi09a]. Profile [FFCN05, HHL07, Tza09]. Profiles [KNA08, MK03, NAS08, SHLT09, VTN09]. Profiling [JLBB07, QHB00]. Prognostic [SHRA06]. Programming [NN09]. Progressive [AE08, Agg01, Asg06, BCK07, KK07b, Mar02a, Sol08, Wan08]. Progressively [BA05a, BK05a, CAB02, Cra02, LWB04, LWB06, LNC09, MR09, MAS05, PP07, Sol06, ZP04]. Projection [Bie07b, Bie07c, CLY08, EK04a, Pre00]. Projections [GO06, Mar08a]. Projective [ADM08]. Proof [Kol07]. Properties [AF08, ABN04, ANV09, BÖ07, BCMRPAEG07, BB01, BM01, BLZ08, Bru08, But06, CM00a, Din01, Dol08, EK04a, FDR08, FR08, FJC07, FV02, FV07, GdT05, HLLL08, HC02, JPB09, KN01b, Lee05a, LS02, LN09, LW02, LC05a, LB05, LHQ07, McE08, MGD08, MWM03, Myn06, NSMP03, NMPS05, NS08, NL07, NRS07a, SL02, SG05a, Sha07, Shi08a, Shi08b, Sho02, SM06b, Usa02, Vau02, Wan06b, WYZ05, XKWW08, ZM09a, ZS06, Aos00, WYY00]. Property [HVW07, Jav09a, KM08a, KP07, Rao04, Roy04, WC07]. Proportion [CSWM09, NS04]. Proportional [DE09, DWB01, Hat09, KG04, LL08, LW07b, OJ06, OJ08, SS07a, YL06b, ZZW07, DM00b]. Proportionality [CL06a, HK06]. Proportions [FSZ07, Fro03, PS07a, RDI01, RT06, SB08, TBM03, VH01, XK08, YC05, Ahm00, BLB00]. Proposal [GIT08]. Protection [YWL08]. Pseudo [KR03, SS02b, CH00]. Pseudo-Likelihood [SS02b]. Psoriatic [TC09]. Pts [ST01c]. Pulak [Che06]. Pure [Jan05b]. Putative [CSP^{+}09].

Quadratic [For05, Güv06, HH09, Hud07, Liu01b, LRG08, RL09, Sol00]. Quality [AWJ05, Bak06, Cha04b, CS04b, HL03, HML04, LHP06, PS01a, Rog06, SA04, Wen02, XL06, ZYZ06]. Quality-Adjusted [CS04b]. Quantification [ACS01b]. Quantifying [Ayi09]. Quantile [AY00, DGZ01, DGZ06, DGR07, EOS08, Ham08, KK00b, Li08a, NS09, OSS05, SAS04, SSD01, Zie06, Zie09]. Quantile-Based [NS09]. Quantiles [BM01, GG05a, GBR01, Par09, ZZ09]. Quantitative [Mon01, Tin01, WCL08, YWL08, ZS02, LZ00, Mon00]. Quantities [HH05]. Quantized [MPV06]. Quantizers [DFP04]. Quasi [CD08, Kak07, Li09a, SC09a, SG05b, SDM08, XKWW08, XSD^{+}08, YZZ08, YYL09, Tan00]. Quasi-Binomial [SDM08]. Quasi-Independent [Kak07]. Quasi-Least [SC09a]. Quasi-Likelihood [CD08, XKWW08, XSD^{+}08, YZZ08, YYL09]. quasi-optimal [Tan00]. Quasi-Residuals [Li09a]. Quaternion [LX09]. Questionnaire [ZF03]. Queue [AM07]. Queueing [IM04, Sta09]. Queues [BP09b]. Quick [Abo05].

Random [ABN07, AN07, BK08, BK09, BW03, BH05, CL01, Chi07, CDL01, CKK04, Dau03, Din01, Dod04, Dol08, EGS05, EEPL08, Ery05, FJJZ05, FJSZ07, GBSW09, GdT05, Gen06, GZ05, GV05, HV03, HWZ06, HPR07, HLW08, HJ01, Hui05, HR03, KC05, KC07, KO05a, KK07a, KLBW07, KL04, KK09, Lee05b, Li01, Li08c, Li09c, Lia08c, LH09, LC05a, Liu01a, LWW03, LNRS07, LBSM09, LVNF07, MLY02, Muk03, NM03, NMPS05, NS08, NM05, NM09, Nor03, OJ06, OJ08, PD07, PR08a, PGTM06, Ram05, Ram08, Rao01b, Rie03, Ros02, ST05a, Sam09, Sch02, SHLT09, ST01c, SSM09, Tin01, Tso09a, Wan05a, Wan07b, XM08, YC02, ZM09b, Zha09, Cho00, Man00a, Sep00, SJ00, SSG00]. Random-Effects [GBSW09, XM08]. Randomized [BBB03, BM08b, Cha04a, Dup08, JM09, KW05, LMR09, OS09a, YWL08]. randomness [CYHL00]. Range [BM08b, FMAA08, HRB01, KV05]. Ranges [CD07, SG05b]. Rank [Bak06, BES04, BC05a, Jeo03, Kra05, MPS04, Oga08, Tau09, Ter02, PB00]. Rank-Order [BES04]. Ranked [ASS05, AO09, BF02, CSWM09, Ghi05, NMS08, OG08, Özt01, SAS04, SASAS06, SM08a, SSM06, TK01, ENS00]. Ranked-Set [BF02, Özt01]. Ranking [AO09, KK00a]. Rankings [CSWM09, QS06, SQ04]. Rao [BB01, CDB06, Gru08, Liu01b, MPP01, Pan02, Tu01]. Rasch [CBKP04, HM04, KC04, MM06]. Rate [AAS01, ASN03, BP09a, Bha01, BL04, BC08a, CG07b, CKM07, Coo09, De 07, GL04, GMKFH08, GN01, GK04b, GGS04, Hat09, KWW04, LOS07, Li02b, Li02a, LL08, LWB03, MLY02, MG09, NR04, OST09, QAT05, RYS09, SK09a, WHZ03, hWlF03, XWW07, YZZ08]. Raters [HLS09]. Rates [FE06, GW01, Lia09a, SL02, Wan05a, Fu,00, SYCZ02]. Ratio [AA09, CDL01, CM01, FW07, Fig07, For05, FWB08, GEGMMI09, HLW08, Jun02, KC05, Kim06a, Lag04, LZZ08a, LVVA04, Men01, NGUT00, PGTM06, QT02, SAS04, Sco07, SB09, SST^{+}09, SBJ^{+}07, Syk00, Tak01, TM03, XLZL07, YL06b, Zha01, ZLY05, Zha08, ZK08, dL07, JS00]. Ratios [Blu08, EEPL08, JBS09, JO05, Nad05, Nad09, dSFM07, PG00]. Rayleigh [BRM00, MAS05, Sol00, Sol06]. Reaches [SK05]. Read [GHMGMPP06]. Reasoning [dC02]. Reassembly [PS01a]. Recapture [DLZ09, HC02, Hug04, SC06, ZLY05]. Recaptures [SC06]. Receiver [LSW05]. Reciprocals [SW04]. Reconciling [De 05, MD07]. Record [AA01, ADP05, AF08, ABN07, AA03, Bak08b, BGAH09, BAA07, Bie07c, Cha07, Cha09, HAA06, Hsi01, KAHS09, Raq04, Raq07, ZS06, LWB03, Raq00, RR02]. Records [Cha07, DLb05, Dem07, DD08, IR07, Jah04, JP08, Kli07, LWB03, LBSM09]. Rectangular [DSPS09]. Recurrence [Bar06, Lia07b, PS01b, ST00]. Recurrent [CCK02, LN05, SC09b]. Reduced [BC05a, CGR09, Oga08]. Reduced-Bias [CGR09]. Reduction [AB01, NA07, Sai01, Coo00]. Redundancy [HW06]. Referees [Ano09]. Reference [Tri08]. Reflections [Rao01a]. Refractory [Kov09]. Region [LX08]. Regions [Hud07, WF05, HAS00, KOF02]. Regression [AA09, Akd01, Akd02, AE03a, AT09, AR09, APdR01, AFdBP05, AAG03, BMN06, BK02, BA07, BHD04, Bar03, BCV01, BÖG09, BZ08, Bel02, BV05b, BB04, Bro02, BM08b, CS07a, CS08a, CHL01, CS09a, CWS01, CL06a, Che05, Chi07, CBKP04, CK04b, Col01, Col03, CS06b, CKK04, DG08, DYKG09, DY08, DM08a, Deb08, DE09, DGRLS03, DWB01, DCN07, DAN08, Dra02, EBF09, FCL06, FW04, FR08, FJC07, FFVF01, FFVFVF03, FP09, GG05b, Gil00, GGT05, Güv06, Hat09, HdMS01, Hod09, HJ01, HK02, Imo03, IH08, ITY01, IT05, IPS08, Jac08, JSV01, Jeo01, KS01, KK03, Kan07, KK02a, KG08, KM05, KS05a, Kha05, KS08, Kha08, KP02, Kna02, KK01, KS04a, KG02, KS04b, LAR07, LBAVM07, Laj06, Led08, LMM01, Li08a, LN09, LW02, LW03]. Regression [LC08, LWW03, Liu04, LZZ08a, LSRD04, Lun02, MXZ08, Mak02, MPP08, MYS07, MI08, Mud03, Mud08, Nai02, Nav08b, NP01, NP08, Oga08, Ora06, PH02, PH03, PB04, PY08, Pat07, Pet09, RS03, RS01, Rei06, RF02, Ryn07, Sad08b, SKA01, Sam03, SO05, SHC02, ST05b, SL07, sS08a, SM09b, SHLT09, ST01b, SSJ02, SW02, SGC09, Sun05a, Sun05b, TOS09, TH06, TIS01, TAS01, TLW08, Tsi01a, Tsi01b, Tsi03, Tsi08, TC04, Tso09a, TS08b, UNN05, UOFC08, UT01, Usa02, VG06, VG09a, WCS03, WLQ09, WD04, WFC01, XY07b, YYK07, YX08, YCL09, YC02, YC03, Yu07, YSZ09, Yua05, ZCL09, ZL06, Zha07a, ZZW07, Zha09, dlCMM03, AENT00, BCV02, Bar00, BL00b, Coo00, CKB00, CNZ00, CBsK00, DR00, DW00, DJ00, EAN00, FC00, GT00, Kim02, KM02, Mar00, MM02, Nor00, Rya00]. regression [SK00, SMS00, SYCZ02, Syk00, VC00, WC00, WYY00]. Regressions [Gil07, Liu03a, Tsi09, WYZ05]. Regressors [HHL03, KM02]. Regret [JP08]. Regular [KH03, SP01]. Regularized [AD02, AKKM09]. Regularly [Sim08]. Regulatory [CSP^{+}09]. Related [ABN04, Bra03, GH04, HVW07, HH05, KMS09, LG03, LR09, Roy04, SA04, SW04, SC02, Vin07, Yeh09, ZT09, CC00, VKT00]. Relation [Bar06, Tan03, ST00]. Relations [LBSM01, PS01b]. Relationship [Jon07, LNP02, Min03, YCW^{+}08]. Relationships [LNRS07, NR04, Wil03]. Relative [DYKG09, Guo03, JO05, KS07, MCY09, TLW08, GCB00, MM02]. Release [LNP01]. Relevant [CV08]. Reliabilities [GCLG06]. Reliability [AFM08, BBL04, BR05, BJM04, DR01, EGS05, ENS00, FV02, FV07, Ghi05, GAMG09, GS05, GK04a, GB01, HZSX06, Han08a, IM04, KP09a, Led06, Lia05, Lia08b, MRT09, Mok05, MXX08, MS07c, NS06, NS09, NSMP03, Pul02, Pul06, SL02, SA01, WHZ03, ZS06, ZYL06, Tsi00]. Rely [GS05]. Remark [Mud08]. Remarks [Bla09, BOR02]. REML [Vas08]. Removal [YL06b]. Renewable [BBK05]. Renewal [AH04b, GS04b, HA01, Pon07, RCX06, SK05]. Repair [KSW09, MS07c]. Repairable [Pul01, Pul02, SM06b, YKC06]. Reparameterisation [Sha02b]. reparametrization [FY00]. Repeated [CS04a, Hat09, IT09, Par01, RvdGM05, SSL08, BTH00]. Replacement [GR09]. Replicated [JW05, SL07, SHA00]. Replications [Koy06]. Reply [Ano03a, Nad09]. Report [Ano04c]. Representation [Jan05b, KS09, NC05, Rod08, Rud02, NC02]. Reproductive [TWW01]. Required [BBK05]. Resampling [DWB01, GZ03]. Research [Gia03, YZ07]. reset [LY00]. Residual [AJ06, AJ09, AB05b, BGNR01, GK04b, Kha05, KMPS09, LCF06, NSMP03, NBA06, PN08, Sad08a, ZQ06, PPL00, SN00]. Residuals [Akd01, Imo03, KP02, Li09b, Li09a, UT01, DM00b]. Resistant [SB03]. Resolution [CDL00, GY05, KLHT06]. Resonance [KMS09]. Respect [BGM08, BC09, Gru04, LWZ07, Xu09]. Response [BBB09, BBB03, BTH00, DGRQ01, FJSZ07, GGVM01, HR03, HR08, JM09, Kan07, KW05, KDCW09, LNP02, Lun02, MXZ08, MW01, OS09a, Par01, PH02, SJS09, Sho02, Sho07, SST^{+}09, Smi05, Tin01, TLF04, WV06, Wan07b, YWL08, BL00a, Cho00, Don00, Gil00, GT00, HAS00, RJ00, Zhe00]. Response-Adaptive [BBB09]. Responses [BBB09, BS06a, Hud07, KRB^{+}09, KDCW09, YW07, GK00, LC00]. Restricted [AT09, Bie07c, CWS01, MKS01, ÖK07c, Özk09, Pan05, TU06, WH00, Wan06a, YZ07, YLX09, ZY07, GO00]. Restriction [Li08a]. Restrictions [EGB06, KM08b, WWJ09, XY07a]. Result [BHH07, RS04]. Results [Agg01, CL08b, Deb08, EOS08, Gil07, GN01, GGS04, HLW08, Hud02, ITV06, KAHS09, KW05, MR05, MOPS08, NRS07b, Pum07, Roy04, Sun02, YLZ09]. Retransformation [GSYB03]. Retrospective [Gur07]. Returns [CGT03]. Revealing [GdT05]. Reversed [GN01, GGS04, KG04, LL08, NSMP03, NBA06]. Review [Car09, Dem07, NS08, Pan05, Pat07, RS01]. Revisited [AA02, CB06, GVMS02, UY04, WJB01]. Revisiting [GPM05, RT06]. Reward [Liu01a, MPT04]. Rewards [Pap04]. Rheumatic [FKB09, TF09]. Rheumatoid [Kov09, She09, ZC09]. Rheumatology [SC09b]. Rice [WW02]. Ridge [Akd02, AT09, AR09, AKS06, AS08, BÖG09, FR08, FP09, Fu,00, KS05a, KS04a, LAR07, ÖK07a, Özk09, PY08, RF02, Ton02, ZM05, ZY07, PK00a]. Riesz [AMP05, Laj06]. Right [Fio08, Hwa05, JIP08, LH09, PS07b, QS06, She05b, She07, SY08, SSD01, YYL09]. Right-Censored [JIP08, QS06, She05b, She07, YYL09]. Right-Censoring [Hwa05, SY08]. Risk [CZL08, CG04, ENL05, Koi06, Kum01, PL09, WW07, WZ07, WWY08, WD04, YYW05, ZY08]. Risk-Efficient [CG04]. Risks [Bas09, EB08, Nag02, Ohy02, SSF04]. RMM [Sho07, Sho02, Sho05]. RMM-Based [Sho05]. Roberts [MPT09]. Robust [AB01, AO09, ALT04, AG03, ALOS09, BC05a, BFZ03, Bla09, Bon08, Fab01, GEGMMI09, Guo00, Hol07, JS06b, MS00, MKS01, SSL08, SCL09, TC09, Tom03, Tsa00, Tso03, TC04, Tso05, Tso09b, ZCL09, Zho01, Had00, PRW00]. Robustness [BD03, DM00a, De 07, Grz03, LS02, LP05, LC05b, MD04, Mic06, MWM03, dlHRB03]. ROC [BRG07, GBSW09, KG00]. Role [GPM05, ST05b]. Root [GT03b, SS06, TWH08, ZK08, NGUT00]. Roots [Iba03]. Rotation [SP08a]. Row [PGP03, Sal06]. Row-Column [PGP03]. Ruin [CZL08, TT03, WWY08, ZY08]. Rule [Cha04a, CWS01, KSS03, LY05, SW02]. Rules [CL01, GM02, JLBB07, SC03]. Run [AH04a, SC03, RJ00]. Run-Length [SC03]. Runs [Ant03, BBQ02, EB07, EK04b, IA09, MPP07, SC03, DLT00].

Sa [Nad09]. Saddle [Ger09]. Saddlepoint [LCL09, XK08, ZK08, ZZ09]. Same [BM08a, GVGP06, MKvdMT05]. Sample [Aos01, BK03, BGM08, CDB06, Chi09, Chi07, CY07, CK04a, CZ07, Dam05, Dut07, Eks02, Fig07, GZ03, GTB07, Grz03, HN03, Iba03, Jun02, Kat01, KK02b, Kim07, KS04b, Lia06, LW02, LC05b, Luc01, MS09, MS06, MCY09, MvdM01a, Muk05b, NMPS05, NS08, NC05, Ozt02, PSZ07, Pat01, Pew02, RCL08, Ric07, Ruk09, SASAS06, SB06, SN06, SP04, SG05b, SW09b, TS09, VHS00, VH01, WS00, Wan09b, WL03, YNI05, YTN07, YA09, ZZ09, Zie06, dL07, CR00, DC00, KK00a, MS00, Neu00, SM00, WYY00]. Samples [AE08, BA05a, BW03, CAB02, Cra01, HN09, HWZ06, JE01, KO05a, Kli07, Lim07, Özt01, PY08, PP07, PR04, SP05, ZGS03]. Sampling [ASS05, AO09, AL05, Ami03, Arn01a, Arn01b, Asp06, BBB03, BF02, Ber05, BD05, Bor03, BCRW02, CSWM09, Dix03, Eid05, Eid06, GR05, GT03a, Ghi05, GN07, GSA08, HL02a, Hui05, KC05, KC07, Kan07, KK09, LD08, Lia09b, LVNF07, MS09, MPG08, Muk03, Muk05b, NMS08, OS09a, OG08, PNJ07, RT04, Rau09, Ros09, RMAM09, Sal06, SAS04, SM08a, SG06, SV08a, SV09, SSM06, TK01, WCL04, Zha01, BLB00, ENS00, Man00b, Zhe00]. Saralees [dSFM07]. Sarcoidosis [CSP^{+}09]. Sarmanov [SL04]. SARS [CHYL08]. Saunders [FL06, Rie03, SLB08, Tsi01a]. Savage [CDL01]. Save [Coo00, ZXM07]. Savits [BS08]. Scalar [Laj06]. Scale [AG03, Bla09, BD05, CL01, CP09, CSV04, CM01, Dan05, EM01, FZ05, FJC07, GYX06, Gra03, Ham06, HKR02, JE01, JC08, KK07a, KF06, LM06, LN05, LSD07, MD07, MS05, NMS09, PK02, SB03, SP04, SG05c, ST07, SVC07, SV08b, Tan06, TK01, ZGS03, AO00, CC02, DM00a, DM00b]. scale-location [DM00b]. Scaled [BA05a, Wan08]. Scan [FWB08]. Scanning [KMS09]. Scans [IA09, SC03]. Scheme [BBB03, CDB06, Dix03, KL04, LD08, LVNF07, Ros09]. Schemes [BCK07, KMPS09, Luc01, Mar02a, Smi05]. Schindler [Neu00]. Schlömilch [Bak08a]. Scleroderma [Li09b]. Score [BCV01, BB01, GGT05, JHL07, Lag04, LZ01, Sil06, Tso09b, Tu01, UOFC08, Wu02, BCV02, CNZ00, CC02].

Scores [FKB09, FBH03, HK06]. Scoring [Pat07, Smi05]. SCRAM [MG09]. Screen [Hsu05]. Screening [CC08, Kim09]. Search [SPG07, SP08a, CDL00]. Seasonal [SA03]. Secant [KF07, Kra05, Vau02]. Second [Aos00, BCV02, CF01, CS06b, Lag04, SJS09, HAS00]. Second-Order [CS06b, SJS09, Aos00, BCV02, HAS00]. Secondary [NMS08]. Sectional [AAGM05]. Sections [DJ08, KKMS07]. Seemingly [FR08, Liu03a, SW02, TAS01]. Segmented [Par01, VG09a]. Selected [CP09, GK05, NMS09, SS07b, AY00, KK00b]. Selecting [ACP03, BC05b, BC08b, FJ09, HL09, Lin04, LH01, MKvdMT05, Tak09, VH04, WL00]. Selection [CS07b, CH09, DY08, De 08, GY05, GN07, JM09, JWSP04, KSS03, KK02a, KP02, Kov09, Luc01, Mal09, MvdM01b, MKvdMT05, MOS04, NWTI03, Pan05, RTT01, Shi01, WC07, Wan09b, WJB01, XY07b, YYK07, YX08, Yi01, dlHRB06, AENT00, KH00, KLK00, Mil00]. Selective [MS07a]. Self [BB04, FRMAA08]. Self-Feeding [BB04]. Self-Similar [FRMAA08]. Seme [CM00b]. Semi [Ani04, BBL04, BJM04, CT04, CKL08, DD04, EB08, GL04, GH04, GMW04, Har04, HLLL08, Huz04, Jaś07, KL04, LOS05, LL05b, LVVA04, Mac08, MPT04, Pap04, RS03, RCPO04, SS04, Sil04, SL09b, VT04, WV06, WL08]. Semi- [RS03]. Semi-Folding [HLLL08]. Semi-Markov [Ani04, BBL04, BJM04, CT04, CKL08, DD04, GL04, GH04, GMW04, Har04, Huz04, Jaś07, KL04, LOS05, LL05b, LVVA04, Mac08, MPT04, Pap04, RCPO04, SS04, Sil04, VT04]. Semi-Parametric [EB08, SL09b, WV06, WL08, Ano04c]. semiconductor [PJ00]. Semiparametric [AS07b, AT09, APdR01, BMN06, BM08a, Che05, CGT02, DY08, DM04, Eid06, GMW04, Hat09, HWL09, Li09b, Lie05, LSRD04, Nai02, OJ08, Pat07, WLQ09, XD09, CFB00, LD00]. Sense [App02]. Sensitivity [GDCOCO06, GVMS07, SVC07, ZYL06]. Sensory [FBH03]. Separable [Shi08a, Shi08b]. Separate [SW02]. Sequence [BMS01, Ery05, EB07, PJ06, SP07, SP09]. Sequences [GM01, IM04, Rao03, Tau09]. Sequential [Ano04c, ACP03, BBB03, Blu08, CZR02, CZ03, CG04, DRB07, GH09, HC01, Hus05, KKP07, LLX09, MT05, RC07, Ros09, Ruk09, Sai00, SP01, Ton05, ZM09a, CAC00]. Serial [BFZ03, Blo02, HWL09, LCC08, PR04]. Serially [ST01b, SSJ02, WYZ05]. Series [ANN08, AJ06, AT01b, BH06, BLZ08, CM08, CC01, CWG06, CCGB03, DGZ01, Dut01, EGMPMRPM05, EOS08, FV02, GM02, Gil07, GH09, JH08, JS06a, KA04, KL02, Led08, McG07, MXX08, MB07c, NC02, NC05, Quo06, RDI01, SA09, Sha01b, Sha09, TWH08, Tri07a, UY04, Wan05b, WL08, Wan09b, Wei09, ZT09, HR00, SHA00]. Sessions [KMS09]. Set [ASS05, AO09, ANMSL06, BF02, CSWM09, Ghi05, JR07, MOS04, NMS08, OG08, Özt01, SAS04, SASAS06, SM08a, SSM06, TK01, VTKB04, ENS00, KK00a]. Sets [BS04, NÚF05, RLÚF05, WCS03]. Setting [KP07, Mud08, Sun05a]. Several [GKFM^{+}08, JE01, Men01, Tso09b, WW07, BA00]. Sex [ZLY05]. Shannon [AF08, Fro03]. Shape [NK05, Pat01, Ram05, Ram08, SPS07, dBSSvV01].

Shaped [AS07a, GM08, CC00]. Shapiro [VG09b]. Share [XL06]. Shared [LL03]. Sharp [Dan05, KO05b, RR02, Raq04]. Sharpening [LH01]. Shelemyahu [Edi09a]. Shelf [FBH03]. Shepp [HVW07]. Shift [SP07, SP09]. Shifted [JJ08, Lia03]. Shiryaev [MPT09]. Short [BK01]. Shortest [FK06]. Shoulder [Zha01]. Shrinkage [ASW06, AR09, Gru04, Gru08, Kha08, RF02, SS06, SS05b, SPS07, ST07]. Sib [HN07, LZ00]. Sib-Sib [HN07]. Sided [De 05, DRB07, Din01, Han08a, MD07, OB05, SG05b, Wan06a, ZGS03, vDK03]. Sign [Dub06, SASAS06]. Signal [KK03, McE08]. Signals [CH09, KMPS09, RTW06, Ric07, WTZL01]. Signed [Bak06, Tau09, Ter02, ZK08]. Signed-Rank [Ter02]. Signed-Rank-Like [Bak06]. Significance [Jac08, MKS01, Rod06]. Significant [SAAAS07]. SIMEX [KG00]. Similar [FRMAA08]. Similarities [Nav08b]. Similarity [AS07c, VTKB04, YC05, pLM00]. Simple [Ahm04, AL05, BZX09, BHD04, Bor03, EGS05, FCL06, GDCOCO06, Güv06, KL05, KC07, KSS03, Kha08, Kit02, MYS07, NP01, Pan02, PH02, RTT01, RZ02, TM03, TH06, MV00, ZFY00]. Simplex [MFPG07]. Simplified [Bru08]. Simulation [AMP05, UP03]. Simulations [LSRD04]. Simultaneous [JWSP04, KOF02, KMPS09, Lia09b, LH01, RTW06, Stȩ07, SSD01, WDRB00]. Simultaneously [CP04]. Sine [Gen07]. Single [Ahm08, BFZ03, BK09, CC09, GR05, MS07c, PS01b, SO05, SHRA06, Tak01, FTV00]. Single-Index [CC09]. Singular [DGLSG02, DGLS03, XY07a, Fu,00]. Sinusoids [RTW06]. Sites [CP04]. Situations [AT01a]. Size [Aos01, AG08, Chi09, CZ07, GH06, Kat01, LC05b, Luc01, MS09, MCY09, Muk05b, NMPS05, NS08, RJ04, RCL08, Ruk09, SW09b, VH01, YNI05, YA09, ZLY05, dBSSvV01, CNZ00, GA00, VHS00]. Size-Biased [GH06]. Sizes [BGM08, Lia06, ZM09b, BA00, WS00]. Skew [Adc07, AN07, AVGQ04, BMW08, CG07a, DMGFRQG07, GTB07, GB01, JBS09, JPB09, JK04a, LBAVM07, LNRS07, MFPP07, MFPG07, MLB09, Pou07, SAVG07, TD07, VLLS06, Pew00, ITY01]. Skew- [JPB09]. Skew-Elliptical [VLLS06]. Skew-Exponential [SAVG07]. Skew-Normal [Adc07, AN07, AVGQ04, DMGFRQG07, GB01, JBS09, JPB09, LBAVM07, LNRS07, MFPP07, MFPG07, MLB09, Pou07, Pew00]. Skew-Symmetric [CG07a]. Skewed [ACS07, VH07, Kim02]. Skewing [KKP03]. Skewness [BBQ07, CCBB09, dHCC01, DZK04, Hol07, KF06, NA07]. Skewness-Kurtosis [DZK04]. Slepian [WS05]. Sliced [Lue08]. Slightly [BS04]. Slope [KS08]. Slopes [Kha08, HAS00]. Slowly [Myn06]. Small [CL08a, Chi09, FSZ07, KK00a, LW02, PR04, SA04, TBM03, Zie06, DLT00]. Small-Sample [Zie06]. Smirnov [KL02]. Smoking [QWSL09]. Smooth [Byo08, PS07b, RC02, SSD01, YC03, Bar00]. Smoothed [GM08, Ham08, TS09]. Smoother [Hwa02]. Smoothing [BCPK05, Bar01, Hua06, HK02, KP02, Led08, LT02, LC08, Yi01, Yua05, ZL06]. Smoothness [GO06]. Software [LNP01]. Sojourn [Pon07, VKT00]. Solution [Kab05, OS09b, PR08b, TTX01, Tza08]. Solutions [FS05, Lee05b].

Solve [Jaś07]. Solving [Gav08, SC02]. Some [AJ06, Agg01, AA01, Akd01, AE03a, AKS06, Asa04, BAA07, BB01, Bla09, BYP07, BM01, BLZ08, Bru08, BOR02, Cha08, CS06b, CL00, Deb08, Fie07, FR08, FS05, Fro07, GdT05, Ger09, GT00, GVMS02, Gru08, GN01, Gup04a, GK04b, GGS04, HLW08, HLLL08, Ili03, JS06a, KAHS09, KW05, LJ06, LNRS07, MT06, MS07b, MR05, MGD08, MOPS08, Muk05a, Muk06, NMPS05, RCX06, RTW06, RSG07, Rya00, SKA01, Shi08a, Shi08b, SDM08, SS05c, Sun02, Sun05b, SM06b, TY00, TSSI02, Usa02, YLX09, YLZ09, Yeh09, ZZ09, GO00, SS00a]. Sources [SDZ01]. Space [DZK04, Mac08, Wan06a, TY00]. Spaced [DATM05, Hat09]. Spaces [Fok08, LVVA04, Pat01]. Spacing [JS07, Mon01]. Spacings [Bie07b, GR09, HWZ06, HLW08, JS09, MvdM01a, XLZL07, YBCH08]. Spanning [BQ05]. Spares [BBK05]. Spatial [AN07, But06, ED08, Eks02, FRMAA08, GSYB03, ITV06, MN08, PGF03, Shi08a, Shi08b, YGKD09]. Spatiotemporal [FMAA08, LMG06]. SPC [ZYZ06]. Spearman [QS06, SQ04]. Special [MS09, SK05]. Species [YC05]. Specification [Byo08, TAS01, Wan09b, Yu07]. Specifications [PX03, YD00]. Specified [ACS01a, DLN08, MS05]. Spectra [McG07]. Spectral [FMAA08, KV05, SHA00, SM06a]. Spectral-Marginal-Based [FMAA08]. spectrum [SSW00]. Spheres [SI02]. Spherical [WSPP09]. Spline [LM04, LZZ08b, UNN05]. Splines [ADS02, Hua06]. Split [Güv05, Jac09, VL06, WSPP08, WSPP09, Wul05, YLZ09]. Split-Plot [Jac09, WSPP08, YLZ09]. Splitting [KF06]. Spondylitis [CSP^{+}09, TC09]. Spot [QAT05]. SPRT [Fra04, Lee05b]. Spurious [Gil07]. Square [Bor03, BM01, CMLB05, Iba03, NC05, SS06]. Square-Root [SS06]. Squared [AE03a, CL08a, DG08, FZ05, FZ06, GHL07, MS07a, QN08, SKA01, SV08b, YZ07, YLX09]. Squares [AA09, DI08, DM08b, HdMS01, Hod09, HL02b, KG08, Kha05, Mon01, SC09a, TT07, Usa02, YSH05, Mil00]. Stability [BS04, Yeh09, Yi01]. Stable [Fan06, Fan09, GdT05, LR01, MBB08, MT06, RS04]. Stage [BBB03, DGZ01, JZ01, Liu01a, Liu03a, Muk05b, NMS08, PP07, Rau09, SS07b, ST07, YYAZ01, Yue04, ZM09a, Aos00, LC00, Man00b, WL00]. Staggered [KWW04]. Standard [Jac08]. Standardization [LR01]. Standardized [LFRLFG07]. Standby [MS07c]. Star [BS08, GM08]. Star-Shaped [GM08]. Starting [Eid09]. Starts [Sam09]. State [AAGM05, DD04, LHP06, Mac08, PD07, SL08a, TC09, CJ04, TY00, Tan00]. Stationary [AH04b, CWG06, GH09, Had09, JH08, KMPS09, LOS05, Mac08, Pak04, Cho00]. Stationary-Excess [Pak04]. Statistic [DSW07, FWB08, FV09, GEGMMI09, Jun02, MPP01, MT05, PC08, Sco07, TH08, TF04, ZN03, Coe00, Neu00, Yan00a]. Statistical [ANV09, BHN05, BN09, BKB02, BV05a, Car09, CB06, CCGB03, Dat03, DLN08, FBH03, GO06, HAA06, HLS09, Hin05, IPS08, KSW09, LNC09, LHQ07, McE08, SP09, TXG03, Tin09, TF09, TT03, YW07, ZBLG05, ZYZ06, dC02]. statistician [Chi00]. Statistics [ADS02, Ahm08, Ahs04, ABN04, AS07c, Ano07, BÖ07, BT07, Bak06, BBQ02, BES04, Bar06, BGAH09, Bar09, BAA07, BK08, BK09, Bie07a, Bie07b, BL09, Cha09, CF05, Cra02, CHB09, CNSV08, Dan05, Edi09b, Fri03, GS04a, Gen07, GKFM^{+}08, Gor09, Gov01, GN03a, Guo03, HN09, HMM03, Hus05, Jah05, JBS09, JS07, JC08, Kak07, KO05a, KO05b, KK07a, Lag04, Laj06, LNRS07, MBS17, Mar05, MPP08, NMPS05, NS08, Nas06, NRS07b, Nav08a, PS01b, PD07, PJH09, Rao01a, Raq04, Raq07, RC07, Ric07, Sad08b, ST05a, SM08a, SM09a, Shi08a, SSM06, ST08, Sul07, TSSI02, TS08a, TA07, TA08, Vex08, WSB09, YTN07, YBCH08, ZK08, GO00, HA00, Hür00, LCL01, PC02, ST00, Wu00]. Status [CS09b]. Stayer [SC09b]. Steel [VHS00]. Stein [Adc07, ACS01a, CWS01, SW02]. Stein-Rule [CWS01, SW02]. Step [AGAAG02, Ami03, BZX09, HMA05, Jan05a]. Step-Down [HMA05]. Step-Loss [Ami03]. Step-Stress [BZX09]. Stepwise [YYK07]. Stochastic [AA01, ASW06, DMGFRQG07, DAN08, FW07, FV09, Gan04, Hod09, HWZ06, KK07a, KAHS09, MJS09, MGD08, MOPS08, NN09, Ora06, PR08b, RF02, ST05b, SS04, SL09b, TD07, VKT00, XP04, XY07a, CP00, SSQ00, WDRB00, XM00]. Stochastically [Gur06]. Stock [CGT03]. Stopped [LL05b]. Stopping [CL01]. Storage [GKB06]. Strategies [MS07a, YWL08, Yue04]. Stratification [Sch05]. Stratified [BW03, Dup08, KC05, KK09, LW07b, NS04, PNJ07, RT04, SG06, SV08a, WH08]. Stratum [Dup08]. Strength [Cra01, DR01, GK04a, GB01, HHL07, Mok05]. Strength-Stress [GB01]. Stress [BZX09, Cra01, DR01, GK04a, GB01, Hsu05, LNP02, Mok05]. Stress-Response [LNP02]. Stress-Strength [Cra01, DR01, GK04a, Mok05]. Strict [KK04b]. Stroke [LCL01]. Strong [FFVFVF03, OSS05, OST09, XWW07, YZZ08]. Strongly [IT09, LH02]. Structural [KA04, KDCW09, Li01, XGZ09]. Structurally [PGP03]. Structure [AS07c, BLK09, GEGMMI09, GH09, MA01, NK08, Pan02, PH02, PB04, Tri07b, XG09, YH08, YBCH08, CBsK00, Shu00]. Structures [YSH05]. Student [FDR08, Kha00, KS08, SN06, Tsi08, VC00]. Student- [Kha00, KS08, Tsi08]. Studentized [WJ06]. Studies [Cha04b, EGS05, EA05, GAMG09, GB01, JS06b, LMR09, RCPO04, SHRA06, sS08a, TF09, XG09, KG00, Shi00, ZFY00]. Study [BR05, GS04b, GDCOCO06, JW05, Kim06b, LMR09, NS04, SA04, Tin09, UP03, ZC09, CM00b]. Sub [Ozt02]. Sub-Sample [Ozt02]. Subcopulas [Car02]. Subgroups [GAB01]. Subject [ED08, PM02, Pul01, sS05a, Zha08]. Subsample [SP04]. Subsamples [SC05]. Subsampling [LM07, PRW00, RvdGM05]. Subscales [HM04]. Subset [GK05, KK02a, MKvdMT05, Pan05, Mil00]. subsets [KH00]. Success [MPP07, NS04]. Successes [Bol07]. Successful [Sha01a]. Successive [RMAM09, SY08, SV09, WSB09]. Such [ACS07]. Sufficiency [IP06]. Sufficient [Dut07, Tza08]. Suitable [GVMS02]. Sultan [Che06]. Sum [CMLB05, Kak07, Kha05, KP02, LL05a, LM07, MK03, Mon01, TT07]. Sums [HPR07, JC08, Nad05, Nad09, PR08a, Ram06, Ram08, SW04, dSFM07]. Sungur [Mud08]. Super [Set04]. Super-efficient [Set04]. Super-powerful [Set04]. Superimposed [WTZL01]. Superiorities [WZ07].

Superpopulation [BA07, Xu09]. Superposition [RCA02]. Supersaturated [But05, CCG09]. Supplementary [Wu02]. Support [LT08, SHC02, YGKD09]. Support-Adjusted [YGKD09]. Supported [FW04]. Supports [GM08]. Supremum [Rao03]. Surface [DGRQ01, SJS09, BL00a, Don00, GT00, HAS00, RJ00]. Surveillance [And09, LLX09]. Survey [KR02, KR03, KML01, NRS07b, PY08, Zie09, Cha00]. Surveys [Smi05]. Survival [AWJ05, AND04, BBL04, BS06b, BDF06, CCD02, CS04b, CDKK05, DW07, EB08, JN03, KD08, LLCNM01, LZ01, LSRD04, PD07, Ros02, sS05a, SAT05, Tu01, WL05, YH06, BRM00, Man00b, SSQ00]. Suspected [Kha08]. Switches [Ani04]. Switching [Lee05a, SP08b]. Symmetric [CG07a, EOOPGB09, EE09, GdT05, HV03, Hod09, JS07, KZ04, MT06, SQ01, TK01, UOFC08, VH07, WQC07, YH08, CR00, TIS01]. symmetrization [PRW00]. Symmetry [BBQ07, Bon08, MW09, NW09, Özt01, SM01, JS00]. Synchronized [Sal03]. Synergism [WW02]. Synthetic [AdL09, GS08]. Synthetic- [AdL09]. System [AH04a, AB05b, BBK05, CN09, DR01, DZK04, FR08, KSW09, MXX08, MB07c, MS07c, Sad08a, SW02, Sta09, ZYL06]. Systematic [Ber05]. Systemic [AF09, FKB09]. Systems [FV02, GCLG06, HRVW07, JT09, MPT04, MPS08, NL07, NRS07a, Pap04, RC07, SM06b, VT04, YKC06].

Table [Nan09, Zhe03]. Tables [AFM08, CL09, CY09, DBA05, FW07, GVP05, RC03, Sco00, WW07, PB00, Smi00b]. Tabulated [Han08b]. Tail [Bur08, CGR09, DSPS09, FDR08, GG05a, Gir04, GPM05, GT03c, Jav09b, Kol07, Li08b, Mar05, PR04, ZBN09]. Tail-Coefficient [Gir04]. Tailed [Gal02, GO05, HT06, HT07]. Tails [FK06, HH05, SvDM07, WWY08]. Tampered [hWlF03]. Tanks [GKB06]. Target [CC08, SAAAS07]. Targets [CP04]. Taylor [CS09a]. TDT [WS00]. Tears [Bro02]. Technique [KK03, LMM01, PD07, RTT01, Sha01b]. Techniques [BCPK05, Dat03, MPS04]. Temporal [ES09, ZT09]. Tendency [CS08b, Fab08, Fab09]. Teratological [HR03]. Terms [ABN07, EBF09, TD07, TS08b]. Terpstra [LZ07]. Test [AS07a, AJ09, Akd02, AO01, BHD04, BA00, Blo02, BS05, Bro07, CC01, CY07, CF05, Dam05, Dub06, Dup08, FW07, Fie07, FWB08, FM06, Gat06, GF06, Gra03, Guo03, GGT05, Güv03, HN05, HA01, HMA05, HKW09, Jac08, Jeo03, Jun02, Kab05, KL02, Kra05, LL02, LL07, LZJ00, LCF06, LZZ08a, MM06, MB07a, Men01, MvdM01a, MWM03, NGYZ09, NM07, Özt01, Ozt02, PS07a, PRVHVA09, Rei06, Rod06, Ryn07, Sak02, SASAS06, She07, SG05b, SBJ^{+}07, Tak01, TTX01, TS08a, TOS09, Tau09, TH08, Ton05, TS07, Tso09b, VG09b, VPS08, Wan06a, WW07, Wil09, Wil05b, Wu02, WW03, YKC06, Zha01, ZYSZ08, vdM08, BLB00, CYHL00, JS00, KCL00, LR00, LCL01, Neu00, NO01, SM00, Tsa00, VHS00, WL00, Yan00a]. Testability [WWJ09]. Testimator [ST07]. Testimators [SPS07]. Testing [Ahm04, BBQ07, BB01, BFZ03, BBB08, CK05, CM09, CKF09, CMNP07, sC07, CB07, CL09, CY09, CH00, CK04b, De 05, DRB07, DE09, Dut07, EE09, FSZ07, GBG06, Gal02, GVMS02, GVGP06, GMO05, HT07, HKR02, HA01, HMM03, Hol07, Hsu05, HS08, HWL09, HJ01, IWC05, JW06, JZ01, JHL07, KKC06, KKSA07, KL02, KZ04, KWW04, Kim06b, KKL06, KS05b, KM08b, KSG03, LL03, Li08a, Li09a, Lia03, Lia05, Lia06, Lia07a, Lia08c, LW03, LW04, LW07a, LWW03, LSW05, LCC08, MS06, MPP08, MD07, NK03, NW09, PR04, RSG07, Ric07, Ruk09, SB03, SSQ00, SM01, TBM03, TWH08, TM07, TAS01, Tso09b, WC08, XK08, ZS06, ZGS03, ZZ01, ZS02, CAC00]. Tests [AGAAG02, ALT04, Art08, BCV01, BS06a, BKB02, Byo08, CZ03, CHC07, Chi09, CK04a, CM01, DV05, Deh07, EH04, Fig07, For09, Fri03, GBS05, GH09, GF08, Gur06, Gur07, HK05, HM02, HN03, Hin09, Imo03, JH01, JH08, JHL07, Kan02, KK02b, LLCNM01, LZ01, LZ07, LNP02, Mar02a, Mar02b, MFPP07, McE08, McG07, Mei05, MT01, MKS01, MT02, MW09, Mur01, OGM06, PSZ07, Que09, RC02, Ram05, Sal03, SN06, SS00a, SB03, SP04, Sil06, SJ04, Sür06, TM03, Tu01, UOFC08, VJ07, VG09a, WW02, Wu06, Yam07, YIS04, Zha07c, dL07, AC00, BCV02, CR00, CNZ00, CC02, GCB00, LY00, Lia00, LCL01, MM02, MMPV01, NGUT00, PB00, WH00, WS00]. th [Bie07c, DLb05, Dem07, Kli07, RR02]. Thank [Ano04f]. Thanks [Ano09]. Their [ALT04, BM01, Cha08, Fab01, Guo03, Har04, Imo03, JLBB07, Jon07, WS05, WQC07, vDK03, CWG06, CM00b]. Them [SC02]. Theorem [AM07, Led06, MYS07, PR08b, ZZ01, Syk00]. Theorems [GH04, GBR01, JS09, MLY02, Smi07]. Theoretic [RS04]. Theoretical [BLZ08, FJJZ05]. Theory [AH04b, Ano07, Arn01a, BL00b, CNSV08, Edi09b, Efr09, LHP06, MBS17, NS06, ON03, SLB08, Shi08a, Zie09, Zhe00]. Therapy [Kov09, LMR09]. Think [GS05]. Third [CGR09, CF05, Guo03, KR08]. Third-Order [CGR09, CF05, KR08]. Thompson [CDB06]. Three [DM08b, FW07, Hsi01, RMK08, TH08, Tza08, Tza09, YZ05, Tsi00]. Three-Parameter [Hsi01, RMK08, TH08, Tza08, Tza09, YZ05, Tsi00]. Three-Parametric [DM08b]. Threshold [ANV09, DGR07, HR03, HR08, Nie05]. Thresholding [DJ09]. Thresholds [SAAAS07]. Tied [CCD02]. Ties [Bro02]. Time [ANN08, AT01b, AH04b, BHN05, BBL04, BH06, Bro09, CM08, CCK02, CC01, CKL07, CWG06, CKL08, CCGB03, DGZ01, DM08a, Dix03, DM04, Dut01, EOS08, GM02, Gil07, GH09, GVW05, Hug04, JH08, KA04, KL02, Led08, LLCNM01, LVVA04, MPT04, McG07, Quo06, RDI01, SS07a, Sha09, sS05a, SGC09, TWH08, Ton05, Tri07a, UY04, Wan05b, WL08, Wan09b, Wei09, YH06, ZT09, ZYSZ08, GA00, HR00, KLK00, MV00, SHA00, SS00b, VKT00]. Time-Dependent [CCK02, DM04, sS05a]. Time-Series [Gil07, WL08]. Time-Varying [SGC09]. Times [Ant03, BDF06, CCK02, Cha07, JN03, Kan02, LWB03, MKvdMT05, Mur01, Pon07, SY08, Sil04, WL05, BRM00]. Toeplitz [Mar07]. Tolerance [RAK08, ZYZ06]. Tolerances [SC05, WC08, PL00]. Tools [CSP^{+}09, SC02, HR00]. Total [AH04a, ASN03, GZ03, PB04, Smi05]. toxicity [WCT00]. Toxicology [RK01]. TRA [ALT04]. Trace [ZM05]. Tracking [BBQ02]. Traditional [KM05]. trails [LD00]. Training [BS04, JR07]. Trait [WCL08, LZ00]. Transcription [CSP^{+}09]. Transect [Eid05, Eid06, Eid09, Zha01]. Transform [KS09, MPS04, NC05, Hür00, NC02]. Transformation [Bak08a, FM06, GSYB03, MB07a, SA01, TSSI02, VH01, LD00]. Transformations [GMO05, MF03, BTH00]. Transformed [Gra03, LXG06, NA09]. Transition [Jaś07, ZLH^{+}09]. Translation [LBSM09]. transmission [WS00]. Transplant [PD07]. Trapping [YL06b]. Trapping-Removal [YL06b]. Treatment [AJ06, AJ09, BJM04, CL06a, CB07, Dup08, Kov09, KS02, LR09, PM02, AKP00, LD00, ZW00]. Treatment-Covariate [CB07]. Treatments [AJ09, BC05b, CZR02, Hus05, LMR09, PLL00]. Tree [BQ05, Lun02, Shi01, TM03]. Tree-Based [Lun02]. Trees [IA09]. Trend [BHH07, CKL07, Sha09, SSJ02, Tsa00]. trending [KM02]. Trends [MG09, Myn06]. Trial [Li09b, Tin09, LCL01]. Trials [Blu08, Bol07, CL08b, FJJZ05, FJSZ07, HML04, HC01, KRB^{+}09, LW07b, LHQ07, LR09, MT05, Ruk09, SA04, XP04, LC00, WCT00, ZW00]. Triangular [HZSX06, Has08, MM00]. Trigonometric [SA09, Zha07a]. Trimmed [JC08, QT02]. Trinomial [KO07]. Triparametric [RACSSCJ05]. Trivariate [Bar06]. True [Tso09a]. Truncated [BCMRPAEG07, BD05, DSC01, DK02, JIP08, JPB09, JHL07, KH02, LOS07, LC05a, OR07, Ohy02, PH03, RCL08, She05b, She07, SSM09, WJ06, CM00a, MD00]. Truncating [HT06]. Truncation [BD05, GT03b, Hwa05, Jav09a, MLY02, OST09, sS05a, SY08, Sun02]. Truncation-Invariance [Jav09a]. TTT [KS09]. TTT-Transform [KS09]. Tukey [FHK07]. Tukey-Type [FHK07]. Tuning [YX08]. Two [ADP05, Bak08b, BBB03, BL07, BCRW02, CS06a, CSM08, DBA05, De 05, Din01, FUOC02, FFCN05, Fri03, GK05, Gen06, GT03c, GN04, GSA08, HI09, HK06, Han08a, HK05, HN03, HWZ06, Hua08, Hus05, Jac09, JW09, JBS09, Kan02, Kha08, KKSA07, KK02b, KKL06, Kim06a, Kim09, Kol07, KK09, KN01b, LM07, Li08c, Li09a, Lin01, Liu01a, Liu03a, MS06, MS07b, MM05, Mon01, Muk05b, Nan09, NL07, NMS08, NM09, OS09a, OS05, OB05, ÖK07c, PF05, PSZ07, PS07a, PP07, PGTM06, Pon07, PO06, QWSL09, Rau09, Rie03, RCA02, RT06, SM09a, SP04, SB08, SS07b, SST^{+}09, Sri02, SS05c, Sta09, Tak01, TS09, Tso03, Wan06a, Wil09, XK08, YYAZ01, YL06a, YZ07, Yue04, ZM09a, ZS02, vDK03, Aos00, Coe00, FL00, LC00, MS00, Neu00, PPL00, PLL00, RJ00, SM00, WL00]. Two-Dimensional [BCRW02, Pon07]. Two-Factor [YL06a, PLL00, RJ00]. Two-Fold [LM07, Li08c]. Two-Latent-Class [QWSL09]. Two-Level [Jac09, OS05]. Two-Parameter [ADP05, FUOC02, FFCN05, GK05, ÖK07c, SM09a]. Two-Part [HK06]. Two-Per-Block [Hua08]. Two-Phase [GSA08]. Two-Sample [HN03, KK02b, MS06, PSZ07, SP04, TS09, MS00, Neu00]. Two-Sided [De 05, Din01, Han08a, OB05, Wan06a, vDK03]. Two-Stage [BBB03, Liu01a, Liu03a, Muk05b, NMS08, PP07, Rau09, SS07b, YYAZ01, Yue04, ZM09a, Aos00, LC00, WL00]. Two-Tail [Kol07]. Two-Way [DBA05, KKSA07, Nan09]. Type [AAS01, Agg06, AO07, AWA05, ASW06, Ars09, BLK09, FUOC02, FHK07, Gir04, Gra03, HN07, Hus05, JK04b, KK04a, KL02, Koi06, Li08b, Liu03b, Liu04, Mok05, QN08, SI02, Tu01, Vex08, YX08, Yeh04, AGHMP05, iHA00, KLK00, ZFY00, AE08, AO00, Asg06, BA05a, BK05a, BZX09, BD05, CAB02, CKM07, Cra02, DRB07, EM01, Has08, HL02a, JC08, LWB06, LNC09, MAS05, PP07, SB06, ST08, Wan08, WL03]. Type-I [BZX09, BD05, PP07, ST08]. Type-II [BA05a, BK05a, CAB02, LWB06, LNC09, MAS05, SB06, Wan08]. Type-III [Has08]. Types [Fok08].

Ultrastructural [LS09, SL07]. UMP [BLB00, NO01]. Unbalanced [CSWM09, Din01, Güv05, Kim06b, LL05a, LM07, Mon01, PB04, Wul05]. Unbiased [AG08, BÖG09, Bur08, GD05, GA05, Güv06, LVNF07, ÖK07a, SM08a, SM09a, SW03, SS07b, SSM06, SDM08, Wul05, YLX09, GA00]. Unbiasedness [TS08b]. Uncertain [HRVW07, KS08]. Uncertainties [Kum01]. Uncertainty [Car09, Pul02, YGKD09]. Unconditional [Fri03]. Under-estimating [GZ03, KH03]. Underground [GKB06]. Underlying [Tso09a, Tso09b]. Understanding [Bar03]. Unequal [BGM08, BS05, Bro09, GN07, Koy06, KSG03, KKT05, Lia06, Mon01, XW08b, ZYSZ08, BA00]. Unequally [DATM05]. Unexpected [GdT05]. Unfolding [Hin05]. Unified [Ant03, KS02, Lee05b, Lin05, Mic06, lTTNlT08, iHA00, LZ00, QTR00]. Uniform [ADS02, Ahs04, BK09, Deb08, De 08, FFVFVF03, GM08, Grz03, Ham08, LH09, LM06, MPP01, Pre08, SL09a, Wan05a, SSQ00]. Uniformity [MvdM01a, QC09]. Unifying [CKK04]. Unimodal [Coo09, Pet09]. Union [AGHMP05, Lee03]. Uniqueness [DZK04, lTTNlT08, hWlF03]. Unit [GT03b, TWH08, NGUT00]. Units [BCRW02, Dan05, MPG08]. Univariate [AA01, AG03, Hud02, Lin01, OJ06, PNJ07, Wu06, YWK01, AKP00, CM00b]. University [PD07]. Unknown [CL01, CBB06, FL04, Nag02, PGF03, PK02, SDZ01, Tso09b, MBS17]. Unmodified [Dut07]. Unobserved [Ayi09]. Unrelated [AL05, FR08, Liu03a, SW02, TAS01]. Unreplicated [Abo05, BS06a, HMA05, KKSA07]. Unrestricted [ÖK07c]. Untransformed [dBSSvV01]. Up-and-Down [Sty09]. Update [Dra02, Nor00]. Updating [CL08b]. Upper [CZL08, Dan05, Sil04, Spu05a, XW08b, YIS04]. Upshifted [NBA06]. Urn [CL08b]. Use [BBQ07, CMNP07, Fab01, HHL03, Imo03, NC05, OS09a, PJ00, Ree07, SB06, SN05, VTKB04, WS05, XP04, Alw00, NC02]. Used [HA01, LZJ00, SS00a]. Useful [IPS08]. Using [ADM08, ASS05, AAGM05, Ami03, Bak08b, BBQ02, BA07, Che08, CCGPO03, DBA05, DATM05, DH06, Eid05, FS05, FBH03, Fro03, GM02, GdT05, GDSPSVP08, GVMS07, Gov01, GSA08, HI09, HM04, HH05, HN09, Hin09, Hsi01, Hus05, JW06, KKC06, KP02, KD08, Kim09, KM08b, KR02, KK09, KS02, LQ05, Li09a, Lie05, Liu03b, LSRD04, Mal09, MP08, MK05, MD04, MF03, MG09, MS05, Mur01, NP01, OS05, PR04, RvdGM05, RS01, RCBM09, SQ04, SAS04, SASAS06, SG04, SM08a, SM09a, SG07, SS03a, SPS07, SV08a, SK09b, SSM06, Smi05, Sol08, ST08, SSM09, TTX01, TN08, Tin01, Tin02, Tsi09, UNN05, VH01, Yan07, Zha07c, BRM00, LZ00, Lu00, MM00, MM02, Mil00, Sai00, WDRB00, Zhe00]. Usual [Tin02].

v.2 [CDL00]. Valid [Dut07]. Validate [SN06]. Validation [APdR01, Spi07, Wan06b, WTZL01]. Validity [AFM08]. Value [AV05, BES04, CG07a, CS06b, GM08, Hsi01, JS06b, Wan06a, WSB09, WW03, WL03, XW08b, Yu07, dlHRB03, RR02, dlHRB00]. Valued [Zho01]. Values [AA01, ADP05, AA03, AAG03, Bak08b, BGAH09, BAA07, Bie07c, CL09, CY09, De 05, GM01, HAA06, HI09, HS09, KAHS09, ZS06, Zha07c, Raq00]. Variability [DMH08, Fab08, HR08, PvdL07, TN08]. Variable [Ami03, BP09a, EA05, Fok08, Guo03, KK03, Kim09, Lia09b, LC05b, MS09, MBPP02, Nad09, RTT01, SM08b, WC07, YYK07, AENT00, CFB00, CBsK00, KH00, WYY00]. Variables [ABN07, BN09, BZ08, BL07, BW03, CMLB05, CD07, CL01, CC09, CCGPO03, CKK04, DBA05, Dod04, EEPL08, FS04, For05, FJC07, GdT05, Gen06, GZ05, GSA08, HV03, HM04, HWZ06, HPR07, HLW08, JBS09, KK07a, KK09, KP08, Lee05b, LX08, Liu01a, LCC08, MCY08, MB07b, Nad05, PGF03, PR08a, PGTM06, Pon07, Ram05, Ram08, Rod05, ST05a, Sch02, SK05, SSM09, TTB08, XY07b, ZL06, dSFM07, CM00a, Don00, KCV00, LY00, PG00, SSG00]. Variance [AL05, Ars09, BK03, Ber05, Bie07c, Che05, Du08, DM04, Eks02, FZ06, FL04, GBSW09, GZ05, Güv05, Güv06, JH01, JW05, KC07, KRB05, KKSA07, KN01a, Kna02, KMPS09, KG02, Las08, Li01, LL05a, LM07, Li08c, LP05, LWW03, Lue08, MPS04, Mar07, Men07, NS06, OS05, PB04, RT04, RW02, Sam09, ST01a, SG07, sS02a, SM09b, SK09b, ST01c, Ton02, TLW08, Vin04, WSPP09, YSH05, PC02]. Variance-Mean [Ars09]. Variances [BRG07, BS05, CM09, CC01, KSG03, KKT05, LZ07, Mak02, Men01, Tso03, Wil09, WW03, XW08b, JS00]. Variant [KP07]. Variate [DMGFRQG07, SMGVMD02, Tri08]. Variates [BK09, Ram06]. Variation [AB01, CK02, For09, HKR02, HK05, LFRLFG07, MCY09, SS03a, TN08, Tso09b, VJ07, WP02]. Variations [Tin01]. Various [Dan05, NÚF05, WWJ09, YX08]. Varying [CS08a, CWG06, FP09, HWL09, LM04, LZZ08b, QL09, Sim08, SGC09, WYY00]. Varying-Coefficient [CS08a, HWL09, LZZ08b, WYY00]. Vector [CY07, CZ07, GS08, GMM01, Kha05, LT08, MKS01, SHC02, FY00, HR00, Kha00]. Vectorial [BBQ07]. Vectors [Aos01, Ery05, LNRS07, Mar08a, NM03, NA09]. Version [ACS01a, Hwa02, KM05, KS04b, RSG07, vdM08]. Versus [Özt01, SW02, MM02]. Via [AKKM09, SA01, ZQ07, AE03b, CKK04, Fal03, HN07, HL09, KO05b, LM06, LC08, Pak04, QTR00, ZQ06]. View [GVGP06]. VII [SI02]. Violated [CL06a]. Visions [Rao01a]. Visualization [SL05]. Viterbi [SL08a]. Vitro [MD03]. Volatility [DI08, SS04, TPS08, CP00, Yan00b]. Volume [Chi09, Shi08a]. vs [LR01].

Waiting [Ant03, Mur01]. Wald [GCB00, Hus05, Lag04]. Wald-Type [Hus05]. Walk [Din01, KL04, Man00a]. Walker [Cho01]. Waring [XZ08]. Warner [BBB03]. Warnings [LB05]. Watson [Fig07, Fro07, GMM01, LR00]. Wavelet [AD02, DJ09, DCN07, Du08, GY05, LH02, LT04, Li02a, ON03, TH06, WS05]. Wavelet-Based [DCN07]. Wavelets [DAN08, UNN05]. Way [Bou09, BH05, DBA05, FW07, KKSA07, KSG03, Nan09, Nor03, Rao01b, SHRA06, Spu05b, Tso03, XM08, Yan00a]. Weak [DLb05, Ham08]. Weakly [GH09, KO05a]. Wear [BN09]. Weathering [CM08]. Weibull [Wan09a, AGAAG02, CC88, Coo09, GG05a, Gir04, Hür00, JN09, LNC09, MRT09, NG05, NE03, NE05, PJ06, TXG03, Tsi00, Tsi02, UP03, hWlF03, YCW^{+}08]. Weighted [Bla09, CMLB05, CS09a, Deh07, DM08b, Hus05, KSG03, LZ01, MGD08, SL09a, SL07, sS03b, SY08, SM06b, TT07, Wan07b, YCL09]. Weights [Asp06, Wan06b]. Weinman [Cra01]. Weiss [Neu00]. Well [BYP07]. Well-Known [BYP07]. Wesołowski [Bou09]. Where [Gav08, Muk05a]. Which [Bie08, Laj06]. Whitney [Ozt02, TS09]. Whittle [LL07]. Wide [IM08, XG09]. Wilcoxon [Ozt02, TS09, Tau09]. Wilcoxon-Signed [Tau09]. Wildlife [Mac02]. Wilk [VG09b]. Wilks [Coe00]. Williams [Tsa00]. Winner [Cha04a]. Winsorized [BK03]. Wise [Las08, PM02, WYY00, Zha09]. Wishart [Bou09, LG03, PK00b]. With-Replacement [GR09]. within [KOF02, SC09a]. within-individuals [KOF02]. Without [Bro02, Gav08, Alw00, Tso09a]. Words [CKL08]. Works [PTY09]. Worth [GK05]. Wrapped [JK04a, RSG07, Pew00]. Wright [HCL00].

XII [JW09].

yield [PJ00]. Yuan [PTY09]. Yule [CCK02, Cho01, XD04].

Z [Ano04a]. Zacks [Edi09a]. Zellner [AWA05, Gru04]. Zero [GGT05, JHL07, KP07, Rod03, Rod06, XD09, Gil00]. Zero-Inflated [Rod03, Rod06, XD09]. Zero-Truncated [JHL07]. Zonal [LX09, She05a]. Zones [JLBB07].

References

Ahmadi:2001:SUS

[AA01] J. Ahmadi and N. R. Arghami. Some univariate stochastic orders on record values. Communications in Statistics: Theory and Methods, 30(1):69–74, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ali:2002:CPR

[AA02] M. A. Ali and M. Y. Ashkar. The calibration problem revisited. Communications in Statistics: Theory and Methods, 31 (10):1733–1741, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

AlZaid:2003:CGD

[AA03] A. A. AlZaid and M. Ahsanullah. A characterization of the Gumbel distribution based on record values. Communications in Statistics: Theory and Methods, 32(11):2101–2108, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akbilgic:2009:NRA

[AA09] Oguz Akbilgic and Eylem Deniz Akinci. A novel regression approach: Least squares ratio. Communications in Statistics: Theory and Methods, 38(9):1539–1545, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Artis:2003:APV

[AAG03] Manuel Artis, Mercedes Ayuso, and Montserrat Guillen. Approximated perfect values in logistic regression for prediction and outlier detection. Communications in Statistics: Theory and Methods, 32(4):841–850, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Albarran:2005:MSM

[AAGM05] Irene Albarran, Mercedes Ayuso, Montserrat Guillén, and Malena Monteverde. A multiple state model for disability using the decomposition of death probabilities and crosssectional data. Communications in Statistics: Theory and Methods, 34(9–10):2063–2075, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Agarwal:2001:GGT

[AAS01] Satish K. Agarwal and Jamal A. Al-Saleh. Generalized gamma type distribution and its hazard rate function. Communications in Statistics: Theory and Methods, 30(2):309– 318, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abraham:2001:VRR

[AB01] Bovas Abraham and Mike Brajac. Variation reduction and robust designs. Communications in Statistics: Theory and Methods, 30(8–9):1951–1962, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alfonsi:2005:NFC

[AB05a] Aurélien Alfonsi and Damiano Brigo. New families of copulas based on periodic functions. Communications in Statistics: Theory and Methods, 34(7):1437–1447, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Asadi:2005:NMR

[AB05b] Majid Asadi and Ismihan Bayramoglu (Bairamov). A note on the mean residual life function of a parallel system. Communications in Statistics: Theory and Methods, 34(2):475– 484, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akhundov:2004:NCP

[ABN04] I. S. Akhundov, N. Balakrishnan, and V. B. Nevzorov. New characterizations by properties of midrange and related statistics. Communications in Statistics: Theory and Methods, 33(12):3133–3143, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akhundov:2007:IRT

[ABN07] I. Akhundov, A. Berred, and V. B. Nevzorov. On the influence of record terms in the addition of independent random variables. Communications in Statistics: Theory and Methods, 36(7):1291–1303, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aboukalam:2005:QEP

[Abo05] M. A. F. Aboukalam. Quick, easy, and powerful analysis of unreplicated factorial designs. Communications in Statistics: Theory and Methods, 34(5):1169–1175, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aubin:2000:BCT

[AC00] Elisete C. Q. Aubin and Gauss M. Cordeiro. Bartlettcorrected tests for normal linear models when the error covariance matrix is nonscalar. Communications in Statistics: Theory and Methods, 29(11):2405–2426, 2000. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aoshima:2003:SPS

[ACP03] Makoto Aoshima, Pinyuen Chen, and S. Panchapakesan. Sequential procedures for selecting the most probable multinomial cell when a nuisance cell is present. Communications in Statistics: Theory and Methods, 32(4):893–906, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnold:2001:MVS

[ACS01a] Barry C. Arnold, Enrique Castillo, and José María Sarabia. A multivariate version of Stein’s identity with applications to moment calculations and estimation of conditionally specified distributions. Communications in Statistics: Theory and Methods, 30(12):2517–2542, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnold:2001:QIC

[ACS01b] Barry C. Arnold, Enrique Castillo, and José María Sarabia. Quantification of incompatibility of conditional and marginal information. Communications in Statistics: Theory and Methods, 30(3):381–395, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnold:2007:DGS

[ACS07] Barry C. Arnold, Enrique Castillo, and José María Sarabia. Distributions with generalized skewed conditionals and mixtures of such distributions. Communications in Statistics: Theory and Methods, 36(8):1493–1503, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Angelini:2002:PCW

[AD02] C. Angelini and D. De Canditiis. Pointwise convergence of the wavelet regularized estimators. Communications in Statistics: Theory and Methods, 31(9):1561–1578, 2002. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adcock:2007:ESL

[Adc07] C. J. Adcock. Extensions of Stein’s lemma for the skewnormal distribution. Communications in Statistics: Theory and Methods, 36(9):1661–1671, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aparisi:2009:DPM

[AdL09] Francisco Aparisi and Marco A. de Luna. The design and performance of the multivariate synthetic- T^{2} control chart. Communications in Statistics: Theory and Methods, 38(2): 173–192, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aggarwal:2008:OFF

[ADM08] M. L. Aggarwal, Lih-Yuan Deng, and Mukta Datta Mazumder. Optimal fractional factorial plans using finite projective geometry. Communications in Statistics: Theory and Methods, 37(8):1258–1265, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmadi:2005:EPT

[ADP05] J. Ahmadi, M. Doostparast, and A. Parsian. Estimation and prediction in a two-parameter exponential distribution based on k-record values under LINEX loss function. Communications in Statistics: Theory and Methods, 34(4):795–805, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Agarwal:2002:LFU

[ADS02] Girdhar G. Agarwal, Rohan J. Dalpatadu, and Ashok K. Singh. Linear functions of uniform order statistics and Bsplines. Communications in Statistics: Theory and Methods, 31(2):181–192, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akdeniz:2003:MSE

[AE03a] Fikri Akdeniz and Hamza Erol. Mean squared error matrix comparisons of some biased estimators in linear regression. Communications in Statistics: Theory and Methods, 32 (12):2389–2413, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Assoudou:2003:BMM

[AE03b] S. Assoudou and B. Essebbar. A Bayesian model for Markov chains via Jeffrey’s prior. Communications in Statistics: Theory and Methods, 32(11):2163–2184, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abo-Eleneen:2008:FIT

[AE08] Z. A. Abo-Eleneen. Fisher information in Type II progressive censored samples. Communications in Statistics: Theory and Methods, 37(5):682–691, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andre:2000:CDV

[AENT00] Carmen D. S. André, Silvia N. Elian, Subhash C. Narula, and Rodrigo A. Tavares. Coefficients of determinations for variable selection in the MSAE regression. Communications in Statistics: Theory and Methods, 29(3):623–642, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmadi:2008:SIP

[AF08] J. Ahmadi and M. Fashandi. Shannon information properties of the endpoints of record coverage. Communications in Statistics: Theory and Methods, 37(4):481–493, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Allen:2009:PAL

[AF09] E. J. Allen and Vernon T. Farewell. Pragmatic analysis of longitudinal data on disease activity in systemic lupus erythematosus. Communications in Statistics: Theory and Methods, 38(18):3369–3388, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Araujo:2005:APD

[AFdBP05] Maria Ivanilde Araújo, Marcelo Fernandes, and Basilio de Bragança Pereira. Alternative procedures to discriminate non nested multivariate linear regression models. Communications in Statistics: Theory and Methods, 34(9–10):2047– 2062, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Allal:2004:AEA

[AFK04] J. Allal, M. Faizi, and A. Kaaouachi. Adaptive R-estimation in autoregressions. Communications in Statistics: Theory and Methods, 33(2):381–395, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andres:2008:CCM

[AFM08] Antonio Martín Andrés and Pedro Femia-Marzo. Chancecorrected measures of reliability and validity in 2 × 2 tables. Communications in Statistics: Theory and Methods, 37(5): 760–772, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arslan:2003:RLS

[AG03] Olcay Arslan and Ali I. Genç. Robust location and scale estimation based on the univariate generalized t(GT) distribution. Communications in Statistics: Theory and Methods, 32(8):1505–1525, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ashbridge:2008:CUE

[AG08] J. Ashbridge and I. B. J. Goudie. Conditionally unbiased estimation of population size under plant-capture. Communications in Statistics: Theory and Methods, 38(1):1–12, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Agapie:2009:CGA

[Aga09] Adriana Agapie. Convergence of the guesstimation algorithm. Communications in Statistics: Theory and Methods, 38(5):711–718, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abdel-Ghaly:2002:MLE

[AGAAG02] Abdalla A. Abdel-Ghaly, Ahmed F. Attia, and Magda M. Abdel-Ghani. The maximum likelihood estimates in step partially accelerated life tests for the Weibull parameters in censored data. Communications in Statistics: Theory and Methods, 31(4):551–573, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aggarwala:2001:PIC

[Agg01] Rita Aggarwala. Progressive interval censoring: Some mathematical results with applications to inference. Communications in Statistics: Theory and Methods, 30(8–9):1921–1935, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aggarwal:2006:BPO

[Agg06] Priyanka Aggarwal. Bayes predictor of one-parameter exponential family type population mean under balanced loss function. Communications in Statistics: Theory and Methods, 35(8):1397–1408, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmed:2008:GMB

[AGH08] S. E. Ahmed, M. N. Goria, and A. Hussein. Gamma mixture: Bimodality, inflexions and L-moments. Communications in Statistics: Theory and Methods, 37(8):1147–1161, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Atienza:2005:CML

[AGHMP05] N. Atienza, J. García-Heras, and J. Muñoz-Pichardo. Consistency of maximum likelihood estimators in finite mixture models of the union of W-type families. Communications in Statistics: Theory and Methods, 34(7):1471–1485, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Hussaini:2001:BPF

[AH01] Essam K. Al-Hussaini. On Bayes prediction of future median. Communications in Statistics: Theory and Methods, 30(7):1395–1410, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abdel-Hameed:2004:OPM

[AH04a] Mohamed S. Abdel-Hameed. Optimal predictive maintenance policies for a deteriorating system: The total discounted cost and the long-run average cost cases. Communications in Statistics: Theory and Methods, 33(3):735–745, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alsmeyer:2004:MRT

[AH04b] Gerold Alsmeyer and Volker Hoefs. Markov renewal theory for stationary (m + 1)-block factors: First passage time and overshoot. Communications in Statistics: Theory and Methods, 33(3):545–568, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmed:2000:CIE

[Ahm00] S. E. Ahmed. Construction of improved estimators of multinomial proportions. Communications in Statistics: Theory and Methods, 29(5–6):1273–1291, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmad:2004:SME

[Ahm04] Ibrahim A. Ahmad. A simple and more efficient new approach to life testing. Communications in Statistics: Theory and Methods, 33(9):2199–2215, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmad:2008:SPM

[Ahm08] Abd El-Baset A. Ahmad. Single and product moments of generalized order statistics from linear exponential distribution. Communications in Statistics: Theory and Methods, 37(8):1162–1172, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahsanullah:2004:CUD

[Ahs04] M. Ahsanullah. A characterization of the uniform distribution by dual generalized order statistics. Communications in Statistics: Theory and Methods, 33(12):2921–2928, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aggarwal:2006:SNS

[AJ06] M. L. Aggarwal and Mithilesh Kumar Jha. Some new series of partially balanced residual treatment effects designs. Communications in Statistics: Theory and Methods, 35(8): 1477–1482, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aggarwal:2009:CRT

[AJ09] M. L. Aggarwal and Mithilesh Kumar Jha. Constructions of residual treatment effects designs for comparing test treatments with a control. Communications in Statistics: Theory and Methods, 38(15):2567–2577, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Avi:2008:CPG

[AJSC08] J. Rodríguez Avi, M. J. Olmo Jiménez, A. Conde Sánchez, and A. J. Sáez Castillo. The _{3}F_{2} with complex parameters as generating function of discrete distribution. Communications in Statistics: Theory and Methods, 37(19):3009–3022, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akdeniz:2001:EAR

[Akd01] Fikri Akdeniz. The examination and analysis of residuals for some biased estimators in linear regression. Communications in Statistics: Theory and Methods, 30(6):1171–1183, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akdeniz:2002:MPT

[Akd02] Fikri Akdeniz. More on the pre-test estimator in ridge regression. Communications in Statistics: Theory and Methods, 31(6):987–994, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Araki:2009:FLD

[AKKM09] Yuko Araki, Sadanori Konishi, Shuichi Kawano, and Hidetoshi Matsui. Functional logistic discrimination via regularized basis expansions. Communications in Statistics: Theory and Methods, 38(16–17):2944–2957, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahn:2000:ETE

[AKP00] Choon Mo Ahn, Byung Chun Kim, and Sang Gue Park. The estimation of treatment effect for the censored bivariate data under the univariate censoring. Communications in Statistics: Theory and Methods, 29(2):319–334, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alkhamisi:2006:SMC

[AKS06] Mahdi Alkhamisi, Ghadban Khalaf, and Ghazi Shukur. Some modifications for choosing ridge parameters. Communications in Statistics: Theory and Methods, 35(11):2005– 2020, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Allaire:2005:SVU

[AL05] Denis Allaire and Louis Laurencelle. The sampling variance of r(X, Y) for unrelated X and Y: A simple algebraic demonstration. Communications in Statistics: Theory and Methods, 33(10):2389–2392, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Attouch:2009:ADR

[ALOS09] Mohammed Attouch, Ali Laksaci, and Elias Ould-Saïd. Asymptotic distribution of robust estimator for functional nonparametric models. Communications in Statistics: Theory and Methods, 38(8):1317–1335, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Allende:2004:RPT

[ALT04] Héctor Allende, Nancy Lacourly, and Soledad Torres. Robust portmanteau TRA tests and their limit distribution. Communications in Statistics: Theory and Methods, 33(8):1899– 1915, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alwan:2000:DEE

[Alw00] Layth C. Alwan. Designing an effective exponential cusum chart without the use of nomographs. Communications in Statistics: Theory and Methods, 29(12):2879–2893, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adam:2006:GCC

[AM06] Maria Adam and John Maroulas. Geometry in canonical correlations. Communications in Statistics: Theory and Methods, 35(12):2263–2273, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Acharya:2007:BMT

[AM07] S. K. Acharya and M. N. Mishra. Bernstein–von Mises theorem for M—M—1 queue. Communications in Statistics: Theory and Methods, 36(1):203–209, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aminzadeh:2003:BEV

[Ami03] M. S. Aminzadeh. Bayesian economic variable acceptancesampling plan using inverse Gaussian model and step-loss function. Communications in Statistics: Theory and Methods, 32(5):961–982, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anh:2005:ESR

[AMP05] V. V. Anh, R. McVinish, and C. Pesee. Estimation and simulation of the Riesz–Bessel distribution. Communications in Statistics: Theory and Methods, 34(9–10):1881–1897, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Allard:2007:NSS

[AN07] Denis Allard and Philippe Naveau. A new spatial skewnormal random field model. Communications in Statistics: Theory and Methods, 36(9):1821–1834, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Antolini:2004:ICD

[AND04] Laura Antolini, Byung-Ho Nam, and Ralph B. D’Agostino. Inference on correlated discrimination measures in survival analysis: A nonparametric approach. Communications in Statistics: Theory and Methods, 33(9):2117–2135, 2004. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andersson:2009:HMM

[And09] E. Andersson. Hotelling’s T^{2} method in multivariate online surveillance: On the delay of an alarm. Communications in Statistics: Theory and Methods, 38(16–17):2625– 2633, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anisimov:2004:AMM

[Ani04] Vladimir V. Anisimov. Averaging in Markov models with fast semi-Markov switches and applications. Communications in Statistics: Theory and Methods, 33(3):517–531, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arias-Nicolas:2006:NDS

[ANMSL06] J. P. Arias-Nicolás, J. Martín, and A. Suárez-Lloréns. The non dominated set in Bayesian decision problems with convex loss functions. Communications in Statistics: Theory and Methods, 35(4):593–607, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abegaz:2008:DCB

[ANN08] Fentaw Abegaz and U. V. Naik-Nimbalkar. Dynamic copulabased Markov time series. Communications in Statistics: Theory and Methods, 37(15):2447–2460, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2000:EBb

[Ano00a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 29(1):ebi, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2000:EBc

[Ano00b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 29(2):ebi, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2000:EBd

[Ano00c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 29(3):ebi, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2000:EBe

[Ano00d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 29(4):ebi, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2000:EBa

[Ano00e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 29(5–6):ebi, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2000:EBf

[Ano00f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 29(7):ebi, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2000:EBg

[Ano00g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 29(8):ebi, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2000:EB999

[Ano00h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 29(9–10):ebi, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610920008832587.

Anonymous:2000:EBh

[Ano00i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 29(11):ebi, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2000:EBi

[Ano00j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 29(12):ebi, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2003:AR

[Ano03a] Anonymous. Author’s reply. Communications in Statistics: Theory and Methods, 32(4):911–912, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2003:Ea

[Ano03b] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 32(5):1085–1086, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2003:Eb

[Ano03c] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 32(5):1087, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2004:BPZ

[Ano04a] Anonymous. Biography of Professor Z. Govindarajulu. Communications in Statistics: Theory and Methods, 33(9):1993– 1996, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2004:LE

[Ano04b] Anonymous. Letter to the Editor. Communications in Statistics: Theory and Methods, 33(11):2869–2870, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2004:MCR

[Ano04c] Anonymous. Mini-conference report: Non/semi-parametric models and sequential analysis. Communications in Statistics: Theory and Methods, 33(9):1997–1999, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2004:Pa

[Ano04d] Anonymous. Preface. Communications in Statistics: Theory and Methods, 33(3):xiii, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2004:Pb

[Ano04e] Anonymous. Preface. Communications in Statistics: Theory and Methods, 33(6):ix, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2004:TYN

[Ano04f] Anonymous. A thank you note. Communications in Statistics: Theory and Methods, 33(9):2001, 2004. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2005:E

[Ano05a] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 34(7):1687, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2005:LEa

[Ano05b] Anonymous. Letter to the Editor. Communications in Statistics: Theory and Methods, 33(1):195–196, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2005:LEb

[Ano05c] Anonymous. Letter to the Editor. Communications in Statistics: Theory and Methods, 33(4):991–992, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2006:E

[Ano06] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 35(6):1157, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2007:ECS

[Ano07] Anonymous. Erratum in Communications in Statistics — Theory and Methods 36(5–8): 1527–1542, for the article “Constrained Bayes and Empirical Bayes Estimation with Balanced Loss Functions”. Communications in Statistics: Theory and Methods, 36(12):2327, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [GKK07].

Anonymous:2009:TR

[Ano09] Anonymous. Thanks to the referees. Communications in Statistics: Theory and Methods, 38(16–17):3265–3266, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Antzoulakos:2003:WTN

[Ant03] Demetrios L. Antzoulakos. Waiting times and number of appearances of runs: A unified approach. Communications in Statistics: Theory and Methods, 32(7):1289–1315, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Amendola:2009:SPT

[ANV09] Alessandra Amendola, Marcella Niglio, and Cosimo Vitale. Statistical properties of threshold models. Communications in Statistics: Theory and Methods, 38(15):2479–2497, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anaya:2000:FOI

[AO00] K. Anaya and F. O’Reilly. Fisher’s observed information matrix for location and scale parameters under Type I censoring. Communications in Statistics: Theory and Methods, 29(7):1527–1537, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anaya:2001:TFI

[AO01] K. Anaya and F. O’Reilly. Test of fit for the inverse Gaussian and gamma distributions under censoring. Communications in Statistics: Theory and Methods, 30(4):757–773, 2001. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akahira:2007:ABK

[AO07] M. Akahira and N. Ohyauchi. The asymptotic bound by the Kiefer-type information inequality and its attainment. Communications in Statistics: Theory and Methods, 36(11):2049– 2059, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alexandridis:2009:RIA

[AO09] Roxana Alexandridis and Omer Ozturk. Robust inference against imperfect ranking in ranked set sampling. Communications in Statistics: Theory and Methods, 38(12):2126– 2143, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aoshima:2000:SOP

[Aos00] Makoto Aoshima. Second-order properties of improved twostage procedure for a multivariate normal distribution. Communications in Statistics: Theory and Methods, 29(3):611– 622, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aoshima:2001:SSD

[Aos01] Makoto Aoshima. Sample size determination for multiple comparisons with components of a linear function of mean vectors. Communications in Statistics: Theory and Methods, 30(8–9):1773–1788, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aneiros-Perez:2001:MCV

[APdR01] Germán Aneiros-Pérez and Alejandro Quintela del Río. Modified cross-validation in semiparametric regression models with dependent errors. Communications in Statistics: Theory and Methods, 30(2):289–307, 2001. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Appel:2002:BEB

[App02] M. J. Appel. Best estimation of a binary outcome in the sense of ks and aroc. Communications in Statistics: Theory and Methods, 31(8):1251–1257, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alheety:2009:CIS

[AR09] M. I. Alheety and T. V. Ramanathan. Confidence interval for shrinkage parameters in ridge regression. Communications in Statistics: Theory and Methods, 38(19):3489–3497, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnold:2001:BBB

[Arn01a] Barry C. Arnold. Beads, bags, Bayes, and the fundamental problem of sampling theory. Communications in Statistics: Theory and Methods, 30(8–9):1963–1967, 2001. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnoldsson:2001:ODB

[Arn01b] Göran Arnoldsson. Optimal designs for binomial sampling under a mixture hypothesis. Communications in Statistics: Theory and Methods, 30(5):897–911, 2001. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arslan:2009:VMM

[Ars09] Olcay Arslan. Variance-mean mixture of Kotz-type distributions. Communications in Statistics: Theory and Methods, 38(2):272–284, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Artes:2008:HTC

[Art08] Rinaldo Artes. Hypothesis tests for covariance analysis models for circular data. Communications in Statistics: Theory and Methods, 37(10):1632–1640, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abebe:2007:ETS

[AS07a] Asheber Abebe and Parminder Singh. An exact test for Ushaped alternatives of location parameters: The exponential distribution case. Communications in Statistics: Theory and Methods, 37(1):37–45, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adams:2007:SED

[AS07b] Robert M. Adams and Robin C. Sickles. Semiparametric efficient distribution free estimation of panel models. Communications in Statistics: Theory and Methods, 36(13):2425– 2442, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Saleh:2007:SSO

[AS07c] Mohammad Fraiwan Al-Saleh. On the similarity structure of order statistics. Communications in Statistics: Theory and Methods, 36(7):1433–1439, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alkhamisi:2008:DRP

[AS08] M. A. Alkhamisi and G. Shukur. Developing ridge parameters for SUR model. Communications in Statistics: Theory and Methods, 37(4):544–564, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Saleh:2002:FMC

[ASA02] Jamal A. Al-Saleh and Satish K. Agarwal. Finite mixture of certain distributions. Communications in Statistics: Theory and Methods, 31(12):2123–2137, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Asadi:2004:SCG

[Asa04] M. Asadi. Some characterizations on generalized Pareto distributions. Communications in Statistics: Theory and Methods, 33(12):2929–2939, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Asgharzadeh:2006:PIE

[Asg06] A. Asgharzadeh. Point and interval estimation for a generalized logistic distribution under progressive Type II censoring. Communications in Statistics: Theory and Methods, 35 (9):1685–1702, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Asha:2003:PMC

[ASN03] G. Asha, P. G. Sankaran, and N. Unnikrishnan Nair. Probability models with constant total failure rate. Communications in Statistics: Theory and Methods, 32(6):1089–1099, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Asparouhov:2006:GML

[Asp06] Tihomir Asparouhov. General multi-level modeling with sampling weights. Communications in Statistics: Theory and Methods, 35(3):439–460, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Saleh:2005:ECC

[ASS05] Mohammad Fraiwan Al-Saleh and Hani M. Samawi. Estimation of the correlation coefficient using bivariate ranked set sampling with application to the bivariate normal distribution. Communications in Statistics: Theory and Methods, 34(4):875–889, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akdeniz:2006:GEM

[ASW06] Fikri Akdeniz, George P. H. Styan, and Hans Joachim Werner. The general expressions for the moments of the stochastic shrinkage parameters of the Liu type estimator. Communications in Statistics: Theory and Methods, 35(3): 423–437, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akkaya:2001:EPA

[AT01a] Aysen Akkaya and Moti L. Tiku. Estimating parameters in autoregressive models in non-normal situations: Asymmetric innovations. Communications in Statistics: Theory and Methods, 30(3):517–536, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akkaya:2001:CTS

[AT01b] Aysen D. Akkaya and M. L. Tiku. Corrigendum: Time series models with asymmetric innovations. Communications in Statistics: Theory and Methods, 30(10):2227–2230, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alavi:2002:NEC

[AT02] A. Alavi and A. Thavaneswaran. Nonparametric estimators for censored correlated data. Communications in Statistics: Theory and Methods, 31(6):977–985, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akkaya:2005:LEM

[AT05] Aysen D. Akkaya and Moti L. Tiku. Letter to the Editor: Modified maximum likelihood estimators for autoregressive models. Communications in Statistics: Theory and Methods, 34(5):1243–1244, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akdeniz:2009:RRE

[AT09] Fıkrı Akdenız and Gülın Tabakan. Restricted ridge estimators of the parameters in semiparametric regression model. Communications in Statistics: Theory and Methods, 38(11): 1852–1869, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnold:2005:VIP

[AV05] Barry C. Arnold and Jose A. Villasenor. On the value of imprecise prior information. Communications in Statistics: Theory and Methods, 34(4):807–822, 2005. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arellano-Valle:2004:NCS

[AVGQ04] Reinaldo B. Arellano-Valle, Héctor W. Gómez, and Fernando A. Quintana. A new class of skew-normal distributions. Communications in Statistics: Theory and Methods, 33(7):1465–1480, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akdeniz:2005:GLT

[AWA05] Fikri Akdeniz, Alan T. K. Wan, and Esra Akdeniz. Generalized Liu type estimators under Zellner’s balanced loss function. Communications in Statistics: Theory and Methods, 34(8):1725–1736, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Almanassra:2005:MES

[AWJ05] Mahmoud Almanassra, Jiantian Wang, and S. Jeyaratnam. Monotonic estimators for the survival function of quality adjusted lifetime. Communications in Statistics: Theory and Methods, 34(5):1217–1231, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adrover:2000:QES

[AY00] Jorge G. Adrover and Victor J. Yohai. Quantile estimation for a selected normal population. Communications in Statistics: Theory and Methods, 29(2):243–262, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aydogdu:2005:PEF

[Ayd05] Halil Aydogdu. A pointwise estimator for the k-fold convolution of a distribution function. Communications in Statistics: Theory and Methods, 34(9–10):1939–1956, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ayis:2009:QIU

[Ayi09] Salma Ayis. Quantifying the impact of unobserved heterogeneity on inference from the logistic model. Communications in Statistics: Theory and Methods, 38(13):2164–2177, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andel:2002:MEP

[AZ02] Jirí Andel and Jitka Zichová. A method for estimating parameter in nonnegative MA(1) models. Communications in Statistics: Theory and Methods, 31(11):2101–2111, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhadary:2000:TEI

[BA00] Madhusudan Bhadary and Mohammed Khorsed Alam. Test for the equality of interclass correlation coefficients under unequal family sizes for several populations. Communications in Statistics: Theory and Methods, 29(4):755–768, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:2005:ISH

[BA05a] N. Balakrishnan and A. Asgharzadeh. Inference for the scaled half-logistic distribution based on progressively Type-II censored samples. Communications in Statistics: Theory and Methods, 34(1):73–87, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bentarzi:2005:CFI

[BA05b] Mohamed Bentarzi and Abdelhakim Aknouche. Calculation of the Fisher information matrix for periodic ARMA models. Communications in Statistics: Theory and Methods, 34(4): 891–903, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bentarzi:2006:LEA

[BA06] Mohamed Bentarzi and Abdelhakim Aknouche. An online estimation algorithm for periodic autoregressive models. Communications in Statistics: Theory and Methods, 35 (8):1495–1512, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bansal:2007:BPH

[BA07] Ashok K. Bansal and Priyanka Aggarwal. Bayes prediction for a heteroscedastic regression superpopulation model using balanced loss function. Communications in Statistics: Theory and Methods, 36(8):1565–1575, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baratpour:2007:SCB

[BAA07] S. Baratpour, J. Ahmadi, and N. R. Arghami. Some characterizations based on entropy of order statistics and record values. Communications in Statistics: Theory and Methods, 36(1):47–57, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bairamov:2007:P

[Bai07] Ismihan Bairamov. Preface. Communications in Statistics: Theory and Methods, 36(7):1271–1272, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bakir:2006:DFQ

[Bak06] Saad T. Bakir. Distribution-free quality control charts based on signed-rank-like statistics. Communications in Statistics: Theory and Methods, 35(4):743–757, 2006. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baker:2008:PAS

[Bak08a] Rose Baker. Probabilistic applications of the schlömilch transformation. Communications in Statistics: Theory and Methods, 37(14):2162–2176, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baklizi:2008:EUR

[Bak08b] Ayman Baklizi. Estimation of Pr(X < Y) using record values in the one and two parameter exponential distributions. Communications in Statistics: Theory and Methods, 37(5): 692–698, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:2006:E

[Bal06] N. Balakrishnan. Editorial. Communications in Statistics: Theory and Methods, 35(12):2145, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:2007:E

[Bal07] N. Balakrishnan. Editorial. Communications in Statistics: Theory and Methods, 36(1):1, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barry:2000:BAL

[Bar00] Daniel Barry. A Bayesian analysis for less smooth departures from polynomial regression models. Communications in Statistics: Theory and Methods, 29(12):2849–2864, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barry:2001:BSO

[Bar01] Daniel Barry. Bayesian smoothing with occasional jumps. Communications in Statistics: Theory and Methods, 30(12): 2707–2719, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barry:2002:BAD

[Bar02] Daniel Barry. A Bayesian analysis for derivative change points. Communications in Statistics: Theory and Methods, 31(8):1335–1348, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barrett:2003:UIM

[Bar03] Bruce E. Barrett. Understanding influence in multivariate regression. Communications in Statistics: Theory and Methods, 32(3):667–680, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baran:2005:CEL

[Bar05] S. Baran. A consistent estimator for linear models with dependent observations. Communications in Statistics: Theory and Methods, 33(10):2469–2486, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barakat:2006:GRR

[Bar06] H. M. Barakat. A general recurrence relation for the moments of bivariate and trivariate order statistics. Communications in Statistics: Theory and Methods, 35(11):1985– 1992, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barakat:2009:NGC

[Bar09] H. M. Barakat. A nonparametric general criterion of asymptotic dependence between order statistics. Communications in Statistics: Theory and Methods, 38(12):1960–1968, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Basu:2009:IAM

[Bas09] Sanjib Basu. Inference about the masking probabilities in the competing risks model. Communications in Statistics: Theory and Methods, 38(16–17):2677–2690, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bera:2001:SOP

[BB01] Anil K. Bera and Yannis Bilias. On some optimality properties of Fisher–Rao score function in testing and estimation.

Communications in Statistics: Theory and Methods, 30(8– 9):1533–1559, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brannas:2004:CHC

[BB04] Kurt Brännäs and Eva Brännäs. Conditional heteroskedasticity in count data regression: Self-feeding activity in fish. Communications in Statistics: Theory and Methods, 33(11): 2745–2758, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bar-Lev:2003:TSS

[BBB03] Shaul K. Bar-Lev, D.Sc., Elizabeta Bobovich, M.A., and Benzion Boukai, Ph.D. A two-stage sequential sampling scheme for Warner’s randomized response model. Communications in Statistics: Theory and Methods, 32(12):2373– 2387, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhattacharya:2008:GFT

[BBB08] Sabyasachi Bhattacharya, Ayanendranath Basu, and Subhadip Bandyopadhyay. Goodness-of-fit testing for exponential polynomial growth curves. Communications in Statistics: Theory and Methods, 38(3):340–363, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bandyopadhyay:2009:KBR

[BBB09] Uttam Bandyopadhyay, Atanu Biswas, and Rahul Bhattacharya. Kernel-based response-adaptive design for continuous responses. Communications in Statistics: Theory and Methods, 38(16–17):2691–2705, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bagdonavicius:2005:NCI

[BBK05] Vilijandas Bagdonavicius, Algimantas Bikelis, and Vytautas Kazakevicius. Nonparametric confidence intervals for the number of spares required for a renewable system. Communications in Statistics: Theory and Methods, 34(5):1203– 1212, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barbu:2004:DTS

[BBL04] Vlad Barbu, Michel Boussemart, and Nikolaos Limnios. Discrete-time semi-Markov model for reliability and survival analysis. Communications in Statistics: Theory and Methods, 33(11):2833–2868, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:2002:ULO

[BBQ02] N. Balakrishnan, M. R. Brito, and A. J. Quiroz. Using large order statistics of runs for tracking a changing Bernoulli probability. Communications in Statistics: Theory and Methods, 31(5):719–732, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:2007:VNS

[BBQ07] N. Balakrishnan, M. R. Brito, and A. J. Quiroz. A vectorial notion of skewness and its use in testing for multivariate symmetry. Communications in Statistics: Theory and Methods, 36(9):1757–1767, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bashir:2005:RRR

[BC05a] Shaheena Bashir and E. M. Carter. Robust reduced rank mixture discriminant analysis. Communications in Statistics: Theory and Methods, 34(1):135–145, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Buzaianu:2005:SAT

[BC05b] Elena Buzaianu and Pinyuen Chen. On selecting among treatments with binomial outcomes. Communications in Statistics: Theory and Methods, 34(6):1247–1264, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brunel:2008:AEH

[BC08a] Elodie Brunel and Fabienne Comte. Adaptive estimation of hazard rate with censored data. Communications in Statistics: Theory and Methods, 37(8):1284–1305, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Buzaianu:2008:CPS

[BC08b] Elena M. Buzaianu and Pinyuen Chen. Curtailment procedure for selecting among Bernoulli populations. Communications in Statistics: Theory and Methods, 37(7):1085–1102, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Buzaianu:2009:PBP

[BC09] Elena M. Buzaianu and Pinyuen Chen. Partitioning Bernoulli populations with respect to a control. Communications in Statistics: Theory and Methods, 38(16–17):2769– 2783, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burkschat:2007:OCO

[BCK07] M. Burkschat, E. Cramer, and U. Kamps. Optimality criteria and optimal schemes in progressive censoring. Communications in Statistics: Theory and Methods, 36(7):1419–1431, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barranco-Chamorro:2007:OAP

[BCMRPAEG07] I. Barranco-Chamorro, J. L. Moreno-Rebollo, A. Pascual-Acosta, and A. Enguix-González. An overview of asymptotic properties of estimators in truncated distributions. Communications in Statistics: Theory and Methods, 36(13):2351– 2366, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bae:2005:STB

[BCPK05] Whasoo Bae, Hyemi Choi, Byeong U. Park, and Choongrak Kim. Smoothing techniques for the bivariate Kaplan–Meier estimator. Communications in Statistics: Theory and Methods, 34(7):1659–1674, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bryant:2002:TDB

[BCRW02] Darryn Bryant, Yanxun Chang, C. A. Rodger, and R. Wei. Two-dimensional balanced sampling plans excluding contiguous units. Communications in Statistics: Theory and Methods, 31(8):1441–1455, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barroso:2001:IST

[BCV01] Lúcia P. Barroso, Gauss M. Cordeiro, and Klaus L. P. Vasconcellos. Improved score tests for von Mises regression models. Communications in Statistics: Theory and Methods, 30 (7):1295–1315, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barroso:2002:SOA

[BCV02] Lúcia P. Barroso, Gauss M. Cordeiro, and Klaus L. P. Vasconcellos. Second-order asymptotics for score tests in heteroskedastic t regression models. Communications in Statistics: Theory and Methods, 31(9):1515–1529, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhar:2003:RBD

[BD03] Lalmohan Bhar and Amitava Dey. Robustness of block designs for diallel crosses against missing data. Communications in Statistics: Theory and Methods, 32(1):193–213, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blumenthal:2005:EST

[BD05] Saul Blumenthal and Ram C. Dahiya. Estimating scale and truncation parameters for the truncated exponential distribution with Type-I censored sampling. Communications in Statistics: Theory and Methods, 34(1):1–21, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bravo:2006:SEM

[BDF06] José I. Bravo, Íñigo De Fuentes, and Arturo J. Fernández. Survival estimation with missing censoring times for the generalized Koziol–Green model. Communications in Statistics: Theory and Methods, 35(2):363–372, 2006. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bansal:2008:AEA

[BDH08] Naveen K. Bansal, Hong Du, and G. G. Hamedani. An application of EM algorithm to a change-point problem. Communications in Statistics: Theory and Methods, 37(13):2010– 2021, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Belitser:2002:CNM

[Bel02] Eduard Belitser. Consistency in nonparametric minimax regression estimation. Communications in Statistics: Theory and Methods, 31(7):1159–1165, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berger:2005:VES

[Ber05] Yvesg G. Berger. A variance estimator for systematic sampling from a deliberately ordered population. Communications in Statistics: Theory and Methods, 34(7):1533–1541, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barakat:2004:CDE

[BES04] H. M. Barakat and M. A. El-Shandidy. Computing the distribution and expected value of the concomitant rank-order statistics. Communications in Statistics: Theory and Methods, 33(11):2575–2594, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barabesi:2002:KEP

[BF02] Lucio Barabesi and Lorenzo Fattorini. Kernel estimators of probability density functions by ranked-set sampling. Communications in Statistics: Theory and Methods, 31(4):597– 610, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berkoun:2003:RTS

[BFZ03] Youcef Berkoun, Hocine Fellag, and Ryszard Zieliński. Robust testing serial correlation in AR(1) processes in the presence of a single additive outlier. Communications in Statistics: Theory and Methods, 32(8):1527–1540, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barakat:2009:ADO

[BGAH09] H. M. Barakat, M. E. Ghitany, and E. K. Al-Hussaini. Asymptotic distributions of order statistics and record values under the Marshall–Olkin parametrization operation. Communications in Statistics: Theory and Methods, 38(13):2267– 2273, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bofinger:2008:PRC

[BGM08] Eve Bofinger, Petra L. Graham, and Kerrie L. Mengersen. Partitioning with respect to a control: Unequal sample sizes case. Communications in Statistics: Theory and Methods, 37 (17):2713–2723, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Belzunce:2001:KER

[BGNR01] F. Belzunce, A. Guillamón, J. Navarro, and J. M. Ruiz. Kernel estimation of residual entropy. Communications in Statistics: Theory and Methods, 30(7):1243–1255, 2001. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bouveyron:2007:HDD

[BGS07] Charles Bouveyron, Stéphane Girard, and Cordelia Schmid. High-dimensional discriminant analysis. Communications in Statistics: Theory and Methods, 36(14):2607–2623, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bond:2001:NCB

[BH01] Marjorie E. Bond, Ph.D. and James J. Higgins, Ph.D. A note on ‘A comparison of Bayes and maximum likelihood estimation of the intraclass correlation coefficient’. Communications in Statistics: Theory and Methods, 30(2):371–380, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burch:2005:OOW

[BH05] Brent D. Burch and Ian R. Harris. Optimal one-way random effects designs for the intraclass correlation based on confidence intervals. Communications in Statistics: Theory and Methods, 34(9–10):2009–2023, 2005. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bibi:2006:EPB

[BH06] Abdelouahab Bibi and Moon-Ho Ringo Ho. Estimation of periodic bilinear time series models. Communications in Statistics: Theory and Methods, 35(9):1745–1756, 2006. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhandary:2000:BEI

[Bha00] Madhusudan K. Bhandary. Bayes estimation of intraclass correlation coefficient. Communications in Statistics: Theory and Methods, 29(1):79–93, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhattacharya:2001:CDC

[Bha01] Bhaskar Bhattacharya. Csiszar divergence from constant failure rate model for grouped data. Communications in Statistics: Theory and Methods, 30(6):1131–1141, 2001. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bar-Hen:2004:SGF

[BHD04] A. Bar-Hen and J. J. Daudin. A simple goodness-of-fit test for parametric regression models. Communications in Statistics: Theory and Methods, 33(6):1357–1370, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bischoff:2007:ARN

[BHH07] Wolfgang Bischoff, Enkelejd Hashorva, and Jürg Hüsler. An asymptotic result for non crossing probabilities of Brownian motion with trend. Communications in Statistics: Theory and Methods, 36(16):2821–2828, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bagdonavicius:2005:SAG

[BHN05] Vilijandas Bagdonavicius, Firoozeh Haghighi, and Mikhail Nikulin. Statistical analysis of general degradation path model and failure time data with multiple failure modes. Communications in Statistics: Theory and Methods, 34(8): 1771–1791, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bieniek:2007:BED

[Bie07a] Mariusz Bieniek. Bounds for expectations of differences of generalized order statistics based on general and life distributions. Communications in Statistics: Theory and Methods, 36(1):59–72, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bieniek:2007:PBE

[Bie07b] Mariusz Bieniek. Projection bounds on expectations of spacings of generalized order statistics from DD and DDA families. Communications in Statistics: Theory and Methods, 36 (7):1343–1357, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bieniek:2007:PMV

[Bie07c] Milena Bieniek. Projection mean-variance bounds on expectations of k th record values from restricted families. Communications in Statistics: Theory and Methods, 36(4): 679–692, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bieniek:2008:FDW

[Bie08] Mariusz Bieniek. On families of distributions for which optimal bounds on expectations of GOS can be derived. Communications in Statistics: Theory and Methods, 37(13):1997– 2009, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishna:2003:PEM

[BJ03] N. Balakrishna and T. M. Jacob. Parameter estimation in minification processes. Communications in Statistics: Theory and Methods, 32(11):2139–2152, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blasi:2004:NTH

[BJM04] A. Blasi, J. Janssen, and R. Manca. Numerical treatment of homogeneous and non-homogeneous semi-Markov reliability models. Communications in Statistics: Theory and Methods, 33(3):697–714, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Byrne:2001:SEC

[BK01] John Byrne and Paul Kabaila. Short exact confidence intervals for the Poisson mean. Communications in Statistics: Theory and Methods, 30(2):257–261, 2001. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bansal:2002:PRP

[BK02] Naveen K. Bansal and Debasis Kundu. Poisson regression with a periodic function. Communications in Statistics: Theory and Methods, 31(7):1123–1136, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:2003:VWM

[BK03] N. Balakrishnan and N. Kannan. Variance of a Winsorized mean when the sample contains multiple outliers. Communications in Statistics: Theory and Methods, 32(1):139–149, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:2005:PIE

[BK05a] N. Balakrishnan and J.-A. Kim. Point and interval estimation for bivariate normal distribution based on progressively Type-II censored data. Communications in Statistics: Theory and Methods, 34(6):1297–1347, 2005. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Byrne:2005:CPC

[BK05b] John Byrne and Paul Kabaila. Comparison of Poisson confidence intervals. Communications in Statistics: Theory and Methods, 34(3):545–556, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beutner:2008:RCR

[BK08] Eric Beutner and Udo Kamps. Random contraction and random dilation of generalized order statistics. Communications in Statistics: Theory and Methods, 37(14):2185–2201, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beutner:2009:IDS

[BK09] Eric Beutner and Udo Kamps. Identical distributions of single variates and random convex combinations of uniform fractional order statistics. Communications in Statistics: Theory and Methods, 38(12):1950–1959, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhaumik:2002:STN

[BKB02] Dulal K. Bhaumik, Ravindra Khattree, and Subhash C. Bagui. Statistical tests for nested designs under partial balance. Communications in Statistics: Theory and Methods, 31 (10):1795–1813, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balkin:2000:NNA

[BL00a] Sandy D. Balkin and Dennis K. J. Lin. A neural network approach to response surface methodology. Communications in Statistics: Theory and Methods, 29(9–10):2215– 2227, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/03610920008832604.

Brown:2000:TIR

[BL00b] G. C. Brown and A. J. Lawrance. Theory and illustration of regression influence diagnostics. Communications in Statistics: Theory and Methods, 29(9–10):2079–2107, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610920008832597.

Boussemart:2004:MDP

[BL04] M. Boussemart and N. Limnios. Markov decision processes with asymptotic average failure rate constraint. Communications in Statistics: Theory and Methods, 33(7):1689–1714, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bilder:2007:MAB

[BL07] Christopher R. Bilder and Thomas M. Loughin. Modeling association between two or more categorical variables that allow for multiple category choices. Communications in Statistics: Theory and Methods, 36(2):433–451, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burkschat:2009:MDC

[BL09] M. Burkschat and B. Lenz. Marginal distributions of the counting process associated with generalized order statistics. Communications in Statistics: Theory and Methods, 38 (12):2089–2106, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blazej:2009:SRA

[Bla09] P. Blazej. Some remarks about robust estimation of the scale parameter in weighted models. Communications in Statistics: Theory and Methods, 38(13):2232–2240, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bar-Lev:2000:UCD

[BLB00] Shaui K. Bar-Lev and Benzion Boukal. On a UMP and curtailed double sampling test for proportions. Communications in Statistics: Theory and Methods, 29(12):2791–2803, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bar-Lev:2009:BDG

[BLK09] Shaul K. Bar-Lev and Abram M. Kagan. Bivariate distributions with Gaussian-type dependence structure. Communications in Statistics: Theory and Methods, 38(16–17):2669– 2676, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blodgett:2002:MIO

[Blo02] Robert J. Blodgett. Measuring improbability of outcomes from a serial dilution test. Communications in Statistics: Theory and Methods, 31(12):2209–2223, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blodgett:2007:PMN

[Blo07] Robert J. Blodgett. The path of the minimum l^{p}-norm estimator for p between 1 and infinity. Communications in Statistics: Theory and Methods, 36(16):2829–2839, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blodgett:2008:FCC

[Blo08] Robert J. Blodgett. Fitting a curve to a confidence interval. Communications in Statistics: Theory and Methods, 37 (17):2681–2686, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blume:2008:HOL

[Blu08] Jeffrey D. Blume. How often likelihood ratios are misleading in sequential trials. Communications in Statistics: Theory and Methods, 37(8):1193–1206, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Braun:2008:STP

[BLZ08] W. John Braun, Wei Li, and Yiqiang Q. Zhao. Some theoretical properties of the geometric and α-series processes. Communications in Statistics: Theory and Methods, 37(9):1483– 1496, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brain:2001:SPQ

[BM01] Carlos W. Brain and Jie Mi. On some properties of the quantiles of the chi-square distribution and their applications to interval estimation. Communications in Statistics: Theory and Methods, 30(8–9):1851–1867, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berred:2008:SEC

[BM08a] Alexandre Berred and Sergey V. Malov. Semiparametric estimation in copulas with the same marginals. Communications in Statistics: Theory and Methods, 38(3):372–392, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bryk:2008:RFD

[BM08b] Artur Bryk and Jan Mielniczuk. Randomized fixed design regression under long-range-dependent errors. Communications in Statistics: Theory and Methods, 37(4):520–531, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishna:2009:NEA

[BM09] N. Balakrishna and Angel Mathew. Nonparametric estimation of the average availability. Communications in Statistics: Theory and Methods, 38(8):1207–1218, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bagdonavicius:2006:SRE

[BMN06] Vilijandas Bagdonavicius, Sergey V. Malov, and Mikhail S. Nikulin. Semiparametric regression estimation in copula models. Communications in Statistics: Theory and Methods, 35(8):1449–1467, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boudjellaba:2001:EIC

[BMS01] H. Boudjellaba, B. MacGibbon, and P. Sawyer. On exact inference for change in a Poisson sequence. Communications in Statistics: Theory and Methods, 30(3):407–434, 2001. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bansal:2008:EBH

[BMW08] Naveen K. Bansal, Mehdi Maadooliat, and Xiaowei Wang. Empirical Bayes and hierarchical Bayes estimation of skew normal populations. Communications in Statistics: Theory and Methods, 37(7):1024–1037, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bagdonavicius:2009:SMA

[BN09] Vilijandas Bagdonavicius and Mikhail Nikulin. Statistical models to analyze failure, wear, fatigue, and degradation data with explanatory variables. Communications in Statistics: Theory and Methods, 38(16–17):3031–3047, 2009. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bairamov:2007:CDT

[BÖ07] Ismihan Bairamov and Tugba Özkal. On characterization of distributions through the properties of conditional expectations of order statistics. Communications in Statistics: Theory and Methods, 36(7):1319–1326, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Batah:2009:CUR

[BÖG09] Feras Sh. M. Batah, M. Revan Özkale, and S. D. Gore. Combining unbiased ridge and principal component regression estimators. Communications in Statistics: Theory and Methods, 38(13):2201–2209, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boland:2007:PDN

[Bol07] Philip J. Boland. The probability distribution for the number of successes in independent trials. Communications in Statistics: Theory and Methods, 36(7):1327–1331, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bondell:2008:REE

[Bon08] Howard D. Bondell. On robust and efficient estimation of the center of symmetry. Communications in Statistics: Theory and Methods, 37(3):318–327, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burbea:2002:SRI

[BOR02] J. Burbea, J. M. Oller, and F. Reverter. Some remarks on the information geometry of the gamma distribution. Communications in Statistics: Theory and Methods, 31(11):1959– 1975, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Borkowski:2003:SLS

[Bor03] John J. Borkowski. Simple latin square sampling ±k designs. Communications in Statistics: Theory and Methods, 32(1): 215–237, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boutouria:2009:CWD

[Bou09] I. Boutouria. Characterization of the Wishart distribution on homogeneous cones in the Bobecka and Wesołowski way. Communications in Statistics: Theory and Methods, 38(15): 2552–2566, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baade:2000:MCD

[BP00] Ingrid A. Baade and Anthony N. Pettitt. Multiple and conditional deletion diagnostics for general linear models. Communications in Statistics: Theory and Methods, 29(8):1899– 1910, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bagkavos:2009:VBN

[BP09a] Dimitrios Bagkavos and Prakash Patil. Variable bandwidths for nonparametric hazard rate estimation. Communications in Statistics: Theory and Methods, 38(7):1055–1078, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boxma:2009:CMM

[BP09b] Onno J. Boxma and David Perry. On the cycle maximum of mountains, dams and queues. Communications in Statistics: Theory and Methods, 38(16–17):2706–2720, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brito:2005:DFM

[BQ05] María R. Brito and Adolfo J. Quiroz. Degree frequencies in the minimal spanning tree and dimension identification. Communications in Statistics: Theory and Methods, 33(1): 99–105, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bekker:2005:RCM

[BR05] A. Bekker and J. J. J. Roux. Reliability characteristics of the Maxwell distribution: A Bayes estimation study. Communications in Statistics: Theory and Methods, 34(11):2169– 2178, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Benhenni:2007:NEA

[BR07] Karim Benhenni and Mustapha Rachdi. Nonparametric estimation of average growth curve with general nonstationary error process. Communications in Statistics: Theory and Methods, 36(6):1173–1186, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bakouch:2009:BBG

[BR09] Hassan S. Bakouch and Miroslav M. Ristić. A bivariate betagamma autoregressive process (BVBGAR(1)). Communications in Statistics: Theory and Methods, 38(7):1113–1131, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brazauskas:2003:IMP

[Bra03] Vytaras Brazauskas. Information matrix for Pareto(IV), Burr, and related distributions. Communications in Statistics: Theory and Methods, 32(2):315–325, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bandos:2007:EBV

[BRG07] Andriy I. Bandos, Howard E. Rockette, and David Gur. Exact bootstrap variances of the area under ROC curve. Communications in Statistics: Theory and Methods, 36(13):2443– 2461, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bekker:2000:GCR

[BRM00] A. Bekker, J. J. J. Roux, and P. J. Mosteit. A generalization of the compound Rayleigh distribution: using a Bayesian method on cancer survival times. Communications in Statistics: Theory and Methods, 29(7):1419–1433, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brostrom:2002:CRT

[Bro02] Göran Broström. Cox regression; ties without tears. Communications in Statistics: Theory and Methods, 31(2):285– 297, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Broemeling:2007:DLT

[Bro07] Lyle D. Broemeling. Detection and localization in test accuracy: A Bayesian perspective. Communications in Statistics: Theory and Methods, 36(8):1555–1564, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brown:2009:DCP

[Bro09] Marlo Brown. Detecting changes in a Poisson process monitored at unequal discrete time intervals. Communications in Statistics: Theory and Methods, 38(16–17):2721–2732, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bruce:2008:SPS

[Bru08] Daniel Bruce. Some properties for a simplified Cox binary model. Communications in Statistics: Theory and Methods, 37(16):2606–2616, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burgess:2003:ODC

[BS03] Leonie Burgess and Deborah J. Street. Optimal designs for 2^{k} choice experiments. Communications in Statistics: Theory and Methods, 32(11):2185–2206, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berhane:2004:SNN

[BS04] Indrias Berhane and C. Srinivasan. Stability of neural networks for slightly perturbed training data sets. Communications in Statistics: Theory and Methods, 33(9):2259–2270, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bristol:2005:MMT

[BS05] David R. Bristol and Nicole Stouffer. A maximum modulus test with unequal variances. Communications in Statistics: Theory and Methods, 33(10):2333–2342, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Basso:2006:DPT

[BS06a] D. Basso and L. Salmaso. A discussion of permutation tests conditional to observed responses in unreplicated 2^{M} full factorial designs. Communications in Statistics: Theory and Methods, 35(1):83–97, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Biswas:2006:KSF

[BS06b] Animikh Biswas and Rajeshwari Sundaram. Kernel survival function estimation based on doubly censored data. Communications in Statistics: Theory and Methods, 35(7):1293– 1307, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bartoszewicz:2008:BSC

[BS08] Jaroslaw Bartoszewicz and Magdalena Skolimowska. Block-Savits characterization and star ordering of exponential mixtures. Communications in Statistics: Theory and Methods, 37(16):2461–2471, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bairamov:2007:DEB

[BT07] Ismihan Bairamov and Halil Tanil. On distributions of exceedances based on generalized order statistics. Communications in Statistics: Theory and Methods, 36(7):1479–1491, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boik:2000:RTR

[BTH00] Robert J. Boik, Charles Todd, and Scott K. Hyde. Response transformations in repeated measures and growth curve models. Communications in Statistics: Theory and Methods, 29(4):699–733, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burch:2008:CET

[Bur08] Brent D. Burch. Comparing equal-tail probability and unbiased confidence intervals for the intraclass correlation coefficient. Communications in Statistics: Theory and Methods, 37(20):3264–3275, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Butler:2005:SLH

[But05] Neil A. Butler. Supersaturated latin hypercube designs. Communications in Statistics: Theory and Methods, 34(2): 417–428, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Butler:2006:PDE

[But06] Neil A. Butler. Properties of designs in experiments on spatial lattices. Communications in Statistics: Theory and Methods, 35(2):337–347, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bordes:2005:SIP

[BV05a] Laurent Bordes and Pierre Vandekerkhove. Statistical inference for partially hidden Markov models. Communications in Statistics: Theory and Methods, 34(5):1081–1104, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Braekers:2005:CRM

[BV05b] Roel Braekers and Noël Veraverbeke. Cox’s regression model under partially informative censoring. Communications in Statistics: Theory and Methods, 34(8):1793–1811, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bosch:2003:OAS

[BW03] V. Bosch and R. Wildner. Optimum allocation of stratified random samples designed for multiple mean estimates and multiple observed variables. Communications in Statistics: Theory and Methods, 32(10):1897–1909, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Byoun:2008:SGF

[Byo08] Soku Byoun. Smooth goodness-of-fit specification tests under the Lagrange multiplier principle. Communications in Statistics: Theory and Methods, 37(3):443–459, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Borzadaran:2007:IMS

[BYP07] G. R. Mohtashami Borzadaran, G. H. Yari, and E. Pasha. Information measures for some well-known families. Communications in Statistics: Theory and Methods, 36(4):669–677, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Batsidis:2008:MLR

[BZ08] A. Batsidis and K. Zografos. Multivariate linear regression model with elliptically contoured distributed errors and monotone missing dependent variables. Communications in Statistics: Theory and Methods, 37(3):349–372, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:2009:ISS

[BZX09] N. Balakrishnan, Li Zhang, and Qihao Xie. Inference for a simple step-stress model with Type-I censoring and lognormally distributed lifetimes. Communications in Statistics: Theory and Methods, 38(10):1690–1709, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cooray:2008:GHN

[CA08] Kahadawala Cooray and Malwane M. A. Ananda. A generalization of the half-normal distribution with applications to lifetime data. Communications in Statistics: Theory and Methods, 37(9):1323–1337, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chandrasekar:2002:EEP

[CAB02] B. Chandrasekar, T. Leo Alexander, and N. Balakrishnan. Equivariant estimation for parameters of exponential distributions based on Type-II progressively censored samples. Communications in Statistics: Theory and Methods, 31(10): 1675–1686, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:2000:SCT

[CAC00] Jessica Y. Chang, Hongshik Ahn, and James J. Chen. On sequential closed testing dose groups with a control. Communications in Statistics: Theory and Methods, 29(5–6):941–956, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Carley:2002:MME

[Car02] Holly Carley. Maximum and minimum extensions of finite subcopulas. Communications in Statistics: Theory and Methods, 31(12):2151–2166, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Carayanni:2009:HUC

[Car09] Vilelmine Carayanni. Handling uncertainty in the cost effectiveness healthcare evaluations. A review of statistical approaches. Communications in Statistics: Theory and Methods, 38(8):1224–1240, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaouche:2006:SIG

[CB06] Ali Chaouche and Jean-Noël Bacro. Statistical inference for the generalized Pareto distribution: Maximum likelihood revisited. Communications in Statistics: Theory and Methods, 35(5):785–802, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2007:TTC

[CB07] Wonsun Chen and Arne Bathke. Testing treatment-covariate interaction in a nonparametric analysis of covariance model. Communications in Statistics: Theory and Methods, 36(6): 1153–1171, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:2006:CMF

[CBB06] Gauss M. Cordeiro, Lúcia P. Barroso, and Denise A. Botter. Covariance matrix formula for generalized linear models with unknown dispersion. Communications in Statistics: Theory and Methods, 35(1):113–120, 2006. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic). See corrigendum [MBS17].

Christensen:2004:LRL

[CBKP04] Karl Bang Christensen, Jakob Bue Bjorner, Svend Kreiner, and Jørgen Holm Petersen. Latent regression in loglinear Rasch models. Communications in Statistics: Theory and Methods, 33(6):1295–1313, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cumsille:2000:EDC

[CBsK00] Francisco Cumsille, Shrikant I. Bangdiwala, Pranab K. sen, and Lawrence L. Kupper. Effect of dichotomizinlg a continuous variable on the model structure in multiple linear regression models. Communications in Statistics: Theory and Methods, 29(3):643–654, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:1988:EWP

[CC88] K. F. Cheng and C. H. Chen. Estimation of the Weibull parameters with grouped data. Communications in Statistics: Theory and Methods, 17(2):325–341, 1988. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic). See correction [Wan09a].

Cutler:2000:ESB

[CC00] D. Richard Cutler and Adele Cutler. Estimation for Ushaped beta distributions: minimum Hellinger distance and related methods. Communications in Statistics: Theory and Methods, 29(7):1487–1509, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2001:NTC

[CC01] Min Chen and Gemai Chen. A nonparametric test of changing conditional variances in autoregressive time series. Communications in Statistics: Theory and Methods, 30(3):557– 578, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cysneiros:2002:IAS

[CC02] Audrey Helen M. A. Cysneiros and Gauss M. Cordeiro. On improving the χ^{2} approximation of score tests in locationscale nonlinear models. Communications in Statistics: Theory and Methods, 31(10):1709–1732, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cai:2008:SMT

[CC08] Weixing Cai and Pinyuen Chen. Screening for multiple target detection. Communications in Statistics: Theory and Methods, 37(12):1930–1947, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2009:ELP

[CC09] Xia Chen and Hengjian Cui. Empirical likelihood for partially linear single-index errors-in-variables model. Communications in Statistics: Theory and Methods, 38(15):2498– 2514, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cavalcanti:2009:ASE

[CCBB09] Alexsandro B. Cavalcanti, Gauss M. Cordeiro, Denise A. Botter, and Lúcia P. Barroso. Asymptotic skewness in exponential family nonlinear models. Communications in Statistics: Theory and Methods, 38(14):2275–2287, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:2006:BAO

[CCC06] Gauss M. Cordeiro, Audrey Helen M. A. Cysneiros, and Francisco José A. Cysneiros. Bartlett adjustments for overdispersed generalized linear models. Communications in Statistics: Theory and Methods, 35(5):937–952, 2006. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:2009:BCM

[CCC09] Gauss M. Cordeiro, Audrey H. M. A. Cysneiros, and Francisco José A. Cysneiros. Bias-corrected maximum likelihood estimators in nonlinear heteroscedastic models. Communications in Statistics: Theory and Methods, 38(15):2463–2478, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chalita:2002:LAD

[CCD02] Liciana V. A. S. Chalita, Enrico A. Colosimo, and Clarice G. B. Demétrio. Likelihood approximations and discrete models for tied survival data. Communications in Statistics: Theory and Methods, 31(7):1215–1229, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chai:2009:GCM

[CCG09] Feng-Shun Chai, Kashinath Chatterjee, and Sudhir Gupta. Generalized E(s^{2}) criterion for multilevel supersaturated designs. Communications in Statistics: Theory and Methods, 38(20):3725–3735, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Contreras-Cristan:2003:SMD

[CCGB03] Alberto Contreras-Cristán and Jose M. González-Barrios. A statistical method for the determination of the appropriate order in a general class of time series models. Communications in Statistics: Theory and Methods, 32(4):875–891, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Contreras-Cristan:2003:IUL

[CCGPO03] Alberto Contreras-Cristán, Eduardo Gutiérrez-Peña, and Federico O’Reilly. Inference using latent variables for mixtures of distributions for censored data with partial identification. Communications in Statistics: Theory and Methods, 32(4):749–774, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:2002:ARF

[CCK02] Chung-Chi Helen Chang, Wenyaw Chan, and Asha S. Kapadia. The analysis of recurrent failure times: the timedependent Yule process approach. Communications in Statistics: Theory and Methods, 31(7):1203–1213, 2002. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chatterjee:2003:PEC

[CD03] Snigdhansu Chatterjee and Samarjit Das. Parameter estimation in conditional heteroscedastic models. Communications in Statistics: Theory and Methods, 32(6):1135–1153, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaganty:2007:RMA

[CD07] N. Rao Chaganty and Yihao Deng. Ranges of measures of association for familial binary variables. Communications in Statistics: Theory and Methods, 36(3):587–598, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:2008:CEE

[CD08] Gauss M. Cordeiro and Clarice G. B. Demétrio. Corrected estimators in extended quasi-likelihood models. Communications in Statistics: Theory and Methods, 37(6):873–880, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaudhuri:2006:FBH

[CDB06] Arijit Chaudhuri, Kajal Dihidar, and Mausumi Bose. On the feasibility of basing Horvitz and Thompson’s estimator on a sample by Rao, Hartley, and Cochran’s scheme. Communications in Statistics: Theory and Methods, 35(12):2239–2244, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chung:2005:BMC

[CDKK05] Younshik Chung, Dipak K. Dey, Misook Kim, and Chansoo Kim. Bayesian model choice in exponential survival models. Communications in Statistics: Theory and Methods, 34 (12):2311–2330, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chatterjee:2000:RVS

[CDL00] K. Chatterjee, Lih-Yuan Deng, and Dennis K. J. Lin. Resolution v.2 search designs. Communications in Statistics: Theory and Methods, 29(5–6):1143–1154, 2000. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chung:2001:ASD

[CDL01] Younshik Chung, Dipak K. Dey, and Sangjeen Lee. Applying the Savage–Dickey density ratio to default Bayes factors, with an illustration to outlier detection in random effects models. Communications in Statistics: Theory and Methods, 30(12):2563–2582, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:2001:LDS

[CF01] P. C. Consul and Felix Famoye. On Lagrangian distributions of the second kind. Communications in Statistics: Theory and Methods, 30(1):165–178, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:2005:TOA

[CF05] Gauss M. Cordeiro and Silvia L. P. Ferrari. Third-order asymptotic distributions of classic test statistics for oneparameter exponential family models. Communications in Statistics: Theory and Methods, 34(5):1041–1055, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2000:SED

[CFB00] Cong Chen, Wayne A. Fuller, and F. Jay Breidt. A semiparametric estimator of the distribution function of a variable measured with error. Communications in Statistics: Theory and Methods, 29(5–6):1293–1310, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cho:2004:RES

[CG04] Hokwon Cho and Z. Govindarajulu. Risk-efficient sequential estimation of the number of multinomial cells. Communications in Statistics: Theory and Methods, 33(9):2159–2174, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:2007:EVD

[CG07a] Sheng-Mao Chang and Marc G. Genton. Extreme value distributions for the skew-symmetric family of distributions. Communications in Statistics: Theory and Methods, 36(9): 1705–1717, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ciuperca:2007:EER

[CG07b] Gabriela Ciuperca and Valerie Girardin. Estimation of the entropy rate of a countable Markov chain. Communications in Statistics: Theory and Methods, 36(14):2543–2557, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cooray:2006:FLD

[CGA06] Kahadawala Cooray, Sumith Gunasekera, and Malwane M. A. Ananda. The folded logistic distribution. Communications in Statistics: Theory and Methods, 35(3):385–393, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Caeiro:2009:RBT

[CGR09] Frederico Caeiro, M. Ivette Gomes, and Lígia Henriques Rodrigues. Reduced-bias tail index estimators under a thirdorder framework. Communications in Statistics: Theory and Methods, 38(7):1019–1040, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cook:2002:SCI

[CGT02] R. J. Cook, V. P. Godambe, and M. E. Thompson. Semiparametric confidence intervals based on estimating functions. Communications in Statistics: Theory and Methods, 31(10):1701–1708, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2003:DSR

[CGT03] John T. Chen, Arjun K. Gupta, and Cas G. Troskie. The distribution of stock returns when the market is up. Communications in Statistics: Theory and Methods, 32(8):1541– 1558, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Choi:2000:TNN

[CH00] Hyun Jip Choi and Chong Sun Hong. Testing non-nested log-linear models with pseudo estimator. Communications in Statistics: Theory and Methods, 29(7):1539–1547, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2009:SPN

[CH09] Pinyuen Chen and Lifang Hsu. A selection procedure for the number of signals in presence of colored noise. Communications in Statistics: Theory and Methods, 38(10):1741–1757, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chan:2000:ODE

[Cha00] Ling-Yau Chan. Optimal designs for experiments with mixtures: a survey. Communications in Statistics: Theory and Methods, 29(9–10):2281–2312, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610920008832607.

Chattopadhyay:2004:RPW

[Cha04a] Gopaldeb Chattopadhyay. Randomized play-the-winner rule for ordered categorical data. Communications in Statistics: Theory and Methods, 33(11):2683–2708, 2004. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chavance:2004:HMI

[Cha04b] M. Chavance. Handling missing items in quality of life studies. Communications in Statistics: Theory and Methods, 33 (6):1371–1383, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Charalambides:2007:EDN

[Cha07] Ch. A. Charalambides. Exact distributions of the number of r-records and the r-record and inter-r-record times. Communications in Statistics: Theory and Methods, 36(7):1305– 1317, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chakraborty:2008:SNM

[Cha08] Subrata Chakraborty. On some new α-modified binomial and Poisson distributions and their applications. Communications in Statistics: Theory and Methods, 37(11):1755– 1769, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Charalambides:2009:DRS

[Cha09] Ch. A. Charalambides. Distribution of record statistics in a q-factorially increasing population. Communications in Statistics: Theory and Methods, 38(12):2042–2055, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cramer:2009:PED

[CHB09] E. Cramer, K. Herle, and N. Balakrishnan. Permanent expansions and distributions of order statistics in the INID case. Communications in Statistics: Theory and Methods, 38(12):2078–2088, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2007:MAP

[CHC07] Lin-An Chen, Hui-Nien Hung, and Chih-Rung Chen. Maximum average-power (MAP) tests. Communications in Statistics: Theory and Methods, 36(12):2237–2249, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2000:GAM

[Che00] Colin Chen. Generalized additive mixed models. Communications in Statistics: Theory and Methods, 29(5–6):1257– 1271, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chechile:2001:BAG

[Che01] Richard A. Chechile. Bayesian analysis of Gumbel distributed data. Communications in Statistics: Theory and Methods, 30(3):485–496, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:2005:VES

[Che05] Fuxia Cheng. On variance estimation in semiparametric regression models. Communications in Statistics: Theory and Methods, 34(8):1737–1742, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2006:MPA

[Che06] Chung-Ho Chen. The modified Pulak and Al-Sultan’s model for determining the optimum process parameters. Communications in Statistics: Theory and Methods, 35(10):1767– 1778, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2008:IME

[Che08] John T. Chen. Inference on the minimum effective dose using binary data. Communications in Statistics: Theory and Methods, 37(13):2124–2135, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chiang:2000:MLS

[Chi00] Chin Long Chiang. My life as a statistician. Communications in Statistics: Theory and Methods, 29(5–6):905–921, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chicken:2007:NRS

[Chi07] Eric Chicken. Nonparametric regression with sample design following a random process. Communications in Statistics: Theory and Methods, 36(10):1915–1934, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chi:2009:DVP

[Chi09] Zhiyi Chi. Data volume and power of multiple tests with small sample size per null. Communications in Statistics: Theory and Methods, 38(16–17):2784–2803, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Castillo:2001:RDL

[CHL01] Enrique Castillo, Ali S. Hadi, and Beatriz Lacruz. Regression diagnostics for the least absolute deviations and the minimax methods. Communications in Statistics: Theory and Methods, 30(6):1197–1225, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Choi:2000:ADM

[Cho00] ByoungSeon Choi. On the asymptotic distributions of mean, autocovariance, autocorrelation, crosscovariance and impulse response estimators of a stationary multidimensional random field. Communications in Statistics: Theory and Methods, 29(8):1703–1724, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Choi:2001:AJD

[Cho01] ByoungSeon Choi. The asymptotic joint distribution of the Yule–Walker estimators of a causal multidimensional AR process. Communications in Statistics: Theory and Methods, 30(4):609–614, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Choi:2005:LE

[Cho05] Jeen-Kap Choi. Letter to the Editor. Communications in Statistics: Theory and Methods, 34(4):997–998, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2005:P

[CHP05] Pinyuen Chen, Lifang Hsu, and S. Panchapakesan. Preface. Communications in Statistics: Theory and Methods, 34(6): 1245, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:2008:AEN

[CHW08] Yi-Ping Chang, Wen-Tao Huang, and Hsiuying Wang. ϵadmissible estimators for normal and Poisson means. Communications in Statistics: Theory and Methods, 37(8):1181– 1192, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2008:NEM

[CHYL08] Feng Chen, Richard M. Huggins, Paul S. F. Yip, and K. F. Lam. Nonparametric estimation of multiplicative counting process intensity functions with an application to the Beijing SARS epidemic. Communications in Statistics: Theory and Methods, 37(2):294–306, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chitturi:2002:NOM

[CJ02] Pallavi Chitturi and Peter W. M. John. Nesting optimal main effects plans and optimal foldover designs. Communications in Statistics: Theory and Methods, 31(10):1777– 1794, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Commenges:2004:MSM

[CJ04] Daniel Commenges and Pierre Joly. Multi-state model for dementia, institutionalization, and death. Communications in Statistics: Theory and Methods, 33(6):1315–1326, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Choi:2002:MDI

[CK02] Byungjin Choi and Keeyoung Kim. Minimum discrimination information estimator of the inverse Gaussian mean with known coefficient of variation. Communications in Statistics: Theory and Methods, 31(1):1–10, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Choi:2004:CMS

[CK04a] Byungjin Choi and Keeyoung Kim. Certain multi sample tests for inverse Gaussian populations. Communications in Statistics: Theory and Methods, 33(7):1557–1576, 2004. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Christensen:2004:TFL

[CK04b] Karl Bang Christensen and Svend Kreiner. Testing the fit of latent regression models. Communications in Statistics: Theory and Methods, 33(6):1341–1356, 2004. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calinski:2005:CTP

[CK05] Tadeusz Caliński and Mirosław Krzyśko. A closed testing procedure for canonical correlations. Communications in Statistics: Theory and Methods, 34(5):1105–1116, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chiu:2007:ACC

[CK07] Jing-Er Chiu and Tsen-I Kuo. Attribute control chart for multivariate Poisson distribution. Communications in Statistics: Theory and Methods, 37(1):146–158, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Copas:2000:CPR

[CKB00] J. B. Copas and A. J. Krebs-Brown. Calibration of predictions in regression. Communications in Statistics: Theory and Methods, 29(9–10):1973–1986, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610920008832590.

Causeur:2009:CFM

[CKF09] D. Causeur, M. Kloareg, and C. Friguet. Control of the FWER in multiple testing under dependence. Communications in Statistics: Theory and Methods, 38(16–17):2733– 2747, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cramer:2004:CLR

[CKK04] E. Cramer, U. Kamps, and C. Keseling. Characterizations via linear regression of ordered random variables: A unifying approach. Communications in Statistics: Theory and Methods, 33(12):2885–2911, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chiang:2007:EIP

[CKL07] Min-Hsien Chiang, Chihwa Kao, and Chih-Hsien Lo. The estimation and inference of a panel cointegration model with a time trend. Communications in Statistics: Theory and Methods, 36(6):1233–1250, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chryssaphinou:2008:DTS

[CKL08] Ourania Chryssaphinou, Margarita Karaliopoulou, and Nikolaos Limnios. On discrete time semi-Markov chains and applications in words occurrences. Communications in Statistics: Theory and Methods, 37(8):1306–1322, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Coffey:2007:PMB

[CKM07] Christopher S. Coffey, John A. Kairalla, and Keith E. Muller. Practical methods for bounding Type I error rate with an internal pilot design. Communications in Statistics: Theory and Methods, 36(11):2143–2157, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cuadras:2000:SOE

[CL00] Caries M. Cuadras and Younes Lahlou. Some orthogonal expansions for the logistic distribution. Communications in Statistics: Theory and Methods, 29(12):2643–2663, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2001:AOS

[CL01] Juei-Chao Chen and Shen-Ming Lee. Approximation to optimal stopping rules for gamma random variables with unknown location and scale parameters. Communications in Statistics: Theory and Methods, 30(4):775–784, 2001. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2006:NET

[CL06a] Y. H. Joshua Chen and G. H. Frank Liu. A note on the estimate of treatment effect from a Cox regression model when the proportionality assumption is violated. Communications in Statistics: Theory and Methods, 35(3):521–526, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cui:2006:P

[CL06b] Lirong Cui and Haijun Li. Preface. Communications in Statistics: Theory and Methods, 35(10):1765–1766, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2008:MSP

[CL08a] Shijie Chen and P. Lahiri. On mean squared prediction error estimation in small area estimation problems. Communications in Statistics: Theory and Methods, 37(11):1792–1798, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Crimaldi:2008:ARG

[CL08b] Irene Crimaldi and Fabrizio Leisen. Asymptotic results for a generalized Pólya urn with “multi-updating” and applications to clinical trials. Communications in Statistics: Theory and Methods, 37(17):2777–2794, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2009:BVT

[CL09] Lee-Shen Chen and Che-Yang Lin. Bayesian p-values for testing independence in 2 × 2 contingency tables. Communications in Statistics: Theory and Methods, 38(10):1635– 1648, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cui:2008:EPD

[CLY08] Xia Cui, Lu Lin, and Guangren Yang. An extended projection data depth and its applications to discrimination. Communications in Statistics: Theory and Methods, 37(14):2276– 2290, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Coffey:2000:PDT

[CM00a] Christopher S. Coffey and Keith E. Muller. Properties of doubly-truncated gamma variables. Communications in Statistics: Theory and Methods, 29(4):851–857, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Coffey:2000:SDT

[CM00b] Christopher S. Coffey and Keith E. Muller. Seme distributions and their implications for an internal pilot study with a univariate linear model. Communications in Statistics: Theory and Methods, 29(12):2677–2691, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:2001:BCL

[CM01] Gauss M. Cordeiro and Lourdes Coral C. Montenegro. Bartlett corrected likelihood ratio tests in location-scale nonlinear models. Communications in Statistics: Theory and Methods, 30(7):1353–1372, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Campos:2005:CGG

[CM05] L. M. B. C. Campos and J. M. G. Marques. On a combination of gamma and generalized error distributions with application to aircraft flight path deviations. Communications in Statistics: Theory and Methods, 33(10):2307–2332, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chan:2008:TSM

[CM08] Victor Chan and William Q. Meeker. Time series modeling of degradation due to outdoor weathering. Communications in Statistics: Theory and Methods, 37(3):408–424, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cao:2009:TEG

[CM09] Li Cao and Thomas Mathew. Testing the equivalence of generalized variances in multivariate bioequivalence. Communications in Statistics: Theory and Methods, 38(16–17):2983– 2991, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Castano-Martinez:2005:DSW

[CMLB05] Antonia Castaño-Martínez and Fernando López-Blázquez. Distribution of a sum of weighted central chi-square variables. Communications in Statistics: Theory and Methods, 34(3):515–524, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chacon:2007:UBF

[CMNP07] J. E. Chacón, J. Montanero, A. G. Nogales, and P. Pérez. On the use of Bayes factor in frequentist testing of a precise hypothesis. Communications in Statistics: Theory and Methods, 36(12):2251–2261, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chou:2001:CCI

[CMY01] Youn-Min Chou, Robert L. Mason, and John C. Young. The control chart for individual observations from a multivariate non-normal distribution. Communications in Statistics: Theory and Methods, 30(8–9):1937–1949, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cha:2009:ABP

[CN09] Ji Hwan Cha and Myung Hwan Na. Accelerated burn-in procedures and system maintenance policies. Communications in Statistics: Theory and Methods, 38(5):719–733, 2009. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cribari-Neto:2003:CIH

[CNG03] Francisco Cribari-Neto and Nila M. S. Galvão. A class of improved heteroskedasticity-consistent covariance matrix estimators. Communications in Statistics: Theory and Methods, 32(10):1951–1980, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cribari-Neto:2007:IUH

[CNSV07] Francisco Cribari-Neto, Tatiene C. Souza, and Klaus L. P. Vasconcellos. Inference under heteroskedasticity and leveraged data. Communications in Statistics: Theory and Methods, 36(10):1877–1888, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See erratum [CNSV08].

Cribari-Neto:2008:EIU

[CNSV08] Francisco Cribari-Neto, Tatiene C. Souza, and Klaus L. P. Vasconcellos. Errata: Inference Under Heteroskedasticity and Leveraged Data, Communications in Statistics, Theory and Methods, 36, 1877–1888, 2007. Communications in Statistics: Theory and Methods, 37(20):3329–3330, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [CNSV07].

Cribari-Neto:2000:SPA

[CNZ00] Francisco Cribari-Neto and Spyeos G. Zarkos. The size and power of analytical and bootstrap corrections to score tests in regression models. Communications in Statistics: Theory and Methods, 29(2):279–289, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Coelho:2000:TAN

[Coe00] Carlos A. Coelho. On two asymptotic normal distributions for the generalized Wilks lambda statistic. Communications in Statistics: Theory and Methods, 29(7):1465–1486, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Collins:2001:FEA

[Col01] Joseph Collins. Functional estimation by asymptotic regression. Communications in Statistics: Theory and Methods, 30 (12):2601–2635, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Collins:2003:NFE

[Col03] Joseph Collins. Nonparametric functional estimation by asymptotic regression. Communications in Statistics: Theory and Methods, 32(6):1169–1195, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cook:2000:SMD

[Coo00] R. Dennis Cook. Save: a method for dimension reduction and graphics in regression. Communications in Statistics: Theory and Methods, 29(9–10):2109–2121, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610920008832598.

Cooray:2009:WPC

[Coo09] Kahadawala Cooray. The Weibull–Pareto composite family with applications to the analysis of unimodal failure rate data. Communications in Statistics: Theory and Methods, 38(11):1901–1915, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chan:2000:LMS

[CP00] Ngai Hang Chan and Giovanni Petris. Long memory stochastic volatility: a Bayesian approach. Communications in Statistics: Theory and Methods, 29(5–6):1367–1378, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2004:DMT

[CP04] Pinyuen Chen and S. Panchapakesan. Detecting multiple targets simultaneously at k sites. Communications in Statistics: Theory and Methods, 33(7):1667–1688, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:2009:SPB

[CP09] Shuenn-Ren Cheng and S. Panchapakesan. Is the selected population the best-location and scale parameter cases. Communications in Statistics: Theory and Methods, 38(10): 1553–1560, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaudhry:2005:IEH

[CQ05] M. A. Chaudhry and Asghar Qadir. Incomplete exponential and hypergeometric functions with applications to the non central χ^{2}-distribution. Communications in Statistics: Theory and Methods, 34(3):525–535, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Carolan:2000:OST

[CR00] A. M. Carolan and J. C. W. Rayner. One sample tests of location for nonnormal symmetric data. Communications in Statistics: Theory and Methods, 29(7):1569–1581, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cramer:2001:ISS

[Cra01] Erhard Cramer. Inference for stress-strength models based on Weinman multivariate exponential samples. Communications in Statistics: Theory and Methods, 30(2):331–346, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cramer:2002:NMP

[Cra02] Erhard Cramer. A note on moments of progressively Type II censored order statistics. Communications in Statistics: Theory and Methods, 31(8):1301–1307, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaganty:2004:NEA

[CS04a] N. Rao Chaganty and Genming Shi. A note on the estimation of autocorrelation in repeated measurements. Communications in Statistics: Theory and Methods, 33(5):1157–1170, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2004:QAS

[CS04b] Pai-Lien Chen and Pranab K. Sen. Quality-adjusted survival estimation with periodic observations: A multistate survival analysis approach. Communications in Statistics: Theory and Methods, 33(6):1327–1339, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:2006:EOM

[CS06a] Yuan-Tsung Chang and Nobuo Shinozaki. Estimation of ordered means of two Poisson distributions. Communications in Statistics: Theory and Methods, 35(11):1993–2003, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:2006:SSO

[CS06b] Gauss M. Cordeiro and Edson Marcos Leal Soares Ramos. Some second-order asymptotics for extreme value linear regression models. Communications in Statistics: Theory and Methods, 35(1):197–207, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cano:2007:OBA

[CS07a] J. A. Cano and D. Salmerón. Objective Bayesian analysis of an exponential regression model with constrained parameters applied to animal digestibility. Communications in Statistics: Theory and Methods, 36(13):2463–2473, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2007:MCM

[CS07b] Pai-Lien Chen and Pranab K. Sen. Markov chain model selection by misclassified model probabilities. Communications in Statistics: Theory and Methods, 36(1):143–153, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cardot:2008:VCF

[CS08a] Hervé Cardot and Pascal Sarda. Varying-coefficient functional linear regression models. Communications in Statistics: Theory and Methods, 37(20):3186–3203, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chavent:2008:CTD

[CS08b] Marie Chavent and Jérôme Saracco. On central tendency and dispersion measures for intervals and hypercubes. Communications in Statistics: Theory and Methods, 37(9):1471– 1482, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:2008:CMF

[CS08c] Gauss M. Cordeiro and Rosangela G. Santana. Covariance matrix formula for exponential family nonlinear models. Communications in Statistics: Theory and Methods, 37 (17):2724–2734, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:2009:COD

[CS09a] Fu-Chuen Chang and Yang-Chan Su. Computing A-optimal designs for weighted polynomial regression by Taylor expansion. Communications in Statistics: Theory and Methods, 38 (10):1622–1634, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2009:MIA

[CS09b] Ling Chen and Jianguo Sun. A multiple imputation approach to the analysis of current status data with the additive hazards model. Communications in Statistics: Theory and Methods, 38(7):1009–1018, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaubey:2008:IOT

[CSM08] Yogendra P. Chaubey, Debaraj Sen, and Satya N. Mishra. Inference on overlap for two inverse Gaussian populations: Equal means case. Communications in Statistics: Theory and Methods, 37(12):1880–1894, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Choi:2009:ABB

[CSP^{+}09] Dongseok Choi, Srilakshmi M. Sharma, Sirichai Pasadhika, Zhixin Kang, Christina A. Harrington, Justine R. Smith, Stephen R. Planck, and James T. Rosenbaum. Application of biostatistics and bioinformatics tools to identify putative transcription factor-gene regulatory network of ankylosing spondylitis and sarcoidosis. Communications in Statistics: Theory and Methods, 38(18):3326–3338, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Collins:2004:MAB

[CSV04] J. R. Collins and D. Szatmari-Voicu. Maximal asymptotic biases of M-estimators of location with preliminary scale estimates. Communications in Statistics: Theory and Methods, 33(8):1877–1886, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2009:URS

[CSWM09] Haiying Chen, Elizabeth A. Stasny, Douglas A. Wolfe, and Steven N. MacEachern. Unbalanced ranked set sampling for estimating a population proportion under imperfect rankings. Communications in Statistics: Theory and Methods, 38 (12):2116–2125, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2004:SMM

[CT04] Pai-Lien Chen and Hsiao-Chuan Tien. Semi-Markov models for multistate data analysis with periodic observations. Communications in Statistics: Theory and Methods, 33(3): 475–486, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:2008:BCG

[CU08] Gauss M. Cordeiro and Margareth C. Toyama Udo. Bias correction in generalized nonlinear models with dispersion covariates. Communications in Statistics: Theory and Methods, 37(14):2219–2225, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calo:2008:FRL

[CV08] Daniela G. Caló and Cinzia Viroli. Finding relevant linear manifolds in classification by Gaussian mixtures. Communications in Statistics: Theory and Methods, 37(19):3040– 3053, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2001:MLF

[CW01] M. H. Chen and P. C. Wang. Multi-level factorial designs with minimum numbers of level changes. Communications in Statistics: Theory and Methods, 30(5):875–885, 2001. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Choi:2006:EPT

[CWG06] Eun-Ha Choi, Wayne A. Woodward, and Henry L. Gray. Euler (p, q) processes and their application to non stationary time series with time varying frequencies. Communications in Statistics: Theory and Methods, 35(12):2245–2262, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaturvedi:2001:SRR

[CWS01] Anoop Chaturvedi, Alan T. K. Wan, and Shri P. Singh. Stein-rule restricted regression estimator in a linear regression model with nonspherical disturbances. Communications in Statistics: Theory and Methods, 30(1):55–68, 2001. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chigira:2007:FSM

[CY07] Hiroaki Chigira and Taku Yamamoto. Finite sample modifications of the Granger non causality test in cointegrated vector autoregressions. Communications in Statistics: Theory and Methods, 36(5):981–1003, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2009:IVT

[CY09] Lee-Shen Chen and Ming-Chung Yang. Improved p-values for testing marginal homogeneity in 2 × 2 contingency tables. Communications in Statistics: Theory and Methods, 38 (10):1649–1663, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chatterjee:2000:AET

[CYHL00] Sangit Chatterjee, Mustafa R. Yilmaz, Mohamed Habibuliah, and Matthew Laudato. An approximate entropy test for randomness. Communications in Statistics: Theory and Methods, 29(3):655–675, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cerutti:2003:EAG

[CZ03] Bernard Cerutti and Toufik Zoubeidi. An efficient allocation in group sequential tests for the linear mixed effects model. Communications in Statistics: Theory and Methods, 32(3): 585–599, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cui:2007:NSS

[CZ07] Wen Cui and Feiqi Zhu. A note on sample size determination for the estimation of the mean vector of a multivariate population. Communications in Statistics: Theory and Methods, 36(8):1607–1610, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2008:EFC

[CZA08] Xinjia Chen, Kemin Zhou, and Jorge L. Aravena. Explicit formula for constructing binomial confidence interval with guaranteed coverage probability. Communications in Statistics: Theory and Methods, 37(8):1173–1180, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2008:UBR

[CZL08] Feng Chen, Jinxia Zhu, and Zehui Li. Upper bounds for the ruin probabilities of the entrance-based risk model. Communications in Statistics: Theory and Methods, 37(16):2634– 2652, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cerutti:2002:AGS

[CZR02] Bernard Cerutti, Toufik Zoubeidi, and David M. Reboussin. Adaptive group sequential procedures for multiple comparison of treatments. Communications in Statistics: Theory and Methods, 31(5):781–793, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Damico:2005:NOS

[Dam05] Joe Damico. A new one-sample test for goodness-of-fit. Communications in Statistics: Theory and Methods, 33(1):181– 193, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Danielak:2005:SUB

[Dan05] Katarzyna Danielak. Sharp upper bounds for expectations of differences of order statistics in various scale units. Communications in Statistics: Theory and Methods, 33(4):787–803, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Doosti:2008:WNS

[DAN08] H. Doosti, M. Afshari, and H. A. Niroumand. Wavelets for nonparametric stochastic regression with mixing stochastic process. Communications in Statistics: Theory and Methods, 37(3):373–385, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Darkhovsky:2009:MEF

[Dar09] Boris Darkhovsky. Minimax estimation of the first derivative by finite number of noisy observations. Communications in Statistics: Theory and Methods, 38(16–17):2804– 2811, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Das:2007:ILG

[Das07] Shubhabrata Das. Information loss in grade point conversion. Communications in Statistics: Theory and Methods, 36 (13):2377–2393, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Datta:2003:STM

[Dat03] Susmita Datta. Statistical techniques for microarray data: A partial overview. Communications in Statistics: Theory and Methods, 32(1):263–280, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dang:2005:MUS

[DATM05] Qianyu Dang, Stewart Anderson, Lingshi Tan, and Sati Mazumdar. Modeling unequally spaced bivariate growth curve data using a Kalman filter approach. Communications in Statistics: Theory and Methods, 34(8):1821–1831, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dauxois:2003:EDE

[Dau03] Jean-Yves Dauxois. Estimating the deviation from exponentiality under random censorship. Communications in Statistics: Theory and Methods, 32(11):2117–2137, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dhar:2006:CBG

[DB06] Sunil K. Dhar and Srinivasan Balaji. On the characterization of a bivariate geometric distribution. Communications in Statistics: Theory and Methods, 35(5):759–765, 2006. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DAmbra:2005:CTW

[DBA05] Luigi D’Ambra, Eric J. Beh, and Pietro Amenta. CATANOVA for two-way contingency tables with ordinal variables using orthogonal polynomials. Communications in Statistics: Theory and Methods, 34(8):1755–1769, 2005. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

deBruin:2001:SSL

[dBSSvV01] R. de Bruin, W. Schaafsma, A. G. M. Steerneman, and G. N. van Vark. Size and shape as linear combinations of untransformed measurements. Communications in Statistics: Theory and Methods, 30(8–9):1695–1714, 2001. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeLeon:2000:OSL

[DC00] A. R. De Leon and K. C. Carriere. On the one-sample location hypothesis for mixed bivariate data. Communications in Statistics: Theory and Methods, 29(11):2573–2581, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

deCristofaro:2002:IRS

[dC02] Rodolfo de Cristofaro. The inductive reasoning in statistical inference. Communications in Statistics: Theory and Methods, 31(7):1079–1089, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Doosti:2007:WBM

[DCN07] Hassan Doosti, Yogendra P. Chaubey, and Hosein Ali Niroumand. On a wavelet-based method of estimating a regression function. Communications in Statistics: Theory and Methods, 36(11):2083–2098, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dragut:2004:CBA

[DD04] A. B. Dragut and M. Dragut. Clustering based approximation procedure for semi-Markov decision processes with incomplete state information. Communications in Statistics: Theory and Methods, 33(3):747–768, 2004. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dembinska:2008:MRD

[DD08] Anna Dembińska and Katarzyna Danielak. On moments of k-records from discrete distributions. Communications in Statistics: Theory and Methods, 37(16):2516–2531, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeLaHorra:2005:RCP

[De 05] Julián De La Horra. Reconciling classical and prior predictive p-values in the two-sided location parameter testing problem. Communications in Statistics: Theory and Methods, 34(3):575–583, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DelaHorra:2007:BRP

[De 07] Julián De la Horra. Bayesian robustness of the positive false discovery rate. Communications in Statistics: Theory and Methods, 36(10):1905–1914, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DelaHorra:2008:BMS

[De 08] Julián De la Horra. Bayesian model selection: Measuring the χ^{2} discrepancy with the uniform distribution. Communications in Statistics: Theory and Methods, 37(9):1412–1424, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Devarajan:2009:TCE

[DE09] Karthik Devarajan and Nader Ebrahimi. Testing for covariate effect in the Cox proportional hazards regression model. Communications in Statistics: Theory and Methods, 38(14):2333–2347, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Debbarh:2008:SUL

[Deb08] Mohammed Debbarh. Some uniform limit results in additive regression model. Communications in Statistics: Theory and Methods, 37(19):3090–3114, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deheuvels:2007:WMT

[Deh07] Paul Deheuvels. Weighted multivariate tests of independence. Communications in Statistics: Theory and Methods, 36(14):2477–2491, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dembinska:2007:RCD

[Dem07] Anna Dembińska. A review on characterizations of discrete distributions based on records and k th records. Communications in Statistics: Theory and Methods, 36(7):1381–1387, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Demirtas:2009:MIL

[Dem09] Hakan Demirtas. Multiple imputation for longitudinal data under a Bayesian multilevel model. Communications in Statistics: Theory and Methods, 38(16–17):2812–2828, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deng:2008:NML

[DF08] Dianliang Deng and Hong-Bin Fang. On nonparametric maximum likelihood estimations of multivariate distribution function based on interval-censored data. Communications in Statistics: Theory and Methods, 38(1):54–74, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Delattre:2004:LDM

[DFP04] Sylvain Delattre, Jean-Claude Fort, and Gilles Pagès. Local distortion and μ-mass of the cells of one dimensional asymptotically optimal quantizers. Communications in Statistics: Theory and Methods, 33(5):1087–1117, 2004. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Draper:2000:API

[DG00] Norman R. Draper and Irwin Guttman. Assessing prior information for designed experiments. Communications in Statistics: Theory and Methods, 29(9–10):2201–2214, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/03610920008832603.

Daimon:2008:MSE

[DG08] Takashi Daimon and Masashi Goto. The mean squared error optimum design criterion for parameter estimation in nonlinear regression models. Communications in Statistics: Theory and Methods, 37(4):508–519, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Diaz-Garcia:2003:DNC

[DGLS03] José A. Díaz-García and Víctor Leiva-Sánchez. Doubly noncentral t and distributions obtained under singular and nonsingular elliptic distributions. Communications in Statistics: Theory and Methods, 32(1):11–32, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Diaz-Garcia:2002:SED

[DGLSG02] José A. Díaz-García, Víctor Leiva-Sánchez, and Manuel Galea. Singular elliptical distribution: Density and applications. Communications in Statistics: Theory and Methods, 31(5):665–681, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Diebolt:2007:ANE

[DGR07] Jean Diebolt, Armelle Guillou, and Pierre Ribereau. Asymptotic normality of extreme quantile estimators based on the peaks-over-threshold approach. Communications in Statistics: Theory and Methods, 36(5):869–886, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Diaz-Garcia:2003:IDE

[DGRLS03] José A. Díaz-García, Manuel Galea Rojas, and Víctor Leiva-Sánchez. Influence diagnostics for elliptical multivariate linear regression models. Communications in Statistics: Theory and Methods, 32(3):625–641, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Diaz-Garcia:2001:AOR

[DGRQ01] José A. Díaz-García and Rogelio Ramos-Quiroga. An approach to optimization in response surface methodology. Communications in Statistics: Theory and Methods, 30(5): 827–835, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeGooijer:2001:MSK

[DGZ01] Jan G. De Gooijer, Ali Gannoun, and Dawit Zerom. Multistage kernel-based conditional quantile prediction in time series. Communications in Statistics: Theory and Methods, 30 (12):2499–2515, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeGooijer:2006:MQP

[DGZ06] Jan G. De Gooijer, Ali Gannoun, and Dawit Zerom. A multivariate quantile predictor. Communications in Statistics: Theory and Methods, 35(1):133–147, 2006. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Demirhan:2006:BAE

[DH06] H. Demirhan and C. Hamurkaroglu. A Bayesian approach to the estimation of expected cell counts by using log-linear models. Communications in Statistics: Theory and Methods, 35(2):325–335, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:2001:SPG

[dHCC01] Hérbetes de H. Cordeiro and Gauss M. Cordeiro. Skewness for parameters in generalized linear models. Communications in Statistics: Theory and Methods, 30(7):1317–1334, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeGregorio:2008:LSV

[DI08] Alessandro De Gregorio and Stefano M. Iacus. Least squares volatility change point estimation for partially observed diffusion processes. Communications in Statistics: Theory and Methods, 37(15):2342–2357, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ding:2001:PMP

[Din01] Keyue Ding. Properties of the minimum point of an unbalanced two-sided random walk. Communications in Statistics: Theory and Methods, 30(11):2393–2414, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Director:2006:AAM

[Dir06] William B. Smith Executive Director. An appreciation of Anant M. Kshirsagar. Communications in Statistics: Theory and Methods, 35(1):1–2, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dixit:2003:EPM

[Dix03] Vaijayanti U. Dixit. Estimation of parameters of mixed failure time distribution based on an extended modified sampling scheme. Communications in Statistics: Theory and Methods, 32(10):1911–1923, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dong:2000:BKL

[DJ00] Jianping Dong and Renfang Jiang. A boundary kernel for local polynomial regression. Communications in Statistics: Theory and Methods, 29(7):1549–1558, 2000. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Durante:2008:ACC

[DJ08] Fabrizio Durante and Piotr Jaworski. Absolutely continuous copulas with given diagonal sections. Communications in Statistics: Theory and Methods, 37(18):2924–2942, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dong:2009:MPE

[DJ09] Jianping Dong and Renfang Jiang. Multinomial probability estimation by wavelet thresholding. Communications in Statistics: Theory and Methods, 38(9):1486–1507, 2009. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deshmukh:2002:BDT

[DK02] S. R. Deshmukh and M. S. Kasture. Bivariate distribution with truncated Poisson marginal distributions. Communications in Statistics: Theory and Methods, 31(4):527–534, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

deLeon:2007:OSL

[dL07] A. R. de Leon. One-sample likelihood ratio tests for mixed data. Communications in Statistics: Theory and Methods, 36(1):129–141, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dembinska:2005:CGD

[DLb05] Anna Dembińska and Fernando López-blázquez. A characterization of geometric distribution through k th weak records. Communications in Statistics: Theory and Methods, 34(12):2345–2351, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

delaCruz-Mesia:2003:BAN

[dlCMM03] Rolando de la Cruz-Mesía and Guillermo Marshall. A Bayesian approach for nonlinear regression models with continuous errors. Communications in Statistics: Theory and Methods, 32(8):1631–1646, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

delaHorra:2000:OPP

[dlHRB00] Julian de la Horra and Maria Teresa Rodriguez-Bernal. Optimality of the posterior predictive p-value based on the posterior odds. Communications in Statistics: Theory and Methods, 29(1):181–192, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

delaHorra:2003:BRP

[dlHRB03] Julian de la Horra and Maria Teresa Rodriguez-Bernal. Bayesian robustness of the posterior predictive p-value. Communications in Statistics: Theory and Methods, 32(8): 1493–1503, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

delaHorra:2006:PDS

[dlHRB06] Julián de la Horra and María Teresa Rodríguez-Bernal. Prior density selection as a particular case of Bayesian model selection: A predictive approach. Communications in Statistics:

Theory and Methods, 35(8):1387–1396, 2006. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ding:2008:SPD

[DLN08] Aidong Adam Ding, Jennifer Lin, and Tianhua Niu. A statistical procedure for detecting highly correlated genes with a pre-specified candidate gene in microarray analysis. Communications in Statistics: Theory and Methods, 37(18):2991– 3007, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deng:2000:CSR

[DLT00] Lih-Yuan Deng, Yingfu Li, and Boxin Tang. Catalogue of small runs nonregular designs from Hadamard matrices with generalized minimum aberration. Communications in Statistics: Theory and Methods, 29(5–6):1379–1395, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Duan:2009:ECD

[DLZ09] Xiaogang Duan, Liping Liu, and Ping Zhao. Estimation of covariate distribution with capture-recapture data. Communications in Statistics: Theory and Methods, 38(20):3705– 3712, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dasiou:2000:RCI

[DM00a] Despina Dasiou and Chronis Moyssiadis. Robustness of confidence intervals for scale parameters based on M-estimators. Communications in Statistics: Theory and Methods, 29(12): 2761–2773, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Derryberry:2000:SLA

[DM00b] De Wayne R. Derryberry and Paul A. Murtaugh. A scalelocation adjustment for proportional hazards deviance residuals. Communications in Statistics: Theory and Methods, 29 (11):2583–2591, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dupuy:2004:EAV

[DM04] J.-F. Dupuy and M. Mesbah. Estimation of the asymptotic variance of semiparametric maximum likelihood estimators in the Cox model with a missing time-dependent covariate. Communications in Statistics: Theory and Methods, 33(6):1385–1401, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Debbarh:2008:ARM

[DM08a] Mohammed Debbarh and Bertrand Maillot. Additive regression model for continuous time processes. Communications in Statistics: Theory and Methods, 37(15):2416–2432, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dubeau:2008:EOW

[DM08b] François Dubeau and Youness Mir. Existence of optimal weighted least squares estimate for three-parametric exponential model. Communications in Statistics: Theory and Methods, 37(9):1383–1398, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dominguez-Molina:2007:MVC

[DMGFRQG07] J. Armando Domínguez-Molina, Graciela González-Farías, Rogelio Ramos-Quiroga, and Arjun K. Gupta. A matrix variate closed skew-normal distribution with applications to stochastic frontier analysis. Communications in Statistics: Theory and Methods, 36(9):1691–1703, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Djauhari:2008:MPV

[DMH08] Maman A. Djauhari, Muhammad Mashuri, and Dyah E. Herwindiati. Multivariate process variability monitoring. Communications in Statistics: Theory and Methods, 37(11): 1742–1754, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dodunekova:2004:CGE

[Dod04] Rossitza Dodunekova. Characterization of the geometric and exponential random variables. Communications in Statistics: Theory and Methods, 33(8):1755–1765, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dolati:2008:DPR

[Dol08] Ali Dolati. On dependence properties of random minima and maxima. Communications in Statistics: Theory and Methods, 38(3):393–399, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Donev:2000:DEV

[Don00] A. N. Donev. Dealing with errors in variables in response surface exploration. Communications in Statistics: Theory and Methods, 29(9–10):2065–2077, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610920008832596.

Damilano:2006:NHL

[DP06] Gabriela Damilano and Pedro Puig. A note on the Hodges– Lehmann estimator and the logistic distribution. Communications in Statistics: Theory and Methods, 35(2):257–261, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dodge:2000:DDR

[DR00] Yadolah Dodge and Valentin Rousson. Direction dependence in a regression line. Communications in Statistics: Theory and Methods, 29(9–10):1957–1972, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610920008832589.

Dewanji:2001:SRU

[DR01] Anup Dewanji and Tummala S. Rao. On system reliability under stress-strength modeling. Communications in Statistics: Theory and Methods, 30(6):1185–1196, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Draper:2002:ARA

[Dra02] Norman R. Draper. Applied regression analysis bibliography update 2000–2001. Communications in Statistics: Theory and Methods, 31(11):2051–2075, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeLaHorra:2007:CPT

[DRB07] Julián De La Horra and María Teresa Rodríguez-Bernal. Computing probabilities of Type I error and Type II error of sequential Bayesian procedures for testing one-sided hypotheses. Communications in Statistics: Theory and Methods, 36(4):731–740, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Drezner:2006:LGB

[Dre06] Zvi Drezner. On the limit of the generalized binomial distribution. Communications in Statistics: Theory and Methods, 35(2):209–221, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dahiya:2001:MLE

[DSC01] Ram C. Dahiya, Paul G. Staneski, and N. Rao Chaganty. Maximum likelihood estimation of parameters of the truncated Cauchy distribution. Communications in Statistics: Theory and Methods, 30(8–9):1737–1750, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

deSa:2007:CSP

[dSFM07] Antonio Mauricio F. L. Miranda de Sá, Leonardo B. Felix, and Eduardo M. A. M. Mendes. Comments on “sums, products, and ratios of non-central beta variables” by saralees nadarajah. Communications in Statistics: Theory and Methods, 36(8):1505–1506, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [Nad05] and reply [Nad09].

Durante:2009:RPB

[DSPS09] Fabrizio Durante, Susanne Saminger-Platz, and Peter Sarkoci. Rectangular patchwork for bivariate copulas and tail dependence. Communications in Statistics: Theory and Methods, 38(15):2515–2527, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dembinska:2007:HMO

[DSW07] A. Dembińska, A. Stepanov, and J. Wesołowski. How many observations fall in a neighborhood of an order statistic? Communications in Statistics: Theory and Methods, 36(5): 851–867, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Du:2008:NWV

[Du08] Xunming Du. A note of wavelet variance. Communications in Statistics: Theory and Methods, 37(5):737–745, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

deUna-Alvarez:2006:CNE

[dUÁRC06] J. de Uña-Álvarez and A. Rodríguez-Casal. Comparing nonparametric estimators for length-biased data. Communications in Statistics: Theory and Methods, 35(5):905–919, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dubnicka:2006:CIM

[Dub06] Suzanne R. Dubnicka. Confidence intervals for the median of a finite population based on the sign test. Communications in Statistics: Theory and Methods, 35(4):551–569, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dupuy:2008:SLT

[Dup08] Jean-François Dupuy. Stratified logrank test of no randomized treatment effect with missing stratum information. Communications in Statistics: Theory and Methods, 37(17): 2760–2776, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dutilleul:2001:MFP

[Dut01] Pierre Dutilleul. Multi-frequential periodogram analysis and the detection of periodic components in time series. Communications in Statistics: Theory and Methods, 30(6):1063– 1098, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dutilleul:2007:NSC

[Dut07] Pierre Dutilleul. A note on sufficient conditions for valid unmodified t testing in correlation analysis with autocorrelated and heteroscedastic sample data. Communications in Statistics: Theory and Methods, 37(1):137–145, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeMartini:2005:CBM

[DV05] D. De Martini and E. Vespa. Copula-based models for the power of independence tests. Communications in Statistics: Theory and Methods, 34(12):2283–2297, 2005. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dodge:2000:IDD

[DW00] Yadolah Dodge and Joe Whittaker. The information for the direction of dependence in L_{1} regression. Communications in Statistics: Theory and Methods, 29(9–10):1945– 1955, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/03610920008832588.

Ding:2007:IBS

[DW07] Aidong Adam Ding and Weijing Wang. Inference for bivariate survival data by copula models adjusted for the boundary effect. Communications in Statistics: Theory and Methods, 36(16):2927–2936, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Do:2001:IBR

[DWB01] Kim-Anh Do, Xuemei Wang, and Bradley M. Broom. Importance bootstrap resampling for proportional hazards regression. Communications in Statistics: Theory and Methods, 30 (10):2173–2188, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeGooijer:2008:MMF

[DY08] Jan G. De Gooijer and Ao Yuan. MDL mean function selection in semiparametric kernel regression models. Communications in Statistics: Theory and Methods, 37(14):2237– 2248, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Daimon:2009:RCM

[DYKG09] Takashi Daimon, Toshihiro Yoshikawa, Tominori Kobayashi, and Masashi Goto. Relative curvature measure for heteroscedastic or non normal nonlinear regression. Communications in Statistics: Theory and Methods, 38(2):193–207, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dudewicz:2004:CUJ

[DZK04] Edward J. Dudewicz, Charles X. Zhang, and Zaven A. Karian. The completeness and uniqueness of Johnson’s system in skewness-kurtosis space. Communications in Statistics: Theory and Methods, 33(9):2097–2116, 2004. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eickhoff:2005:GLL

[EA05] Jens C. Eickhoff and Yasuo Amemiya. Generalized linear latent variable modeling for multi-group studies. Communications in Statistics: Theory and Methods, 34(9–10):1991– 2008, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Elian:2000:IMR

[EAN00] Silvia N. Elian, Carmen D. S. André, and Subhash C. Narula. Influence measure for the L_{1} regression. Communications in Statistics: Theory and Methods, 29(4):837–849, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eryilmaz:2007:RBS

[EB07] Serkan Eryilmaz and Muhammet Bekçi. Runs in a bivariate sequence over the first coordinate. Communications in Statistics: Theory and Methods, 36(7):1389–1395, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Escarela:2008:FSP

[EB08] Gabriel Escarela and Russell J. Bowater. Fitting a semiparametric mixture model for competing risks in survival data. Communications in Statistics: Theory and Methods, 37(2):277–293, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Esteve:2009:ITD

[EBF09] Anna Esteve, Eva Boj, and Josep Fortiana. Interaction terms in distance-based regression. Communications in Statistics: Theory and Methods, 38(19):3498–3509, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ecker:2008:BSM

[ED08] Mark D. Ecker and Victor De Oliveira. Bayesian spatial modeling of housing prices subject to a localized externality. Communications in Statistics: Theory and Methods, 37 (13):2066–2078, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Editor:2009:GCP

[Edi09a] Nitis Mukhopadhyay Guest Editor. A glimpse at the career of Professor shelemyahu zacks. Communications in Statistics: Theory and Methods, 38(16–17):2623–2624, 2009. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Editor:2009:GER

[Edi09b] Nitis Mukhopadhyay Guest Editor. Guest editorial: Recent advances in theory and applications of statistics. Communications in Statistics: Theory and Methods, 38(16–17):2621– 2622, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ennis:2009:HTE

[EE09] Daniel M. Ennis and John M. Ennis. Hypothesis testing for equivalence defined on symmetric open intervals. Communications in Statistics: Theory and Methods, 38(11):1792– 1803, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ennis:2008:CBP

[EEPL08] Daniel M. Ennis, John M. Ennis, Joseph Palen, and Richard E. Lampe. Confidence bounds for positive ratios of normal random variables. Communications in Statistics: Theory and Methods, 37(3):307–317, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Efromovich:2009:MTB

[Efr09] Sam Efromovich. Multiwavelets: Theory and bioinformatic applications. Communications in Statistics: Theory and Methods, 38(16–17):2829–2842, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Erdely:2006:CFA

[EGB06] Arturo Erdely and José M. González-Barrios. On the construction of families of absolutely continuous copulas with given restrictions. Communications in Statistics: Theory and Methods, 35(4):649–659, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Enguix-Gonzalez:2005:IAP

[EGMPMRPM05] A. Enguix-González, J. M. Muñoz-Pichardo, J. L. Moreno-Rebollo, and R. Pino-Mejías. Influence analysis in principal component analysis through power-series expansions. Communications in Statistics: Theory and Methods, 34(9– 10):2025–2046, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ebneshahrashoob:2005:RAP

[EGS05] Morteza Ebneshahrashoob, Tangan Gao, and Milton Sobel. Random accessibility as a parallelism to reliability studies on simple graphs. Communications in Statistics: Theory and Methods, 34(6):1423–1436, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Evans:2004:GFT

[EH04] Scott R. Evans, Ph.D. and David W. Hosmer, Jr., Ph.D. Goodness of fit tests in mixed effects logistic models characterized by clustering. Communications in Statistics: Theory and Methods, 33(5):1139–1155, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Elmore:2004:ECC

[EHT04] Ryan T. Elmore, Thomas P. Hettmansperger, and Hoben Thomas. Estimating component cumulative distribution functions in finite mixture models. Communications in Statistics: Theory and Methods, 33(9):2075–2086, 2004. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eidous:2005:IKE

[Eid05] Omar M. Eidous. On improving kernel estimators using line transect sampling. Communications in Statistics: Theory and Methods, 34(4):931–941, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eidous:2006:SML

[Eid06] Omar Eidous. A semiparametric model for line transect sampling. Communications in Statistics: Theory and Methods, 35(7):1211–1221, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eidous:2009:KMS

[Eid09] Omar Eidous. Kernel method starting with half-normal detection function for line transect density estimation. Communications in Statistics: Theory and Methods, 38(14):2366– 2378, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Evangelaras:2004:ALP

[EK04a] H. Evangelaras and C. Koukouvinos. Another look at projection properties of Hadamard matrices. Communications in Statistics: Theory and Methods, 33(7):1607–1620, 2004. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Evangelaras:2004:CAM

[EK04b] H. Evangelaras and C. Koukouvinos. Combined arrays with minimum number of runs and maximum estimation efficiency. Communications in Statistics: Theory and Methods, 33(7):1621–1628, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ekstrom:2002:NEV

[Eks02] Magnus Ekström. Nonparametric estimation of the variance of sample means based on nonstationary spatial data. Communications in Statistics: Theory and Methods, 31(10):1743– 1775, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eugene:2002:BND

[ELF02] Nicholas Eugene, Carl Lee, and Felix Famoye. Beta-normal distribution and its applications. Communications in Statistics: Theory and Methods, 31(4):497–512, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Escobar:2001:AEF

[EM01] Luis A. Escobar and William Q. Meeker. The asymptotic equivalence of the Fisher information matrices for Type I and Type II censored data from location-scale families. Communications in Statistics: Theory and Methods, 30(10):2211– 2225, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

El-Nouty:2005:PNA

[ENL05] Charles El-Nouty and Rémi Lancar. The presmoothed Nelson–Aalen estimator in the competing risk model. Communications in Statistics: Theory and Methods, 33(1):135– 151, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

El-Neweihi:2000:REB

[ENS00] Emad El-Neweihi and Bikas K. Sinha. Reliability estimation based on ranked set sampling. Communications in Statistics: Theory and Methods, 29(7):1583–1595, 2000. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Elal-Olivero:2009:NCN

[EOOPGB09] David Elal-Olivero, Juan F. Olivares-Pacheco, Héctor W. Gómez, and Heleno Bolfarine. A new class of non negative distributions generated by symmetric distributions. Communications in Statistics: Theory and Methods, 38(7):993– 1008, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ezzahrioui:2008:ARN

[EOS08] M’hamed Ezzahrioui and Elias Ould-Saïd. Asymptotic results of a nonparametric conditional quantile estimator for functional time series. Communications in Statistics: Theory and Methods, 37(17):2735–2759, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eryilmaz:2005:CSI

[Ery05] Serkan Eryilmaz. Concomitants in a sequence of independent nonidentically distributed random vectors. Communications in Statistics: Theory and Methods, 34(9–10):1925– 1933, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ellis:2009:TDB

[ES09] Steven P. Ellis and Jonathan W. Stewart. Temporal dependence and bias in meta-analysis. Communications in Statistics: Theory and Methods, 38(15):2453–2462, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fabian:2001:ICT

[Fab01] Zdenek Fabián. Induced cores and their use in robust parametric estimation. Communications in Statistics: Theory and Methods, 30(3):537–555, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fabian:2008:NMC

[Fab08] Zdenek Fabián. New measures of central tendency and variability of continuous distributions. Communications in Statistics: Theory and Methods, 37(2):159–174, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fabian:2009:CIN

[Fab09] Zdenek Fabián. Confidence intervals for a new characteristic of central tendency of distributions. Communications in Statistics: Theory and Methods, 38(11):1804–1814, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Falk:2003:DMA

[Fal03] Michael Falk. Domain of multivariate attraction via angular decomposition. Communications in Statistics: Theory and Methods, 32(11):2109–2116, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fan:2006:PES

[Fan06] Zhaozhi Fan. Parameter estimation of stable distributions. Communications in Statistics: Theory and Methods, 35(2): 245–255, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fan:2009:MDE

[Fan09] Zhaozhi Fan. Minimum-distance estimator for stable exponent. Communications in Statistics: Theory and Methods, 38(4):511–528, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Freitas:2003:SMS

[FBH03] Marta Afonso Freitas, Wagner Borges, and Linda Lee Ho. A statistical model for shelf life estimation using sensory evaluations scores. Communications in Statistics: Theory and Methods, 32(8):1559–1589, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Farrell:2000:LIB

[FC00] P. J. Farrell and N. G. Cadigan. Local influence in binary regression models, and its correspondence with global influence. Communications in Statistics: Theory and Methods, 29(2):349–368, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fan:2006:BEN

[FCL06] Tsai-Hung Fan, Kuo-Ching Chang, and Chung-Bow Lee. Bayesian estimation of the number of change points in simple linear regression models. Communications in Statistics: Theory and Methods, 35(4):689–710, 2006. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Finner:2008:ATP

[FDR08] Helmut Finner, Thorsten Dickhaus, and Markus Roters. Asymptotic tail properties of Student’s t-distribution. Communications in Statistics: Theory and Methods, 37(2):175– 179, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Finkelstein:2006:MFR

[FE06] Maxim Finkelstein and Veronica Esaulova. On mixture failure rates ordering. Communications in Statistics: Theory and Methods, 35(11):1943–1955, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ferrari:2005:APL

[FFCN05] Silvia L. P. Ferrari, Michel Ferreira Da Silva, and Francisco Cribari-Neto. Adjusted profile likelihood for two-parameter exponential family models. Communications in Statistics: Theory and Methods, 34(2):257–276, 2005. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Francisco-Fernandez:2001:LPR

[FFVF01] M. Francisco-Fernández and J. M. Vilar-Fernández. Local polynomial regression estimation with correlated errors. Communications in Statistics: Theory and Methods, 30(7): 1271–1293, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Francisco-Fernandez:2003:USC

[FFVFVF03] M. Francisco-Fernández, J. M. Vilar-Fernández, and J. A. Vilar-Fernández. On the uniform strong consistency of local polynomial regression under dependence conditions. Communications in Statistics: Theory and Methods, 32(12):2415– 2440, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fang:2000:FBD

[FFvR00] Kai-Tai Fang, Hong-Bin Fang, and Dietrich von Rosen. A family of bivariate distributions with non-elliptical contours. Communications in Statistics: Theory and Methods, 29(8): 1885–1898, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feddag:2003:GEE

[FGM03] Mohand-Larbi Feddag, Ion Grama, and Mounir Mesbah. Generalized estimating equations (GEE) for mixed logistic models. Communications in Statistics: Theory and Methods, 32(4):851–874, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fischer:2007:TTD

[FHK07] Matthias Fischer, Armin Horn, and Ingo Klein. Tukeytype distributions in the context of financial data. Communications in Statistics: Theory and Methods, 36(1):23–35, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fuh:2003:EPA

[FHL03] Cheng-Der Fuh, Inchi Hu, and Shih-Kuei Lin. Empirical performance and asset pricing in hidden Markov models. Communications in Statistics: Theory and Methods, 32(12):2477– 2512, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fierro:2007:THS

[Fie07] Raúl Fierro. Test of homogeneity for some population models based on counting processes. Communications in Statistics: Theory and Methods, 37(1):46–54, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Figueiredo:2007:MSL

[Fig07] Adelaide Figueiredo. Multi-sample likelihood ratio tests based on bipolar Watson distributions defined on the hypersphere. Communications in Statistics: Theory and Methods, 36(4):815–820, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fiori:2008:MKR

[Fio08] Anna Maria Fiori. Measuring kurtosis by right and left inequality orders. Communications in Statistics: Theory and Methods, 37(17):2665–2680, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fan:2009:EPP

[FJ09] Tsai-Hung Fan and Wen-Kuei Jian. Expected posterior priors in selecting the largest mean of the exponential distributions. Communications in Statistics: Theory and Methods, 38(10):1561–1575, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fotopoulos:2007:RPA

[FJC07] Stergios B. Fotopoulos, Venkata K. Jandhyala, and Kim Heng Chen. Regression properties for asymmetric generalized scale mixtures of multivariate Gaussian variables. Communications in Statistics: Theory and Methods, 36(3): 613–628, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fedorov:2005:MCC

[FJJZ05] Valerii Fedorov, Byron Jones, C. Matthew Jones, and Anatoly Zhigljavsky. Multi-center clinical trials with random enrollment: Theoretical approximations. Communications in Statistics: Theory and Methods, 34(4):955–985, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fedorov:2007:DAR

[FJSZ07] Valerii Fedorov, Byron Jones, Vippal Savani, and Anatoly Zhigljavsky. Deriving approximations in a random effects model for multicenter clinical trials with binary response. Communications in Statistics: Theory and Methods, 36(3): 629–644, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ferentinos:2006:MSE

[FK06] K. K. Ferentinos and K. X. Karakostas. More on shortest and equal tails confidence intervals. Communications in Statistics: Theory and Methods, 35(5):821–829, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fransen:2009:DSA

[FKB09] Jaap Fransen, Arthur Kavanaugh, and George F. Borm. Desirability scores for assessing multiple outcomes in systemic rheumatic diseases. Communications in Statistics: Theory and Methods, 38(18):3461–3471, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feigin:2000:CLD

[FL00] Paul D. Feigin and Yan P. Lumelskii. On confidence limits for the difference of two binomial parameters. Communications in Statistics: Theory and Methods, 29(1):131–141, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fedorov:2004:PEM

[FL04] Valerii V. Fedorov and Sergei L. Leonov. Parameter estimation for models with unknown parameters in variance. Communications in Statistics: Theory and Methods, 33(11):2627– 2657, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

From:2006:EPB

[FL06] Steven G. From and Linxiong Li. Estimation of the parameters of the Birnbaum–Saunders distribution. Communications in Statistics: Theory and Methods, 35(12):2157–2169, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Flynn:2005:MPL

[Fly05] Michael R. Flynn. On the moments of the 4-parameter lognormal distribution. Communications in Statistics: Theory and Methods, 34(4):745–751, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Freeman:2006:ETH

[FM06] Jade Freeman and Reza Modarres. Efficiency of t-test and Hotelling’s T^{2}-test after Box–Cox transformation. Communications in Statistics: Theory and Methods, 35(6):1109– 1122, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frias:2008:SMB

[FMAA08] M. P. Frías, M. D. Ruiz Medina, F. J. Alonso, and J. M. Angulo. Spectral-marginal-based estimation of spatiotemporal long-range dependence. Communications in Statistics: Theory and Methods, 38(1):103–114, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fokoue:2008:LVM

[Fok08] Ernest Fokoué. Latent variable models in heterogeneous spaces for observations of mixed types. Communications in Statistics: Theory and Methods, 38(1):42–53, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Forchini:2005:DRQ

[For05] G. Forchini. The distribution of a ratio of quadratic forms in noncentral normal variables. Communications in Statistics: Theory and Methods, 34(5):999–1008, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Forkman:2009:ETC

[For09] Johannes Forkman. Estimator and tests for common coefficients of variation in normal distributions. Communications in Statistics: Theory and Methods, 38(2):233–251, 2009. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fountain:2000:CCC

[Fou00] Robert L. Fountain. A class of closeness criteria. Communications in Statistics: Theory and Methods, 29(8):1865–1883, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fitzgerald:2001:AMC

[FP01] Mark Fitzgerald and Rick Picard. Accelerated Monte Carlo for particle dispersion. Communications in Statistics: Theory and Methods, 30(11):2459–2471, 2001. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frouin:2009:CRF

[FP09] Robert Frouin and Bruno Pelletier. Consistency of ridge function fields for varying nonparametric regression. Communications in Statistics: Theory and Methods, 38(8):1272– 1283, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Firinguetti:2008:SAP

[FR08] Luis Firinguetti and Hernán Rubio. Some asymptotic properties of ridge regression in a system of seemingly unrelated regression equations. Communications in Statistics: Theory and Methods, 37(15):2433–2446, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Franzen:2004:SFL

[Fra04] Stefan Franzén. SPRT fixed length confidence intervals. Communications in Statistics: Theory and Methods, 33(2): 305–319, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frick:2003:TSU

[Fri03] Hans Frick. Two statistics for unconditional exact tests in matched-pairs design. Communications in Statistics: Theory and Methods, 32(9):1789–1799, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frias:2008:PES

[FRMAA08] M. P. Frías, M. D. Ruiz-Medina, F. J. Alonso, and J. M. Angulo. Parameter estimation of self-similar spatial covariogram models. Communications in Statistics: Theory and Methods, 37(7):1011–1023, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

From:2003:CIG

[Fro03] Steven G. From. Confidence intervals for Gini’s diversity measure and Shannon’s entropy using adjusted proportions. Communications in Statistics: Theory and Methods, 32(5): 935–954, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

From:2007:SNB

[Fro07] Steven G. From. Some new bounds on the probability of extinction of a Galton–Watson process with numerical comparisons. Communications in Statistics: Theory and Methods, 36(10):1993–2009, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Finos:2004:NMF

[FS04] Livio Finos and Luigi Salmaso. Nonparametric multi-focus analysis for categorical variables. Communications in Statistics: Theory and Methods, 33(8):1931–1941, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frankowski:2005:SSB

[FS05] Krzysztof S. Frankowski and Milton Sobel. Solutions of some birthday problems using Dirichlet integrals. Communications in Statistics: Theory and Methods, 34(6):1393–1410, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fang:2007:IEB

[FSZ07] Xiang Fang, Walter W. Stroup, and Shunpu Zhang. Improved empirical Bayes estimation in group testing procedure for small proportions. Communications in Statistics: Theory and Methods, 36(16):2937–2944, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Finkelstein:2000:CCI

[FTV00] Mark Finkelstein, Howard G. Tucker, and Jerry Alan Veeh. Conservative confidence intervals for a single parameter. Communications in Statistics: Theory and Methods, 29(8): 1911–1928, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fu:2000:RES

[Fu,00] Wenjiang J. Fu, Ph.D. Ridge estimator in singular design with application to age-period-cohort analysis of disease rates. Communications in Statistics: Theory and Methods, 29(2):263–278, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fu:2005:NPE

[Fu05] Wenjiang J. Fu. A note on prediction error with collinearity. Communications in Statistics: Theory and Methods, 34(3): 619–623, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ferrari:2002:BTC

[FUOC02] Silvia L. P. Ferrari, Miguel A. Uribe-Opazo, and Gauss M. Cordeiro. Bartlett-type corrections for two-parameter exponential family models. Communications in Statistics: Theory and Methods, 31(6):901–924, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Franco:2002:RPS

[FV02] Manuel Franco and Juana María Vivo. Reliability properties of series and parallel systems from bivariate exponential models. Communications in Statistics: Theory and Methods, 31(12):2349–2360, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Franco:2007:GMG

[FV07] Manuel Franco and Juana María Vivo. Generalized mixtures of gamma and exponentials and reliability properties of the maximum from friday and patil bivariate exponential model. Communications in Statistics: Theory and Methods, 36(11):2011–2025, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Franco:2009:SAC

[FV09] Manuel Franco and Juana-María Vivo. Stochastic aging classes for the maximum statistic from friday and patil bivariate exponential distribution family. Communications in Statistics: Theory and Methods, 38(6):902–911, 2009. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fang:2004:BMS

[FW04] Zhide Fang and Douglas P. Wiens. Bayesian minimally supported D-optimal designs for an exponential regression model. Communications in Statistics: Theory and Methods, 33(5):1187–1204, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feng:2007:LRT

[FW07] Yanqin Feng and Jinde Wang. Likelihood ratio test against stochastic order in three way contingency tables. Communications in Statistics: Theory and Methods, 37(1):81–96, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fraker:2008:NPL

[FWB08] Shannon E. Fraker, William H. Woodall, and Howard S. Burkom. A note on the Poisson likelihood ratio test statistic for Kulldorff’s scan methods. Communications in Statistics: Theory and Methods, 37(7):998–1001, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fujioka:2000:CEM

[FY00] Teruo Fujioka and Takerni Yanagimoto. Constructing estimators of a mean vector through orthogonal reparametrization and differential equations. Communications in Statistics: Theory and Methods, 29(8):1851–1863, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Farsipour:2005:EGS

[FZ05] N. Sanjari Farsipour and H. Zakerzadeh. Estimation of a gamma scale parameter under asymmetric squared log error loss. Communications in Statistics: Theory and Methods, 34 (5):1127–1135, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Farsipour:2006:EGV

[FZ06] N. Sanjari Farsipour and H. Zakerzadeh. Estimation of generalized variance under an asymetric loss function “squared log error”. Communications in Statistics: Theory and Methods, 35(4):571–581, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goudie:2000:CUE

[GA00] I. B. J. Goudie and J. Ashbridge. A conditionally-unbiased estimator of population size based on plant-capture in continuous time. Communications in Statistics: Theory and Methods, 29(11):2605–2619, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goudie:2005:CUE

[GA05] I. B. J. Goudie and J. Ashbridge. A conditionally unbiased estimator for the equal-catchability model. Communications in Statistics: Theory and Methods, 34(7):1543–1553, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghitany:2007:MOE

[GAAA07] M. E. Ghitany, F. A. Al-Awadhi, and L. A. Alkhalfan. Marshall–Olkin extended Lomax distribution and its application to censored data. Communications in Statistics: Theory and Methods, 36(10):1855–1866, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garcia-Alfaro:2001:CCS

[GAB01] K. H. Garcia-Alfaro and H. Bolfarine. Comparative calibration with subgroups. Communications in Statistics: Theory and Methods, 30(10):2057–2078, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gallagher:2002:TLD

[Gal02] Colin M. Gallagher. Testing for linear dependence in heavytailed data. Communications in Statistics: Theory and Methods, 31(4):611–623, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghitany:2009:RSL

[GAMG09] M. E. Ghitany, D. K. Al-Mutairi, and Ramesh C. Gupta. Reliability studies of the log-exponential inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 38(19):3532–3543, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gani:2004:SCM

[Gan04] J. Gani. Stochastic compartmental models for diabetes. Communications in Statistics: Theory and Methods, 33(9): 2015–2026, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gatto:2006:BTC

[Gat06] Riccardo Gatto. A bootstrap test for circular data. Communications in Statistics: Theory and Methods, 35(2):281–292, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gautier:2006:AIM

[Gau06] Antony Gautier. Asymptotic inefficiency of mean-correction on parameter estimation for a periodic first-order autoregressive model. Communications in Statistics: Theory and Methods, 35(11):2083–2106, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gavriliadis:2008:MIP

[Gav08] P. N. Gavriliadis. Moment information for probability distributions, without solving the moment problem. I: Where is the mode? Communications in Statistics: Theory and Methods, 37(5):671–681, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2001:RSS

[GB01] Ramesh C. Gupta and Nicole Brown. Reliability studies of the skew-normal distribution and its application to a strength-stress model. Communications in Statistics: Theory and Methods, 30(11):2427–2445, 2001. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Galea:2006:HTF

[GBG06] Manuel Galea, Heleno Bolfarine, and Patricia Giménez. Hypothesis testing in functional comparative calibration models. Communications in Statistics: Theory and Methods, 35 (11):2021–2033, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gonzalez-Barrios:2001:CTQ

[GBR01] Jose M. González-Barrios and Raúl Rueda. On convergence theorems for quantiles. Communications in Statistics: Theory and Methods, 30(5):943–955, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gill:2005:TCE

[GBS05] Paramjit S. Gill, Sarath G. Banneheka, and Tim B. Swartz. Tests concerning equicorrelation matrices with grouped normal data. Communications in Statistics: Theory and Methods, 34(4):857–873, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gallas:2009:FRE

[GBSW09] Brandon D. Gallas, Andriy Bandos, Frank W. Samuelson, and Robert F. Wagner. A framework for random-effects ROC analysis: Biases with the bootstrap and other variance estimators. Communications in Statistics: Theory and Methods, 38(15):2586–2603, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gimenez:2000:ARE

[GCB00] Patricia Gimenez, Enrico A. Colosimo, and Heleno Bolfarine. Asymptotic relative efficiency of Wald tests in measurement error models. Communications in Statistics: Theory and Methods, 29(3):549–564, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guo:2006:RS

[GCLG06] Yanlan Guo, Lirong Cui, Jinlin Li, and Song Gao. Reliabilities for (n, f, k(i, j)) and ⟨n, f, k(i, j)⟩ systems. Communications in Statistics: Theory and Methods, 35(10):1779–1789, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:2005:ODB

[GD05] Himadri Ghosh and Ashish Das. Optimal designs for best linear unbiased prediction in diallel crosses. Communications in Statistics: Theory and Methods, 34(7):1579–1586, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guo:2006:GPP

[GD06] Renkuan Guo and Tim Dunne. Grey predictive process control charts. Communications in Statistics: Theory and Methods, 35(10):1857–1868, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez-Deniz:2006:SMS

[GDCOCO06] E. Gómez-Déniz, E. Calderín-Ojeda, and I. Cabrera-Ortega. A simple method to study sensitivity of BMP’s. Communications in Statistics: Theory and Methods, 35(4):583–591, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez-Deniz:2008:UBH

[GDSPSVP08] E. Gómez-Déniz, J. M. Sarabia, J. M. Pérez-Sánchez, and F. J. Vázquez-Polo. Using a Bayesian hierarchical model for fitting automobile claim frequency data. Communications in Statistics: Theory and Methods, 37(9):1425–1435, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garel:2005:RSU

[GdT05] Bernard Garel, Ludovic d’Estampes, and Dag Tjøstheim. Revealing some unexpected dependence properties of linear combinations of stable random variables using symmetric covariation. Communications in Statistics: Theory and Methods, 33(4):769–786, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garcia-Escudero:2009:RMR

[GEGMMI09] L. A. García-Escudero, A. Gordaliza, C. Matrán, and A. Mayo-Iscar. A robust maximal F -ratio statistic to detect clusters structure. Communications in Statistics: Theory and Methods, 38(5):682–694, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Genc:2006:NDL

[Gen06] Ali I. Genç. A note on the distribution of linear functions of two ordered correlated normal random variables. Communications in Statistics: Theory and Methods, 35(6):989–995, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Genc:2007:OSS

[Gen07] Ali I. Genç. On order statistics from sine distribution. Communications in Statistics: Theory and Methods, 36(12):2169– 2176, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gerhold:2009:AAS

[Ger09] Stefan Gerhold. Asymptotic analysis of some discrete distributions by the saddle point method. Communications in Statistics: Theory and Methods, 38(2):262–271, 2009. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Grane:2006:AGF

[GF06] Aurea Grane and Josep Fortiana. An adaptive goodness-of-fit test. Communications in Statistics: Theory and Methods, 35(6):1141–1155, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Grane:2008:KLB

[GF08] Aurea Grané and Josep Fortiana. Karhunen–Loève basis in goodness-of-fit tests decomposition: An evaluation. Communications in Statistics: Theory and Methods, 37(19):3144– 3163, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gardes:2005:EEQ

[GG05a] Laurent Gardes and Stéphane Girard. Estimating extreme quantiles of Weibull tail distributions. Communications in Statistics: Theory and Methods, 34(5):1065–1080, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gijbels:2005:DDD

[GG05b] I. Gijbels and A.-C. Goderniaux. Data-driven discontinuity detection in derivatives of a regression function. Communications in Statistics: Theory and Methods, 33(4):851–871, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2004:SCR

[GGS04] Rameshwar D. Gupta, Ramesh C. Gupta, and P. G. Sankaran. Some characterization results based on factorization of the (reversed) hazard rate function. Communications in Statistics: Theory and Methods, 33(12):3009–3031, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2005:STZ

[GGT05] Pushpa Lata Gupta, Ramesh C. Gupta, and Ram C. Tripathi. Score test for zero inflated generalized Poisson regression model. Communications in Statistics: Theory and Methods, 33(1):47–64, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez:2001:RLE

[GGVM01] E. Gómez, M. A. Gómez-Villegas, and J. M. Marín. Response to letter to the Editor. Communications in Statistics: Theory and Methods, 30(5):989–992, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Glynn:2004:FCL

[GH04] Peter W. Glynn and Peter J. Haas. On functional central limit theorems for semi-Markov and related processes. Communications in Statistics: Theory and Methods, 33(3): 487–506, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghorbel:2006:IPA

[GH06] M. Ghorbel and T. Huillet. Interval partitioning arising from size-biased picking. Communications in Statistics: Theory and Methods, 35(7):1173–1187, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gombay:2009:STC

[GH09] Edit Gombay and Lajos Horváth. Sequential tests and change detection in the covariance structure of weakly stationary time series. Communications in Statistics: Theory and Methods, 38(16–17):2872–2883, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghitany:2005:REB

[Ghi05] M. E. Ghitany. On reliability estimation based on ranked set sampling. Communications in Statistics: Theory and Methods, 34(5):1213–1216, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garren:2007:IEP

[GHL07] Steven T. Garren, Elizabeth R. Hume, and Glen R. Leppert. Improved estimation of the Pareto location parameter under squared error loss. Communications in Statistics: Theory and Methods, 36(1):97–102, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garca-Heras:2006:IMB

[GHMGMPP06] J. García-Heras, J. Muñoz-García, J. M. Muñoz-Pichardo, and L. Pardo. Influence measures based on Cressie–Read divergence measures in multivariate linear model. Communications in Statistics: Theory and Methods, 35(11):2055– 2073, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giancristofaro:2003:NCA

[Gia03] Rosa Arboretti Giancristofaro. A new conjoint analysis procedure with application to marketing research. Communications in Statistics: Theory and Methods, 32(11):2271–2283, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gilchrist:2000:RMD

[Gil00] Robert Gilchrist. Regression models for data with a non-zero probability of a zero response. Communications in Statistics: Theory and Methods, 29(9–10):1987–2003, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610920008832591.

Giles:2007:SRT

[Gil07] David E. A. Giles. Spurious regressions with time-series data: Further asymptotic results. Communications in Statistics: Theory and Methods, 36(5):967–979, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Girard:2004:HTE

[Gir04] Stéphane Girard. A Hill type estimator of the Weibull tailcoefficient. Communications in Statistics: Theory and Methods, 33(2):205–234, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gzyl:2008:PNB

[GIT08] H. Gzyl, P. L. Novi Inverardi, and A. Tagliani. A proposal for a new bound for discrete distributions. Communications in Statistics: Theory and Methods, 37(14):2149–2161, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2000:IER

[GK00] A. K. Gupta and D. G. Kabe. On the impact of equicorrelated responses on multicollinearity. Communications in Statistics: Theory and Methods, 29(7):1565–1568, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2003:CGG

[GK03] Rameshwar D. Gupta and Debasis Kundu. Closeness of gamma and generalized exponential distribution. Communications in Statistics: Theory and Methods, 32(4):705–721, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guo:2004:NAI

[GK04a] Huizhen Guo and K. Krishnamoorthy. New approximate inferential methods for the reliability parameter in a stressstrength model: The normal case. Communications in Statistics: Theory and Methods, 33(7):1715–1731, 2004. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2004:SCD

[GK04b] Ramesh C. Gupta and S. N. U. A. Kirmani. Some characterization of distributions by functions of failure rate and mean residual life. Communications in Statistics: Theory and Methods, 33(12):3115–3131, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gangopadhyay:2005:EAW

[GK05] Aditi Kar Gangopadhyay and Somesh Kumar. Estimating average worth of the selected subset from two-parameter exponential populations. Communications in Statistics: Theory and Methods, 34(12):2257–2267, 2005. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gill:2006:DAL

[GKB06] Ryan S. Gill, Jerome P. Keating, and Michael I. Baron. Detecting abrupt leaks in blended underground storage tanks. Communications in Statistics: Theory and Methods, 35(4): 727–742, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Glueck:2008:FCB

[GKFM^{+}08] D. H. Glueck, A. Karimpour-Fard, J. Mandel, L. Hunter, and K. E. Muller. Fast computation by block permanents of cumulative distribution functions of order statistics from several populations. Communications in Statistics: Theory and Methods, 37(18):2815–2824, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:2007:CBE

[GKK07] Malay Ghosh, Myung Joon Kim, and Dalho Kim. Constrained Bayes and empirical Bayes estimation with balanced loss functions. Communications in Statistics: Theory and Methods, 36(8):1527–1542, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See erratum [Ano07].

Girardin:2004:ERM

[GL04] Valerie Girardin and Nikolaos Limnios. Entropy rate and maximum entropy methods for countable semi-Markov chains. Communications in Statistics: Theory and Methods, 33(3):609–622, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gao:2005:NAD

[GLW05] Shaogen Gao, Guoying Li, and Dongqian Wang. A new approach for detecting multivariate outliers. Communications in Statistics: Theory and Methods, 34(8):1857–1865, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Glueck:2001:EVS

[GM01] Deborah H. Glueck and Keith E. Muller. On the expected values of sequences of functions. Communications in Statistics: Theory and Methods, 30(2):363–369, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gamerman:2002:BAE

[GM02] Dani Gamerman and Ajax R. B. Moreira. Bayesian analysis of econometric time series models using hybrid integration rules. Communications in Statistics: Theory and Methods, 31(1):49–72, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Girard:2008:SEV

[GM08] Stéphane Girard and Ludovic Menneteau. Smoothed extreme value estimators of non-uniform point processes boundaries with application to star-shaped supports estimation. Communications in Statistics: Theory and Methods, 37(6):881–897, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Glueck:2008:EPF

[GMKFH08] D. H. Glueck, K. E. Muller, A. Karimpour-Fard, and L. Hunter. Expected power for the false discovery rate with independence. Communications in Statistics: Theory and Methods, 37(12):1855–1866, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gonzalez:2001:EOD

[GMM01] M. González, M. Molina, and M. Mota. Estimation of the offspring distribution and the mean vector for a bisexual Galton–Watson process. Communications in Statistics: Theory and Methods, 30(3):497–516, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gracia-Medrano:2005:TTF

[GMO05] L. Gracia-Medrano and F. O’Reilly. Transformations for testing the fit of the inverse-Gaussian distribution. Communications in Statistics: Theory and Methods, 33(4):919–924, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gonzalez-Manteiga:2004:NME

[GMPG04] W. González-Manteiga and A. Pérez-González. Nonparametric mean estimation with missing data. Communications in Statistics: Theory and Methods, 33(2):277–303, 2004. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Greenwood:2004:EES

[GMW04] Priscilla E. Greenwood, Ursula U. Müller, and Wolfgang Wefelmeyer. Efficient estimation for semiparametric semi-Markov processes. Communications in Statistics: Theory and Methods, 33(3):419–435, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2001:SRR

[GN01] Rameshwar D. Gupta and Asok K. Nanda. Some results on reversed hazard rate ordering. Communications in Statistics: Theory and Methods, 30(11):2447–2457, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gulati:2003:GFS

[GN03a] Sneh Gulati and Jordan Neus. Goodness of fit statistics for the exponential distribution when the data are grouped. Communications in Statistics: Theory and Methods, 32(3): 681–700, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2003:LE

[GN03b] Arjun Gupta and Saralees Nadarajah. Letter to the Editor. Communications in Statistics: Theory and Methods, 32(4): 907–909, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gudi:2004:DTE

[GN04] S. V. Gudi and B. N. Nagnur. Deficiency of two estimators in one-parameter exponential family of distributions. Communications in Statistics: Theory and Methods, 33(8):1779– 1800, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2005:MBN

[GN05] Arjun K. Gupta and Saralees Nadarajah. On the moments of the beta normal distribution. Communications in Statistics: Theory and Methods, 33(1):1–13, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Greco:2007:ISU

[GN07] Luigi Greco and Stefania Naddeo. Inverse sampling with unequal selection probabilities. Communications in Statistics: Theory and Methods, 36(5):1039–1048, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gajek:2000:IGO

[GO00] L. Gajek and A. Okolewski. Inequalities for generalized order statistics from some restricted family of distributions. Communications in Statistics: Theory and Methods, 29(11):2427– 2438, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2005:ALT

[GO05] Ramesh C. Gupta and S. H. Ong. Analysis of long-tailed count data by Poisson mixtures. Communications in Statistics: Theory and Methods, 34(3):557–573, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garcia:2006:SPS

[GO06] Gloria García and Josep M. Oller. On the smoothness of projections in statistical models. Communications in Statistics: Theory and Methods, 35(8):1375–1385, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2008:P

[GO08] Ramesh Gupta and S. H. Ong. Preface. Communications in Statistics: Theory and Methods, 37(11):1687, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gonen:2000:BAI

[Gön00] Mithat Gönen. A Bayesian analysis of the intraclass correlations in the mixed linear model. Communications in Statistics: Theory and Methods, 29(7):1451–1464, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goroncy:2009:LBP

[Gor09] Agnieszka Goroncy. Lower bounds on positive L-statistics. Communications in Statistics: Theory and Methods, 38(12): 1989–2002, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Govindarajulu:2001:CDE

[Gov01] Z. Govindarajulu. Characterization of double exponential distribution using moments of order statistics. Communications in Statistics: Theory and Methods, 30(11):2355–2372, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gut:2003:GLI

[GP03] Allan Gut and Valentin V. Petrov. Generalized laws of the iterated logarithm. Communications in Statistics: Theory and Methods, 32(9):1685–1693, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomes:2005:RRJ

[GPM05] M. Ivette Gomes, Hugo Pereira, and M. Cristina Miranda. Revisiting the role of the jackknife methodology in the estimation of a positive tail index. Communications in Statistics: Theory and Methods, 34(2):319–335, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gadre:2005:EMA

[GR05] M. P. Gadre and R. N. Rattihalli. Exact multi-attribute acceptance single sampling plan. Communications in Statistics: Theory and Methods, 33(1):165–180, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goroncy:2009:LBS

[GR09] Agnieszka Goroncy and Tomasz Rychlik. Lower bounds on spacings based on with-replacement model. Communications in Statistics: Theory and Methods, 38(12):1982–1988, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Graneri:2003:TGF

[Gra03] Jorge Graneri. χ^{2}-type goodness of fit test based on transformed empirical processes for location and scale families.

Communications in Statistics: Theory and Methods, 32(10): 2073–2087, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gruber:2004:ESE

[Gru04] Marvin H. J. Gruber. The efficiency of shrinkage estimators with respect to Zellner’s balanced loss function. Communications in Statistics: Theory and Methods, 33(2):235–249, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gruber:2008:SAR

[Gru08] Marvin H. J. Gruber. Some applications of the Rao distance to shrinkage estimators. Communications in Statistics: Theory and Methods, 37(2):180–193, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Grzegorzewski:2003:REU

[Grz03] Przemyslaw Grzegorzewski. Robustness of estimation in the uniform distribution under dependencies in a sample. Communications in Statistics: Theory and Methods, 32(6):1155– 1168, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gather:2004:CDL

[GS04a] Ursula Gather and Gábor Székely. Characterizations of distributions by linear forms of order statistics. Communications in Statistics: Theory and Methods, 33(12):2913–2919, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gertsbakh:2004:RFI

[GS04b] I. Gertsbakh and Y. Shpungin. Renewal function and interval availability: A numerical Monte-Carlo study. Communications in Statistics: Theory and Methods, 33(3):639–649, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Graver:2005:YMR

[GS05] Jack Graver and Milton Sobel. You May rely on the reliability polynomial for much more than you might think. Communications in Statistics: Theory and Methods, 34(6):1411– 1422, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghute:2008:MSC

[GS08] V. B. Ghute and D. T. Shirke. A multivariate synthetic control chart for monitoring process mean vector. Communications in Statistics: Theory and Methods, 37(13):2136–2148, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2008:EMT

[GSA08] Sat Gupta, Javid Shabbir, and Shabbir Ahmad. Estimation of median in two-phase sampling using two auxiliary variables. Communications in Statistics: Theory and Methods, 37(11):1815–1822, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez-Sanchez-Manzano:2008:MEP

[GSMGVM08] E. Gómez-Sánchez-Manzano, M. A. Gómez-Villegas, and J. M. Marín. Multivariate exponential power distributions as mixtures of normal distributions with Bayesian applications. Communications in Statistics: Theory and Methods, 37(6):972–985, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gannoun:2003:ATR

[GSYB03] Ali Gannoun, Jérome Saracco, Ao Yuan, and George E. Bonney. On adaptive transformation–retransformation estimate of conditional spatial median. Communications in Statistics: Theory and Methods, 32(10):1981–2011, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gilmour:2000:SPA

[GT00] Steven G. Gilmour and Luzia A. Trinca. Some practical advice on polynomial regression analysis from blocked response surface designs. Communications in Statistics: Theory and Methods, 29(9–10):2157–2180, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610920008832601.

Geweke:2003:NSD

[GT03a] John Geweke and Hisashi Tanizaki. Note on the sampling distribution for the Metropolis–Hastings algorithm. Communications in Statistics: Theory and Methods, 32(4):775–789, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goldman:2003:ADU

[GT03b] Elena Goldman and Hiroki Tsurumi. Asymptotic distribution of a unit root process under double truncation. Communications in Statistics: Theory and Methods, 32(10):2059– 2071, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goria:2003:BTP

[GT03c] M. N. Goria and A. Tagliani. Bounds on the tail probability and absolute difference between two distributions. Communications in Statistics: Theory and Methods, 32(3):519–532, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez:2007:LSI

[GTB07] Héctor W. Gómez, Francisco J. Torres, and Heleno Bolfarine. Large-sample inference for the epsilon-skew- t distribution. Communications in Statistics: Theory and Methods, 36(1): 73–81, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez:2008:LE

[GTB08] Héctor W. Gómez, Francisco J. Torres, and Heleno Bolfarine. Letter to the Editor. Communications in Statistics: Theory and Methods, 37(6):986, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guillou:2000:BCI

[Gui00] Armelle Guillou. Bootstrap confidence intervals for the Pareto index. Communications in Statistics: Theory and Methods, 29(1):211–226, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guo:2000:REC

[Guo00] Jiin-Huarng Guo. Robust estimation for the coefficient of a first order autoregressive process. Communications in Statistics: Theory and Methods, 29(1):55–66, 2000. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guo:2003:FCC

[Guo03] Jiin-Huarng Guo. Four correlation coefficients with a third blocking variable: Their efficacy, relative efficiency, and test statistics. Communications in Statistics: Theory and Methods, 32(9):1835–1858, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2004:SAM

[Gup04a] Ramesh C. Gupta. On some association measures in the multivariate normal distribution. Communications in Statistics: Theory and Methods, 33(11):2817–2832, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2004:P

[Gup04b] Ramesh C. Gupta. Preface. Communications in Statistics: Theory and Methods, 33(12):2871, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2006:BFD

[Gup06] Sudhir Gupta. Balanced factorial designs for cDNA microarray experiments. Communications in Statistics: Theory and Methods, 35(8):1469–1476, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gurevich:2006:NAC

[Gur06] Gregory Gurevich. Nonparametric AMOC changepoint tests for stochastically ordered alternatives. Communications in Statistics: Theory and Methods, 35(5):887–903, 2006. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gurevich:2007:RPT

[Gur07] Gregory Gurevich. Retrospective parametric tests for homogeneity of data. Communications in Statistics: Theory and Methods, 36(16):2841–2862, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guven:2003:TIF

[Güv03] Bilgehan Güven. Test of independence in the Farlie-Gumbel-Morgenstern distribution. Communications in Statistics: Theory and Methods, 32(9):1753–1765, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guven:2005:IVC

[Güv05] Bilgehan Güven. Inference for variance components in a mixed model for unbalanced split plot design. Communications in Statistics: Theory and Methods, 34(1):101–112, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guven:2006:MVQ

[Güv06] Bilgehan Güven. Minimum variance quadratic unbiased estimation for the variance components in simple linear regression with onefold nested error. Communications in Statistics: Theory and Methods, 35(7):1309–1318, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goos:2005:EIO

[GV05] Peter Goos and Martina Vandebroek. Estimating the intercept in an orthogonally blocked experiment when the block effects are random. Communications in Statistics: Theory and Methods, 33(4):873–890, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez-Villegas:2006:COS

[GVGP06] Miguel A. Gómez-Villegas and Beatriz González-Pérez. A condition to obtain the same decision in the homogeneity testing problem from the frequentist and Bayesian point of view. Communications in Statistics: Theory and Methods, 35(12):2211–2222, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez-Villegas:2002:SBA

[GVMS02] Miguel A. Gómez-Villegas, Paloma Maín, and Luis Sanz. A suitable Bayesian approach in testing point null hypothesis: Some examples revisited. Communications in Statistics: Theory and Methods, 31(2):201–217, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez-Villegas:2007:SAG

[GVMS07] Miguel A. Gómez-Villegas, Paloma Maín, and Rosario Susi. Sensitivity analysis in Gaussian Bayesian networks using a divergence measure. Communications in Statistics: Theory and Methods, 36(3):523–539, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez-Villegas:2005:BAC

[GVP05] Miguel A. Gómez-Villegas and Beatriz González Pérez. Bayesian analysis of contingency tables. Communications in Statistics: Theory and Methods, 34(8):1743–1754, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gray:2005:NDA

[GVW05] Henry L. Gray, Chu-Ping C. Vijverberg, and Wayne A. Woodward. Nonstationary data analysis by time deformation. Communications in Statistics: Theory and Methods, 34(1):163–192, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2001:DCP

[GW01] Ramesh C. Gupta and Robin Warren. Determination of change points of non-monotonic failure rates. Communications in Statistics: Theory and Methods, 30(8–9):1903–1920, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghorai:2005:DBR

[GY05] J. K. Ghorai and Dong Yu. Data-based resolution selection in positive wavelet density estimation. Communications in Statistics: Theory and Methods, 33(10):2393–2408, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gaudoin:2006:CIS

[GYX06] Olivier Gaudoin, Bo Yang, and Min Xie. Confidence intervals for the scale parameter of the power-law process. Communications in Statistics: Theory and Methods, 35(8):1525–1538, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gaffke:2003:RAU

[GZ03] Norbert Gaffke and Andreas Zöllner. A resampling approach for under-estimating a finite population total from a censored sample. Communications in Statistics: Theory and Methods, 32(12):2305–2320, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghezzi:2005:ICV

[GZ05] Daniel J. Ghezzi and Shelemyahu Zacks. Inference on the common variance of correlated normal random variables. Communications in Statistics: Theory and Methods, 34(7): 1517–1531, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Habibullah:2000:EPP

[HA00] Mohamed Habibullah and M. Ahsanullah. Estimation of parameters of a Pareto distribution by generalized order statistics. Communications in Statistics: Theory and Methods, 29 (7):1597–1609, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hendi:2001:TNB

[HA01] M. I. Hendi and A. M. Abouammoh. Testing new better than renewal used life distributions based on U-test. Communications in Statistics: Theory and Methods, 30(10):2135–2147, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Habibi:2006:SEE

[HAA06] A. Habibi, N. R. Arghami, and J. Ahmadi. Statistical evidence in experiments and in record values. Communications in Statistics: Theory and Methods, 35(11):1971–1983, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haddad:2000:REF

[Had00] John N. Haddad. On robust estimation in the first order autoregressive processes. Communications in Statistics: Theory and Methods, 29(1):45–54, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haddad:2009:PEG

[Had09] John N. Haddad. Preliminary estimation in Gaussian stationary processes. Communications in Statistics: Theory and Methods, 38(5):742–747, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamdan:2006:CAI

[Ham06] Hasan Hamdan. Characterizing and approximating infinite scale mixtures of normals. Communications in Statistics: Theory and Methods, 35(3):407–413, 2006. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamadouche:2008:WHC

[Ham08] D. Hamadouche. Weak hölder convergence of smoothed uniform quantile process. Communications in Statistics: Theory and Methods, 37(10):1497–1505, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Han:2008:TSB

[Han08a] Ming Han. Two-sided M-Bayesian credible limits method of reliability parameters and its applications. Communications in Statistics: Theory and Methods, 37(10):1659–1670, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanayama:2008:APC

[Han08b] Nobutane Hanayama. Age period cumulative environment model for analyzing (age, period)-tabulated data on breast cancer deaths. Communications in Statistics: Theory and Methods, 37(11):1770–1782, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Harlamov:2004:CSM

[Har04] B. P. Harlamov. Continuous semi-Markov processes and their applications. Communications in Statistics: Theory and Methods, 33(3):569–589, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huda:2000:MOD

[HAS00] S. Huda and A. A. Al-Shiha. On D- and E-minimax optimal designs for estimating the axial slopes of a second-order response surface over hypercubic regions. Communications in Statistics: Theory and Methods, 29(8):1827–1849, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hashorva:2008:MDA

[Has08] Enkelejd Hashorva. On the max-domain of attraction of Type-III elliptical triangular arrays. Communications in Statistics: Theory and Methods, 37(10):1543–1551, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hattori:2009:RDS

[Hat09] Satoshi Hattori. Regression diagnostics of the semiparametric proportional rate model for irregularly spaced repeated measurements. Communications in Statistics: Theory and Methods, 38(4):542–559, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsiao:2001:BDS

[HC01] Chin-Fu Hsiao and Murray K. Clayton. Bayes discrete sequential boundaries for clinical trials. Communications in Statistics: Theory and Methods, 30(7):1381–1394, 2001. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huggins:2002:APO

[HC02] Richard M. Huggins and Anne Chao. Asymptotic properties of an optimal estimating function approach to the analysis of mark recapture data. Communications in Statistics: Theory and Methods, 31(4):575–595, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Han:2000:BCL

[HCL00] JeongHye Han, JoongJae Cho, and Choon Seong Leem. Bootstrap confidence limits for Wright’s C_{5}. Communications in Statistics: Theory and Methods, 29(3):485–505, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ho:2001:GLS

[HdMS01] Linda Lee Ho and Julio da Motta Singer. Generalized least squares methods for bivariate Poisson regression. Communications in Statistics: Theory and Methods, 30(2):263–277, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hashorva:2005:ETR

[HH05] Enkelejd Hashorva and Jürg Hüsler. Estimation of tails and related quantities using the number of near-extremes. Communications in Statistics: Theory and Methods, 34(2): 337–349, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2009:ODQ

[HH09] Miao-Kuan Huang and Mong-Na Lo Huang. D_{s}-optimal designs for quadratic log contrast model for experiments with mixtures. Communications in Statistics: Theory and Methods, 38(10):1607–1621, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hayes:2003:UMG

[HHL03] Kathy J. Hayes, Joe Hirschberg, and Jenny Lye. The use of multivariate generated regressors in the presence of heteroskedasticity. Communications in Statistics: Theory and Methods, 32(8):1607–1629, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

He:2007:ADP

[HHL07] Yanling He, Wenzheng Huang, and Hua Liang. Axiomatic development of profile likelihoods as the strength of evidence for composite hypotheses. Communications in Statistics: Theory and Methods, 36(15):2695–2706, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamaz:2009:CTE

[HI09] Abdelghani Hamaz and Mohamed Ibazizen. Comparison of two estimation methods of missing values using Pitmancloseness criterion. Communications in Statistics: Theory and Methods, 38(13):2210–2213, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hinich:2005:NMS

[Hin05] Melvin J. Hinich. A new method for statistical multidimensional unfolding. Communications in Statistics: Theory and Methods, 34(12):2299–2310, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hinich:2009:FAG

[Hin09] Melvin Hinich. Falsifying ARCH/GARCH models using bispectral based tests. Communications in Statistics: Theory and Methods, 38(4):529–541, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huh:2001:RPT

[HJ01] Myung-Hoe Huh and Myoungshic Jhun. Random permutation testing in multiple linear regression. Communications in Statistics: Theory and Methods, 30(10):2023–2032, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hawkes:2006:PEC

[HJ06] A. G. Hawkes and A. Jalali. Parameter estimation and censored lives. Communications in Statistics: Theory and Methods, 35(10):1791–1802, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hwang:2002:DSE

[HK02] Ruey-Ching Hwang and Mei-Lien Ken. Double smoothing estimation of the multivariate regression function in nonparametric regression. Communications in Statistics: Theory and Methods, 31(3):419–434, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Handa:2005:TEP

[HK05] B. R. Handa and N. S. Kambo. Tests for equality of parameters of two exponential distributions with common known coefficient of variation. Communications in Statistics: Theory and Methods, 34(11):2147–2155, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Han:2006:TPM

[HK06] Cong Han and Richard Kronmal. Two-part models for analysis of Agatston scores with possible proportionality constraints. Communications in Statistics: Theory and Methods, 35(1):99–111, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Handa:2002:TSP

[HKR02] B. R. Handa, N. S. Kambo, and C. D. Ravindran. Testing of scale parameter of the exponential distribution with known coefficient of variation: Conditional approach. Communications in Statistics: Theory and Methods, 31(1):73–86, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hoque:2009:PPT

[HKW09] Zahirul Hoque, Shahjahan Khan, and Jacek Wesołowski. Performance of preliminary test estimator under LINEX loss function. Communications in Statistics: Theory and Methods, 38(2):252–261, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2002:IBS

[HL02a] Wen-Tao Huang and Yu-Pin Lin. An improved Bayesian sampling plan for exponential population with Type I censoring. Communications in Statistics: Theory and Methods, 31(11):2003–2025, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hudson:2002:GLS

[HL02b] Richard Hudson and Charles R. Lawoko. Generalised least squares (gls) estimation of the difference parameter in long memory (arfima) processes. Communications in Statistics: Theory and Methods, 31(9):1629–1646, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hajlaoui:2003:EQC

[HL03] Mekki Hajlaoui and Mohamed Limam. Economical quality control with measurement error. Communications in Statistics: Theory and Methods, 32(11):2255–2270, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2009:SBP

[HL09] Wen-Tao Huang and Yao-Tsung Lai. Selecting the best process based on capability index via empirical Bayes approach. Communications in Statistics: Theory and Methods, 38(10):1576–1588, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2008:SPS

[HLLL08] Pi-Hsiang Huang, Pen-Hwang Liau, Chien-Shiang Liau, and Chien-Te Lee. Some properties of semi-folding designs. Communications in Statistics: Theory and Methods, 37(8):1245– 1257, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hedayat:2009:SAA

[HLS09] A. S. Hedayat, Congrong Lou, and Bikas K. Sinha. A statistical approach to assessment of agreement involving multiple raters. Communications in Statistics: Theory and Methods, 38(16–17):2899–2922, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2008:SNR

[HLW08] Taizhong Hu, Qingshu Lu, and Songqiao Wen. Some new results on likelihood ratio orderings for spacings of heterogeneous exponential random variables. Communications in Statistics: Theory and Methods, 37(16):2506–2515, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Henze:2002:GFT

[HM02] Norbert Henze and Simos G. Meintanis. Goodness-of-fit tests based on a new characterization of the exponential distribution. Communications in Statistics: Theory and Methods, 31(9):1479–1497, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hardouin:2004:CBV

[HM04] J.-B. Hardouin and M. Mesbah. Clustering binary variables in subscales using an extended Rasch model and Akaike information criterion. Communications in Statistics: Theory and Methods, 33(6):1277–1294, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Holm:2005:STE

[HMA05] Sture Holm, Sigyn Mark, and Tobias Adolfsson. A step-down test for effects in unreplicated factorial designs. Communications in Statistics: Theory and Methods, 34(2):405–416, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hassairi:2005:IDE

[HMK05] Adelhamid Hassairi, Afif Masmoudi, and Célestin C. Kokonendji. Implicit distributions and estimation. Communications in Statistics: Theory and Methods, 34(2):245–252, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See comments [Muk06].

Hosmane:2004:QLD

[HML04] Balakrishna Hosmane, David Morris, and Charles Locke. Quality of life: Design and analyses issues in clinical trials. Communications in Statistics: Theory and Methods, 33 (6):1429–1443, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hobza:2003:LDS

[HMM03] Tomás Hobza, Isabel Molina, and Domingo Morales. Likelihood divergence statistics for testing hypotheses in familial data. Communications in Statistics: Theory and Methods, 32(2):415–434, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2003:MID

[HMN03] Taizhong Hu, Min Ma, and Asok K. Nanda. Moment inequalities for discrete ageing families. Communications in Statistics: Theory and Methods, 32(1):61–90, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hartley:2001:EFC

[HN01] Andrew M. Hartley and Dayanand N. Naik. Estimation of familial correlations under autoregressive circular covariance. Communications in Statistics: Theory and Methods, 30(8– 9):1811–1828, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Henze:2003:TST

[HN03] Norbert Henze and Ya. Yu. Nikitin. Two-sample tests based on the integrated empirical process. Communications in Statistics: Theory and Methods, 32(9):1767–1788, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Heimann:2005:ECD

[HN05] Günter Heimann and Georg Neuhaus. On the exact conditional distribution for Peto’s combined test for carcinogenicity assays. Communications in Statistics: Theory and Methods, 33(10):2535–2572, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Helu:2007:ESS

[HN07] Amal Helu and Dayanand N. Naik. Estimation of sib-sib correlation via a Kotz-type density function. Communications in Statistics: Theory and Methods, 36(5):1021–1029, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

He:2009:CEU

[HN09] Qinying He and H. N. Nagaraja. Correlation estimation using concomitants of order statistics from bivariate normal samples. Communications in Statistics: Theory and Methods, 38(12):2003–2015, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hodoshima:2009:ACM

[Hod09] Jiro Hodoshima. The asymptotic covariance matrix of the least squares estimator in the stochastic linear regression model: The case of elliptically symmetric distribution. Communications in Statistics: Theory and Methods, 38(5):661– 674, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Holgersson:2007:RTS

[Hol07] H. E. T. Holgersson. Robust testing for skewness. Communications in Statistics: Theory and Methods, 36(3):485–498, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2007:LDS

[HPR07] Zhishui Hu, Valentin V. Petrov, and John Robinson. On large deviations of sums of independent random variables. Communications in Statistics: Theory and Methods, 36(10): 1981–1992, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Harvill:2000:ILI

[HR00] Janc L. Harvill and Bonnic K. Ray. An investigation of lag identification tools for vector nonlinear time series. Communications in Statistics: Theory and Methods, 29(8):1677– 1701, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hunt:2003:TDR

[HR03] Daniel Hunt and Shesh N. Rai. A threshold dose-response model with random effects in teratological experiments. Communications in Statistics: Theory and Methods, 32(7): 1439–1457, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hunt:2008:IRV

[HR08] Daniel L. Hunt and Shesh N. Rai. Interlitter response variability in a threshold dose-response model. Communications in Statistics: Theory and Methods, 37(14):2304–2314, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsu:2001:LRD

[HRB01] Nan-Jung Hsu, Bonnie K. Ray, and F. Jay Breidt. Longrange dependent common factor models: a Bayesian approach. Communications in Statistics: Theory and Methods, 30(6):1047–1061, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hilgert:2007:IEC

[HRVW07] Nadine Hilgert, Vivien Rossi, Jean-Pierre Vila, and Vérène Wagner. Identification, estimation, and control of uncertain dynamic systems: A nonparametric approach. Communications in Statistics: Theory and Methods, 36(14):2509–2525, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hebert:2005:CLE

[HS05] Jaimie L. Hebert and Stephen M. Scariano. Comparing location estimators for exponential mixtures under Pitman’s measure of closeness. Communications in Statistics: Theory and Methods, 33(1):29–46, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2008:EET

[HS08] Jianhua Hu and Sarah Shi. On the expressions of estimability in testing general linear hypotheses. Communications in Statistics: Theory and Methods, 37(5):782–790, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Harel:2009:IOH

[HS09] Ofer Harel and Jeffrey Stratton. Inferences on the outfluence — how do missing values impact your analysis? Communications in Statistics: Theory and Methods, 38(16–17):2884– 2898, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Habibi:2005:CPD

[HSAN05] R. Habibi, S. M. Sadooghi-Alvandi, and A. R. Nematollahi. Change point detection in a general class of distributions. Communications in Statistics: Theory and Methods, 34(9– 10):1935–1938, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsieh:2001:BPM

[Hsi01] Ping-Hung Hsieh. On Bayesian predictive moments of next record value using three-parameter gamma priors. Communications in Statistics: Theory and Methods, 30(4):729–738, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsu:2005:GTG

[Hsu05] Lifang Hsu. Group testing with a goal in estimating the number of defects under imperfect environmental stress screen levels. Communications in Statistics: Theory and Methods, 34(6):1363–1377, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Han:2006:TEM

[HT06] Sier Han and Zheng Tian. Truncating estimation for the mean change-point in heavy-tailed dependent observations. Communications in Statistics: Theory and Methods, 35(1): 43–52, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Han:2007:BTC

[HT07] Sier Han and Zheng Tian. Bootstrap testing for changes in persistence with heavy-tailed innovations. Communications in Statistics: Theory and Methods, 36(12):2289–2299, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2006:SSB

[Hua06] Chunfeng Huang. Smoothing splines with boundary correction. Communications in Statistics: Theory and Methods, 35(6):1101–1107, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2008:EMC

[Hua08] Pi-Hsiang Huang. The efficient methods to construct twoper-block design in 2^{k} plans. Communications in Statistics: Theory and Methods, 38(3):431–438, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hudson:2002:NRC

[Hud02] Suzanne Hudson. New results for a class of univariate distributions. Communications in Statistics: Theory and Methods, 31(2):239–247, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huda:2007:MDD

[Hud07] S. Huda. Minimax design for the difference between estimated responses for the quadratic model over hypercubic regions. Communications in Statistics: Theory and Methods, 36(6):1225–1231, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huggins:2004:CRE

[Hug04] Richard Huggins. Capture-recapture experiments with time heterogeneous capture probabilities in continuous time: A counting process approach. Communications in Statistics: Theory and Methods, 33(5):1221–1238, 2004. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huillet:2005:SFA

[Hui05] Thierry Huillet. Sampling formulae arising from random Dirichlet populations. Communications in Statistics: Theory and Methods, 34(5):1019–1040, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hurlimann:2000:GLT

[Hür00] Werner Hürlimann. General location transform of the order statistics from the exponential, Pareto and Weibull, with application to maximum likelihood estimation. Communications in Statistics: Theory and Methods, 29(11):2535–2545, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hussein:2005:SCT

[Hus05] Abdulkadir Hussein. Sequential comparison of two treatments using weighted Wald-type statistics. Communications in Statistics: Theory and Methods, 34(7):1631–1641, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huzurbazar:2004:MMF

[Huz04] Aparna V. Huzurbazar. Multistate models, flowgraph models, and semi-Markov processes. Communications in Statistics: Theory and Methods, 33(3):457–474, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamedani:2003:CSR

[HV03] G. G. Hamedani and Hans Volkmer. A characterization of symmetric random variables. Communications in Statistics: Theory and Methods, 32(4):723–728, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamedani:2007:CPR

[HVW07] G. G. Hamedani, Hans Volkmer, and Jacek Wesołowski. Characterization problems related to the Shepp property. Communications in Statistics: Theory and Methods, 36(6): 1049–1057, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hou:2006:GAB

[HW06] Fujun Hou and Qizong Wu. Genetic algorithm-based redundancy optimization problems in fuzzy framework. Communications in Statistics: Theory and Methods, 35(10):1931– 1941, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hwang:2002:NVL

[Hwa02] Ruey-Ching Hwang. A new version of the local constant M-smoother. Communications in Statistics: Theory and Methods, 31(5):833–848, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hwang:2005:MPL

[Hwa05] Y.-T. Hwang. Moments of the product-limit estimator under left-truncation and right-censoring. Communications in Statistics: Theory and Methods, 34(8):1833–1850, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2009:TSC

[HWL09] Xuemei Hu, Zhizhong Wang, and Feng Liu. Testing serial correlation in semiparametric varying-coefficient partially linear models. Communications in Statistics: Theory and Methods, 38(13):2145–2163, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2003:UML

[hWlF03] Rong hua Wang and He liang Fei. Uniqueness of the maximum likelihood estimate of the Weibull distribution tampered failure rate model. Communications in Statistics: Theory and Methods, 32(12):2321–2338, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2006:SCD

[HWZ06] Taizhong Hu, Feng Wang, and Zegang Zhu. Stochastic comparisons and dependence of spacings from two samples of exponential random variables. Communications in Statistics: Theory and Methods, 35(6):979–988, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hossain:2000:CME

[HZ00] Anwar M. Hossain and William J. Zimmer. Comparisons of methods of estimation for a Pareto distribution of the first kind. Communications in Statistics: Theory and Methods, 29(4):859–878, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Han:2006:RAF

[HZSX06] Bo Tang Han, Cai Bo Zhang, Chang Sen Sun, and Chun Jie Xu. Reliability analysis of flexible manufacturing cells based on triangular fuzzy number. Communications in Statistics: Theory and Methods, 35(10):1897–1907, 2006. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Inoue:2009:DRS

[IA09] Kiyoshi Inoue and Sigeo Aki. Distributions of runs and scans on higher-order Markov trees. Communications in Statistics: Theory and Methods, 38(5):621–641, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ibazizen:2003:BMS

[Iba03] Mohamed Ibazizen. Bias and mean square error of the sample roots under the contaminated Gaussian model. Communications in Statistics: Theory and Methods, 32(5):955–959, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Imon:2008:IMO

[IH08] A. H. M. Rahmatullah Imon and Ali S. Hadi. Identification of multiple outliers in logistic regression. Communications in Statistics: Theory and Methods, 37(11):1697–1709, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Han:2000:UAB

[iHA00] Sung il Han and Sigeo Aki. A unified approach to binomialtype distributions of order k. Communications in Statistics: Theory and Methods, 29(8):1929–1943, 2000. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koike:2002:IK

[iK02] Ken ichi Koike. On the inequality of kshirsagar. Communications in Statistics: Theory and Methods, 31(9):1617–1627, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iliopoulos:2003:SNE

[Ili03] George Iliopoulos. Some new estimators of the binomial parameter n. Communications in Statistics: Theory and Methods, 32(7):1361–1372, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iyer:2004:CBS

[IM04] Srikanth K. Iyer and D. Manjunath. Correlated bivariate sequences for queueing and reliability applications. Communications in Statistics: Theory and Methods, 33(2):331–350, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Izady:2008:DWE

[IM08] Navid Izady and Hashem Mahlooji. Developing a wide easyto-generate class of bivariate copulas. Communications in Statistics: Theory and Methods, 37(12):1919–1929, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Imon:2003:RRM

[Imo03] A. H. M. Rahmatullah Imon. Regression residuals, moments, and their use in tests for normality. Communications in Statistics: Theory and Methods, 32(5):1021–1034, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Isotalo:2006:LPS

[IP06] Jarkko Isotalo and Simo Puntanen. Linear prediction sufficiency for new observations in the general Gauss–Markov model. Communications in Statistics: Theory and Methods, 35(6):1011–1023, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Isotalo:2008:UMD

[IPS08] Jarkko Isotalo, Simo Puntanen, and George P. H. Styan. A useful matrix decomposition and its statistical applications in linear regression. Communications in Statistics: Theory and Methods, 37(9):1436–1457, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Impert:2007:CRM

[IR07] Agnieszka Impert and Tomasz Rychlik. Comparison of records and maxima. Communications in Statistics: Theory and Methods, 36(7):1397–1408, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Inverardi:2003:MED

[IT03] P. L. Novi Inverardi and A. Tagliani. Maximum entropy density estimation from fractional moments. Communications in Statistics: Theory and Methods, 32(2):327–345, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Islam:2005:MLR

[IT05] M. Qamarul Islam and Moti L. Tiku. Multiple linear regression model under nonnormality. Communications in Statistics: Theory and Methods, 33(10):2443–2467, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iqbal:2009:CSB

[IT09] Ijaz Iqbal and M. H. Tahir. Circular strongly balanced repeated measurements designs. Communications in Statistics: Theory and Methods, 38(20):3686–3696, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Illig:2006:ARS

[ITV06] Aude Illig and Benoît Truong-Van. Asymptotic results for spatial ARMA models. Communications in Statistics: Theory and Methods, 35(4):671–688, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Islam:2001:NRS

[ITY01] M. Qamarul Islam, Moti L. Tiku, and F. Yildirim. Nonnormal regression. I. Skew distributions. Communications in Statistics: Theory and Methods, 30(6):993–1020, 2001. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ip:2005:TNL

[IWC05] Wai-Cheung Ip, Hueng Wong, and Min Chen. Testing normality for linear AR(p) models. Communications in Statistics: Theory and Methods, 33(4):891–908, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jackson:2008:SLM

[Jac08] Dan Jackson. The significance level of meta-regression’s standard hypothesis test. Communications in Statistics:

Theory and Methods, 37(10):1576–1590, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jacroux:2009:MAT

[Jac09] Mike Jacroux. Minimum aberration two-level fractional factorial split-plot designs with hard to change factors. Communications in Statistics: Theory and Methods, 38(9):1337– 1349, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jaheen:2004:EBI

[Jah04] Zeinhum F. Jaheen. Empirical Bayes inference for generalized exponential distribution based on records. Communications in Statistics: Theory and Methods, 33(8):1851–1861, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jaheen:2005:EBG

[Jah05] Zeinhum F. Jaheen. Estimation based on generalized order statistics from the Burr model. Communications in Statistics: Theory and Methods, 34(4):785–794, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Janardan:2001:LE

[Jan01] K. G. Janardan. Letter to the Editor. Communications in Statistics: Theory and Methods, 30(12):2721–2722, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jang:2002:MEN

[Jan02] Dae-Heung Jang. Measures for evaluating non-orthogonality of experimental designs. Communications in Statistics: Theory and Methods, 31(2):249–260, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jan:2005:SCI

[Jan05a] Show-Li Jan. Step contrasts for identifying the minimum effective dose. Communications in Statistics: Theory and Methods, 34(1):45–57, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Janardan:2005:IRD

[Jan05b] K. G. Janardan. Integral representation of a distribution associated with a pure birth process. Communications in Statistics: Theory and Methods, 34(11):2097–2103, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jaskiewicz:2007:FPA

[Jaś07] Anna Jaśkiewicz. A fixed point approach to solve the average cost optimality equation for semi-Markov decision processes with Feller transition probabilities. Communications in Statistics: Theory and Methods, 36(14):2559–2575, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Javid:2009:CTI

[Jav09a] Amir Ahmadi Javid. Copulas with truncation-invariance property. Communications in Statistics: Theory and Methods, 38(20):3756–3771, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Javid:2009:LTD

[Jav09b] Amir Ahmadi Javid. Limiting tail dependence copulas. Communications in Statistics: Theory and Methods, 38(20):3772– 3781, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jamalizadeh:2008:GBC

[JB08] A. Jamalizadeh and N. Balakrishnan. On a generalization of bivariate Cauchy distribution. Communications in Statistics: Theory and Methods, 37(4):469–474, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jamalizadeh:2009:DRT

[JBS09] A. Jamalizadeh, N. Balakrishnan, and A. Sheikhy. Distributions of ratios of two correlated skew-normal variables and of ratios of two linear functions of order statistics from bivariate normal distribution. Communications in Statistics: Theory and Methods, 38(12):2107–2115, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

John:2008:LTS

[JC08] Thomas T. John and Pinyuen Chen. L^{1}-limit of trimmed sums of order statistics from location-scale distributions with applications to Type II censored data analysis. Communications in Statistics: Theory and Methods, 37(16):2488–2505, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jin:2001:CSP

[JE01] Chun Jin and Abdulaziz Elfessi. On the common scale parameter of several Pareto populations in censored samples. Communications in Statistics: Theory and Methods, 30(3): 451–462, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jeong:2001:NAE

[Jeo01] Jong-Hyeon Jeong. A note on asymptotic efficiency of a regression coefficient parameter under ordinal logistic regression model. Communications in Statistics: Theory and Methods, 30(7):1257–1269, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jeong:2003:ELR

[Jeo03] Jong-Hyeon Jeong. Efficiency of log-rank test under dependent censorship. Communications in Statistics: Theory and Methods, 32(6):1197–1211, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jackson:2001:EPT

[JH01] J. Michael Jackson and John E. Hewett. Evaluation of power of tests involving covariates under nonconstant variance. Communications in Statistics: Theory and Methods, 30 (10):2005–2021, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jahan:2008:BBG

[JH08] Nusrat Jahan and Jane L. Harvill. Bispectral-based goodness-of-fit tests of Gaussianity and linearity of stationary time series. Communications in Statistics: Theory and Methods, 37(20):3216–3227, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jung:2007:STT

[JHL07] Byoung Cheol Jung, Sang Moon Han, and Jungbok Lee. Score tests for testing independence in the zero-truncated bivariate Poisson models. Communications in Statistics: Theory and Methods, 36(3):599–611, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jacome:2008:PEL

[JIP08] M. A. Jácome and M. C. Iglesias-Pérez. Presmoothed estimation with left-truncated and right-censored data. Communications in Statistics: Theory and Methods, 37(18):2964– 2983, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jimenez:2008:NMC

[JJ08] Fernando Jiménez and Pedro Jodrá. A note on the moments and computer generation of the shifted Gompertz distribution. Communications in Statistics: Theory and Methods, 38(1):75–89, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jammalamadaka:2004:NFW

[JK04a] Sreenivasa Rao Jammalamadaka and Tomasz J. Kozubowski. New families of wrapped distributions for modeling skew circular data. Communications in Statistics: Theory and Methods, 33(9):2059–2074, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Janssen:2004:BTB

[JK04b] Arnold Janssen and Michael Kunz. Brownian type boundary crossing probabilities for piecewise linear boundary functions. Communications in Statistics: Theory and Methods, 33(7):1445–1464, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jeske:2007:CRI

[JLBB07] Daniel R. Jeske, Zheng Liu, Elizabeth Bent, and James Borneman. Classification rules that include neutral zones and their application to microbial community profiling. Communications in Statistics: Theory and Methods, 36(10): 1965–1980, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jones:2000:BGM

[JLR00] G. Jones, C. D. Lai, and J. C. W. Rayner. A bivariate gamma mixture distribution. Communications in Statistics: Theory and Methods, 29(12):2775–2790, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jia:2009:AHR

[JM09] Feiyi Jia and Gary C. McDonald. Analyzing hybrid randomized response data with a binomial selection procedure. Communications in Statistics: Theory and Methods, 38(6): 784–807, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jensen:2003:ESF

[JN03] Uwe Jensen and Andreas Narr. An estimator for the survival function on the basis of observable censoring times. Communications in Statistics: Theory and Methods, 32(10):2043– 2052, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jose:2009:WDA

[JN09] K. K. Jose and Shanoja R. Naik. On the q-Weibull distribution and its applications. Communications in Statistics: Theory and Methods, 38(6):912–926, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jeong:2005:EDH

[JO05] Jong-Hyeon Jeong and David Oakes. Effects of different hazard ratios on asymptotic relative efficiency of estimates from Cox’s model. Communications in Statistics: Theory and Methods, 34(2):429–448, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

John:2000:BAC

[Joh00] Peter W. M. John. Breaking alias chains in fractional factorials. Communications in Statistics: Theory and Methods, 29 (9–10):2143–2155, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/03610920008832600.

Jones:2007:CDD

[Jon07] M. C. Jones. On a class of distributions defined by the relationship between their density and distribution functions. Communications in Statistics: Theory and Methods, 36(10):1835–1843, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jozani:2008:PRM

[JP08] Mohammad Jafari Jozani and Ahmad Parsian. Posterior regret Γ-minimax estimation and prediction with applications on k-records data under entropy loss function. Communications in Statistics: Theory and Methods, 37(14):2202–2212, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jamalizadeh:2009:TLS

[JPB09] A. Jamalizadeh, R. Pourmousa, and N. Balakrishnan. Truncated and limited skew-normal and skew- t distributions: Properties and an illustration. Communications in Statistics: Theory and Methods, 38(16–17):2653–2668, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jorgensen:2005:GH

[JR05] Bent Jørgensen and Jeevanantham Rajeswaran. A generalization of Hotelling’s T^{2}. Communications in Statistics: Theory and Methods, 34(11):2179–2195, 2005. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Juutilainen:2007:MMD

[JR07] Ilmari Juutilainen and Juha Röning. A method for measuring distance from a training data set. Communications in Statistics: Theory and Methods, 36(14):2625–2639, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jokiel-Rokita:2009:MPE

[JRM09] Alicja Jokiel-Rokita and Ryszard Magiera. Minimax prediction of the empirical distribution function. Communications in Statistics: Theory and Methods, 38(11):1776–1791, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jiang:2000:LRT

[JS00] Guoyong Jiang and Sanat K. Sarkar. The likelihood ratio test for the homogeneity of the variances in a covariance matrix with block compound symmetry. Communications in Statistics: Theory and Methods, 29(5–6):1155–1178, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jayasree:2006:SND

[JS06a] G. Jayasree and R. J. R. Swamy. Some new discrete probability distributions derived from power series distributions. Communications in Statistics: Theory and Methods, 35(9): 1555–1567, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jung:2006:RRE

[JS06b] Inkyung Jung and Pranab Kumar Sen. Robust R-estimation of a consensus value from multi-center studies. Communications in Statistics: Theory and Methods, 35(11):2127–2144, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jimenez:2007:ADS

[JS07] Raúl Jiménez and Yongzhao Shao. Asymptotic distribution for symmetric spacing statistics. Communications in Statistics: Theory and Methods, 36(1):37–46, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jimenez:2009:LTD

[JS09] Raúl Jiménez and Yongzhao Shao. Limit theorems for ϕdivergences based on k-spacings. Communications in Statistics: Theory and Methods, 38(5):695–710, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Janssen:2001:ELR

[JSV01] Paul Janssen, Jan Swanepoel, and Noël Veraverbeke. Efficiency of linear regression estimators based on presmoothing. Communications in Statistics: Theory and Methods, 30 (10):2079–2097, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jou:2009:EDM

[JT09] Bo-Yan Jou and Sheng-Tsaing Tseng. An enhanced double MEWMA controller for drifted MIMO systems. Communications in Statistics: Theory and Methods, 38(10):1730–1740, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jung:2002:IFL

[Jun02] Kang-Mo Jung. Influence function of the likelihood ratio test statistic for multivariate normal sample. Communications in Statistics: Theory and Methods, 31(8):1273–1281, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jiang:2005:AME

[JW05] Guoyong Jiang and Jianrong Wu. An alternative method for estimating the variance components in a replicated crossover study. Communications in Statistics: Theory and Methods, 34(1):113–125, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jing:2006:TEM

[JW06] Ping Jing and Jinfang Wang. Testing the equality of multivariate distributions using the bootstrap and integrated empirical processes. Communications in Statistics: Theory and Methods, 35(4):661–670, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jalali:2009:MLE

[JW09] Assad Jalali and Alan J. Watkins. On maximum likelihood estimation for the two parameter Burr XII distribution. Communications in Statistics: Theory and Methods, 38 (11):1916–1926, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johannsen:2004:SSF

[JWSP04] David A. Johannsen, Edward J. Wegman, Jeffrey L. Solka, and Carey E. Priebe. Simultaneous selection of features and metric for optimal nearest neighbor classification. Communications in Statistics: Theory and Methods, 33(9):2137–2157, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jones:2001:CCM

[JZ01] C. Matthew Jones and Anatoly A. Zhigljavsky. Comparison of costs for multi-stage group testing methods in the pharmaceutical industry. Communications in Statistics: Theory and Methods, 30(10):2189–2209, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kezim:2004:BAS

[KA04] Boualem Kezim and Zahia Abdelli. A Bayesian analysis of a structural change in the parameters of a time series. Communications in Statistics: Theory and Methods, 33(8):1863– 1876, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kabaila:2005:APT

[Kab05] Paul Kabaila. Assessment of a preliminary F -test solution to the Behrens–Fisher problem. Communications in Statistics: Theory and Methods, 34(2):291–302, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khaledi:2009:SNR

[KAHS09] Baha-Eldin Khaledi, Fariba Amiripour, Taizhong Hu, and S. Roohollah Shojaei. Some new results on stochastic comparisons of record values. Communications in Statistics: Theory and Methods, 38(12):2056–2066, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kakizawa:2007:AED

[Kak07] Yoshihide Kakizawa. Asymptotic expansions for the distributions of maximum and sum of quasi-independent Hotelling’s T^{2} statistics under non normality. Communications in Statistics: Theory and Methods, 37(1):97–120, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kang:2002:TET

[Kan02] Shin-Soo Kang. Tests for equality of two distributions with correlated interval censored life-times. Communications in Statistics: Theory and Methods, 31(12):2183–2196, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kang:2007:GSA

[Kan07] Sung Won Kang. A Gibbs sampling algorithm for a changing regression model with pooled binary response data. Communications in Statistics: Theory and Methods, 36(2):349–373, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Katsis:2001:COS

[Kat01] Athanassios Katsis. Calculating the optimal sample size for binomial populations. Communications in Statistics: Theory and Methods, 30(4):665–678, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kawamura:2004:DPG

[Kaw04] T. Kawamura. Derivation of the power and generalized inverse Gaussian distributions based on the entropy maximization principle. Communications in Statistics: Theory and Methods, 33(7):1481–1490, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2002:IEN

[KBC02] Byung Hwee Kim, Hoh Yoo Baek, and In Hong Chang. Improved estimators of the natural parameters in continuous multiparameter exponential families. Communications in Statistics: Theory and Methods, 31(1):11–29, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kreiner:2004:ALD

[KC04] Svend Kreiner and Karl B. Christensen. Analysis of local dependence and multidimensionality in graphical loglinear Rasch models. Communications in Statistics: Theory and Methods, 33(6):1239–1276, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kadilar:2005:NRE

[KC05] Cem Kadilar and Hulya Cingi. A new ratio estimator in stratified random sampling. Communications in Statistics: Theory and Methods, 34(3):597–602, 2005. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kadilar:2007:IVE

[KC07] Cem Kadilar and Hulya Cingi. Improvement in variance estimation in simple random sampling. Communications in Statistics: Theory and Methods, 36(11):2075–2081, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kadilar:2008:EPM

[KC08] Cem Kadilar and Hulya Cingi. Estimators for the population mean in the case of missing data. Communications in Statistics: Theory and Methods, 37(14):2226–2236, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2000:CTP

[KCL00] Soohwa Kim, Sinsup Cho, and Sangyeol Lee. On the Cusum test for parameter changes in garch(1, 1) models. Communications in Statistics: Theory and Methods, 29(2):445–462, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kowalski:2000:NMC

[KCV00] Scott Kowalski, John A. Cornell, and G. Geoffrey Vining. A new model and class of designs for mixture experiments with process variables. Communications in Statistics: Theory and Methods, 29(9–10):2255–2280, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610920008832606.

Kim:2008:MMS

[KD08] Sungduk Kim and Dipak K. Dey. Modeling multilevel survival data using frailty models. Communications in Statistics: Theory and Methods, 37(11):1734–1741, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2009:BSE

[KDCW09] Sungduk Kim, Sonali Das, Ming-Hui Chen, and Nicholas Warren. Bayesian structural equations modeling for ordinal response data with missing responses and missing covariates. Communications in Statistics: Theory and Methods, 38(16– 17):2748–2768, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Keen:2001:PA

[KE01] Kevin J. Keen and Robert C. Elston. A problem in ascertainment. Communications in Statistics: Theory and Methods, 30(8–9):1615–1631, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kemp:2001:BGD

[Kem01] Adrienne W. Kemp. The q-beta-geometric distribution as a model for fecundability. Communications in Statistics: Theory and Methods, 30(11):2373–2384, 2001. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kemp:2004:CDL

[Kem04] Adrienne W. Kemp. Classes of discrete lifetime distributions. Communications in Statistics: Theory and Methods, 33(12):3069–3093, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Klein:2006:SSS

[KF06] Ingo Klein and Matthias Fischer. Skewness by splitting the scale parameter. Communications in Statistics: Theory and Methods, 35(7):1159–1171, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Klein:2007:NKO

[KF07] Ingo Klein and Matthias Fischer. A note on the kurtosis ordering of the generalized secant hyperbolic distribution. Communications in Statistics: Theory and Methods, 37(1): 1–7, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2000:SAM

[KG00] Jonghyeon Kim and Leon J. Gleser. SIMEX approaches to measurement error in ROC studies. Communications in Statistics: Theory and Methods, 29(11):2473–2491, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulasekera:2002:VEN

[KG02] K. B. Kulasekera and Colin Gallagher. Variance estimation in nonparametric multiple regression. Communications in Statistics: Theory and Methods, 31(8):1373–1383, 2002. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kundu:2004:CPR

[KG04] Debasis Kundu and Rameshwar D. Gupta. Characterizations of the proportional (reversed) hazard model. Communications in Statistics: Theory and Methods, 33(12):3095–3102, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kayhan:2008:NAL

[KG08] Yasemin Kayhan and Suleyman Gunay. A new approach to least median of squares and regression through the origin. Communications in Statistics: Theory and Methods, 37(5): 773–781, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2000:ISV

[KH00] Choongrak Kim and Soonyoung Hwang. Influential subsets on the variable selection. Communications in Statistics: Theory and Methods, 29(2):335–347, 2000. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Komori:2002:PEB

[KH02] Yoshio Komori and Hideo Hirose. Parameter estimation based on grouped or continuous data for truncated exponential distributions. Communications in Statistics: Theory and Methods, 31(6):889–900, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Keribin:2003:APE

[KH03] C. Keribin and D. Haughton. Asymptotic probabilities of over-estimating and under-estimating the order of a model in general regular families. Communications in Statistics: Theory and Methods, 32(7):1373–1404, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khan:2000:IEM

[Kha00] Shahjahan Khan. Improved estimation of the mean vector for Student-t model. Communications in Statistics: Theory and Methods, 29(3):507–527, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khan:2005:PDR

[Kha05] Shahjahan Khan. Predictive distribution of regression vector and residual sum of squares for normal multiple regression model. Communications in Statistics: Theory and Methods, 33(10):2423–2441, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khan:2008:SEI

[Kha08] Shahjahan Khan. Shrinkage estimators of intercept parameters of two simple regression models with suspected equal slopes. Communications in Statistics: Theory and Methods, 37(2):247–260, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2000:MOD

[Kim00] Myung Geun Kim. Multivariate outliers and decompositions of Mahalanobis distance. Communications in Statistics: Theory and Methods, 29(7):1511–1526, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2002:BRC

[Kim02] Hea-Jung Kim. Binary regression with a class of skewed t link models. Communications in Statistics: Theory and Methods, 31(10):1863–1886, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2006:RTF

[Kim06a] Hea-Jung Kim. On the ratio of two folded normal distributions. Communications in Statistics: Theory and Methods, 35(6):965–977, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2006:HTP

[Kim06b] Jonghyeon Kim. Hypothesis testing problems in an unbalanced longitudinal ophthalmology study. Communications in Statistics: Theory and Methods, 35(3):461–476, 2006. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2007:MCM

[Kim07] Hea-Jung Kim. A Monte Carlo method for estimating prediction limit for the arithmetic mean of lognormal sample. Communications in Statistics: Theory and Methods, 36(11):2159– 2167, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2009:COO

[Kim09] Hea-Jung Kim. Classification of observations into one of two artificially dichotomized classes by using a normal screening variable. Communications in Statistics: Theory and Methods, 38(5):607–620, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kitsos:2002:SLC

[Kit02] Christos P. Kitsos. The simple linear calibration problem as an optimal experimental design. Communications in Statistics: Theory and Methods, 31(7):1167–1177, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Keison:2000:SSS

[KK00a] P. I. Keison and K. E. Kemp. Small sample set estimation of a baseline ranking. Communications in Statistics: Theory and Methods, 29(1):19–43, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:2000:QES

[KK00b] Somesh Kumar and Aditi Kar. Quantile estimation for a selected normal population. Communications in Statistics: Theory and Methods, 29(2):437–444, 2000. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kocherlakota:2001:RBP

[KK01] Subrahmaniam Kocherlakota and Kathleen Kocherlakota. Regression in the bivariate Poisson distribution. Communications in Statistics: Theory and Methods, 30(5):815–825, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kashid:2002:MGC

[KK02a] D. N. Kashid and S. R. Kulkarni. A more general criterion for subset selection in multiple linear regression. Communications in Statistics: Theory and Methods, 31(5):795–811, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2002:IPT

[KK02b] Seong W. Kim and Dal Ho Kim. Intrinsic priors for twosample tests in normal populations. Communications in Statistics: Theory and Methods, 31(7):1091–1105, 2002. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kalagonda:2003:DMC

[KK03] A. A. Kalagonda and S. R. Kulkarni. Diagnosis of multivariate control chart signal based on dummy variable regression technique. Communications in Statistics: Theory and Methods, 32(8):1665–1684, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kemp:2004:FMC

[KK04a] Adrienne W. Kemp and C. David Kemp. Factorial moment characterizations for certain binomial-type distributions. Communications in Statistics: Theory and Methods, 33(12):3059–3068, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kokonendji:2004:SAD

[KK04b] Célestin C. Kokonendji and Mohamed Khoudar. On strict arcsine distribution. Communications in Statistics: Theory and Methods, 33(5):993–1006, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khaledi:2007:SOO

[KK07a] Baha-Eldin Khaledi and Subhash Kochar. Stochastic orderings of order statistics of independent random variables with different scale parameters. Communications in Statistics: Theory and Methods, 36(7):1441–1449, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kus:2007:EPP

[KK07b] Coskun Kus and Mehmet F. Kaya. Estimation for the parameters of the Pareto distribution under progressive censoring. Communications in Statistics: Theory and Methods, 36(7):1359–1365, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koyuncu:2009:FEP

[KK09] Nursel Koyuncu and Cem Kadilar. Family of estimators of population mean using two auxiliary variables in stratified random sampling. Communications in Statistics: Theory and Methods, 38(14):2398–2417, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kallioras:2006:TFG

[KKC06] Athanasios G. Kallioras, Ioannis A. Koutrouvelis, and George C. Canavos. Testing the fit of gamma distributions using the empirical moment generating function. Communications in Statistics: Theory and Methods, 35(3):527–540, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2006:IPT

[KKL06] Dal Ho Kim, Sang Gil Kang, and Woo Dong Lee. Intrinsic priors for testing two normal means with intrinsic Bayes factors. Communications in Statistics: Theory and Methods, 35(1):63–81, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kannan:2007:CCI

[KKM07] Nandini Kannan, Jerome P. Keating, and Robert L. Mason. A comparison of classical and inverse estimators in the calibration problem. Communications in Statistics: Theory and Methods, 36(1):83–95, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Klement:2007:CCH

[KKMS07] Erich Peter Klement, Anna Kolesárová, Radko Mesiar, and Carlo Sempi. Copulas constructed from horizontal sections. Communications in Statistics: Theory and Methods, 36(16): 2901–2911, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2003:SGJ

[KKP03] Choongrak Kim, Woochul Kim, and Byeong U. Park. Skewing and generalized jackknifing in kernel density estimation. Communications in Statistics: Theory and Methods, 32 (11):2153–2162, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2007:ASP

[KKP07] Hee Soo Kim, Young Sub Kwon, and Dong Ho Park. Adaptive sequential preventive maintenance policy and Bayesian consideration. Communications in Statistics: Theory and Methods, 36(6):1251–1269, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kharrati-Kopaei:2007:NMT

[KKSA07] M. Kharrati-Kopaei and S. M. Sadooghi-Alvandi. A new method for testing interaction in unreplicated two-way analysis of variance. Communications in Statistics: Theory and Methods, 36(15):2787–2803, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:2005:NPE

[KKT05] Somesh Kumar, Ajay Kumar, and Yogesh Mani Tripathi. A note on the Pitman estimator of ordered normal means when the variances are unequal. Communications in Statistics: Theory and Methods, 34(11):2115–2122, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2002:KST

[KL02] Youngjin Kim and Sangyeol Lee. On the Kolmogorov– Smirnov type test for testing nonlinearity in time series. Communications in Statistics: Theory and Methods, 31(2): 299–309, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Korolyuk:2004:SMR

[KL04] V. S. Korolyuk and N. Limnios. Semi-Markov random walk in Poisson approximation scheme. Communications in Statistics: Theory and Methods, 33(3):507–516, 2004. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kabaila:2005:SME

[KL05] Paul Kabaila and Chris J. Lloyd. A simple measure of the efficiency of a Buehler confidence limit. Communications in Statistics: Theory and Methods, 34(4):767–774, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Konwerska:2007:IPM

[KLBW07] Małgorzata Konwerska, Fernando López-Blázquez, and Jacek Wesołowski. Identification of product measures by random choice of marginals. Communications in Statistics: Theory and Methods, 36(7):1465–1478, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kuwada:2006:GOP

[KLHT06] Masahide Kuwada, Shujie Lu, Yoshifumi Hyodo, and Eiji Taniguchi. GA-optimal partially balanced fractional 2^{m1+m2} factorial designs of resolution R({00, 10, 01, 20}|Ω) with 2 ≤ m_{1}, m_{2} ≤ 4. Communications in Statistics: Theory and Methods, 35(11):2035–2053, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Klimczak:2007:BBT

[Kli07] Monika Klimczak. Best bounds of k th records from bounded samples. Communications in Statistics: Theory and Methods, 36(7):1451–1464, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kirn:2000:BMS

[KLK00] Dal Ho Kirn, Woo Dong Lee, and Sang Gil Kang. Bayesian model selection for life time data under type II censoring. Communications in Statistics: Theory and Methods, 29 (12):2865–2878, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kramer:2002:OBA

[KM02] Walter Krämer and Francesc Marmol. OLS-based asymptotic inference in linear regression models with trending regressors and AR(p)-disturbances. Communications in Statistics: Theory and Methods, 31(2):261–270, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Keating:2005:PNC

[KM05] Jerome P. Keating and Robert L. Mason. Pitman nearness comparison of the traditional estimator of the coefficient of determination and its adjusted version in linear regression models. Communications in Statistics: Theory and Methods, 34(2):367–374, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kokonendji:2008:PCD

[KM08a] Célestin C. Kokonendji and Dhafer Malouche. A property of count distributions in the Hinde–Demétrio family. Communications in Statistics: Theory and Methods, 37(12):1823– 1834, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kost:2008:TEM

[KM08b] James T. Kost and Michael P. McDermott. Testing equality of means under order restrictions using multiple contrasts. Communications in Statistics: Theory and Methods, 37(19):3029–3039, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kuk:2001:EPC

[KML01] Anthony Y. C. Kuk, T. K. Mak, and W. K. Li. Estimation procedures for categorical survey data with nonignorable nonresponse. Communications in Statistics: Theory and Methods, 30(4):643–663, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Knoth:2009:MSS

[KMPS09] Sven Knoth, Manuel Cabral Morais, António Pacheco, and Wolfgang Schmid. Misleading signals in simultaneous residual schemes for the mean and variance of a stationary process. Communications in Statistics: Theory and Methods, 38 (16–17):2923–2943, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kao:2009:EDE

[KMS09] Ming-Hung Kao, Abhyuday Mandal, and John Stufken. Efficient designs for event-related functional magnetic resonance imaging with multiple scanning sessions. Communications in Statistics: Theory and Methods, 38(16–17):3170– 3182, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kleffe:2001:MNE

[KN01a] Jürgen Kleffe and Ragnar Norberg. Minimum norm estimation of variance components for life insurance data. Communications in Statistics: Theory and Methods, 30(8–9):1591– 1603, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kundu:2001:APT

[KN01b] Debasis Kundu and Swagata Nandi. On asymptotic properties of a two dimensional frequency estimator. Communications in Statistics: Theory and Methods, 30(8–9):1561–1577, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Knapp:2002:VEE

[Kna02] Guido Knapp. Variance estimation in the error components regression model. Communications in Statistics: Theory and Methods, 31(9):1499–1514, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kazemzadeh:2008:PMP

[KNA08] R. B. Kazemzadeh, R. Noorossana, and A. Amiri. Phase I monitoring of polynomial profiles. Communications in Statistics: Theory and Methods, 37(10):1671–1686, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaluszka:2005:BSW

[KO05a] M. Kaluszka and A. Okolewski. Bounds for L-statistics from weakly dependent samples of random length. Communications in Statistics: Theory and Methods, 34(9–10):1899– 1910, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaluszka:2005:SBG

[KO05b] M. Kaluszka and A. Okolewski. Sharp bounds for generalized order statistics via logarithmic moments. Communications in Statistics: Theory and Methods, 34(9–10):1911– 1923, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khang:2007:NGL

[KO07] T. F. Khang and S. H. Ong. A new generalization of the logarithmic distribution arising from the inverse trinomial distribution. Communications in Statistics: Theory and Methods, 36(1):3–21, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kanda:2002:SCR

[KOF02] Takashi Kanda, Megu Ohtaki, and Yasunori Fujikoshi. Simultaneous confidence regions in an extended growth curve model with k hierarchical within-individuals design matrices. Communications in Statistics: Theory and Methods, 31 (9):1605–1616, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koike:2006:IBT

[Koi06] Ken-Ichi Koike. An integral Bhattacharyya type bound for the Bayes risk. Communications in Statistics: Theory and Methods, 35(12):2185–2195, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kolassa:2007:PAE

[Kol07] John E. Kolassa. A proof of the asymptotic equivalence of two-tail probability approximations. Communications in Statistics: Theory and Methods, 36(2):221–228, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kovalchik:2009:BAT

[Kov09] Stephanie Kovalchik. A Bayesian approach to therapy selection in the treatment of refractory rheumatoid arthritis. Communications in Statistics: Theory and Methods, 38 (18):3406–3418, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koyano:2006:MEB

[Koy06] Hitoshi Koyano. Minimax empirical Bayes estimators in multivariate mixed linear models with unequal replications. Communications in Statistics: Theory and Methods, 35(1): 121–131, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2002:RSP

[KP02] Tae Yoon Kim and Cheolyong Park. Regression smoothing parameter selection using cross residuals sum. Communications in Statistics: Theory and Methods, 31(12):2275–2287, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulkarni:2007:NVB

[KP07] H. V. Kulkarni and B. B. Patkure. A new variant of the bivariate setting the clock back to zero property. Communications in Statistics: Theory and Methods, 36(13):2339– 2349, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kozubowski:2008:MBD

[KP08] Tomasz J. Kozubowski and Anna K. Panorska. A mixed bivariate distribution connected with geometric maxima of exponential variables. Communications in Statistics: Theory and Methods, 37(18):2903–2923, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krishna:2009:BGD

[KP09a] Hare Krishna and Pramendra Singh Pundir. A bivariate geometric distribution with applications to reliability. Communications in Statistics: Theory and Methods, 38(7):1079– 1093, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kundu:2009:EPG

[KP09b] Debasis Kundu and Biswabrata Pradhan. Estimating the parameters of the generalized exponential distribution in presence of hybrid censoring. Communications in Statistics: Theory and Methods, 38(12):2030–2041, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kovacevic:2002:LLM

[KR02] Milorad S. Kovacevic and Shesh N. Rai. Log-linear modelling of change using longitudinal survey data. Communications in Statistics: Theory and Methods, 31(10):1815–1835, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kovacevic:2003:PML

[KR03] Milorad S. Kovacević and Shesh N. Rai. A pseudo maximum likelihood approach to multilevel modelling of survey data. Communications in Statistics: Theory and Methods, 32(1): 103–121, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kundhi:2008:TOB

[KR08] Gubhinder Kundhi and Paul Rilstone. The third-order bias of nonlinear estimators. Communications in Statistics: Theory and Methods, 37(16):2617–2633, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kravchuk:2005:RTL

[Kra05] O. Y. Kravchuk. Rank test of location optimal for hyperbolic secant distribution. Communications in Statistics: Theory and Methods, 34(7):1617–1630, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kraus:2008:INC

[Kra08] David Kraus. Identifying nonproportional covariates in the Cox model. Communications in Statistics: Theory and Methods, 37(4):617–625, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kageyama:2005:CNP

[KRB05] Sanpei Kageyama, Sarita Rudra, and Shakti Banerjee. Constructions of nested pairwise efficiency and variance balanced designs. Communications in Statistics: Theory and Methods, 34(7):1571–1578, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kapur:2009:ECB

[KRB^{+}09] Kush Kapur, Anindya Roy, Dulal K. Bhaumik, Robert D. Gibbons, Nicole A. Lazar, John A. Sweeney, Subhash Aryal, and Dave Patterson. Estimation and classification of BOLD responses over multiple trials. Communications in Statistics: Theory and Methods, 38(16–17):3099–3113, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaciranlar:2001:CLE

[KS01] Selahattin Kaçiranlar and Sadullah Sakallioglu. Combining the Liu estimator and the principal component regression estimator. Communications in Statistics: Theory and Methods, 30(12):2699–2705, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kshirsagar:2002:UTA

[KS02] Anant M. Kshirsagar and Kirk Steinhorst. A unified treatment of the analysis of experimental designs using eigenvectors. Communications in Statistics: Theory and Methods, 31(3):435–447, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kubokawa:2004:IEB

[KS04a] Tatsuya Kubokawa and Muni S. Srivastava. Improved empirical Bayes ridge regression estimators under multicollinearity. Communications in Statistics: Theory and Methods, 33 (8):1943–1973, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:2004:PNC

[KS04b] Manoj Kumar and V. K. Srivastava. Pitman nearness and concentration probability comparisons of the sample coefficient of determination and its adjusted version in linear regression models. Communications in Statistics: Theory and Methods, 33(7):1629–1641, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khalaf:2005:CRP

[KS05a] Ghadban Khalaf and Ghazi Shukur. Choosing ridge parameter for regression problems. Communications in Statistics: Theory and Methods, 34(5):1177–1182, 2005. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kollo:2005:ETP

[KS05b] Tõnu Kollo and Muni S. Srivastava. Estimation and testing of parameters in multivariate Laplace distribution. Communications in Statistics: Theory and Methods, 33(10):2363– 2387, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kabaila:2007:REP

[KS07] Paul Kabaila and Khreshna Syuhada. The relative efficiency of prediction intervals. Communications in Statistics: Theory and Methods, 36(15):2673–2686, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khan:2008:ESL

[KS08] Shahjahan Khan and A. K. Md. E. Saleh. Estimation of slope for linear regression model with uncertain prior information and Student-t error. Communications in Statistics: Theory and Methods, 37(16):2564–2581, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kurz:2009:BRT

[KS09] Stefan Kurz and Winfried Stute. The Bahadur representation of the TTT-transform. Communications in Statistics: Theory and Methods, 38(16–17):3183–3189, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulinskaya:2003:PAT

[KSG03] E. Kulinskaya, R. G. Staudte, and H. Gao. Power approximations in testing for unequal means in a one-way ANOVA weighted for unequal variances. Communications in Statistics: Theory and Methods, 32(12):2353–2371, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Karlis:2003:SRS

[KSS03] Dimitris Karlis, Gilbert Saporta, and Antonis Spinakis. A simple rule for the selection of principal components. Communications in Statistics: Theory and Methods, 32(3):643– 666, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ke:2009:ASG

[KSW09] Jau-Chuan Ke, Zheng-Long Su, and Kuo-Hsiung Wang. An availability system with general repair distribution: Statistical inference. Communications in Statistics: Theory and Methods, 38(9):1508–1525, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumbhakar:2001:RPU

[Kum01] Subal C. Kumbhakar. Risk preferences under price uncertainties and production risk. Communications in Statistics: Theory and Methods, 30(8–9):1715–1735, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Karagrigoriou:2005:EAS

[KV05] A. Karagrigoriou and F. Vonta. On efficient AR spectral estimation for long-range predictions. Communications in Statistics: Theory and Methods, 34(7):1587–1603, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2005:SNR

[KW05] Jong-Min Kim and William D. Warde. Some new results on the multinomial randomized response model. Communications in Statistics: Theory and Methods, 34(4):847–856, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2004:TCH

[KWW04] Dong-Yun Kim, Michael Woodroofe, and Yanhong Wu. Testing for a change in the hazard rate with staggered entry. Communications in Statistics: Theory and Methods, 33(9): 2041–2058, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2004:SLE

[KZ04] Kyoungmi Kim and Mai Zhou. Symmetric location estimation/testing by empirical likelihood. Communications in Statistics: Theory and Methods, 33(9):2233–2243, 2004. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lagos:2004:ASO

[Lag04] Bernardo M. Lagos. About second order local power of the likelihood ratio, Wald and score statistics in first order autoregressive and moving average models. Communications in Statistics: Theory and Methods, 33(2):367–379, 2004. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lajmi:2006:SSW

[Laj06] S. Lajmi. Scalar statistics which have constant regression on the mean of a Riesz distribution. Communications in Statistics: Theory and Methods, 35(11):2075–2082, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Labra:2007:ALI

[LAR07] Filidor V. Labra, Reiko Aoki, and Francisco Rojas. An application of the local influence diagnostics to ridge regression under elliptical model. Communications in Statistics: Theory and Methods, 36(4):767–779, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Last:2008:CPE

[Las08] Michael Last. Change point estimation with independent observations and piece-wise continuous variance function. Communications in Statistics: Theory and Methods, 37(5): 722–736, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lindkvist:2005:APE

[LB05] Håkan Lindkvist and Yuri Belyaev. Asymptotic properties of estimators in a model of life data with warnings. Communications in Statistics: Theory and Methods, 34(2):461–474, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lynch:2007:EPD

[LB07] Caroline Lynch and Daniel Barry. Estimating probability distributions by observing betting practices. Communications in Statistics: Theory and Methods, 37(1):8–17, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lachos:2007:LBI

[LBAVM07] Víctor H. Lachos, Heleno Bolfarine, Reinaldo B. Arellano-Valle, and Lourdes C. Montenegro. Likelihood-based inference for multivariate skew-normal regression models. Communications in Statistics: Theory and Methods, 36(9):1769– 1786, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lopez-Blazquez:2001:EAR

[LBSM01] F. López-Blázquez and B. Salamanca-Miño. Exact and approximated relations between negative hypergeometric and negative binomial probabilities. Communications in Statistics: Theory and Methods, 30(5):957–967, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lopez-Blazquez:2009:RTR

[LBSM09] F. López-Blázquez and B. Salamanca-Miño. Random translation of records from geometric distributions. Communications in Statistics: Theory and Methods, 38(12):2067–2077, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2009:MIM

[LBZ09] Nianjun Liu, Richard Bucala, and Hongyu Zhao. Modeling informatively missing genotypes in haplotype analysis. Communications in Statistics: Theory and Methods, 38(18):3445– 3460, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2000:OTS

[LC00] Shang P. Lin and T. Timothy Chen. Optimal two-stage designs for Phase II clinical trials with differentiation of complete and partial responses. Communications in Statistics: Theory and Methods, 29(5–6):923–940, 2000. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2005:PBD

[LC05a] Pei-Sheng Lin and Murray K. Clayton. Properties of binary data generated from a truncated Gaussian random field. Communications in Statistics: Theory and Methods, 34(3): 537–544, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2005:RVS

[LC05b] Yu-Chang Lin and Chao-Yu Chou. Robustness of the variable sample size and control limit^{−}X chart to non normality. Communications in Statistics: Theory and Methods, 34(3): 721–743, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2008:AOL

[LC08] Kuo-Chin Lin and Yi-Ju Chen. Assessing ordinal logistic regression models via nonparametric smoothing. Communications in Statistics: Theory and Methods, 37(6):917–930, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2008:TSC

[LCC08] Feng Liu, Gemai Chen, and Min Chen. Testing serial correlation in partial linear errors-in-variables models based on empirical likelihood. Communications in Statistics: Theory and Methods, 37(12):1905–1918, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2006:NTP

[LCF06] Xiaohu Li, Wenqing Cao, and Xiuying Feng. A new test procedure for decreasing mean residual life. Communications in Statistics: Theory and Methods, 35(12):2171–2183, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2001:PTO

[LCL01] Mei Lu, Gary Chase, and Shuhui Li. Permutation tests and other test statistics for ill-behaved data: experience of the NINDS t-PA Stroke Trial. Communications in Statistics: Theory and Methods, 30(7):1481–1496, 2001. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2009:SAP

[LCL09] Tai-Fang C. Lu, Jen-Loung Chen, and Jau-Yau Lu. Saddlepoint approximations for a product and its application. Communications in Statistics: Theory and Methods, 38(7): 953–963, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liao:2000:CST

[LD00] Qiming Liao and Victor De Gruttola. A class of semiparametric transformation frailty models with application to estimating treatment effects and heterogeneity in AIDS clinical trails. Communications in Statistics: Theory and Methods, 29(5–6):1039–1057, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lahiri:2008:LLM

[LD08] Soumi Lahiri and Sunil K. Dhar. Log-linear modeling under generalized inverse sampling scheme. Communications in Statistics: Theory and Methods, 37(8):1237–1244, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ledoux:2006:PLT

[Led06] James Ledoux. A Poisson limit theorem for reliability models based on Markov chains. Communications in Statistics: Theory and Methods, 35(1):173–196, 2006. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ledoux:2007:FEA

[Led07] James Ledoux. Filtering and the EM-algorithm for the Markovian arrival process. Communications in Statistics: Theory and Methods, 36(14):2577–2593, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ledolter:2008:STS

[Led08] Johannes Ledolter. Smoothing time series with local polynomial regression on time. Communications in Statistics: Theory and Methods, 37(6):959–971, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2003:NBP

[Lee03] Min-Young Lee. New bounds of the probability of a union. Communications in Statistics: Theory and Methods, 32(11): 2089–2099, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2005:PPN

[Lee05a] Oesook Lee. Probabilistic properties of a nonlinear ARMA process with Markov switching. Communications in Statistics: Theory and Methods, 34(1):193–204, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2005:USI

[Lee05b] Yoon-Dong Lee. Unified solutions of integral equations of SPRT for exponential random variables. Communications in Statistics: Theory and Methods, 33(1):65–74, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2007:E

[Lee07] Andy Lee. Erratum. Communications in Statistics: Theory and Methods, 36(8):1655, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leung:2002:PLL

[Leu02] Chi-Ying Leung. Performance of the location linear discriminant function under across-location heteroscedasticity. Communications in Statistics: Theory and Methods, 31(6):1031– 1044, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lopez-Fidalgo:2007:OST

[LFRLFG07] J. López-Fidalgo, M. J. Rivas-López, and B. Fernández-Garzón. A-optimality standardized through the coefficient of variation. Communications in Statistics: Theory and Methods, 36(4):781–792, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2003:NWD

[LG03] Kaican Li and Zhi Geng. The noncentral Wishart distribution and related distributions. Communications in Statistics: Theory and Methods, 32(1):33–45, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2001:SSI

[LH01] W. Liu and A. J. Hayter. Selecting and sharpening inferences in simultaneous inference with a Bayesian approach. Communications in Statistics: Theory and Methods, 30(1): 135–145, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2002:SCW

[LH02] Sungho Lee and Dug Hun Hong. On a strongly consistent wavelet density estimator for the deconvolution problem. Communications in Statistics: Theory and Methods, 31 (8):1259–1272, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liau:2007:CEC

[LH07] Pen-Hwang Liau and Pi-Hsiang Huang. Clear effects in the 2^{n−p} combined design. Communications in Statistics: Theory and Methods, 36(4):793–799, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:2009:EBE

[LH09] Tachen Liang and Wen-Tao Huang. Empirical Bayes estimation for uniform distributions with random right censoring. Communications in Statistics: Theory and Methods, 38 (20):3713–3724, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2006:EME

[LHP06] Songlin Liu, Zheng Haiqi, and Xu Ping. Evaluation method for equipment state quality based on extension theory. Communications in Statistics: Theory and Methods, 35(10):1883– 1895, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2007:SPM

[LHQ07] Jun Liu, Weili He, and Hui Quan. Statistical properties of a modified accelerated design for Phase I cancer clinical trials. Communications in Statistics: Theory and Methods, 36(15):2741–2759, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liao:2000:OFF

[LI00] C. T. Liao and H. K. Iyer. Optimal 2^{n−p} fractional factorial designs for dispersion effects under a location-dispersion model. Communications in Statistics: Theory and Methods, 29(4):823–835, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2001:NDM

[Li01] Heng Li. Nested designs with multivariate measurement: an illustration of the structural approach to random effects multivariate analysis of variance. Communications in Statistics: Theory and Methods, 30(5):887–895, 2001. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2002:HRE

[Li02a] Linyuan Li. Hazard rate estimation for censored data by wavelet methods. Communications in Statistics: Theory and Methods, 31(6):943–960, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2002:BEB

[Li02b] Tze Fen Li. Bayes empirical Bayes approach to estimation of the failure rate in exponential distribution. Communications in Statistics: Theory and Methods, 31(9):1457–1465, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2008:NTE

[Li08a] Bo Li. Nonparametric testing of an exclusion restriction in quantile regression. Communications in Statistics: Theory and Methods, 37(18):2877–2889, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2008:DMD

[Li08b] Haijun Li. Duality of the multivariate distributions of Marshall–Olkin type and tail dependence. Communications in Statistics: Theory and Methods, 37(11):1721–1733, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2008:CIL

[Li08c] Xinmin Li. Confidence intervals for linear function of variance components in two-fold nested random effects model. Communications in Statistics: Theory and Methods, 37(2): 204–212, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2009:TET

[Li09a] Fang Li. Testing for the equality of two autoregressive functions using quasi-residuals. Communications in Statistics: Theory and Methods, 38(9):1404–1421, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2009:SRA

[Li09b] Jialiang Li. Semiparametric residuals and analysis for a scleroderma clinical trial. Communications in Statistics: Theory and Methods, 38(18):3339–3350, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2009:KED

[Li09c] Jiexiang Li. Kernel estimators for distribution functions on dependent random fields. Communications in Statistics: Theory and Methods, 38(20):3678–3685, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:2000:EBT

[Lia00] Tachen Liang. Empirical Bayes tests in a positive exponential family with partial information on priors. Communications in Statistics: Theory and Methods, 29(3):565–577, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:2003:EBT

[Lia03] TaChen Liang. On empirical Bayes testing for a location parameter in a shifted gamma distribution. Communications in Statistics: Theory and Methods, 32(1):123–138, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:2005:EBT

[Lia05] Tachen Liang. On empirical Bayes testing for reliability. Communications in Statistics: Theory and Methods, 34(3): 661–670, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:2006:EBT

[Lia06] Tachen Liang. Empirical Bayes testing for the mean lifetime of exponential distributions: Unequal sample sizes case. Communications in Statistics: Theory and Methods, 35(8): 1409–1428, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:2007:EBT

[Lia07a] Tachen Liang. Empirical Bayes testing for double exponential distributions. Communications in Statistics: Theory and Methods, 36(8):1543–1553, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liao:2007:AMR

[Lia07b] Jason J. Z. Liao. An accelerated model for recurrence data. Communications in Statistics: Theory and Methods, 36(4): 821–831, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lian:2008:BEB

[Lia08a] Heng Lian. Bayes and empirical Bayes inference in changepoint problems. Communications in Statistics: Theory and Methods, 38(3):419–430, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:2008:EBE

[Lia08b] Tachen Liang. Empirical Bayes estimation of reliability in a positive exponential family. Communications in Statistics: Theory and Methods, 37(13):2052–2065, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:2008:EBT

[Lia08c] Tachen Liang. On empirical Bayes testing in a discrete life model with random censoring. Communications in Statistics: Theory and Methods, 37(15):2369–2381, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lian:2009:RCP

[Lia09a] Heng Lian. On rates of convergence for posterior distributions under misspecification. Communications in Statistics: Theory and Methods, 38(11):1893–1900, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:2009:SVS

[Lia09b] Tachen Liang. Simultaneous variable sampling inspection for finite populations. Communications in Statistics: Theory and Methods, 38(10):1589–1606, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liebscher:2005:SDE

[Lie05] Eckhard Liebscher. Semiparametric density estimators using copulas. Communications in Statistics: Theory and Methods, 34(1):59–71, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lim:2007:EEF

[Lim07] Johan Lim. Estimation of the entropy functional from dependent samples. Communications in Statistics: Theory and Methods, 36(8):1577–1589, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2001:NCT

[Lin01] Chien-Tai Lin. Nonparametric classification on two univariate distributions. Communications in Statistics: Theory and Methods, 30(2):319–330, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2004:PIA

[Lin04] Fei Lin. A probability inequality and its applications in selecting best normal means. Communications in Statistics: Theory and Methods, 33(2):321–330, 2004. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2005:ADU

[Lin05] Hung-Chin Lin. The asymptotic distributions of unified process capability indices. Communications in Statistics: Theory and Methods, 33(4):909–918, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2001:TSA

[Liu01a] Jeng-Fu Liu. Two-stage approximation of expected reward for gamma random variables. Communications in Statistics: Theory and Methods, 30(7):1471–1480, 2001. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2001:AEA

[Liu01b] Zhijun Liu. Asymptotic efficiency of analysis of Rao’s quadratic entropy. Communications in Statistics: Theory and Methods, 30(8–9):1497–1506, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2003:TSE

[Liu03a] Jinshan Liu. Two-stage estimators of seemingly unrelated regressions with elliptical distribution. Communications in Statistics: Theory and Methods, 32(6):1251–1263, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2003:ULT

[Liu03b] Kejian Liu. Using Liu-type estimator to combat collinearity. Communications in Statistics: Theory and Methods, 32 (5):1009–1020, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2004:MLT

[Liu04] Kejian Liu. More on Liu-type estimator in linear regression. Communications in Statistics: Theory and Methods, 33 (11):2723–2733, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Laradji:2006:IAS

[LJ06] A. Laradji and Anwar H. Joarder. Inequalities among some measures of location. Communications in Statistics: Theory and Methods, 35(11):1963–1970, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2001:CCD

[LL01] Hua Lin and Andy H. Lee. A conditional Cook’s distance to assess influence in autoregressive models. Communications in Statistics: Theory and Methods, 30(7):1373–1380, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Laib:2002:CHT

[LL02] Naâmane Laib and Djamal Louani. On the conditional homoscedasticity test in autoregressive model with arch error. Communications in Statistics: Theory and Methods, 31(7):1179–1202, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2003:THF

[LL03] Sungim Lee and Sangyeol Lee. Testing heterogeneity for frailty distribution in shared frailty model. Communications in Statistics: Theory and Methods, 32(11):2245–2253, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2005:CIS

[LL05a] Xinmin Li and Guoying Li. Confidence intervals on sum of variance components with unbalanced designs. Communications in Statistics: Theory and Methods, 34(4):833–845, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Limnios:2005:HIS

[LL05b] Nikolaos Limnios and Mei-Ling Ting Lee. Hoeffding’s inequality for stopped martingales and semi-Markov processes. Communications in Statistics: Theory and Methods, 34(3): 713–720, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2007:TPC

[LL07] Taewook Lee and Sangyeol Lee. Test for parameter change in linear processes based on Whittle’s estimator. Communications in Statistics: Theory and Methods, 36(11):2129–2141, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2008:MMP

[LL08] Xiaohu Li and Zhouping Li. A mixture model of proportional reversed hazard rate. Communications in Statistics: Theory and Methods, 37(18):2953–2963, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leffondre:2001:ONT

[LLCNM01] Karen Leffondré, Joseph Lellouch, Catherine Com-Nougué, and Thierry Moreau. Optimality of nonparametric tests for association between survival time and continuous covariates. Communications in Statistics: Theory and Methods, 30(5): 913–929, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2009:MDC

[LLN09] Sangyeol Lee, Youngmi Lee, and Okyoung Na. Monitoring distributional changes in autoregressive models. Communications in Statistics: Theory and Methods, 38(16–17):2969– 2982, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lai:2009:BAS

[LLX09] Tze Leung Lai, Tongwei Liu, and Haipeng Xing. A Bayesian approach to sequential surveillance in exponential families. Communications in Statistics: Theory and Methods, 38(16– 17):2958–2968, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2004:LAB

[LM04] Yiqiang Lu and Shisong Mao. Local asymptotics for B-spline estimators of the varying coefficient model. Communications in Statistics: Theory and Methods, 33(5):1119–1138, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lillo:2006:BEU

[LM06] Rosa E. Lillo and Miguel Martín. On the Bayesian estimation for the uniform scale parameter via a functional equation. Communications in Statistics: Theory and Methods, 35 (12):2197–2210, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2007:CIS

[LM07] Xinmin Li and Zhaowei Meng. Confidence intervals on sum of variance components in unbalanced two-fold nested designs with equal subsampling. Communications in Statistics: Theory and Methods, 36(2):387–397, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2006:SMP

[LMG06] Euikyoo Lee, Myung-Sang Moon, and Richard F. Gunst. A spatiotemporal model for pollutant concentrations. Communications in Statistics: Theory and Methods, 35(3):489–497, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2001:GTD

[LMM01] Baibing Li, Elaine B. Martin, and A. Julian Morris. A graphical technique for detecting influential cases in regression analysis. Communications in Statistics: Theory and Methods, 30(3):463–483, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lachenbruch:2009:DET

[LMR09] Peter A. Lachenbruch, Frederick W. Miller, and Lisa G. Rider. On determining the effects of therapy on disease damage in non randomized studies with multiple treatments: A study of juvenile myositis. Communications in Statistics: Theory and Methods, 38(18):3268–3281, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Louzada-Neto:2005:HSI

[LN05] Francisco Louzada-Neto. A hybrid scale intensity model for recurrent event data. Communications in Statistics: Theory and Methods, 33(1):119–133, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:2009:APE

[LN09] Han-Ying Liang and Si-Li Niu. Asymptotic properties of error density estimator in regression model under α-mixing assumptions. Communications in Statistics: Theory and Methods, 38(6):761–783, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2009:SIT

[LNC09] Chien-Tai Lin, Hon Keung Tony Ng, and Ping Shing Chan. Statistical inference of Type-II progressively hybrid censored data with Weibull lifetimes. Communications in Statistics: Theory and Methods, 38(10):1710–1729, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2001:OSR

[LNP01] Chong Hyung Lee, Kyung Hyun Nam, and Dong Ho Park. Optimal software release policy based on Markovian perfect debugging model. Communications in Statistics: Theory and Methods, 30(11):2329–2342, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Louzada-Neto:2002:ALT

[LNP02] Francisco Louzada-Neto and Gleici Castro Perdoná. Accelerated lifetime tests with a log-non-linear stress-response relationship. Communications in Statistics: Theory and Methods, 31(1):129–146, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Loperfido:2007:SRB

[LNRS07] Nicola Loperfido, Jorge Navarro, Jose M. Ruiz, and Carlos J. Sandoval. Some relationships between skew-normal distributions and order statistics from exchangeable normal random vectors. Communications in Statistics: Theory and Methods, 36(9):1719–1733, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Limnios:2005:EES

[LOS05] N. Limnios, B. Ouhbi, and A. Sadek. Empirical estimator of stationary distribution for semi-Markov processes. Communications in Statistics: Theory and Methods, 34(4):987–995, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lemdani:2007:ABH

[LOS07] Mohamed Lemdani and Elias Ould-Saïd. Asymptotic behavior of the hazard rate kernel estimator under truncated and censored data. Communications in Statistics: Theory and Methods, 36(1):155–173, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Loubes:2008:PIP

[Lou08] Jean-Michel Loubes. l^{1} Penalty for ill-posed inverse problems. Communications in Statistics: Theory and Methods, 37(9):1399–1411, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2005:NRN

[LP05] Jyh-Jiuan Lin and Nabendu Pal. A note on robustness of normal variance estimators under t-distributions. Communications in Statistics: Theory and Methods, 34(5):1117–1126, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lallich:2007:P

[LP07] Stéphane Lallich and Dominique Pastor. Preface. Communications in Statistics: Theory and Methods, 36(14):2475– 2476, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2005:LCM

[LQ05] Ji-Hyun Lee and Bahjat F. Qaqish. A latent changepoint model using a generalized estimating equations approach. Communications in Statistics: Theory and Methods, 34(5): 1233–1242, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lobus:2000:LPD

[LR00] Jörg-Uwe Löbus and Lutz Ritter. The limiting power of the Durbin–Watson test. Communications in Statistics: Theory and Methods, 29(12):2665–2676, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

LePage:2001:CVS

[LR01] Raoul LePage and Michal Ryznar. Conditioning vs. standardization for contrasts on errors attracted to stable laws.

Communications in Statistics: Theory and Methods, 30(8– 9):1829–1850, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2008:NEF

[LR08] I-Li Lu and Donald St. P. Richards. Natural exponential families and generalized hypergeometric measures. Communications in Statistics: Theory and Methods, 37(16):2472– 2487, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2009:AMI

[LR09] Rong Liu and Viswanathan Ramakrishnan. Application of multiple imputation in analysis of data from clinical trials with treatment related dropouts. Communications in Statistics: Theory and Methods, 38(20):3666–3677, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2008:PMA

[LRG08] Xu-Qing Liu, Jian-Ying Rong, and Feng Gao. Permutation matrix with application to optimal invariant quadratic prediction in finite populations. Communications in Statistics: Theory and Methods, 37(18):2825–2835, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lefante:2002:RPL

[LS02] John J. Lefante, Jr. and Arvind K. Shah. Robustness properties of lognormal confidence intervals for lognormal and gamma distributed data. Communications in Statistics: Theory and Methods, 31(11):1939–1957, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liau:2009:CIE

[LS09] Pen-Hwang Liau and Shalabh. Confidence interval estimation in ultrastructural model. Communications in Statistics: Theory and Methods, 38(5):675–681, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2007:ELS

[LSD07] Xuewen Lu, R. S. Singh, and A. F. Desmond. Estimation for location-scale models with censored data. Communications in Statistics: Theory and Methods, 36(6):1069–1089, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lorino:2004:CSR

[LSRD04] T. Lorino, M. Sanaa, S. Robin, and J.-J. Daudin. Comparison of semiparametric regression models for correlated survival data using simulations. Communications in Statistics: Theory and Methods, 33(8):1975–1991, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2005:NEH

[LSW05] Aiyi Liu, Enrique F. Schisterman, and Chengqing Wu. Nonparametric estimation and hypothesis testing on the partial area under receiver operating characteristic curves. Communications in Statistics: Theory and Methods, 34(9–10):2077– 2088, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2002:SCE

[LT02] Chin-Shang Li and Jeremy M. G. Taylor. Smoothing covariate effects in cure models. Communications in Statistics: Theory and Methods, 31(3):477–493, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2004:CWD

[LT04] Sungho Lee and Robert Lee Taylor. On an L_{1} consistent wavelet density estimator for the deconvolution problem. Communications in Statistics: Theory and Methods, 33 (8):1917–1929, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2008:NSV

[LT08] Sungho Lee and Robert L. Taylor. A note on support vector density estimation for the deconvolution problem. Communications in Statistics: Theory and Methods, 37(3):328–336, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lyles:2001:APA

[LTHK01] Robert H. Lyles, Douglas J. Taylor, John J. Hanfelt, and Lawrence L. Kupper. An alternative parametric approach for discrete missing data problems. Communications in Statistics: Theory and Methods, 30(10):1969–1988, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tian:2008:UMC

[lTTNlT08] Guo liang Tian, Ming Tan, Kai Wang Ng, and Man lai Tang. A unified method for checking compatibility and uniqueness for finite discrete conditional distributions. Communications in Statistics: Theory and Methods, 38(1):115–129, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2000:ICI

[Lu00] Wang-Shu Lu. Improved confidence intervals for a binomial parameter using the Bayesian method. Communications in Statistics: Theory and Methods, 29(12):2835–2847, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Luceno:2001:SSS

[Luc01] Alberto Luceño. Selection of sample size for discrete feedback dead-band control schemes. Communications in Statistics: Theory and Methods, 30(4):679–689, 2001. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lue:2008:SAV

[Lue08] Heng-Hui Lue. Sliced average variance estimation for censored data. Communications in Statistics: Theory and Methods, 37(20):3276–3286, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lund:2002:TBR

[Lun02] Ulric J. Lund. Tree-based regression for a circular response. Communications in Statistics: Theory and Methods, 31(9): 1549–1560, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lumelskii:2007:GCR

[LVNF07] Ya. Lumelskii, V. Voinov, M. Nikulin, and P. Feigin. On a generalization of the classical random sampling scheme and unbiased estimation. Communications in Statistics: Theory and Methods, 36(4):693–705, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Luque-Vasquez:2004:TRE

[LVVA04] Fernando Luque-Vásquez and Oscar Vega-Amaya. Time and ratio expected average cost optimality for semi-Markov control processes on Borel spaces. Communications in Statistics: Theory and Methods, 33(3):715–734, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2002:SSP

[LW02] Ming Lin and Laisheng Wei. The small sample properties of the principal components estimator for regression coefficients. Communications in Statistics: Theory and Methods, 31(2):271–283, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2003:THN

[LW03] Jin-Guan Lin and Bo-Cheng Wei. Testing for heteroscedasticity in nonlinear regression models. Communications in Statistics: Theory and Methods, 32(1):171–192, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2004:THC

[LW04] Jin-Guan Lin and Bo-Cheng Wei. Testing for heteroscedasticity and/or correlation in nonlinear models with correlated errors. Communications in Statistics: Theory and Methods, 33(2):251–275, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2007:THA

[LW07a] Jin-Guan Lin and Bo-Cheng Wei. Testing for heteroscedasticity and/or autocorrelation in longitudinal mixed effect nonlinear models with AR(1) errors. Communications in Statistics: Theory and Methods, 36(3):567–586, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2007:DSC

[LW07b] Ivy Liu and Dong Q. Wang. Diagnostics for stratified clinical trials in proportional odds models. Communications in Statistics: Theory and Methods, 36(1):211–220, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2003:PEL

[LWB03] Chien-Tai Lin, Sam J. S. Wu, and N. Balakrishnan. Parameter estimation for the linear hazard rate distribution based on records and inter-record times. Communications in Statistics: Theory and Methods, 32(4):729–748, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2004:IEP

[LWB04] Chien-Tai Lin, Sam J. S. Wu, and N. Balakrishnan. Interval estimation of parameters of log-gamma distribution based on progressively censored data. Communications in Statistics: Theory and Methods, 33(11):2595–2626, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2006:ILG

[LWB06] Chien-Tai Lin, Sam J. S. Wu, and N. Balakrishnan. Inference for log-gamma distribution based on progressively Type-II censored data. Communications in Statistics: Theory and Methods, 35(7):1271–1292, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2003:THC

[LWW03] Ying-An Liu, Bo-Cheng Wei, and Hai-Bin Wang. Testing of homogeneity for correlation and variance in nonlinear regression models with DBL(p, 0, 1) random errors. Communications in Statistics: Theory and Methods, 32(12):2441–2463, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2007:ALE

[LWZ07] Chang-Yu Lu, Jian-Hong Wu, and Bin-Yan Zhang. Admissible linear estimators with respect to inequality constraints. Communications in Statistics: Theory and Methods, 36(5): 961–966, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2008:ELC

[LX08] Gaorong Li and Liugen Xue. Empirical likelihood confidence region for the parameter in a partially linear errors-in-variables model. Communications in Statistics: Theory and Methods, 37(10):1552–1564, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2009:ZPH

[LX09] Fei Li and Yifeng Xue. Zonal polynomials and hypergeometric functions of quaternion matrix argument. Communications in Statistics: Theory and Methods, 38(8):1184–1206, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2006:CCT

[LXG06] J. Y. Liu, M. Xie, and T. N. Goh. CUSUM chart with transformed exponential data. Communications in Statistics: Theory and Methods, 35(10):1829–1843, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leung:2000:HER

[LY00] Siu Fai Leung and Shihti Yu. How effective are the reset tests for omitted variables. Communications in Statistics: Theory and Methods, 29(4):879–902, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2005:BEB

[LY05] Tze Fen Li and Tseng-Chang Yen. A Bayes empirical Bayes decision rule for classification. Communications in Statistics: Theory and Methods, 34(5):1137–1149, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2000:MQT

[LZ00] Zhaohai Li and Heping Zhang. Mapping quantitative trait loci using both extremely discordant and concordant sib pairs: a unified approach for meta-analysis. Communications in Statistics: Theory and Methods, 29(5–6):1115–1127, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leton:2001:EBS

[LZ01] E. Letón and P. Zuluaga. Equivalence between score and weighted tests for survival curves. Communications in Statistics: Theory and Methods, 30(4):591–608, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leton:2007:NVT

[LZ07] E. Letón and P. Zuluaga. A note on the variances of the tests of Kendall, jonckheere, and terpstra. Communications in Statistics: Theory and Methods, 36(5):927–937, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2000:TNB

[LZJ00] Xiaohu Li, Chengkui Zhong, and Bingyi Jing. Test for new better than used in convex ordering. Communications in Statistics: Theory and Methods, 29(12):2751–2760, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2008:ELR

[LZZ08a] Yukun Liu, Changliang Zou, and Runchu Zhang. Empirical likelihood ratio test for a change-point in linear regression model. Communications in Statistics: Theory and Methods, 37(16):2551–2563, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2008:PSE

[LZZ08b] Yiqiang Lu, Riquan Zhang, and Liping Zhu. Penalized spline estimation for varying-coefficient models. Communications in Statistics: Theory and Methods, 37(14):2249–2261, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Montopoli:2001:ADC

[MA01] George Montopoli and Donald A. Anderson. The analysis of discrete choice experiments with correlated error structure. Communications in Statistics: Theory and Methods, 30(4): 615–626, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ma:2006:MAM

[Ma06] Jun Ma. Multiplicative algorithms for maximum penalized likelihood inversion with non negative constraints and generalized error distributions. Communications in Statistics: Theory and Methods, 35(5):831–848, 2006. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mack:2002:BCC

[Mac02] Y. P. Mack. Bias-corrected confidence intervals for wildlife abundance estimation. Communications in Statistics: Theory and Methods, 31(7):1107–1122, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Macci:2008:LDE

[Mac08] Claudio Macci. Large deviations for empirical estimators of the stationary distribution of a semi-Markov process with finite state space. Communications in Statistics: Theory and Methods, 37(19):3077–3089, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mak:2002:MEV

[Mak02] Tak K. Mak. Modeling and estimating variances in regression. Communications in Statistics: Theory and Methods, 31(3):351–365, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Makkonen:2008:BCP

[Mak08] Lasse Makkonen. Bringing closure to the plotting position controversy. Communications in Statistics: Theory and Methods, 37(3):460–467, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Malouche:2009:MGM

[Mal09] Dhafer Malouche. Mixed graphical model selection using Holm’s procedure. Communications in Statistics: Theory and Methods, 38(9):1453–1464, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Man:2000:MLE

[Man00a] K. S. Man. Maximum likelihood estimation for a nearly random walk model. Communications in Statistics: Theory and Methods, 29(3):677–697, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Manda:2000:BAM

[Man00b] Samuel O. M. Manda. A Bayesian analysis of multivariate survival data from multi-stage cluster sampling. Communications in Statistics: Theory and Methods, 29(4):769–782, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marohn:2000:EEL

[Mar00] Michael Falk Frank Marohn. Efficient estimation in local parametric regression analysis. Communications in Statistics: Theory and Methods, 29(11):2363–2389, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marohn:2002:CGP

[Mar02a] Frank Marohn. A characterization of generalized Pareto distributions by progressive censoring schemes and goodness-of-fit tests. Communications in Statistics: Theory and Methods, 31(7):1055–1065, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marohn:2002:CLM

[Mar02b] Frank Marohn. A comment on locally most powerful tests in the presence of nuisance parameters. Communications in Statistics: Theory and Methods, 31(3):337–349, 2002. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marohn:2005:TIE

[Mar05] F. Marohn. Tail index estimation in models of generalized order statistics. Communications in Statistics: Theory and Methods, 34(5):1057–1064, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marin:2007:EVC

[Mar07] Jean-Michel Marin. Estimation of variance components for a linear Toeplitz model. Communications in Statistics: Theory and Methods, 36(12):2273–2288, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marsh:2008:CIP

[Mar08a] Patrick Marsh. Conditional information in projections of Gaussian vectors. Communications in Statistics: Theory and Methods, 38(3):332–339, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maruyama:2008:MMK

[Mar08b] Yosihito Maruyama. A measure of multivariate kurtosis with principal components. Communications in Statistics: Theory and Methods, 37(13):2116–2123, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mousa:2005:BPP

[MAS05] M. A. M. Ali Mousa and S. A. Al-Sagheer. Bayesian prediction for progressively Type-II censored data from the Rayleigh model. Communications in Statistics: Theory and Methods, 34(12):2353–2361, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meintanis:2007:DTT

[MB07a]  Simos G. Meintanis and Yiannis Bassiakos.  Data-

transformation and test of fit for the generalized Pareto hypothesis. Communications in Statistics: Theory and Methods, 36(4):833–849, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mollah:2007:PIV

[MB07b] Md. Nurul Haque Mollah and S. K. Bhattacharjee. Predictive influence of variables in a multivariate distribution in presence of perfect multicollinearity. Communications in Statistics: Theory and Methods, 37(1):121–136, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukhopadhyay:2007:BAM

[MB07c] Chiranjit Mukhopadhyay and Sanjib Basu. Bayesian analysis of masked series system lifetime data. Communications in Statistics: Theory and Methods, 36(2):329–348, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marcheselli:2008:PED

[MBB08] M. Marcheselli, A. Baccini, and L. Barabesi. Parameter estimation for the discrete stable family. Communications in Statistics: Theory and Methods, 37(6):815–830, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maravelakis:2002:ICV

[MBPP02] P. E. Maravelakis, S. Bersimis, J. Panaretos, and S. Psarakis. Identifying the out of control variable in a multivariate control chart. Communications in Statistics: Theory and Methods, 31(12):2391–2408, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Magalhaes:2017:CCM

[MBS17] Tiago M. Magalhães, Denise A. Botter, and Mônica C. Sandoval. Corrigendum to: “Covariance matrix formula for generalized linear models with unknown dispersion” by G. M. Cordeiro, L. P. Barroso, and D. A. Botter [Communications in Statistics — Theory and Methods (2006) 35(1), 113–120]. Communications in Statistics: Theory and Methods, 46(21):10702–10704, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1242745. See [CBB06].

McElroy:2008:SPM

[McE08] Tucker McElroy. Statistical properties of model-based signal extraction diagnostic tests. Communications in Statistics: Theory and Methods, 37(4):591–616, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McGee:2007:TMP

[McG07] Monnie McGee. Tests for multiple peaks in the spectra of categorical time series. Communications in Statistics: Theory and Methods, 36(16):2891–2900, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mason:2008:IVC

[MCY08] Robert L. Mason, Youn-Min Chou, and John C. Young. Identifying variables contributing to outliers in Phase I. Communications in Statistics: Theory and Methods, 37(7): 1103–1118, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mason:2009:MVM

[MCY09] Robert L. Mason, Youn-Min Chou, and John C. Young. Monitoring variation in a multivariate process when the dimension is large relative to the sample size. Communications in Statistics: Theory and Methods, 38(6):939–951, 2009. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McClean:2000:NML

[MD00] Sally McClean and Colum Devine. Nonparametric maximum likelihood estimation for artificially truncated absence data. Communications in Statistics: Theory and Methods, 29 (11):2439–2457, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Molinari:2003:VFI

[MD03] Nicolas Molinari and Jean-Pierre Daurès. In vitro fertilization with immune patients. Communications in Statistics: Theory and Methods, 32(9):1859–1870, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Micheas:2004:MLR

[MD04] Athanasios C. Micheas and Dipak K. Dey. On measuring loss robustness using maximum a posteriori estimate. Communications in Statistics: Theory and Methods, 33(5):1069–1085, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Micheas:2007:RBF

[MD07] Athanasios C. Micheas and Dipak K. Dey. Reconciling Bayesian and frequentist evidence in the one-sided scale parameter testing problem. Communications in Statistics: Theory and Methods, 36(6):1123–1138, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meintanis:2005:COT

[Mei05] Simos G. Meintanis. A class of omnibus tests for the Laplace distribution based on the empirical characteristic function. Communications in Statistics: Theory and Methods, 33(4): 925–948, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meintanis:2008:PMA

[Mei08] Simos G. Meintanis. A powerful method of assessing the fit of the lognormal distribution. Communications in Statistics: Theory and Methods, 37(12):1948–1958, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mentz:2001:LRT

[Men01] Raul P. Mentz. Likelihood ratio test of equality of several variances in the intraclass correlation model. Communications in Statistics: Theory and Methods, 30(1):75–91, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menzefricke:2002:ECC

[Men02] Ulrich Menzefricke. On the evaluation of control chart limits based on predictive distributions. Communications in Statistics: Theory and Methods, 31(8):1423–1440, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menzefricke:2007:CCG

[Men07] Ulrich Menzefricke. Control charts for the generalized variance based on its predictive distribution. Communications in Statistics: Theory and Methods, 36(5):1031–1038, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Minh:2003:UBT

[MF03] Do Le (Paul) Minh and Nicholas R. Farnum. Using bilinear transformations to induce probability distributions. Communications in Statistics: Theory and Methods, 32(1):1–9, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mateu-Figueras:2007:SND

[MFPG07] Glòria Mateu-Figueras and Vera Pawlowsky-Glahn. The skew-normal distribution on the simplex. Communications in Statistics: Theory and Methods, 36(9):1787–1802, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mateu-Figueras:2007:GFT

[MFPP07] Glòria Mateu-Figueras, Pedro Puig, and Arthur Pewsey. Goodness-of-fit tests for the skew-normal distribution when the parameters are estimated from the data. Communications in Statistics: Theory and Methods, 36(9):1735–1755, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mishra:2009:ESR

[MG09] Kaushal K. Mishra and Sujit K. Ghosh. Estimation of SCRAM rate trends in nuclear power plants using hierarchical Bayes models. Communications in Statistics: Theory and Methods, 38(16–17):2856–2871, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Misra:2008:PSA

[MGD08] Neeraj Misra, Nitin Gupta, and Ishwari Dutt Dhariyal. Preservation of some aging properties and stochastic orders by weighted distributions. Communications in Statistics: Theory and Methods, 37(5):627–644, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mohammed:2008:BER

[MI08] Noura Ait Mohammed and Mohamed Ibazizen. Bayesian estimation of the regression model with autocorrelated errors under contamination. Communications in Statistics: Theory and Methods, 37(15):2382–2389, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Micheas:2006:UAP

[Mic06] Athanasios C. Micheas. A unified approach to prior and loss robustness. Communications in Statistics: Theory and Methods, 35(2):309–323, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miller:2000:ALS

[Mil00] Alan J. Miller. Another look at subset selection using linear least squares. Communications in Statistics: Theory and Methods, 29(9–10):2005–2017, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610920008832592.

Minkova:2002:GCD

[Min02] Leda D. Minkova. A generalization of the classical discrete distributions. Communications in Statistics: Theory and Methods, 31(6):871–888, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Minami:2003:MEI

[Min03] Mihoko Minami, Ph.D. A multivariate extension of inverse Gaussian distribution derived from inverse relationship. Communications in Statistics: Theory and Methods, 32(12):2285–2304, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miao:2009:DIE

[MJS09] Yu Miao, Yiwen Jiang, and Si Shen. Deviation inequalities for the estimator of linear parameter in stochastic processes. Communications in Statistics: Theory and Methods, 38(6): 888–901, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mentz:2003:SPM

[MK03] Graciela B. Mentz and Anant M. Kshirsagar. Sum of profiles model with exchangeably distributed errors. Communications in Statistics: Theory and Methods, 32(8):1591–1605, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mentz:2005:CUG

[MK05] Graciela B. Mentz and Anant M. Kshirsagar. Classification using growth curves. Communications in Statistics: Theory and Methods, 33(10):2487–2502, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mudholkar:2001:RTS

[MKS01] Govind S. Mudholkar, James Kost, and Perla Subbaiah. Robust tests for the significance of orthant restricted mean vector. Communications in Statistics: Theory and Methods, 30 (8–9):1789–1810, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Misra:2005:SSP

[MKvdMT05] Neeraj Misra, Somesh Kumar, Edward C. van der Meulen, and Yogesh Mani Tripathi. A subset selection procedure for selecting the exponential population having the longest mean lifetime when the guarantee times are the same. Communications in Statistics: Theory and Methods, 34(7):1555–1569, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Montenegro:2009:LIA

[MLB09] Lourdes C. Montenegro, Victor H. Lachos, and Heleno Bolfarine. Local influence analysis for skew-normal linear mixed models. Communications in Statistics: Theory and Methods, 38(4):484–496, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

MacGibbon:2002:LTA

[MLY02] Brenda MacGibbon, Jiang Lu, and Hassan Younes. Limit theorems for asymptotically minimax estimation of a distribution with increasing failure rate under a random mixed censorship/truncation model. Communications in Statistics: Theory and Methods, 31(8):1309–1333, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Martin:2000:BIT

[MM00] Gael M. Martin and Vance L. Martin. Bayesian inference in the triangular cointegration model using a Jeffreys prior. Communications in Statistics: Theory and Methods, 29(8): 1759–1785, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McKenzie:2002:CTA

[MM02] C. R. McKenzie and Michael McAleer. Comparing tests of autoregressive versus moving average errors in regression models using Bahadur’s asymptotic relative efficiency. Communications in Statistics: Theory and Methods, 31(8):1349– 1371, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moerbeek:2005:OED

[MM05] Mirjam Moerbeek and Cora J. M. Maas. Optimal experimental designs for multilevel logistic models with two binary predictors. Communications in Statistics: Theory and Methods, 34(5):1151–1167, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Martynov:2006:GFT

[MM06] G. Martynov and M. Mesbah. Goodness of fit test and latent distribution estimation in the mixed Rasch model. Communications in Statistics: Theory and Methods, 35(5):921–935, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menendez:2001:APD

[MMPV01] M. Menéndez, D. Morales, L. Pardo, and I. Vajda. Approximations to powers of ϕ-disparity goodness-of-fit tests. Communications in Statistics: Theory and Methods, 30(1): 105–134, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marin:2008:SMM

[MN08] J. M. Marín and C. Nieto. Spatial matching of multiple configurations of points with a bioinformatics application. Communications in Statistics: Theory and Methods, 37(12):1977– 1995, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mokhlis:2005:RSS

[Mok05] Nahed A. Mokhlis. Reliability of a stress-strength model with Burr type III distributions. Communications in Statistics: Theory and Methods, 34(7):1643–1657, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Monlezun:2000:AFP

[Mon00] Charles J. Monlezun. Addendum to fitting polynomials to main effects in multi factor experiments: A note concerning polynomial coefficients when all factors are quantitative. Communications in Statistics: Theory and Methods, 29 (12):2915–2919, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Monlezun:2001:ODI

[Mon01] Charles J. Monlezun. Orthogonally decomposing the interaction sum of squares in two factor experiments when both factors are quantitative: Unbalanced data and unequal spacing. Communications in Statistics: Theory and Methods, 30 (7):1411–1433, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Morais:2008:ECM

[MOPS08] Manuel Cabral Morais, Yarema Okhrin, António Pacheco, and Wolfgang Schmid. EWMA charts for multivariate output: Some stochastic ordering results. Communications in Statistics: Theory and Methods, 37(16):2653–2663, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Munoz:2004:JSM

[MOS04] Laura Muñoz, Pilar Olave, and Manuel Salvador. Joint selection of the model and the information set in heteroskedastic dynamic models. Communications in Statistics: Theory and Methods, 33(11):2659–2682, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mandal:2008:OMD

[MP08] N. K. Mandal and Manisha Pal. Optimum mixture design using deficiency criterion. Communications in Statistics: Theory and Methods, 37(10):1565–1575, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Macci:2009:CED

[MP09] Claudio Macci and Lea Petrella. Censored exponential data: Large deviations for MLEs and posterior distributions. Communications in Statistics: Theory and Methods, 38(15):2435– 2452, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mandal:2008:ISP

[MPG08] B. N. Mandal, Rajender Parsad, and V. K. Gupta. IPPS sampling plans excluding adjacent units. Communications in Statistics: Theory and Methods, 37(16):2532–2550, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menendez:2001:RSA

[MPP01] M. L. Menéndez, J. A. Pardo, and L. Pardo. Rao’s statistic for the analysis of uniform association in cross-classifications. Communications in Statistics: Theory and Methods, 30(12): 2655–2681, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Makri:2007:PIP

[MPP07] Frosso S. Makri, Andreas N. Philippou, and Zaharias M. Psillakis. Pólya, inverse Pólya, and circular Pólya distributions of order k for l-overlapping success runs. Communications in Statistics: Theory and Methods, 36(4):657–668, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menendez:2008:PDS

[MPP08] Maria Luisa Menéndez, Julio Angel Pardo, and Leandro Pardo. Phi-divergence statistics for testing linear hypotheses in logistic regression models. Communications in Statistics: Theory and Methods, 37(4):494–507, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mansouri:2004:ART

[MPS04] H. Mansouri, R. L. Paige, and J. G. Surles. Aligned rank transform techniques for analysis of variance and multiple comparisons. Communications in Statistics: Theory and Methods, 33(9):2217–2232, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Monteiro:2008:OAS

[MPS08] Magda Monteiro, Isabel Pereira, and Manuel G. Scotto. Optimal alarm systems for count processes. Communications in Statistics: Theory and Methods, 37(19):3054–3076, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McClean:2004:DTR

[MPT04] Sally McClean, Aleka A. Papadopoulou, and George Tsaklidis. Discrete time reward models for homogeneous semi-Markov systems. Communications in Statistics: Theory and Methods, 33(3):623–638, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moustakides:2009:NCC

[MPT09] George V. Moustakides, Aleksey S. Polunchenko, and Alexander G. Tartakovsky. Numerical comparison of CUSUM and Shiryaev–Roberts procedures for detecting changes in distributions. Communications in Statistics: Theory and Methods, 38(16–17):3225–3239, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Morales:2006:EEC

[MPV06] Domingo Morales, Leandro Pardo, and Igor Vajda. On efficient estimation in continuous models based on finitely quantized observations. Communications in Statistics: Theory and Methods, 35(9):1629–1653, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Misquitta:2005:SRN

[MR05] P. Misquitta and Frits H. Ruymgaart. Some results on nonparametric calibration. Communications in Statistics: Theory and Methods, 34(7):1605–1616, 2005. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Madi:2009:BIG

[MR09] Mohamed T. Madi and Mohammad Z. Raqab. Bayesian inference for the generalized exponential distribution based on progressively censored data. Communications in Statistics: Theory and Methods, 38(12):2016–2029, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moala:2009:NPD

[MRT09] Fernando Antonio Moala, Josemar Rodrigues, and Vera Lucia D. Tomazella. A note on the prior distributions of Weibull parameters for the reliability function. Communications in Statistics: Theory and Methods, 38(7):1041–1054, 2009. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mudholkar:2000:RAH

[MS00] Govind S. Mudholkar and Deo Kumar Srivastava. Robust analogs of Hotelling’s two-sample T^{2}. Communications in Statistics: Theory and Methods, 29(12):2717–2749, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Murat:2003:MCD

[MS03] Małgorzata Murat and Dominik Szynal. Moments of certain deformed probability distributions. Communications in Statistics: Theory and Methods, 32(2):291–313, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moschopoulos:2005:BIS

[MS05] Panagis Moschopoulos and Naijun Sha. Bayesian inference of scale parameters in exponential family using conditionally specified priors. Communications in Statistics: Theory and Methods, 34(2):303–318, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marozzi:2006:MBA

[MS06] Marco Marozzi and Luigi Salmaso. Multivariate bi-aspect testing for the two-sample location problem. Communications in Statistics: Theory and Methods, 35(3):477–488, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Matsuura:2007:OBS

[MS07a] Shun Matsuura and Nobuo Shinozaki. Optimal binning strategies under squared error loss in selective assembly with measurement error. Communications in Statistics: Theory and Methods, 36(16):2863–2876, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meeden:2007:SBM

[MS07b] Glen Meeden and Dawn Sargent. Some Bayesian methods for two auditing problems. Communications in Statistics: Theory and Methods, 36(15):2727–2740, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Munoli:2007:ERC

[MS07c] S. B. Munoli and M. D. Suranagi. Estimation of reliability for ( k + 1) component lightly loaded standby system with single repair. Communications in Statistics: Theory and Methods, 36(1):193–202, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mahadik:2009:SVS

[MS09] Shashibhushan B. Mahadik and Digambar T. Shirke. A special variable sample size and sampling interval chart. Communications in Statistics: Theory and Methods, 38(8):1284– 1299, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mandal:2000:ACO

[MT00] Saumendranath Mandal and Ben Torsney. Algorithms for the construction of optimizing distributions. Communications in Statistics: Theory and Methods, 29(5–6):1219–1231, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mudholkar:2001:NDE

[MT01] Govind S. Mudholkar and Lili Tian. On the null distributions of the entropy tests for the Gaussian and inverse Gaussian models. Communications in Statistics: Theory and Methods, 30(8–9):1507–1520, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mudholkar:2002:NDE

[MT02] Govind S. Mudholkar and Lili Tian. On the null distributions of the entropy tests for the Gaussian and inverse Gaussian models*. Communications in Statistics: Theory and Methods, 31(2):311–324, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moser:2005:NIM

[MT05] Barry K. Moser and Joshua M. Tebbs. A new interim monitoring statistic for group sequential clinical trials. Communications in Statistics: Theory and Methods, 33(1):153–164, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Matsui:2006:SIN

[MT06] Muneya Matsui and Akimichi Takemura. Some improvements in numerical evaluation of symmetric stable density and its derivatives. Communications in Statistics: Theory and Methods, 35(1):149–172, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Muddapur:2003:DDR

[Mud03] M. V. Muddapur. On directional dependence in a regression line. Communications in Statistics: Theory and Methods, 32 (10):2053–2057, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Muddapur:2008:RPN

[Mud08] M. V. Muddapur. A remark on the paper “A Note on Directional Dependence in Regression Setting”, by Engin A. Sungur. Communications in Statistics: Theory and Methods, 37(3):386–387, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [Sun05a].

Mukhopadhyay:2002:EFP

[Muk02] Parimal Mukhopadhyay. On estimating functions in the presence of a nuisance parameter. Communications in Statistics: Theory and Methods, 31(1):31–36, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukherjee:2003:SDI

[Muk03] Bhramar Mukherjee. On sampling designs for integral estimation of a random process. Communications in Statistics: Theory and Methods, 32(8):1647–1663, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukhopadhyay:2005:SOP

[Muk05a] N. Mukhopadhyay. On some one-parameter families where maximal invariants do not distinguish hypothesized models. Communications in Statistics: Theory and Methods, 34(1): 23–36, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukhopadhyay:2005:NAD

[Muk05b] Nitis Mukhopadhyay. A new approach to determine the pilot sample size in two-stage sampling. Communications in Statistics: Theory and Methods, 34(6):1275–1295, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukhopadhyay:2006:SCH

[Muk06] Nitis Mukhopadhyay. Some comments on hassairi et al.’s “Implicit Distributions and Estimation”. Communications in Statistics: Theory and Methods, 35(2):293–297, 2006. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [HMK05].

Mukhopadhyay:2007:OEF

[Muk07] Parimal Mukhopadhyay. On optimal estimating functions in the presence of nuisance parameters. Communications in Statistics: Theory and Methods, 36(10):1867–1876, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Muralidharan:2001:TEA

[Mur01] K. Muralidharan. Tests for exponentiality against gamma alternatives using normalized waiting times. Communications in Statistics: Theory and Methods, 30(3):397–405, 2001. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meeden:2000:SHM

[MV00] Glen Meeden and Stephen Vardeman. A simple hidden Markov model for Bayesian modeling with time dependent data. Communications in Statistics: Theory and Methods, 29(8):1801–1826, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moerbeek:2001:OED

[MVB01] Mirjam Moerbeek, Gerard J. P. Van Breukelen, and Martijn P. F. Berger. Optimal experimental designs for multilevel models with covariates. Communications in Statistics: Theory and Methods, 30(12):2683–2697, 2001. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Misra:2001:NTU

[MvdM01a] Neeraj Misra and Edward C. van der Meulen. A new test of uniformity based on overlapping sample spacings. Communications in Statistics: Theory and Methods, 30(7):1435–1470, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Misra:2001:EFS

[MvdM01b] Neeraj Misra and Edward C. van der Meulen. On estimation following selection from nonregular distributions. Communications in Statistics: Theory and Methods, 30(12):2543– 2561, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Manatunga:2001:AFA

[MW01] Amita K. Manatunga and John M. Williamson. Assessing familial aggregation with an ordinal response. Communications in Statistics: Theory and Methods, 30(4):627–641, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mudholkar:2009:SPF

[MW09] Govind S. Mudholkar and Hongyue Wang. R-symmetry and the power functions of the tests for inverse Gaussian means. Communications in Statistics: Theory and Methods, 38(13):2178–2186, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mudholkar:2003:RPP

[MWM03] Govind S. Mudholkar, Gregory E. Wilding, and William L. Mietlowski. Robustness properties of the Pitman–Morgan test. Communications in Statistics: Theory and Methods, 32(9):1801–1816, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mu:2008:ISR

[MXX08] Weiyan Mu, Xingzhong Xu, and Shifeng Xiong. Inference on system reliability for independent series components. Communications in Statistics: Theory and Methods, 38(3):409– 418, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ma:2008:IRB

[MXZ08] Jianjun Ma, Xingzhong Xu, and Junlong Zhao. Inverse regression in binary response LDV model. Communications in Statistics: Theory and Methods, 37(2):233–246, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mehmet:2007:COB

[MY07] Yilmaz Mehmet and Tuncer Yalçin. Convex ordering for bivariate distributions. Communications in Statistics: Theory and Methods, 36(3):465–472, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mynbaev:2006:APO

[Myn06] Kairat T. Mynbaev. Asymptotic properties of OLS estimates in autoregressions with bounded or slowly growing deterministic trends. Communications in Statistics: Theory and Methods, 35(3):499–520, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miao:2007:CLT

[MYS07] Yu Miao, Guangyu Yang, and Luming Shen. The central limit theorem for LS estimator in simple linear EV regression models. Communications in Statistics: Theory and Methods, 36(12):2263–2272, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nasser:2006:EIF

[NA06] Mohammed Nasser and Mesbahul Alam. Estimators of influence function. Communications in Statistics: Theory and Methods, 35(1):21–32, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Niaki:2007:SRA

[NA07] Seyed Taghi Akhavan Niaki and Babak Abbasi. Skewness reduction approach in multi-attribute process monitoring. Communications in Statistics: Theory and Methods, 36(12): 2313–2325, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Niaki:2009:MMA

[NA09] Seyed Taghi Akhavan Niaki and Babak Abbasi. Monitoring multi-attribute processes based on NORTA inverse transformed vectors. Communications in Statistics: Theory and Methods, 38(7):964–979, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nadarajah:2005:SPR

[Nad05] Saralees Nadarajah. Sums, products, and ratios of noncentral beta variables. Communications in Statistics: Theory and Methods, 34(1):89–100, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See comments [dSFM07] and reply [Nad09].

Nadarajah:2009:RCS

[Nad09] Saralees Nadarajah. Reply to “Comments on ‘Sums, Products, and Ratios of Non-Central Beta Variables’ by Saralees Nadarajah” by Antonio Mauricio F. L. Miranda de Sa, Leonardo B. Felix, and Eduardo M. A. M. Mendes. Communications in Statistics: Theory and Methods, 38(14):2428– 2433, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [Nad05, dSFM07].

Nagai:2002:AMD

[Nag02] Yoshimitsu Nagai. Analysis of masked data in a dependent competing risks model under unknown environment. Communications in Statistics: Theory and Methods, 31(9):1579– 1593, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Naito:2002:SRM

[Nai02] Kanta Naito. Semiparametric regression with multiplicative adjustment. Communications in Statistics: Theory and Methods, 31(12):2289–2309, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nandram:2009:BIC

[Nan09] Balgobin Nandram. Bayesian inference of the cell probabilities of a two-way categorical table under non ignorability. Communications in Statistics: Theory and Methods, 38(16– 17):3015–3030, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nassar:2006:CMG

[Nas06] Manal M. Nassar. On the characterization of a mixture of generalized power function distributions by conditional expectation of order statistics. Communications in Statistics: Theory and Methods, 35(11):1957–1962, 2006. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Noorossana:2008:MAL

[NAS08] R. Noorossana, A. Amiri, and P. Soleimani. On the monitoring of autocorrelated linear profiles. Communications in Statistics: Theory and Methods, 37(3):425–442, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Navarro:2008:CPC

[Nav08a] Jorge Navarro. Characterizations by power contractions of order statistics. Communications in Statistics: Theory and Methods, 37(7):987–997, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Navidi:2008:PRC

[Nav08b] William Navidi. Poisson regression and the case-crossover design: Similarities and differences. Communications in Statistics: Theory and Methods, 37(2):213–220, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nanda:2006:URM

[NBA06] Asok K. Nanda, Subarna Bhattacharjee, and S. S. Alam. On upshifted reversed mean residual life order. Communications in Statistics: Theory and Methods, 35(8):1513–1523, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nandi:2002:SRD

[NC02] S. B. Nandi and S. Choudhury. Series representation of doubly noncentral t distribution by use of the Mellin integral transform. Communications in Statistics: Theory and Methods, 31(12):2139–2149, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nandi:2005:SRN

[NC05] S. B. Nandi and S. Choudhury. Series representation of non null distribution of the square of sample multiple correlation coefficient by use of the Mellin integral transform. Communications in Statistics: Theory and Methods, 34(3):679–685, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nassar:2003:EWD

[NE03] Manal M. Nassar and Fathy H. Eissa. On the exponentiated Weibull distribution. Communications in Statistics: Theory and Methods, 32(7):1317–1336, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nassar:2005:BEE

[NE05] M. M. Nassar and Fathy H. Eissa. Bayesian estimation for the exponentiated Weibull model. Communications in Statistics: Theory and Methods, 33(10):2343–2362, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Neuhauser:2000:ETS

[Neu00] Markus Neuhäuser. An exact two-sample test based on the Baumgartner–Weiss–Schindler statistic and a modification of Lepage’s test. Communications in Statistics: Theory and Methods, 29(1):67–78, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ng:2000:NPI

[Ng00] Vee Ming Ng. A note on predictive inference for multivariate elliptically contoured distributions. Communications in Statistics: Theory and Methods, 29(3):477–483, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nadarajah:2004:CBD

[NG04] Saralees Nadarajah and Arjun K. Gupta. Characterizations of the beta distribution. Communications in Statistics: Theory and Methods, 33(12):2941–2957, 2004. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nadarajah:2005:MEW

[NG05] Saralees Nadarajah and Arjun K. Gupta. On the moments of the exponentiated Weibull distribution. Communications in Statistics: Theory and Methods, 34(2):253–256, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nakatsuma:2000:RTU

[NGUT00] Teruo Nakatsuma, Elena Gouskova, Jinichi Uemura, and Hiroki Tsurumi. Ratio tests of a unit root. Communications in Statistics: Theory and Methods, 29(11):2547–2571, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ning:2009:MBT

[NGYZ09] Wei Ning, Arjun K. Gupta, Chang Yu, and Sanguo Zhang. A moment-based test for homogeneity in finite mixture models. Communications in Statistics: Theory and Methods, 38 (9):1371–1382, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nieto:2005:MBT

[Nie05] Fabio H. Nieto. Modeling bivariate threshold autoregressive processes in the presence of missing data. Communications in Statistics: Theory and Methods, 34(4):905–930, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nelson:2003:TPM

[NK03] P. I. Nelson and K. E. Kemp. Testing for the presence of a maverick judge. Communications in Statistics: Theory and Methods, 32(4):807–826, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagatsuka:2005:PES

[NK05] Hideki Nagatsuka and Toshinari Kamakura. Parameter estimation of the shape parameter of the Castillo–Hadi model. Communications in Statistics: Theory and Methods, 33(1): 15–27, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nishimura:2008:IGS

[NK08] Tomoaki Nishimura and Fumiyasu Komaki. The information geometric structure of generalized empirical likelihood estimators. Communications in Statistics: Theory and Methods, 37(12):1867–1879, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Navarro:2007:OPS

[NL07] Jorge Navarro and Chin-Diew Lai. Ordering properties of systems with two dependent components. Communications in Statistics: Theory and Methods, 36(3):645–655, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nadarajah:2003:PMM

[NM03] Saralees Nadarajah and Kosto Mitov. Product moments of multivariate random vectors. Communications in Statistics: Theory and Methods, 32(1):47–60, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Njuho:2005:ALM

[NM05] Peter M. Njuho and George A. Milliken. Analysis of linear models with one factor having both fixed and random levels. Communications in Statistics: Theory and Methods, 34(9– 10):1979–1989, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Noda:2007:ATP

[NM07] Kazuo Noda and Etsuo Miyaoka. Admissibility of test procedures for linear hypotheses in general linear model. Communications in Statistics: Theory and Methods, 36(1):117–128, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Njuho:2009:ALM

[NM09] Peter M. Njuho and George A. Milliken. Analysis of linear models with two factors having both fixed and random levels. Communications in Statistics: Theory and Methods, 38 (14):2348–2365, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nanda:2005:SPO

[NMPS05] Asok K. Nanda, Neeraj Misra, Prasanta Paul, and Harshinder Singh. Some properties of order statistics when the sample size is random. Communications in Statistics: Theory and Methods, 34(11):2105–2113, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nematollahi:2008:TSC

[NMS08] N. Nematollahi, M. Salehi M., and R. Aliakbari Saba. Twostage cluster sampling with ranked set sampling in the secondary sampling frame. Communications in Statistics: Theory and Methods, 37(15):2404–2415, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nematollahi:2009:ESP

[NMS09] N. Nematollahi and F. Motamed-Shariati. Estimation of the scale parameter of the selected gamma population under the entropy loss function. Communications in Statistics: Theory and Methods, 38(2):208–221, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nezhad:2009:BIS

[NN09] Mohammad Saber Fallah Nezhad and Seyed Taghi Akhavan Niaki. A Bayesian inference and stochastic dynamic programming approach to determine the best binomial distribution. Communications in Statistics: Theory and Methods, 38(14):2379–2397, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Noda:2001:UIT

[NO01] Kazuo Noda and Hideo Ono. On UMP invariant F -test procedures in a general linear model. Communications in Statistics: Theory and Methods, 30(10):2099–2115, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Norman:2000:ARA

[Nor00] K. Draper Norman. Applied regression analysis bibliography update. Communications in Statistics: Theory and Methods, 29(9–10):2313–2341, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/03610920008832608.

Norell:2003:AEU

[Nor03] Lennart Norell. ANOVA estimators under imbalance in the one-way random model. Communications in Statistics: Theory and Methods, 32(3):601–623, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ndlovu:2001:CUP

[NP01] P. Ndlovu and J. Preater. Calibration using a piecewise simple linear regression model. Communications in Statistics: Theory and Methods, 30(2):229–242, 2001. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ng:2008:CIL

[NP08] K. H. Ng and A. H. Pooi. Calibration intervals in linear regression models. Communications in Statistics: Theory and Methods, 37(11):1688–1696, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nelsen:2001:BBD

[NQMRLÚF01] Roger B. Nelsen, José Juan Quesada-Molina, Joseé Antonio Rodriíguez-Lallena, and Manuel Úbeda-Flores. Bounds on bivariate distribution functions with given margins and measures of association. Communications in Statistics: Theory and Methods, 30(6):1155–1162, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Navarro:2004:CRB

[NR04] Jorge Navarro and Jose M. Ruiz. Characterizations from relationships between failure rate functions and conditional moments. Communications in Statistics: Theory and Methods, 33(12):3159–3171, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagaraja:2009:P

[NR09] H. N. Nagaraja and Mohammad Z. Raqab. Preface. Communications in Statistics: Theory and Methods, 38(12):1949, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Navarro:2007:PCS

[NRS07a] J. Navarro, Jose M. Ruiz, and Carlos J. Sandoval. Properties of coherent systems with dependent components. Communications in Statistics: Theory and Methods, 36(1):175–191, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Navarro:2007:OSO

[NRS07b] Jorge Navarro, Tomasz Rychlik, and Moshe Shaked. Are the order statistics ordered? A survey of recent results. Communications in Statistics: Theory and Methods, 36(7):1273– 1290, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nanthakumar:2004:EPS

[NS04] A. Nanthakumar and K. Selvavel. Estimation of proportion of success from a stratified population: A comparative study. Communications in Statistics: Theory and Methods, 33(9):2245–2257, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nair:2006:CCD

[NS06] N. Unnikrishnan Nair and K. K. Sudheesh. Characterization of continuous distributions by variance bound and its implications to reliability modeling and catastrophe theory. Communications in Statistics: Theory and Methods, 35(7):1189– 1199, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nanda:2008:POA

[NS08] Asok K. Nanda and Moshe Shaked. Partial ordering and aging properties of order statistics when the sample size is random: A brief review. Communications in Statistics: Theory and Methods, 37(11):1710–1720, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nair:2009:QBR

[NS09] N. Unnikrishnan Nair and P. G. Sankaran. Quantile-based reliability analysis. Communications in Statistics: Theory and Methods, 38(2):222–232, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nanda:2003:RPR

[NSMP03] Asok K. Nanda, Harshinder Singh, Neeraj Misra, and Prasanta Paul. Reliability properties of reversed residual lifetime. Communications in Statistics: Theory and Methods, 32(10):2031–2042, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nelsen:2005:CBS

[NÚF05] Roger B. Nelsen and Manuel Úbeda-Flores. A comparison of bounds on sets of joint distribution functions derived from various measures of association. Communications in Statistics: Theory and Methods, 33(10):2299–2305, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ngatchou-Wandji:2009:TSE

[NW09] Joseph Ngatchou-Wandji. Testing symmetry of the error distribution in nonlinear heteroscedastic models. Communications in Statistics: Theory and Methods, 38(9):1465–1485, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Noorossana:2002:EDM

[NWA02] R. Noorossana, William H. Woodall, and S. Amiriparian. On the economic design of multivariate control charts. Communications in Statistics: Theory and Methods, 31(9):1665– 1673, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Noda:2003:IAE

[NWTI03] Kazuo Noda, Jinglong Wang, Rinya Takahashi, and Masashi Itoh. Improvement to AIC as estimator of Kullback–Leibler information for linear model selection. Communications in Statistics: Theory and Methods, 32(11):2207–2225, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oruc:2005:GCT

[OB05] Özlem Ege Oruç and Ismihan Bairamov. On the general class of two-sided power distribution. Communications in Statistics: Theory and Methods, 34(5):1009–1017, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ozdemir:2008:NFI

[OG08] Yaprak Arzu Ozdemir and Fikri Gokpinar. A new formula for inclusion probabilities in median-ranked set sampling. Communications in Statistics: Theory and Methods, 37(13):2022–2033, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ogasawara:2008:AER

[Oga08] Haruhiko Ogasawara. Asymptotic expansion in reduced rank regression under normality and nonnormality. Communications in Statistics: Theory and Methods, 37(7):1051–1070, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

OReilly:2006:CDG

[OGM06] Federico O’Reilly and Leticia Gracia-Medrano. On the conditional distribution of goodness-of-fit tests. Communications in Statistics: Theory and Methods, 35(3):541–549, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ohyauchi:2002:CBR

[Ohy02] Nao Ohyauchi. Comparison of the Bayes risks of estimators for a family of truncated normal distributions. Communications in Statistics: Theory and Methods, 31(5):699–718, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Omori:2006:IRE

[OJ06] Yasuhiro Omori and Richard A. Johnson. The influence of random effects on univariate and bivariate discrete proportional hazards models. Communications in Statistics: Theory and Methods, 35(9):1757–1764, 2006. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Omori:2008:ESB

[OJ08] Yasuhiro Omori and Richard A. Johnson. Efficient semiparametric Bayesian estimation of multivariate discrete proportional hazards model with random effects. Communications in Statistics: Theory and Methods, 38(1):29–41, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ozkale:2007:CUR

[ÖK07a] M. Revan Özkale and Selahattin Kaçiranlar. Comparisons of the unbiased ridge estimation to the other estimations. Communications in Statistics: Theory and Methods, 36(4): 707–723, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ozkale:2007:POC

[ÖK07b] M. Revan Özkale and Selahattin Kaçiranlar. A predictionoriented criterion for choosing the biasing parameter in Liu estimation. Communications in Statistics: Theory and Methods, 36(10):1889–1903, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ozkale:2007:RUT

[ÖK07c] M. Revan Özkale and Selahattin Kaçiranlar. The restricted and unrestricted two-parameter estimators. Communications in Statistics: Theory and Methods, 36(15):2707–2725, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ochiai:2003:ATM

[ON03] Toshimitsu Ochiai and Kanta Naito. Asymptotic theory for the multiscale wavelet density derivative estimator. Communications in Statistics: Theory and Methods, 32(10):1925– 1950, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

OReilly:2007:FIT

[OR07] Federico O’Reilly and Raúl Rueda. Fiducial inferences for the truncated exponential distribution. Communications in Statistics: Theory and Methods, 36(12):2207–2212, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oral:2006:BRS

[Ora06] E. Oral. Binary regression with stochastic covariates. Communications in Statistics: Theory and Methods, 35(8):1429– 1447, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oono:2005:EEV

[OS05] Youhei Oono and Nobuo Shinozaki. Estimation of error variance in the analysis of experiments using two-level orthogonal arrays. Communications in Statistics: Theory and Methods, 33(1):75–98, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Odumade:2009:EUT

[OS09a] Oluseun Odumade and Sarjinder Singh. Efficient use of two decks of cards in randomized response sampling. Communications in Statistics: Theory and Methods, 38(4):439–446, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ong:2009:DDA

[OS09b] S. H. Ong and K. Shimizu. A discrete distribution arising as a solution of a linear difference equation: Extension of the non central negative binomial distribution. Communications in Statistics: Theory and Methods, 38(6):927–938, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ould-Said:2005:SAQ

[OSS05] Elias Ould-Saïd and Ourida Sadki. Strong approximation of quantile function for strong mixing and censored processes. Communications in Statistics: Theory and Methods, 34(7):1449–1459, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ould-Said:2009:SCR

[OST09] Elias Ould-Saïd and Abdelkader Tatachak. Strong consistency rate for the kernel mode estimator under strong mixing hypothesis and left truncation. Communications in Statistics: Theory and Methods, 38(8):1154–1169, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ozkale:2009:CRE

[Özk09] M. Revan Özkale. Comment on ridge estimation to the restricted linear model. Communications in Statistics: Theory and Methods, 38(7):1094–1097, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Orelien:2002:APM

[OZMC02] Jean G. Orelien, Jun Zhai, Richard Morris, and Rich Cohn. An approach to performing multiple comparisons with a control in Gee models. Communications in Statistics: Theory and Methods, 31(1):87–105, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ozturk:2001:NTS

[Özt01] Ömer Öztürk. A nonparametric test of symmetry versus asymmetry for ranked-set samples. Communications in Statistics: Theory and Methods, 30(10):2117–2133, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ozturk:2002:SSM

[Ozt02] Omer Ozturk. Sub-sample Mann–Whitney–Wilcoxon test. Communications in Statistics: Theory and Methods, 31(12): 2361–2377, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pakes:2004:CPL

[Pak04] Anthony G. Pakes. Characterization of probability laws via length-biasing and stationary-excess operations. Communications in Statistics: Theory and Methods, 33(12):2975– 2992, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pakyari:2009:NAB

[Pak09] Reza Pakyari. A note on asymptotic behavior of the nonparametric density estimators in multivariate mixtures. Communications in Statistics: Theory and Methods, 38(8): 1219–1223, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pan:2002:IOI

[Pan02] Jian-Xin Pan. Influential observation identification in the growth curve model with Rao’s simple covariance structure. Communications in Statistics: Theory and Methods, 31(5): 813–831, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Panchapakesan:2005:RSS

[Pan05] S. Panchapakesan. Restricted subset selection procedures for normal means: A brief review with a fresh look at the classical formulations of Bechhofer and Gupta. Communications in Statistics: Theory and Methods, 34(6):1265–1273, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Papadopoulou:2004:ERN

[Pap04] Aleka A. Papadopoulou. Economic rewards in nonhomogeneous semi-Markov systems. Communications in Statistics: Theory and Methods, 33(3):681–696, 2004. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2001:SDR

[Par01] Taesung Park. Segmented dose-response models for repeated measures data. Communications in Statistics: Theory and Methods, 30(10):2045–2056, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parzen:2009:QCQ

[Par09] Emanuel Parzen. Quantiles, conditional quantiles, confidence quantiles for p, Logodds(p). Communications in Statistics: Theory and Methods, 38(16–17):3048–3058, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patrangenaru:2001:NLS

[Pat01] Vic Patrangenaru. New large sample and bootstrap methods on shape spaces in high level analysis of natural images. Communications in Statistics: Theory and Methods, 30(8– 9):1675–1693, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patilea:2007:SRM

[Pat07] Valentin Patilea. Semiparametric regression models with applications to scoring: A review. Communications in Statistics: Theory and Methods, 36(14):2641–2653, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pavlenko:2003:FIH

[Pav03] Tatjana Pavlenko. Feature informativeness in highdimensional discriminant analysis. Communications in Statistics: Theory and Methods, 32(2):459–474, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Permutt:2000:NLR

[PB00] Thomas Permutt and Vance W. Berger. A new look at rank tests in ordered-2 × k contingency tables. Communications in Statistics: Theory and Methods, 29(5–6):989–1003, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2004:CIT

[PB04] Dong Joon Park and Richard K. Burdick. Confidence intervals on total variance in a regression model with an unbalanced onefold nested error structure. Communications in Statistics: Theory and Methods, 33(11):2735–2743, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Putt:2002:EAV

[PC02] Mary E. Putt and Vernon M. Chinchilli. Estimating the asymptotic variance of generalized L-statistics. Communications in Statistics: Theory and Methods, 31(5):733–751, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pan:2008:SBM

[PC08] Jianmin Pan and Jiahua Chen. U-statistic based modified information criterion for change point problems. Communications in Statistics: Theory and Methods, 37(17):2687– 2712, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pelletier:2007:NBR

[PD07] Ronald P. Pelletier and George T. Diderrich. A note on Breiman’s random forest data mining technique and conventional Cox modeling of survival statistics: The case of the phantom “induct” covariate in the Ohio State University kidney transplant database. Communications in Statistics: Theory and Methods, 36(10):1953–1964, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pettersson:2009:URE

[Pet09] Kjell Pettersson. On unimodal regression in the exponential family. Communications in Statistics: Theory and Methods, 38(9):1526–1538, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pewsey:2000:WSN

[Pew00] Arthur Pewsey. The wrapped skew-normal distribution on the circle. Communications in Statistics: Theory and Methods, 29(11):2459–2472, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pewsey:2002:LSI

[Pew02] Arthur Pewsey. Large-sample inference for the general halfnormal distribution. Communications in Statistics: Theory and Methods, 31(7):1045–1054, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pewsey:2004:ILB

[Pew04] Arthur Pewsey. Improved likelihood based inference for the general half-normal distribution. Communications in Statistics: Theory and Methods, 33(2):197–204, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Papaioannou:2005:TFF

[PF05] Takis Papaioannou and Kosmas Ferentinos. On two forms of Fisher’s measure of information. Communications in Statistics: Theory and Methods, 34(7):1461–1470, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pham-Gia:2000:DRI

[PG00] T. Pham-Gia. Distributions of the ratios of independent beta variables and applications. Communications in Statistics: Theory and Methods, 29(12):2693–2715, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peraza:2003:IVS

[PGF03] Felipe Peraza and Graciela González-Farías. Indicator variables in spatial prediction when Σ is unknown. Communications in Statistics: Theory and Methods, 32(3):555–571, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pewsey:2007:P

[PGF07] Arthur Pewsey and Graciela González-Farías. Preface. Communications in Statistics: Theory and Methods, 36(9):1657– 1659, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parsad:2003:SIR

[PGP03] Rajender Parsad, V. K. Gupta, and N. S. G. Prasad. Structurally incomplete row-column designs. Communications in Statistics: Theory and Methods, 32(1):239–261, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pichardo:2000:IAM

[PGPG00] J. M. Muñoz Pichardo, J. Muñoz García, J. M. Fernández Ponce, and M. D. Jiménez Garnero. Influence analysis in multivariate linear general models. Communications in Statistics: Theory and Methods, 29(3):529–547, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pham-Gia:2006:DRT

[PGTM06] T. Pham-Gia, N. Turkkan, and E. Marchand. Density of the ratio of two normal random variables and applications. Communications in Statistics: Theory and Methods, 35(9):1569– 1591, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2002:CIM

[PH02] Dong Joon Park and Hyun Mi Hwang. Confidence intervals for the mean response in the simple linear regression model with balanced nested error structure. Communications in Statistics: Theory and Methods, 31(1):107–117, 2002. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2003:RDC

[PH03] Jinho Park and Jinsoo Hwang. Regression depth with censored and truncated data. Communications in Statistics: Theory and Methods, 32(5):997–1008, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2000:UCD

[PJ00] Kwang-Su Park and Chi-Hyuck Jun. Use of contagious distributions in the semiconductor yield models considering cluster effect. Communications in Statistics: Theory and Methods, 29(1):1–17, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandya:2006:BEC

[PJ06] Mayuri Pandya and P. N. Jani. Bayesian estimation of change point in inverse Weibull sequence. Communications in Statistics: Theory and Methods, 35(12):2223–2237, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Provost:2009:MBA

[PJH09] Serge B. Provost, Min Jiang, and Hyung-Tae Ha. Momentbased approximations of probability mass functions with applications involving order statistics. Communications in Statistics: Theory and Methods, 38(12):1969–1981, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paul:2000:MRA

[PK00a] Sofia Paul and Andre I. Khuri. Modified ridge analyses under nonstandard conditions. Communications in Statistics: Theory and Methods, 29(9–10):2181–2200, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610920008832602.

Pensky:2000:LEB

[PK00b] Marianna Pensky and Kalpana Kirtane. Linear empirical Bayes estimation in the case of the Wishart distribution. Communications in Statistics: Theory and Methods, 29(8):1787–1799, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Petropoulos:2002:CIE

[PK02] Constantinos Petropoulos and Stavros Kourouklis. A class of improved estimators for the scale parameter of an exponential distribution with unknown location. Communications in Statistics: Theory and Methods, 31(3):325–335, 2002. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2006:NFI

[PK06] Sangun Park and Chul Eung Kim. A note on the Fisher information in exponential distribution. Communications in Statistics: Theory and Methods, 35(1):13–19, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pearn:2000:ECI

[PL00] W. L. Pearn and G. H. Lin. Estimating capability index c_{pk} for processes with asymmetric tolerances. Communications in Statistics: Theory and Methods, 29(11):2593–2604, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pan:2009:DNP

[PL09] Jeh-Nan Pan and Chun-Yi Lee. Development of a new process incapability index with an application to the evaluation of manufacturing risk. Communications in Statistics: Theory and Methods, 38(8):1133–1153, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peng:2000:MCP

[PLL00] Jianan Peng, Chu-In Charles Lee, and Lin Liu. A multiple comparisons procedure for detecting differences between treatments and a control in two-factor experiments. Communications in Statistics: Theory and Methods, 29(5–6):1197– 1217, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2000:DFA

[pLM00] Jen pei Liu and Mi-Chia Ma. On difference factor in assessment of dissolution similarity. Communications in Statistics: Theory and Methods, 29(5–6):1089–1113, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Prescott:2002:EPW

[PM02] P. Prescott and R. A. Mansson. Efficiency of pair-wise treatment comparisons in incomplete block experiments subject to the loss of a block of observations. Communications in Statistics: Theory and Methods, 31(3):449–462, 2002. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pensky:2000:ELE

[PN00] M. Pensky and P. Ni. Extended linear empirical Bayes estimation. Communications in Statistics: Theory and Methods, 29(3):579–592, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Poursaeed:2008:MPM

[PN08] M. H. Poursaeed and A. R. Nematollahi. On the mean past and the mean residual life under double monitoring. Communications in Statistics: Theory and Methods, 37(7):1119– 1133, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2007:CAU

[PNJ07] Hyeonah Park, Seongryong Na, and Jongwoo Jeon. Compromise allocation in univariate stratified sampling. Communications in Statistics: Theory and Methods, 36(2):265–271, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Puza:2006:TBP

[PO06] Borek D. Puza and Terence J. O’Neill. Two birthday problems. Communications in Statistics: Theory and Methods, 35(3):415–422, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Polansky:2005:NED

[Pol05] Alan M. Polansky. Nonparametric estimation of distribution functions of nonstandard mixtures. Communications in Statistics: Theory and Methods, 34(8):1711–1724, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pons:2007:EDF

[Pon07] Odile Pons. Estimation for the distribution function of oneand two-dimensional censored variables or sojourn times of Markov renewal processes. Communications in Statistics: Theory and Methods, 36(14):2527–2542, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Portela:2008:CDD

[Por08] J. Portela. Clustering discrete data through the multinomial mixture model. Communications in Statistics: Theory and Methods, 37(20):3250–3263, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pourahmadi:2007:SNA

[Pou07] Mohsen Pourahmadi. Skew-normal ARMA models with nonlinear heteroscedastic predictors. Communications in Statistics: Theory and Methods, 36(9):1803–1819, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:2007:MLE

[PP07] Nandita W. Patel and M. N. Patel. Maximum likelihood estimation in compound geometric failure model with changing parameters from Type-I two-stage progressively censored and group censored samples. Communications in Statistics: Theory and Methods, 36(13):2367–2375, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2000:AEB

[PPL00] Hyo-Ii Park, Jung-Sub Park, and Seung-Joo Lee. Asymptotic equivalence between two empirical mean residual life processes. Communications in Statistics: Theory and Methods, 29(11):2523–2533, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peiris:2004:NTS

[PR04] Shelton Peiris and C. R. Rao. A note on testing for serial correlation in large number of small samples using tail probability approximations. Communications in Statistics: Theory and Methods, 33(8):1767–1777, 2004. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Petrov:2008:LDS

[PR08a] Valentin V. Petrov and John Robinson. Large deviations for sums of independent non identically distributed random variables. Communications in Statistics: Theory and Methods, 37(18):2984–2990, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Prieto-Rumeau:2008:SAE

[PR08b] Tomás Prieto-Rumeau. Stochastic algorithms for the estimation of an optimal solution of a LP problem. Convergence and central limit theorem. Communications in Statistics: Theory and Methods, 37(20):3308–3318, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Prescott:2000:PDM

[Pre00] Philip Prescott. Projection designs for mixture experiments in orthogonal blocks. Communications in Statistics: Theory and Methods, 29(9–10):2229–2253, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610920008832605.

Prescott:2004:CDM

[Pre04] Philip Prescott. A class of designs for mixture experiments based on augmented pairs. Communications in Statistics: Theory and Methods, 33(11):2759–2785, 2004. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Prescott:2008:NUD

[Pre08] Philip Prescott. Nearly uniform designs for mixture experiments. Communications in Statistics: Theory and Methods, 37(13):2095–2115, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Perez-Rodriguez:2009:GFT

[PRVHVA09] Paulino Pérez-Rodríguez, Humberto Vaquera-Huerta, and José A. Villaseñor-Alva. A goodness-of-fit test for the Gumbel distribution based on Kullback–Leibler information. Communications in Statistics: Theory and Methods, 38(6): 842–855, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Politis:2000:SSR

[PRW00] Dimitris N. Politis, Joseph P. Romano, and Michael Wolf. Subsampling, symmetrization, and robust interpolation. Communications in Statistics: Theory and Methods, 29(8): 1741–1757, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parmet:2001:QID

[PS01a] Yisrael Parmet and David M. Steinberg. Quality improvement from disassembly-reassembly experiments. Communications in Statistics: Theory and Methods, 30(5):969–985, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pawlas:2001:RRS

[PS01b] Piotr Pawlas and Dominik Szynal. Recurrence relations for single and product moments of generalized order statistics from Pareto, generalized Pareto, and Burr distributions. Communications in Statistics: Theory and Methods, 30(4): 739–746, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Provost:2005:IAF

[PS05] Serge B. Provost and Deepak Sanjel. Inference about the first-order autoregressive coefficient. Communications in Statistics: Theory and Methods, 34(5):1183–1201, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parmet:2007:TDB

[PS07a] Yisrael Parmet and Edna Schechtman. On a test of the difference between two binomial proportions. Communications in Statistics: Theory and Methods, 36(5):887–895, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peng:2007:CBL

[PS07b] Liang Peng and Shan Sun. Comparisons between local linear estimator and kernel smooth estimator for a smooth distribution based on MSE under right censoring. Communications in Statistics: Theory and Methods, 36(2):297–312, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2007:CTS

[PSZ07] Do-Hwan Park, Jianguo Sun, and Xingqiu Zhao. A class of two-sample nonparametric tests for panel count data. Communications in Statistics: Theory and Methods, 36(8):1611– 1625, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Panchapakesan:2009:P

[PTE09] S. Panchapakesan, Sheng-Tsaing Tseng, and Ming-Chung Yang Guest Editors. Preface. Communications in Statistics: Theory and Methods, 38(10):1547, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Panchapakesan:2009:DYH

[PTY09] S. Panchapakesan, Sheng-Tsaing Tseng, and Ming-Chung Yang. Deng-Yuan Huang: Life and works. Communications in Statistics: Theory and Methods, 38(10):1548–1552, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pulcini:2001:PFF

[Pul01] Gianpaolo Pulcini. On the prediction of future failures for a repairable equipment subject to overhauls. Communications in Statistics: Theory and Methods, 30(4):691–706, 2001. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pulcini:2002:AMU

[Pul02] Gianpaolo Pulcini. Accounting for model uncertainty in reliability analysis of mechanical repairable equipment. Communications in Statistics: Theory and Methods, 31(5):849–869, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pulcini:2006:BDR

[Pul06] G. Pulcini. A bivariate distribution for the reliability analysis of failure data in presence of a forewarning or primer event. Communications in Statistics: Theory and Methods, 35(11):2107–2126, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pumo:2007:ARM

[Pum07] B. Pumo. Asymptotic results for the MPL estimators of the contact process of order p. Communications in Statistics: Theory and Methods, 36(14):2493–2508, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pena:2007:DMV

[PvdL07] Daniel Peña and Angelika van der Linde. Dimensionless measures of variability and dependence for multivariate continuous distributions. Communications in Statistics: Theory and Methods, 36(10):1845–1854, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Perakis:2003:PCI

[PX03] Michael Perakis and Evdokia Xekalaki. On a process capability index for asymmetric specifications. Communications in Statistics: Theory and Methods, 32(7):1459–1492, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2008:RRE

[PY08] Mingue Park and Min Yang. Ridge regression estimation for survey samples. Communications in Statistics: Theory and Methods, 37(4):532–543, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qian:2005:BMG

[QAT05] Xufeng Qian, Rie Ashizawa, and Hiroki Tsurumi. Bayesian, MLE, and GMM estimation of a spot rate model. Communications in Statistics: Theory and Methods, 34(12):2221– 2233, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qin:2009:LBU

[QC09] Hong Qin and Kashinath Chatterjee. Lower bounds for the uniformity pattern of asymmetric fractional factorials. Communications in Statistics: Theory and Methods, 38(9):1383– 1392, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Quinn:2000:NLI

[QHB00] Shannon L. Quinn, Thomas J. Harris, and David W. Bacon. Notes on likelihood intervals and profiling. Communications in Statistics: Theory and Methods, 29(1):109–129, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qingguo:2009:ANE

[QL09] Tang Qingguo and Cheng Longsheng. Asymptotic normality of M-estimators for varying coefficient models with longitudinal data. Communications in Statistics: Theory and Methods, 38(9):1422–1440, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qin:2008:KTL

[QN08] Min Qin and Tapan K. Nayak. Kshirsagar-type lower bounds for mean squared error of prediction. Communications in Statistics: Theory and Methods, 37(6):861–872, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Quade:2006:CCR

[QS06] D. Quade and I. A. Salama. Concordance of complete or right-censored rankings based on Spearman’s footrule. Communications in Statistics: Theory and Methods, 35(6):1059– 1069, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qin:2002:ELR

[QT02] Gengsheng Qin and Min Tsao. Empirical likelihood ratio confidence interval for the trimmed mean. Communications in Statistics: Theory and Methods, 31(12):2197–2208, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qu:2000:UAE

[QTR00] Yinsheng Qu, Ming Tan, and Lisa Rybicki. A unified approach to estimating association measures via a joint generalized linear model for paired binary data. Communications in Statistics: Theory and Methods, 29(1):143–156, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Quessy:2009:TMI

[Que09] Jean-François Quessy. Tests of multivariate independence for ordinal data. Communications in Statistics: Theory and Methods, 38(19):3510–3531, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Quoreshi:2006:BTS

[Quo06] A. M. M. Shahiduzzaman Quoreshi. Bivariate time series modeling of financial count data. Communications in Statistics: Theory and Methods, 35(7):1343–1358, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qin:2009:TLC

[QWSL09] Li Qin, Lisa A. Weissfeld, Changyu Shen, and Michele D. Levine. A two-latent-class model for smoking cessation data with informative dropouts. Communications in Statistics: Theory and Methods, 38(15):2604–2619, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodriguez-Avi:2003:EPG

[RACSSCJ03] J. Rodríguez-Avi, A. Conde-Sánchez, A. J. Sáez-Castillo, and M. J. Olmo Jiménez. Estimation of parameters in Gaussian hypergeometric distributions. Communications in Statistics: Theory and Methods, 32(6):1101–1118, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodriguez-Avi:2005:EIC

[RACSSCJ05] J. Rodríguez-Avi, A. Conde-Sánchez, A. J. Sáez-Castillo, and M. J. Olmo Jiménez. Estimation and inference in complex triparametric Pearson distributions. Communications in Statistics: Theory and Methods, 34(2):277–289, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodriguez-Avi:2007:GHP

[RACSSCOJ07] J. Rodríguez-Avi, A. Conde-Sánchez, A. J. Sáez-Castillo, and M. J. Olmo-Jiménez. Gaussian hypergeometric probability distributions for fitting discrete data. Communications in Statistics: Theory and Methods, 36(3):453–463, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rahimov:2008:ANE

[Rah08] I. Rahimov. Asymptotically normal estimators for the offspring mean in the branching process with immigration. Communications in Statistics: Theory and Methods, 38(1): 13–28, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Razmkhah:2008:NCI

[RAK08] M. Razmkhah, J. Ahmadi, and B. Khatib. Nonparametric confidence intervals and tolerance limits based on minima and maxima. Communications in Statistics: Theory and Methods, 37(10):1525–1542, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ramanayake:2005:TCP

[Ram05] Asoka Ramanayake. Tests for a change point in the shape parameter of gamma random variables. Communications in Statistics: Theory and Methods, 33(4):821–833, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ramsay:2006:DSC

[Ram06] Colin M. Ramsay. The distribution of sums of certain I.I.D. Pareto variates. Communications in Statistics: Theory and Methods, 35(3):395–405, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ramsay:2008:DSD

[Ram08] Colin M. Ramsay. The distribution of sums of I.I.D. Pareto random variables with arbitrary shape parameter. Communications in Statistics: Theory and Methods, 37(14):2177– 2184, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:2001:SRP

[Rao01a] C. Radhakrishna Rao. Statistics: Reflections on the past and visions for the future. Communications in Statistics: Theory and Methods, 30(11):2235–2257, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:2001:NEO

[Rao01b] Poduri S. R. S. Rao. Nonnegative estimators for the one-way random effects model. Communications in Statistics: Theory and Methods, 30(8–9):1605–1613, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:2003:MIS

[Rao03] B. L. S. Prakasa Rao. Moment inequalities for supremum of empirical processes for ϕ-mixing sequences. Communications in Statistics: Theory and Methods, 32(9):1695–1701, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:2004:BLM

[Rao04] B. L. S. Prakasa Rao. On a bivariate lack of memory property under binary associative operation. Communications in Statistics: Theory and Methods, 33(12):3103–3114, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raqab:2000:MRV

[Raq00] Mohammad Z. Raqab. On the moments of record values. Communications in Statistics: Theory and Methods, 29(7):1631–1647, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raqab:2004:SBE

[Raq04] Mohammad Z. Raqab. Sharp bounds on the error in approximating the means of record statistics by inverse hazard functions. Communications in Statistics: Theory and Methods, 33(7):1527–1539, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raqab:2007:IEC

[Raq07] Mohammad Z. Raqab. Inequalities for expected current record statistics. Communications in Statistics: Theory and Methods, 36(7):1367–1380, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rau:2009:BEB

[Rau09] Re-Bin Rau. Bayes and empirical Bayes two-stage procedures for sampling inspection. Communications in Statistics: Theory and Methods, 38(6):871–887, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ravishanker:2001:DGA

[Rav01] Nalini Ravishanker. Differential geometry of ARFIMA processes. Communications in Statistics: Theory and Methods, 30(8–9):1889–1902, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Radouane:2002:NST

[RC02] Ouadia Radouane and Emmanuelle Crétois. Neyman smooth tests for the generalized Pareto distribution. Communications in Statistics: Theory and Methods, 31(7):1067–1078, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Requena:2003:MPC

[RC03] F. Requena and N. Martín Ciudad. The maximum probability 2 × c contingency tables and the maximum probability points of the multivariate hypergeometric distribution. Communications in Statistics: Theory and Methods, 32(9):1737– 1752, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Revathy:2007:EEP

[RC07] S. Amala Revathy and B. Chandrasekar. Equivariant estimation of parameters based on sequential order statistics from (1, 3) and (2, 3) systems. Communications in Statistics: Theory and Methods, 36(3):541–548, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodrigues:2002:BAS

[RCA02] Josemar Rodrigues, Juan E. R. Cid, and Jorge A. Achcar. Bayesian analysis for the superposition of two dependent nonhomogeneous Poisson processes. Communications in Statistics: Theory and Methods, 31(9):1467–1478, 2002. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Retout:2009:DON

[RCBM09] Sylvie Retout, Emmanuelle Comets, Caroline Bazzoli, and France Mentré. Design optimization in nonlinear mixed effects models using cost functions: Application to a joint model of infliximab and methotrexate pharmacokinetics. Communications in Statistics: Theory and Methods, 38(18): 3351–3368, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ren:2008:SSP

[RCL08] Shiquan Ren, Haitao Chu, and Shenghan Lai. Sample size and power calculations for left-truncated normal distribution. Communications in Statistics: Theory and Methods, 37(6):847–860, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ruiz-Castro:2004:SMM

[RCPO04] Juan Eloy Ruiz-Castro and Rafael Pérez-Ocón. A semi-Markov model in biomedical studies. Communications in Statistics: Theory and Methods, 33(3):437–455, 2004. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ran:2006:SAN

[RCX06] Lun Ran, Lirong Cui, and M. Xie. Some analytical and numerical bounds on the renewal function. Communications in Statistics: Theory and Methods, 35(10):1815–1827, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ravishanker:2001:CTS

[RDI01] Nalini Ravishanker, Dipak K. Dey, and Malini Iyengar. Compositional time series analysis of mortality proportions. Communications in Statistics: Theory and Methods, 30(11):

2281–2291, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Reed:2007:BLM

[Ree07] William J. Reed. Brownian–Laplace motion and its use in financial modelling. Communications in Statistics: Theory and Methods, 36(3):473–484, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Reif:2006:PCC

[Rei06] J. Reif. Pitman closeness in classes of general pre-test estimators and regression estimators. Communications in Statistics: Theory and Methods, 35(2):263–279, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rubio:2002:DSS

[RF02] Hernán Rubio and Luis Firinguetti. The distribution of stochastic shrinkage parameters in ridge regression. Communications in Statistics: Theory and Methods, 31(9):1531– 1547, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ramanayake:2004:ECM

[RG04] Asoka Ramanayake and Arjun K. Gupta. Epidemic change model for the exponential family. Communications in Statistics: Theory and Methods, 33(9):2175–2198, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Richman:2007:SES

[Ric07] Joshua S. Richman. Sample entropy statistics and testing for order in complex physiological signals. Communications in Statistics: Theory and Methods, 36(5):1005–1019, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rieck:2003:CTR

[Rie03] James R. Rieck. A comparison of two random number generators for the Birnbaum–Saunders distribution. Communications in Statistics: Theory and Methods, 32(5):929–934, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rohan:2000:ERO

[RJ00] V. M. Rohan and G. Jones. Efficient run orders for a twofactor response surface experiment on a correlated process. Communications in Statistics: Theory and Methods, 29(3): 593–609, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Reed:2004:DPL

[RJ04] William J. Reed and Murray Jorgensen. The double Paretolognormal distribution — a new parametric model for size distributions. Communications in Statistics: Theory and Methods, 33(8):1733–1753, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Razzaghi:2001:EEA

[RK01] Mehdi Razzaghi and Ralph L. Kodell. An extension of the em algorithm for optimization of constrained likelihood: an application in toxicology. Communications in Statistics: Theory and Methods, 30(11):2317–2327, 2001. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rong:2009:CFO

[RL09] Jian-Ying Rong and Xu-Qing Liu. On the closed form of optimal invariant quadratic predictors in finite populations. Communications in Statistics: Theory and Methods, 38(15):2541–2551, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodriguez-Lallena:2005:BPB

[RLÚF05] José Antonio Rodríguez-Lallena and Manuel Úbeda-Flores. Best-possible bounds on sets of multivariate distribution functions. Communications in Statistics: Theory and Methods, 33(4):805–820, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rueda:2009:MEU

[RMAM09] M. Rueda, S. Martínez, A. Arcos, and J. F. Muñoz. Mean estimation under successive sampling with calibration estimators. Communications in Statistics: Theory and Methods, 38(6):808–827, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raqab:2008:ETP

[RMK08] Mohammad Z. Raqab, Mohamed T. Madi, and Debasis Kundu. Estimation of P (Y < X) for the three-parameter generalized exponential distribution. Communications in Statistics: Theory and Methods, 37(18):2854–2864, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raats:2006:MMD

[RMvdG06] V. M. Raats, J. J. A. Moors, and B. B. van der Genugten. A mixed model for double checking fallible auditors. Communications in Statistics: Theory and Methods, 35(9):1723– 1743, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodrigues:2003:BAZ

[Rod03] Josemar Rodrigues. Bayesian analysis of zero-inflated distributions. Communications in Statistics: Theory and Methods, 32(2):281–289, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodrigues:2005:EFM

[Rod05] Josemar Rodrigues. An entropy frailty model for dependent variables. Communications in Statistics: Theory and Methods, 33(4):979–990, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodrigues:2006:FBS

[Rod06] Josemar Rodrigues. Full Bayesian significance test for zeroinflated distributions. Communications in Statistics: Theory and Methods, 35(2):299–307, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodrigues:2008:MRM

[Rod08] Josemar Rodrigues. Mixture representation of the maximum entropy density through Archimedean copulas. Communications in Statistics: Theory and Methods, 37(18):2943–2952, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rogerson:2006:FDC

[Rog06] Peter A. Rogerson. Formulas for the design of CUSUM quality control charts. Communications in Statistics: Theory and Methods, 35(2):373–383, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rossa:2002:ESF

[Ros02] Agnieszka Rossa. On the estimation of survival function under random censorship. Communications in Statistics: Theory and Methods, 31(6):961–975, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rossa:2009:NAK

[Ros09] Agnieszka Rossa. The Nelson–Aalen and Kaplan–Meier estimators under a sequential sampling scheme. Communications in Statistics: Theory and Methods, 38(16–17):3077– 3098, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roy:2003:DND

[Roy03] Dilip Roy. The discrete normal distribution. Communications in Statistics: Theory and Methods, 32(10):1871–1883, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roy:2004:GLM

[Roy04] Dilip Roy. A generalization of the lack of memory property and related characterization results. Communications in Statistics: Theory and Methods, 33(12):3145–3158, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raqab:2002:SBM

[RR02] Mohammad Z. Raqab and Tomasz Rychlik. Sharp bounds for the mean of k th record value. Communications in Statistics: Theory and Methods, 31(11):1927–1937, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rancel:2001:RDU

[RS01] M. Mercedes Suárez Rancel and Miguel A. González Sierra. Regression diagnostic using local influence: a review. Communications in Statistics: Theory and Methods, 30(5):799– 813, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rafajlowicz:2003:EDS

[RS03] E. Rafajlowicz and R. Schwabe. Experimental design for (semi-)local regression. Communications in Statistics: Theory and Methods, 32(5):1035–1055, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:2004:CSL

[RS04] C. R. Rao and D. N. Shanbhag. Characterizations of stable laws based on a number theoretic result. Communications in Statistics: Theory and Methods, 33(12):2873–2884, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:2007:WVS

[RSG07] A. V. Dattatreya Rao, I. Ramabhadra Sarma, and S. V. S. Girija. On wrapped version of some life testing models. Communications in Statistics: Theory and Methods, 36(11):2027– 2035, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ramamurthy:2003:CAC

[RSS03] K. G. Ramamurthy, Arvind Seth, and Mohammad Khanjari Sadegh. A combinatorial approach for component importance. Communications in Statistics: Theory and Methods, 32(2):497–507, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:2004:EFJ

[RT04] J. N. K. Rao and M. Tausi. Estimating function jackknife variance estimators under stratified multistage sampling. Communications in Statistics: Theory and Methods, 33(9):2087–2095, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roths:2006:RBC

[RT06] Scott A. Roths and Joshua M. Tebbs. Revisiting Beal’s confidence intervals for the difference of two binomial proportions. Communications in Statistics: Theory and Methods, 35(9):1593–1609, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rifo:2009:FBA

[RT09] Laura L. R. Rifo and Soledad Torres. Full Bayesian analysis for a class of jump-diffusion models. Communications in Statistics: Theory and Methods, 38(8):1262–1271, 2009. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rech:2001:SVS

[RTT01] Gianluigi Rech, Timo Teräsvirta, and Rolf Tschernig. A simple variable selection technique for nonlinear models. Communications in Statistics: Theory and Methods, 30(6):1227– 1241, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:2006:SEN

[RTW06] C. R. Rao, K. W. Tam, and Y. Wu. On simultaneous estimation of the number of signals and frequencies of multiple sinusoids when some observations are missing. Communications in Statistics: Theory and Methods, 35(12):2147–2155, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rudas:2002:CRL

[Rud02] Tamás Rudas. Canonical representation of log-linear models. Communications in Statistics: Theory and Methods, 31 (12):2311–2323, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rukhin:2009:STG

[Ruk09] Andrew L. Rukhin. Sequential testing to guarantee the necessary sample size in clinical trials. Communications in Statistics: Theory and Methods, 38(16–17):3114–3122, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raats:2005:GMR

[RvdGM05] V. M. Raats, B. B. van der Genugten, and J. J. A. Moors. A general model for repeated audit controls using monotone subsampling. Communications in Statistics: Theory and Methods, 33(4):949–977, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Romano:2002:ENC

[RW02] Joseph P. Romano and Michael Wolf. Explicit nonparametric confidence intervals for the variance with guaranteed coverage. Communications in Statistics: Theory and Methods, 31(8):1231–1250, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ryan:2000:SIL

[Rya00] Thomas P. Ryan. Some issues in logistic regression. Communications in Statistics: Theory and Methods, 29(9–10):2019– 2032, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/03610920008832593.

Rynkiewicz:2007:ETM

[Ryn07] J. Rynkiewicz. Estimation and test for multi-dimensional regression models. Communications in Statistics: Theory and Methods, 36(14):2655–2671, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Riggs:2009:LBC

[RYS09] Kent Riggs, Dean Young, and James Stamey. Likelihoodbased confidence intervals for complementary Poisson rate parameters with misclassified data. Communications in Statistics: Theory and Methods, 38(2):159–172, 2009. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rossa:2002:SIK

[RZ02] Agnieszka Rossa and Ryszard Zieliński. A simple improvement of the Kaplan–Meier estimator. Communications in Statistics: Theory and Methods, 31(1):147–158, 2002. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sansgiry:2001:REL

[SA01] Prashant S. Sansgiry and Olcay Akman. Reliability estimation via length-biased transformation. Communications in Statistics: Theory and Methods, 30(11):2473–2479, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shaarawy:2003:BIS

[SA03] Samir M. Shaarawy and Sherif S. Ali. Bayesian identification of seasonal autoregressive models. Communications in Statistics: Theory and Methods, 32(5):1067–1084, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sebille:2004:EHR

[SA04] Véronique Sébille and Jean-Louis Auget. Evaluating healthrelated quality of life: A case-study of differential item functioning analysis in small trials. Communications in Statistics: Theory and Methods, 33(6):1403–1428, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shaarawy:2008:BIM

[SA08] Samir M. Shaarawy and Sherif S. Ali. Bayesian identification of multivariate autoregressive processes. Communications in Statistics: Theory and Methods, 37(5):791–802, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saadi:2009:EPD

[SA09] Nora Saadi and Smail Adjabi. On the estimation of the probability density by trigonometric series. Communications in Statistics: Theory and Methods, 38(19):3583–3595, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soltani:2007:TMA

[SAAAS07] A. R. Soltani, S. A. Al-Awadhi, and W. M. Al-Shemeri. On thresholds of moving averages with given on-target significant levels. Communications in Statistics: Theory and Methods, 36(14):2595–2606, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sadegh:2008:MPM

[Sad08a] Mohammad Khanjari Sadegh. Mean past and mean residual life functions of a parallel system with nonidentical components. Communications in Statistics: Theory and Methods, 37(7):1134–1145, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sadegh:2008:NRO

[Sad08b] Mohammad Khanjari Sadegh. A note on the regression of order statistics. Communications in Statistics: Theory and Methods, 37(4):475–480, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saibal:2000:SEE

[Sai00] Chattopadhay Saibal. Sequential estimation of exponential location parameter using an asymmetric loss function. Communications in Statistics: Theory and Methods, 29(4): 783–795, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sain:2001:BRE

[Sai01] Stephan R. Sain. Bias reduction and elimination with kernel estimators. Communications in Statistics: Theory and Methods, 30(8–9):1869–1888, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sakaori:2002:NTE

[Sak02] F. Sakaori. A nonparametric test for the equality of dependent correlation coefficients under normality. Communications in Statistics: Theory and Methods, 31(12):2379–2389, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salmaso:2003:SPT

[Sal03] Luigi Salmaso. Synchronized permutation tests in 2^{k} factorial designs. Communications in Statistics: Theory and Methods, 32(7):1419–1437, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salehi:2006:ACR

[Sal06] Mohammad M. Salehi. Adaptive cluster row and column elimination sampling +(1) design. Communications in Statistics: Theory and Methods, 35(2):349–362, 2006. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samuelsen:2003:MCR

[Sam03] Sven Ove Samuelsen. Mimicking Cox-regression. Communications in Statistics: Theory and Methods, 32(1):151–169, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sampath:2009:FPV

[Sam09] S. Sampath. Finite population variance estimation under LSS with multiple random starts. Communications in Statistics: Theory and Methods, 38(19):3596–3607, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samawi:2004:BRS

[SAS04] Hani M. Samawi and Mohammad Fraiwan Al-Saleh. On bivariate ranked set sampling for distribution and quantile estimation and quantile interval estimation using ratio estimator. Communications in Statistics: Theory and Methods, 33(8):1801–1819, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samawi:2006:BST

[SASAS06] Hani M. Samawi, Mohammad F. Al-Saleh, and Obaid Al-Saidy. Bivariate sign test for one-sample bivariate location model using ranked set sample. Communications in Statistics: Theory and Methods, 35(6):1071–1083, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Silva:2005:BAA

[SAT05] Giovani L. Silva and M. Antónia Amaral-Turkman. Bayesian analysis of an additive survival model with frailty. Communications in Statistics: Theory and Methods, 33(10):2517– 2533, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salinas:2007:ESE

[SAVG07] Hugo S. Salinas, Reinaldo B. Arellano-Valle, and Héctor W. Gómez. The extended skew-exponential power distribution and its derivation. Communications in Statistics: Theory and Methods, 36(9):1673–1689, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shetty:2003:COR

[SB03] I. D. Shetty and Sharada V. Bhat. A class of outlier resistant distribution-free tests for testing homogeneity against ordered scale alternatives. Communications in Statistics:

Theory and Methods, 32(2):347–379, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sanjel:2006:UOP

[SB06] Deepak Sanjel and N. Balakrishnan. Use of orthogonal polynomial approximations for inference in exponential distribution based on K-sample doubly Type-II censored data. Communications in Statistics: Theory and Methods, 35(9):1671– 1683, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shi:2008:BCI

[SB08] Lei Shi and Peng Bai. Bayesian confidence interval for the difference of two proportions in the matched-paired design. Communications in Statistics: Theory and Methods, 37 (13):2034–2051, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shi:2009:BCI

[SB09] Lei Shi and Peng Bai. Bayesian confidence interval for the ratio of marginal probabilities in the matched-pair design. Communications in Statistics: Theory and Methods, 38(8):1300–1316, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:2007:CPH

[SBJ^{+}07] Deo Kumar Srivastava, James M. Boyett, Carl W. Jackson, Xin Tong, and Shesh N. Rai. A comparison of permutation Hotelling’s T^{2} test and log-ratio test for analyzing compositional data. Communications in Statistics: Theory and Methods, 36(2):415–431, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Street:2001:ODP

[SBM01] Deborah J. Street, David S. Bunch, and Beverley J. Moore. Optimal designs for 2^{k} paired comparison experiments. Communications in Statistics: Theory and Methods, 30(10): 2149–2171, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saw:2002:LED

[SBT02] Sutaip L. C. Saw, Uditha Balasooriya, and Khye Chong Tan. The log-eig distribution: a new probability model for lifetime data. Communications in Statistics: Theory and Methods, 31 (11):1913–1926, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sobel:2002:DRP

[SC02] Milton Sobel and Aaron Childs. Dirichlet-related probability problems and new tools for solving them. Communications in Statistics: Theory and Methods, 31(11):2027–2050, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shmueli:2003:RLD

[SC03] Galit Shmueli and Ayala Cohen. Run-length distribution for control charts with runs and scans rules. Communications in Statistics: Theory and Methods, 32(2):475–495, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shu:2005:EPC

[SC05] M. H. Shu and K. S. Chen. Estimating process capability indices based on subsamples for asymmetric tolerances. Communications in Statistics: Theory and Methods, 34(2): 485–505, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sisson:2006:NBA

[SC06] S. A. Sisson and Y. V. Chan. A note on Bayesian analyses of capture-recapture data with perfect recaptures. Communications in Statistics: Theory and Methods, 35(1):53–62, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2007:EBT

[sC07] Lee shen Chen. Empirical Bayes testing for a nonexponential family distribution. Communications in Statistics: Theory and Methods, 36(11):2061–2074, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sabo:2009:AQL

[SC09a] Roy T. Sabo and N. Rao Chaganty. Adaptation of quasileast squares to estimate correlations within a nuclear family.

Communications in Statistics: Theory and Methods, 38(16– 17):3059–3076, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sutradhar:2009:BMS

[SC09b] Rinku Sutradhar and Richard J. Cook. A bivariate moverstayer model for interval-censored recurrent event data: Application to joint damage in rheumatology. Communications in Statistics: Theory and Methods, 38(18):3389–3405, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scagliarini:2002:EAD

[Sca02] Michele Scagliarini. Estimation of C_{p} for autocorrelated data and measurement errors. Communications in Statistics: Theory and Methods, 31(9):1647–1664, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schwarz:2002:CIG

[Sch02] Wolf Schwarz. On the convolution of inverse Gaussian and exponential random variables. Communications in Statistics: Theory and Methods, 31(12):2113–2121, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schechtman:2005:SMI

[Sch05] Edna Schechtman. Stratification: Measuring and inference. Communications in Statistics: Theory and Methods, 34(11): 2133–2145, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Song:2009:RCI

[SCL09] Wheyming Tina Song, Chia-Jung Chang, and Sin-Long Liu. Robust confidence intervals for the Bernoulli parameter. Communications in Statistics: Theory and Methods, 38 (19):3544–3560, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scott:2000:TCM

[Sco00] W. F. Scott. Tables of the Cramér–von Mises distributions. Communications in Statistics: Theory and Methods, 29(1): 227–235, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scollnik:2006:IGP

[Sco06] David P. M. Scollnik. On the intervened generalized Poisson distribution. Communications in Statistics: Theory and Methods, 35(6):953–963, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scott:2007:ADL

[Sco07] W. F. Scott. On the asymptotic distribution of the likelihood ratio statistic. Communications in Statistics: Theory and Methods, 36(2):273–281, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinha:2008:SAU

[SDM08] Bikas K. Sinha, Kishore K. Das, and Sujay K. Mukhoti. On some aspects of unbiased estimation of parameters in quasibinomial distributions. Communications in Statistics: Theory and Methods, 37(19):3023–3028, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shaarawy:2001:BCA

[SDZ01] Samir M. Shaarawy, Ahmed L. Daif, and Alyaa R. Zahran. Bayesian classification of ARMA sources with unknown order. Communications in Statistics: Theory and Methods, 30 (6):1143–1154, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sepanski:2000:RCP

[Sep00] Jungsywan H. Sepanski. On a random coefficient probit model. Communications in Statistics: Theory and Methods, 29(11):2493–2505, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sethuraman:2004:SEE

[Set04] Jayaram Sethuraman. Are super-efficient estimators superpowerful? Communications in Statistics: Theory and Methods, 33(9):2003–2013, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sankaran:2004:CUL

[SG04] P. G. Sankaran and R. P. Gupta. Characterizations using local dependence function. Communications in Statistics:

Theory and Methods, 33(12):2959–2974, 2004. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sankaran:2005:GCD

[SG05a] P. G. Sankaran and Rameshwar D. Gupta. A general class of distributions: Properties and applications. Communications in Statistics: Theory and Methods, 34(11):2089–2095, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2005:OST

[SG05b] Parminder Singh and Amar Nath Gill. A one sided test based on sample quasi ranges. Communications in Statistics: Theory and Methods, 33(4):835–849, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:2005:PEP

[SG05c] Tanuja Srivastava and Himani Gupta. Pitman estimator to Pareto scale parameters. Communications in Statistics: Theory and Methods, 34(12):2235–2255, 2005. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shabbir:2006:NEP

[SG06] Javid Shabbir and Sat Gupta. A new estimator of population mean in stratified sampling. Communications in Statistics: Theory and Methods, 35(7):1201–1209, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shabbir:2007:IVE

[SG07] Javid Shabbir and Sat Gupta. On improvement in variance estimation using auxiliary information. Communications in Statistics: Theory and Methods, 36(12):2177–2185, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sankaran:2008:AMB

[SG08] P. G. Sankaran and V. L. Gleeja. Association measures for bivariate lifetime data. Communications in Statistics: Theory and Methods, 37(20):3228–3249, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sun:2009:MRM

[SGC09] Liuquan Sun, Shaojun Guo, and Min Chen. Marginal regression model with time-varying coefficients for panel data. Communications in Statistics: Theory and Methods, 38(8): 1241–1261, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saavedra:2000:SAR

[SHA00] P. Saavedra, C. N. Hernández, and J. Artiles. Spectral analysis with replicated time series. Communications in Statistics: Theory and Methods, 29(11):2343–2362, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shanmugam:2001:PSP

[Sha01a] Ramalingam Shanmugam. Predicting a “successful” prevention of an epidemic. Communications in Statistics: Theory and Methods, 30(1):93–103, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shanmugam:2001:PMG

[Sha01b] Ramalingam Shanmugam. Product moment generating function technique and its applications to modified power series distributions. Communications in Statistics: Theory and Methods, 30(11):2385–2392, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shao:2002:MLE

[Sha02a] Quanxi Shao. Maximum likelihood estimation for generalised logistic distributions. Communications in Statistics: Theory and Methods, 31(10):1687–1700, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shao:2002:RME

[Sha02b] Quanxi Shao. A reparameterisation method for embedded models. Communications in Statistics: Theory and Methods, 31(5):683–697, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shams:2007:PCC

[Sha07] S. Shams. Properties of certain class of mixture distributions. Communications in Statistics: Theory and Methods, 36(2):

229–233, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shao:2009:NTE

[Sha09] Q. Shao. Nonparametric trend estimation for periodic autoregressive time series. Communications in Statistics: Theory and Methods, 38(14):2418–2427, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Seok:2002:PIS

[SHC02] Kyungha Seok, Changha Hwang, and Daehyeon Cho. Prediction intervals for support vector machine regression. Communications in Statistics: Theory and Methods, 31(10):1887– 1898, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sheena:2005:OPE

[She05a] Y. Sheena. Order-preserving estimators and an inequality on the integration of zonal polynomial. Communications in Statistics: Theory and Methods, 34(7):1503–1516, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2005:NBE

[She05b] Pao-Sheng Shen. Nonparametric bivariate estimation with left-truncated and right-censored data. Communications in Statistics: Theory and Methods, 34(7):1675–1685, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2007:GCT

[She07] Pao-Sheng Shen. A general class of test procedures for left-truncated and right-censored data. Communications in Statistics: Theory and Methods, 36(16):2913–2925, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sheng:2009:BMM

[She09] Xiaoming Sheng. A Bayesian model for measurement errors in diagnosis of rheumatoid arthritis. Communications in Statistics: Theory and Methods, 38(18):3419–3431, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shieh:2000:GML

[Shi00] Gwowen Shieh. General multivariate linear models for longitudinal studies. Communications in Statistics: Theory and Methods, 29(4):735–753, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shimodaira:2001:MCL

[Shi01] Hidetoshi Shimodaira. Multiple comparisons of loglikelihoods and combining nonnested models with applications to phylogenetic tree selection. Communications in Statistics: Theory and Methods, 30(8–9):1751–1772, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shitan:2008:CFI

[Shi08a] Mahendran Shitan. Corrigendum: Fractionally Integrated Separable Spatial Autoregressive (FISSAR) Model and Some of Its Properties, Communications in Statistics — Theory and Methods, Volume 37 Issue No. 8, January 2008, Pages 1266–1273. Communications in Statistics: Theory and Methods, 38(1):156–158, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [Shi08b].

Shitan:2008:FIS

[Shi08b] Mahendran Shitan. Fractionally integrated separable spatial autoregressive (FISSAR) model and some of its properties. Communications in Statistics: Theory and Methods, 37 (8):1266–1273, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See corrigendum [Shi08a].

Shiau:2009:MNP

[SHLT09] Jyh-Jen Horng Shiau, Hsiang-Ling Huang, Shuo-Hui Lin, and Ming-Ye Tsai. Monitoring nonlinear profiles with random effects by nonparametric regression. Communications in Statistics: Theory and Methods, 38(10):1664–1679, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shore:2002:RMM

[Sho02] Haim Shore. Response modeling methodology (RMM)- exploring the properties of the implied error distribution. Communications in Statistics: Theory and Methods, 31(12):

2225–2249, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shore:2005:ARB

[Sho05] Haim Shore. Accurate RMM-based approximations for the CDF of the normal distribution. Communications in Statistics: Theory and Methods, 34(3):507–513, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shore:2007:CGL

[Sho07] Haim Shore. Comparison of generalized Lambda distribution (GLD) and response modeling methodology (RMM) as general platforms for distribution fitting. Communications in Statistics: Theory and Methods, 36(15):2805–2819, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sauerbrei:2006:EBA

[SHRA06] Willi Sauerbrei, Norbert Holländer, Richard D. Riley, and Douglas G. Altman. Evidence-based assessment and application of prognostic markers: The long way from single studies to meta-analysis. Communications in Statistics: Theory and Methods, 35(7):1333–1342, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shults:2000:MCS

[Shu00] Justine Shults. Modeling the correlation structure of data that have multiple levels of association. Communications in Statistics: Theory and Methods, 29(5–6):1005–1015, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shimizu:2002:PTV

[SI02] Kunio Shimizu and Kazuyuki Iida. Pearson Type VII distributions on spheres. Communications in Statistics: Theory and Methods, 31(4):513–526, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Silvestrov:2004:UBE

[Sil04] Dmitrii Silvestrov. Upper bounds for exponential moments of hitting times for semi-Markov processes. Communications in Statistics: Theory and Methods, 33(3):533–543, 2004. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Silva:2006:NIA

[Sil06] J. M. C. Santos Silva. A note on influence assessment in score tests. Communications in Statistics: Theory and Methods, 35(7):1243–1256, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Simic:2008:ABR

[Sim08] Slavko Simic. On asymptotic behavior of regularly varying moments for discrete laws. Communications in Statistics: Theory and Methods, 38(1):130–137, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Song:2000:ACI

[SJ00] Peter X. K. Song and Wenxin Jiang. Assessing conditional independence for log-linear Poisson models with random effects. Communications in Statistics: Theory and Methods, 29(5–6):1233–1245, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sim:2004:NTM

[SJ04] Songyong Sim and Richard A. Johnson. New tests for multivariate ordered alternatives. Communications in Statistics: Theory and Methods, 33(9):2027–2039, 2004. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sarika:2009:SOR

[SJS09] Sarika, Seema Jaggi, and V. K. Sharma. Second-order response surface model with neighbor effects. Communications in Statistics: Theory and Methods, 38(9):1393–1403, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:2000:EML

[SK00] D. M. Smith and M. G. Kenward. Extensions of multiple linear regression. Communications in Statistics: Theory and Methods, 29(9–10):2033–2053, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610920008832594.

Sane:2001:EMI

[SK01] Manisha Sane and A. V. Kharshikar. Effect of missing an influential covariate. Communications in Statistics: Theory and Methods, 30(5):837–853, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Song:2005:DRV

[SK05] Ruiguang Song and John M. Karon. Distributions of renewal variables when renewal process reaches a special event. Communications in Statistics: Theory and Methods, 34(8):1813– 1819, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:2006:NPC

[SK06] Ananda Sen and Devadatta Kulkarni. A note on a path counting formula. Communications in Statistics: Theory and Methods, 35(1):3–11, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sarhan:2009:GLF

[SK09a] Ammar M. Sarhan and Debasis Kundu. Generalized linear failure rate distribution. Communications in Statistics: Theory and Methods, 38(5):642–660, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2009:CEU

[SK09b] Housila P. Singh and Namrata Karpe. A class of estimators using auxiliary information for estimating finite population variance in presence of measurement errors. Communications in Statistics: Theory and Methods, 38(5):734–741, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sakallioglu:2001:MSE

[SKA01] Sadullah Sakallioglu, Selabattin Kaçiranlar, and Fikri Akdeniz. Mean squared error comparisons of some biased regression estimators. Communications in Statistics: Theory and Methods, 30(2):347–361, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sadek:2002:APM

[SL02] Amr Sadek and Nikolaos Limnios. Asymptotic properties for maximum likelihood estimators for reliability and failure rates of Markov chains. Communications in Statistics: Theory and Methods, 31(10):1837–1861, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shubina:2004:MAC

[SL04] Maria Shubina and Mei-Ling Ting Lee. On maximum attainable correlation and other measures of dependence for the Sarmanov family of bivariate distributions. Communications in Statistics: Theory and Methods, 33(5):1031–1052, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sobel:2005:DVP

[SL05] Marc Sobel and Longin Jan Latecki. Data visualization by pairwise distortion minimization. Communications in Statistics: Theory and Methods, 34(6):1379–1391, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shalabh:2007:CER

[SL07] Shalabh and Pen-Hwang Liau. Consistent estimation of the regression coefficient through weighted arithmetic mean of the inconsistent estimators in replicated ultrastructural model. Communications in Statistics: Theory and Methods, 36(5):955–960, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saidane:2008:EBV

[SL08a] Mohamed Saidane and Christian Lavergne. An EM-based Viterbi approximation algorithm for mixed-state latent factor models. Communications in Statistics: Theory and Methods, 37(17):2795–2814, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

SenGupta:2008:LIM

[SL08b] Ashis SenGupta and Arnab Kumar Laha. A likelihood integrated method for exploratory graphical analysis of change point problem with directional data. Communications in Statistics: Theory and Methods, 37(11):1783–1791, 2008.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Seck:2009:UCN

[SL09a] Cheikh Tidiane Seck and Gane Samb Lo. Uniform convergence of the non-weighted poverty measures. Communications in Statistics: Theory and Methods, 38(20):3697–3704, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Su:2009:SPE

[SL09b] Yuxia Su and Lu Lin. Semi-parametric estimation for forward-backward stochastic differential equations. Communications in Statistics: Theory and Methods, 38(11):1759– 1775, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sanhueza:2008:GBS

[SLB08] Antonio Sanhueza, Víctor Leiva, and N. Balakrishnan. The generalized Birnbaum–Saunders distribution and its theory, methodology, and application. Communications in Statistics: Theory and Methods, 37(5):645–670, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2008:ADI

[SLZ08] Xinmei Shen, Zhengyan Lin, and Caiya Zhang. Asymptotic distributions of innovation density estimators in linear processes. Communications in Statistics: Theory and Methods, 37(14):2262–2275, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:2000:TBT

[SM00] Kanwar Sen and Sunil Kumar Mathur. A test for bivariate two sample location problem. Communications in Statistics: Theory and Methods, 29(2):417–436, 2000. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Szekely:2001:CMA

[SM01] G. J. Székely and T. F. Móri. A characteristic measure of asymmetry and its application for testing diagonal symmetry. Communications in Statistics: Theory and Methods, 30 (8–9):1633–1639, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saura:2006:EMF

[SM06a] F. Saura and J. Mateu. Estimating mark functions through spectral analysis for marked point patterns. Communications in Statistics: Theory and Methods, 35(5):861–885, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sunoj:2006:SPW

[SM06b] S. M. Sunoj and S. S. Maya. Some properties of weighted distributions in the context of repairable systems. Communications in Statistics: Theory and Methods, 35(2):223–228, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sengupta:2008:UEE

[SM08a] Samindranath Sengupta and Sujay Mukhuti. Unbiased estimation of P (X > Y) for exponential populations using order statistics with application in ranked set sampling. Communications in Statistics: Theory and Methods, 37(6):898–916, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sobel:2008:NAE

[SM08b] Marc Sobel and Bud Mishra. A new approach to the estimation of inter-variable correlation. Communications in Statistics: Theory and Methods, 37(15):2315–2330, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sengupta:2009:UED

[SM09a] S. Sengupta and Sujay Mukhuti. Unbiased estimation of the distribution function of a two-parameter exponential population using order statistics. Communications in Statistics: Theory and Methods, 38(15):2578–2585, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2009:EVF

[SM09b] Silian Shen and Changlin Mei. Estimation of the variance function in heteroscedastic linear regression models. Communications in Statistics: Theory and Methods, 38(7):1098– 1112, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sanchez-Manzano:2002:MVG

[SMGVMD02] Eusebio Gómez Sánchez-Manzano, Miguel Angel Gómez-Villegas, and Juan-Miguel Marín-Diazaraque. A matrix variate generalization of the power exponential family of distributions. Communications in Statistics: Theory and Methods, 31(12):2167–2182, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:2000:E

[Smi00a] William B. Smith. Editorial. Communications in Statistics: Theory and Methods, 29(1):v, 2000. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:2000:IOT

[Smi00b] William B. Smith. Inference with overlapping 2 × 2 tables. Communications in Statistics: Theory and Methods, 29(1): 237–241, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:2005:DSS

[Smi05] Narelle F. Smith. The design of scoring schemes for surveys using the block total response method. Communications in Statistics: Theory and Methods, 34(11):2157–2168, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:2007:ITF

[Smi07] Murray D. Smith. Invariance theorems for Fisher information. Communications in Statistics: Theory and Methods, 36 (12):2213–2222, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stute:2000:NMC

[SMS00] W. Stute, W. González Manteiga, and C. Sánchez Sellero. Nonparametric model checks in censored regression. Communications in Statistics: Theory and Methods, 29(7):1611– 1629, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sankaran:2000:CMR

[SN00] P. G. Sankaran and N. Unnikrishnan Nair. Conditional mean residual life functions. Communications in Statistics: Theory and Methods, 29(7):1663–1675, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sungur:2005:DCU

[SN05] Engin A. Sungur and Peh Ng. A decomposition of copulas and its use. Communications in Statistics: Theory and Methods, 34(12):2269–2282, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schucany:2006:PGF

[SN06] William R. Schucany and H. K. Tony Ng. Preliminary goodness-of-fit tests for normality do not validate the onesample Student t. Communications in Statistics: Theory and Methods, 35(12):2275–2286, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Satoh:2005:NMR

[SO05] Kenichi Satoh and Megu Ohtaki. A note on multiple regression for single index model. Communications in Statistics: Theory and Methods, 33(10):2409–2422, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Satoh:2006:NGC

[SO06] Kenichi Satoh and Megu Ohtaki. Nonparametric growth curve model with local linear approximation. Communications in Statistics: Theory and Methods, 35(4):641–648, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soliman:2000:CLQ

[Sol00] Ahmed A. Soliman. Comparison of LINEX and quadratic Bayes estimators for the Rayleigh distribution. Communications in Statistics: Theory and Methods, 29(1):95–107, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soliman:2006:EFM

[Sol06] Ahmed A. Soliman. Estimators for the finite mixture of Rayleigh model based on progressively censored data. Communications in Statistics: Theory and Methods, 35(5):803– 820, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soliman:2008:EPM

[Sol08] Ahmed A. Soliman. Estimations for Pareto model using general progressive censored data and asymmetric loss. Communications in Statistics: Theory and Methods, 37(9):1353– 1370, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shoemaker:2001:SBC

[SP01] O. J. Shoemaker and P. K. Pathak. The sequential bootstrap: a comparison with regular bootstrap. Communications in Statistics: Theory and Methods, 30(8–9):1661–1674, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sarabia:2002:CLC

[SP02] José María Sarabia and Marta Pascual. A class of Lorenz curves based on linear exponential loss functions. Communications in Statistics: Theory and Methods, 31(6):925–942, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sanchez:2003:IMO

[SP03a] María Jesús Sánchez and Daniel Peña. The identification of multiple outliers in ARIMA models. Communications in Statistics: Theory and Methods, 32(6):1265–1287, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Svetliza:2003:DNN

[SP03b] Carolina F. Svetliza and Gilberto A. Paula. Diagnostics in nonlinear negative binomial models. Communications in Statistics: Theory and Methods, 32(6):1227–1250, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shetty:2004:NNC

[SP04] I. D. Shetty and Parameshwar V. Pandit. A note on a new class of distribution-free tests for the two-sample scale problem based on subsample medians. Communications in Statistics: Theory and Methods, 33(9):2271–2280, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sothinathan:2005:BMO

[SP05] N. Sothinathan and L. I. Pettit. Bayes methods for outliers in binomial samples. Communications in Statistics: Theory and Methods, 34(2):351–366, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shah:2007:BES

[SP07] J. B. Shah and M. N. Patel. Bayes estimation of shift point in geometric sequence. Communications in Statistics: Theory and Methods, 36(6):1139–1151, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2008:SGR

[SP08a] G. N. Singh and Kumari Priyanka. Search of good rotation patterns to improve the precision of estimates at current occasion. Communications in Statistics: Theory and Methods, 37(3):337–348, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spezia:2008:BIF

[SP08b] Luigi Spezia and Roberta Paroli. Bayesian inference and forecasting in dynamic neural networks with fully Markov switching ARCH noises. Communications in Statistics: Theory and Methods, 37(13):2079–2094, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shah:2009:BES

[SP09] J. B. Shah and M. N. Patel. Bayes estimation of shift point in normal sequence and its application to statistical process control. Communications in Statistics: Theory and Methods, 38(4):560–572, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Satpati:2007:EBD

[SPG07] Subrata Kumar Satpati, Rajender Parsad, and V. K. Gupta. Efficient block designs for dependent observations — a computer-aided search. Communications in Statistics: Theory and Methods, 36(6):1187–1223, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spitzner:2007:CCV

[Spi07] Dan J. Spitzner. Constructive cross-validation in linear prediction. Communications in Statistics: Theory and Methods, 36(5):939–953, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2007:STS

[SPS07] D. C. Singh, Gyan Prakash, and P. Singh. Shrinkage testimators for the shape parameter of Pareto distribution using LINEX loss function. Communications in Statistics: Theory and Methods, 36(4):741–753, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spurrier:2002:ILB

[Spu02] John D. Spurrier. An improved lower bound for Lehmann’s Lambda. Communications in Statistics: Theory and Methods, 31(2):193–199, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spurrier:2005:IUB

[Spu05a] John D. Spurrier. Improved upper bounds for Hollander’s Mu and Lehmann’s Lambda. Communications in Statistics: Theory and Methods, 34(8):1851–1855, 2005. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spurrier:2005:MCB

[Spu05b] John D. Spurrier. Multiple comparisons with the best control in a one-way layout. Communications in Statistics: Theory and Methods, 34(3):651–660, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salama:2001:SF

[SQ01] Ibrahim A. Salama and Dana Quade. The symmetric footrule. Communications in Statistics: Theory and Methods, 30(6):1099–1109, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salama:2004:AAC

[SQ04] I. A. Salama and D. Quade. Agreement among censored rankings using Spearman’s footrule. Communications in Statistics: Theory and Methods, 33(8):1837–1850, 2004. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastav:2002:CGG

[Sri02] Sudesh K. Srivastav. On construction of generalized group divisible designs with two groups. Communications in Statistics: Theory and Methods, 31(4):639–647, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:2000:TEA

[SS00a] Kanwar Sen and Preeti Wanti Srivastava. Tests for exponentiality against new better than old in expectation and new better than some used in expectation alternatives. Communications in Statistics: Theory and Methods, 29(1):157–180, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shinde:2000:EPM

[SS00b] Ramkrishna L. Shinde and Ashok Shanubhogue. Estimation of parameters and the mean life of a mixed failure time distribution. Communications in Statistics: Theory and Methods, 29(11):2621–2642, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sutradhar:2001:CBB

[SS01] Santosh C. Sutradhar and Brajendra Sutradhar. On classification of a bivariate binary observation. Communications in Statistics: Theory and Methods, 30(11):2259–2279, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2002:NHE

[sS02a] Pao sheng Shen. A note on the homogenetic estimate for the variance of the Kaplan–Meier estimate. Communications in Statistics: Theory and Methods, 31(9):1595–1603, 2002. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sutradhar:2002:PLI

[SS02b] Brajendra C. Sutradhar and Sanjoy K. Sinha. On pseudolikelihood inference in the binary longitudinal mixed model.

Communications in Statistics: Theory and Methods, 31(3): 397–417, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shanmugam:2003:IVI

[SS03a] Ramalingam Shanmugam and Jagbir Singh. Index of variation with illustration using molecular data. Communications in Statistics: Theory and Methods, 32(2):509–517, 2003. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2003:PLE

[sS03b] Pao sheng Shen. The product-limit estimate as an inverseprobability-weighted average. Communications in Statistics: Theory and Methods, 32(6):1119–1133, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Silvestrov:2004:PPS

[SS04] Dmitrii Silvestrov and Fredrik Stenberg. A pricing process with stochastic volatility controlled by a semi-Markov process. Communications in Statistics: Theory and Methods, 33(3):591–608, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2005:ESF

[sS05a] Pao sheng Shen. Estimation of survival function subject to time-dependent covariates and left truncation. Communications in Statistics: Theory and Methods, 33(10):2503–2516, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2005:BSE

[SS05b] Housila P. Singh and Sharad Saxena. Bayesian and shrinkage estimation of process capability index C_{p}. Communications in Statistics: Theory and Methods, 34(1):205–228, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastav:2005:SMC

[SS05c] Sudesh K. Srivastav and Arti Shankar. Some methods of constructing generalized group divisible designs with two groups. Communications in Statistics: Theory and Methods, 34(1):127–133, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saxena:2006:OSS

[SS06] Sharad Saxena and Housila P. Singh. From ordinary to shrinkage square-root estimators. Communications in Statistics: Theory and Methods, 35(6):1037–1058, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sankaran:2007:PHM

[SS07a] P. G. Sankaran and V. N. Sreeja. Proportional hazards model for multivariate failure time data. Communications in Statistics: Theory and Methods, 36(8):1627–1641, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sill:2007:ETS

[SS07b] Michael W. Sill and Allan R. Sampson. Extension of a twostage conditionally unbiased estimator of the selected population to the bivariate normal case. Communications in Statistics: Theory and Methods, 36(4):801–813, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2008:HRC

[sS08a] Pao sheng Shen. Hazards regression for cancer studies. Communications in Statistics: Theory and Methods, 37(20):3319– 3328, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Siburg:2008:GC

[SS08b] Karl Friedrich Siburg and Pavel A. Stoimenov. Gluing copulas. Communications in Statistics: Theory and Methods, 37 (19):3124–3134, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shaarawy:2007:BIM

[SSA07] Samir M. Shaarawy, Emad E. A. Soliman, and Sherif S. Ali. Bayesian identification of moving average models. Communications in Statistics: Theory and Methods, 36(12):2301– 2312, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sun:2001:SQP

[SSD01] Yanqing Sun, Shan Sun, and Yuanan Diao. Smooth quantile processes from right censored data and construction of simultaneous confidence bands. Communications in Statistics: Theory and Methods, 30(4):707–727, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sun:2004:AHM

[SSF04] Jianguo Sun, Liuquan Sun, and Nancy Flournoy. Additive hazards model for competing risks analysis of the case-cohort design. Communications in Statistics: Theory and Methods, 33(2):351–366, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Song:2000:CBM

[SSG00] Ruiguand Song, Paul C. Schlecht, and Jensen H. Groff. A composite binomial model derived from correlated random variables. Communications in Statistics: Theory and Methods, 29(5–6):1179–1195, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Song:2002:NEL

[SSJ02] Seuck Heun Song, Dietmar Stemann, and Byoung Cheol Jung. A note on estimating linear trend in a regression model with serially correlated error components. Communications in Statistics: Theory and Methods, 31(8):1385–1398, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sanhueza:2008:RPN

[SSL08] Antonio Sanhueza, Pranab K. Sen, and Víctor Leiva. A robust procedure in nonlinear models for repeated measurements. Communications in Statistics: Theory and Methods, 38(1):138–155, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinha:2006:UED

[SSM06] Bikas K. Sinha, Samindranath Sengupta, and Sujay Mukhuti. Unbiased estimation of the distribution function of an exponential population using order statistics with application in ranked set sampling. Communications in Statistics: Theory and Methods, 35(9):1655–1670, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sunoj:2009:CLD

[SSM09] S. M. Sunoj, P. G. Sankaran, and S. S. Maya. Characterizations of life distributions using conditional expectations of doubly (interval) truncated random variables. Communications in Statistics: Theory and Methods, 38(9):1441–1452, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2000:TUS

[SSQ00] Harshinder Singh, R. S. Singh, and Shixian Qian. Testing uniform stochastic ordering for survival functions. Communications in Statistics: Theory and Methods, 29(7):1649– 1661, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2009:ERT

[SST^{+}09] Sarjinder Singh, Housila P. Singh, Rajesh Tailor, Jack Allen, and Marcin Kozak. Estimation of ratio of two finitepopulation means in the presence of non response. Communications in Statistics: Theory and Methods, 38(19):3608– 3621, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shi:2000:NME

[SSW00] Dylan Shi, Junjie Song, and Chi Wang. A note on maximum entropy spectrum estimation. Communications in Statistics: Theory and Methods, 29(5–6):1311–1325, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samuel:2000:IFR

[ST00] Philip Samuel and P. Yageens Thoma. An improved form of a recurrence relation on the product moments of order statistics. Communications in Statistics: Theory and Methods, 29 (7):1559–1564, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Senoglu:2001:AVE

[ST01a] Birdal Senoglu and Moti L. Tiku. Analysis of variance in experimental design with nonnormal error distributions. Communications in Statistics: Theory and Methods, 30(7):1335– 1352, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Song:2001:ECO

[ST01b] Seuck Heun Song and Götz Trenkler. On the efficiency of the Cochrane–Orcutt estimator in the serially correlated error components regression model for panel data. Communications in Statistics: Theory and Methods, 30(2):195–207, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:2001:EPE

[ST01c] Rakesh Srivastava and Vilpa Tanna. An estimation procedure for error variance incorporating pts for random effects model under LINEX loss function. Communications in Statistics: Theory and Methods, 30(12):2583–2599, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sajeevkumar:2005:AOS

[ST05a] N. K. Sajeevkumar and P. Yageen Thomas. Applications of order statistics of independent nonidentically distributed random variables in estimation. Communications in Statistics: Theory and Methods, 34(4):775–783, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shalabh:2005:ELR

[ST05b] Shalabh and H. Toutenburg. Estimation of linear regression models with missing data: The role of stochastic linear constraints. Communications in Statistics: Theory and Methods, 34(2):375–387, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:2007:DSS

[ST07] Rakesh Srivastava and Vilpa Tanna. Double stage shrinkage testimator of the scale parameter of an exponential life model under general entropy loss function. Communications in Statistics: Theory and Methods, 36(2):283–295, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sreekumar:2008:EPT

[ST08] N. V. Sreekumar and P. Yageen Thomas. Estimation of the parameters of Type-I generalized logistic distribution using order statistics. Communications in Statistics: Theory and Methods, 37(10):1506–1524, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stadje:2009:QST

[Sta09] Wolfgang Stadje. A queueing system with two parallel lines, cost-conscious customers, and jockeying. Communications in Statistics: Theory and Methods, 38(16–17):3158–3169, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stepniak:2007:SCL

[Stȩ07] Czesław Stȩpniak. Simultaneous canonization of linear models. Communications in Statistics: Theory and Methods, 36 (13):2405–2412, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stylianou:2009:DED

[Sty09] Mario Stylianou. Derivation of exact distributions following an up-and-down design. Communications in Statistics: Theory and Methods, 38(16–17):3190–3197, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sultan:2007:OSG

[Sul07] Khalaf S. Sultan. Order statistics from the generalized exponential distribution and applications. Communications in Statistics: Theory and Methods, 36(7):1409–1418, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sungur:2002:SRT

[Sun02] Engin A. Sungur. Some results on truncation dependence invariant class of copulas. Communications in Statistics: Theory and Methods, 31(8):1399–1422, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sungur:2005:NDD

[Sun05a] Engin A. Sungur. A note on directional dependence in regression setting. Communications in Statistics: Theory and Methods, 34(9–10):1957–1965, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See remark [Mud08].

Sungur:2005:SOC

[Sun05b] Engin A. Sungur. Some observations on copula regression functions. Communications in Statistics: Theory and Methods, 34(9–10):1967–1978, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Surucu:2006:GFT

[Sür06] Baris Sürücü. Goodness-of-fit tests for multivariate distributions. Communications in Statistics: Theory and Methods, 35(7):1319–1331, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2008:FEP

[SV08a] Housila P. Singh and Gajendra K. Vishwakarma. A family of estimators of population mean using auxiliary information in stratified sampling. Communications in Statistics: Theory and Methods, 37(7):1038–1050, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Szatmari-Voicu:2008:MAM

[SV08b] D. Szatmari-Voicu. Minimax asymptotic mean-squared-error of L-estimators of scale parameter. Communications in Statistics: Theory and Methods, 37(5):709–721, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2009:GPE

[SV09] Housila P. Singh and Gajendra K. Vishwakarma. A general procedure for estimating population mean in successive sampling. Communications in Statistics: Theory and Methods, 38(2):293–308, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Szatmari-Voicu:2007:ESM

[SVC07] D. Szatmari-Voicu and J. R. Collins. M-estimators of scale with minimum gross errors sensitivity. Communications in Statistics: Theory and Methods, 36(11):2037–2048, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2007:NAD

[SvDM07] Amita Singh, J. René van Dorp, and Thomas A. Mazzuchi. A novel asymmetric distribution with power tails. Communications in Statistics: Theory and Methods, 36(2):235–252, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:2002:SVS

[SW02] Viren K. Srivastava and Alan T. K. Wan. Separate versus system methods of Stein-rule estimation in seemingly unrelated regression models. Communications in Statistics: Theory and Methods, 31(11):2077–2099, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shi:2003:MUM

[SW03] Ning-Zhong Shi and Dehui Wang. Median unbiased and maximum likelihood estimations of ARCH(0, 1) coefficient. Communications in Statistics: Theory and Methods, 32(5): 1057–1066, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Seshadri:2004:MDR

[SW04] Vanamamalai Seshadri and Jacek Wesołowski. Martingales defined by reciprocals of sums and related characterizations. Communications in Statistics: Theory and Methods, 33(12): 2993–3007, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schick:2009:IDE

[SW09a] Anton Schick and Wolfgang Wefelmeyer. Improved density estimators for invertible linear processes. Communications in Statistics: Theory and Methods, 38(16–17):3123– 3147, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Solanky:2009:AOS

[SW09b] Tumulesh K. S. Solanky and Yuefeng Wu. On approximate optimality of the sample size for the partition problem. Communications in Statistics: Theory and Methods, 38(16– 17):3148–3157, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swift:2009:CCI

[Swi09] Michael Bruce Swift. Comparison of confidence intervals for a Poisson mean — further considerations. Communications in Statistics: Theory and Methods, 38(5):748–759, 2009. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2008:IPW

[SY08] Pao-Sheng Shen and Ya-Fang Yan. An inverse-probability-weighted approach to the estimation of successive duration times under left-truncation and right-censoring. Communications in Statistics: Theory and Methods, 37(16):2595– 2605, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sun:2002:CRE

[SYCZ02] Xiaoqian Sun, Jinhong You, Gemai Chen, and Xian Zhou. Convergence rates of estimators in partial linear regression models with MA(∞) error process. Communications in Statistics: Theory and Methods, 31(12):2251–2273, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sykes:2000:REF

[Syk00] Alan M. Sykes. Ratio estimates and Fieller’s theorem in regression modelling. Communications in Statistics: Theory and Methods, 29(9–10):2055–2063, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610920008832595.

Savani:2006:EEP

[SZ06] V. Savani and A. A. Zhigljavsky. Efficient estimation of parameters of the negative binomial distribution. Communications in Statistics: Theory and Methods, 35(5):767–783, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tavangar:2007:GPD

[TA07] M. Tavangar and M. Asadi. Generalized Pareto distributions characterized by generalized order statistics. Communications in Statistics: Theory and Methods, 36(7):1333–1341, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tavangar:2008:CGP

[TA08] Mahdi Tavangar and Majid Asadi. On a characterization of generalized Pareto distribution based on generalized order statistics. Communications in Statistics: Theory and Methods, 37(9):1347–1352, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Takeuchi:2001:LRT

[Tak01] Hironori Takeuchi. On the likelihood ratio test for a single model against the mixture of two known densities. Communications in Statistics: Theory and Methods, 30(5):931–942, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Takada:2009:SBC

[Tak09] Yoshikazu Takada. Selecting the best component of a multivariate normal population. Communications in Statistics: Theory and Methods, 38(16–17):3198–3212, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanizaki:2000:NNG

[Tan00] Hisashi Tanizaki. Nonlinear and non-Gaussian state estimation: A quasi-optimal estimator. Communications in Statistics: Theory and Methods, 29(12):2805–2834, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanaka:2003:RBF

[Tan03] Hidekazu Tanaka. On a relation between a family of distributions attaining the Bhattacharyya bound and that of linear combinations of the distributions from an exponential family. Communications in Statistics: Theory and Methods, 32 (10):1885–1896, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanaka:2006:LSP

[Tan06] Hidekazu Tanaka. Location and scale parameter family of distributions attaining the Bhattacharyya bound. Communications in Statistics: Theory and Methods, 35(9):1611–1628, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tarsitano:2005:EGL

[Tar05] Agostino Tarsitano. Estimation of the generalized Lambda distribution parameters for grouped data. Communications in Statistics: Theory and Methods, 34(8):1689–1709, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Timm:2001:TMS

[TAS01] Neil H. Timm and Ali A. Al-Subaihi. Testing model specification in seemingly unrelated regression models. Communications in Statistics: Theory and Methods, 30(4):579–590, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taufer:2009:WSR

[Tau09] Emanuele Taufer. Wilcoxon-signed rank test for long memory sequences. Communications in Statistics: Theory and Methods, 38(16–17):3240–3248, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Torabi:2007:FLA

[TB07] Hamzeh Torabi and Javad Behboodian. Fuzzy leastabsolutes estimates in linear models. Communications in Statistics: Theory and Methods, 36(10):1935–1944, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tebbs:2003:EBG

[TBM03] Joshua M. Tebbs, Christopher R. Bilder, and Barry K. Moser. An empirical Bayes group-testing approach to estimating small proportions. Communications in Statistics: Theory and Methods, 32(5):983–995, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsou:2004:PRR

[TC04] Tsung-Shan Tsou and Kuang-Fu Cheng. Parametric robust regression analysis of contaminated data. Communications in Statistics: Theory and Methods, 33(8):1887–1898, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tolusso:2009:RES

[TC09] David Tolusso and Richard J. Cook. Robust estimation of state occupancy probabilities for interval-censored multistate data: An application involving spondylitis in psoriatic arthritis. Communications in Statistics: Theory and Methods, 38(18):3307–3325, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tchumtchoua:2007:BES

[TD07] Sylvie Tchumtchoua and Dipak K. Dey. Bayesian estimation of stochastic frontier models with multivariate skew t error terms. Communications in Statistics: Theory and Methods, 36(5):907–916, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Terpstra:2000:MEA

[Ter00] Jeffrey T. Terpstra. The median estimate of the autoregressive location parameter. Communications in Statistics: Theory and Methods, 29(8):1725–1739, 2000. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Terpstra:2001:MEA

[Ter01] Jeffrey T. Terpstra. The median estimate of the autoregressive location parameter. Communications in Statistics: Theory and Methods, 30(4):785–786, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Terpstra:2002:GSR

[Ter02] Jeffrey T. Terpstra. Generalized signed-rank estimators for autoregression parameters. Communications in Statistics: Theory and Methods, 31(3):367–395, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tonda:2004:AEN

[TF04] Tetsuji Tonda and Yasunori Fujikoshi. Asymptotic expansion of the null distribution of LR statistic for multivariate linear hypothesis when the dimension is large. Communications in Statistics: Theory and Methods, 33(5):1205–1220, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tom:2009:SMI

[TF09] Brian D. M. Tom and Vernon T. Farewell. Statistical methods for individual-level data in cohort mortality studies of rheumatic diseases. Communications in Statistics: Theory and Methods, 38(18):3472–3487, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tian:2006:ODC

[TH06] Yongge Tian and Agnes M. Herzberg. D-optimal designs for a combined simple linear and Haar-wavelet regression model. Communications in Statistics: Theory and Methods, 35(6):1085–1091, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tolikas:2008:ADG

[TH08] Konstantinos Tolikas and Saeed Heravi. The Anderson– Darling goodness-of-fit test statistic for the three-parameter lognormal distribution. Communications in Statistics: Theory and Methods, 37(19):3135–3143, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tinsson:2001:PVM

[Tin01] Walter Tinsson. Prediction of the variations of the mean response by using experimental design with quantitative factors and random block effects. Communications in Statistics: Theory and Methods, 30(2):209–228, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tinsson:2002:PEF

[Tin02] W. Tinsson. Parameter estimation for the fixed block effects model using a usual design. Communications in Statistics: Theory and Methods, 31(6):995–1017, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ting:2009:PSC

[Tin09] Naitee Ting. Practical and statistical considerations in designing an early phase II osteoarthritis clinical trial: A case study. Communications in Statistics: Theory and Methods, 38(18):3282–3296, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:2001:NRI

[TIS01] Moti L. Tiku, M. Qamarul Islam, and A. Sevtap Selçuk. Nonnormal regression. II. Symmetric distributions. Communications in Statistics: Theory and Methods, 30(6):1021– 1045, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiwari:2001:RSS

[TK01] Ram C. Tiwari and Paul H. Kvam. Ranked set sampling from location-scale families of symmetric distributions. Communications in Statistics: Theory and Methods, 30(8–9):1641– 1659, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tommasi:2004:MDP

[TLF04] Chiara Tommasi and Jesús López-Fidalgo. Minimax designs for a parametrization of binary response models. Communications in Statistics: Theory and Methods, 33(11):2787– 2798, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tong:2008:RED

[TLW08] Tiejun Tong, Anna Liu, and Yuedong Wang. Relative errors of difference-based variance estimators in nonparametric regression. Communications in Statistics: Theory and Methods, 37(18):2890–2902, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tian:2003:LRT

[TM03] Lili Tian and Govind S. Mudholkar. The likelihood ratio tests for homogeneity of inverse Gaussian means under simple order and simple tree order. Communications in Statistics: Theory and Methods, 32(4):791–805, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tian:2007:TEO

[TM07] Lili Tian and Govind S. Mudholkar. Testing equality of order constrained inverse Gaussian means. Communications in Statistics: Theory and Methods, 36(5):897–906, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taye:2008:MFV

[TN08] Girma Taye and Peter Njuho. Monitoring field variability using confidence interval for coefficient of variation. Communications in Statistics: Theory and Methods, 37(6):831–846, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Toma:2003:REP

[Tom03] Aida Toma. Robust estimators for the parameters of multivariate lognormal distribution. Communications in Statistics: Theory and Methods, 32(7):1405–1417, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tong:2002:GRE

[Ton02] Hengqing Tong. Generalized ridge estimate of parameters in variance component model. Communications in Statistics: Theory and Methods, 31(1):119–128, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tong:2005:SMH

[Ton05] Y. L. Tong. A sequential multi-hypothesis test for the mean function of a discrete-time Gaussian process. Communications in Statistics: Theory and Methods, 34(6):1349–1362, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taniguchi:2009:PTE

[TOS09] Masanobu Taniguchi, Hiroaki Ogata, and Hiroshi Shiraishi. Preliminary test estimation for regression models with longmemory disturbance. Communications in Statistics: Theory and Methods, 38(16–17):3213–3224, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tripsiannis:2002:MGP

[TPP02] Gregory A. Tripsiannis, Andreas N. Philippou, and Afroditi A. Papathanasiou. Multivariate generalized Pólya distribution of order k. Communications in Statistics: Theory and Methods, 31(11):1899–1912, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tripsiannis:2003:MGD

[TPP03] Gregory A. Tripsiannis, Andreas N. Philippou, and Afroditi A. Papathanasiou. Multivariate generalized distributions of order k. Communications in Statistics: Theory and Methods, 32(9):1725–1735, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Thavaneswaran:2008:DKO

[TPS08] A. Thavaneswaran, S. Peiris, and Jagbir Singh. Derivation of kurtosis and option pricing formulas for popular volatility models with applications in finance. Communications in Statistics: Theory and Methods, 37(11):1799–1814, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Triantafyllopoulos:2007:CDT

[Tri07a] K. Triantafyllopoulos. Convergence of discount time series dynamic linear models. Communications in Statistics: Theory and Methods, 36(11):2117–2127, 2007. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Triantafyllopoulos:2007:NCS

[Tri07b] K. Triantafyllopoulos. A note on the canonical structure of multivariate dynamic linear models. Communications in Statistics: Theory and Methods, 36(3):563–565, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Triantafyllopoulos:2008:RPM

[Tri08] K. Triantafyllopoulos. Reference priors for matrix-variate dynamic linear models. Communications in Statistics: Theory and Methods, 37(6):947–958, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Towhidi:2007:NGF

[TS07] M. Towhidi and M. Salmanpour. A new goodness-of-fit test based on the empirical characteristic function. Communications in Statistics: Theory and Methods, 36(15):2777–2785, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taneichi:2008:ADT

[TS08a] Nobuhiro Taneichi and Yuri Sekiya. Approximations of the distributions of test statistics for homogeneity of a product multinomial model. Communications in Statistics: Theory and Methods, 37(10):1610–1631, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsurumi:2008:BIR

[TS08b] Hiroki Tsurumi and Haruhiko Shimizu. Bayesian inference of the regression model with error terms following the exponential power distribution and unbiasedness of the LAD estimator. Communications in Statistics: Theory and Methods, 37(2):221–232, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tasdan:2009:SMW

[TS09] Feridun Tasdan and Gerald Sievers. Smoothed Mann– Whitney–Wilcoxon procedure for two-sample location problem. Communications in Statistics: Theory and Methods, 38(6):856–870, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsai:2000:RWT

[Tsa00] Kao-Tai Tsai. Robust Williams trend test. Communications in Statistics: Theory and Methods, 29(5–6):1327–1346, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsionas:2000:PAP

[Tsi00] Efthymios G. Tsionas. Posterior analysis, prediction and reliability in three-parameter Weibull distributions. Communications in Statistics: Theory and Methods, 29(7):1435– 1449, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsionas:2001:BIB

[Tsi01a] Efthymios G. Tsionas. Bayesian inference in Birnbaum– Saunders regression. Communications in Statistics: Theory and Methods, 30(1):179–193, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsionas:2001:BMP

[Tsi01b] Efthymios G. Tsionas. Bayesian multivariate Poisson regression. Communications in Statistics: Theory and Methods, 30(2):243–255, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsionas:2001:EIF

[Tsi01c] Efthymios G. Tsionas. Exact inference in four-parameter generalized gamma distributions. Communications in Statistics: Theory and Methods, 30(4):747–756, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsionas:2002:BAF

[Tsi02] Efthymios G. Tsionas. Bayesian analysis of finite mixtures of Weibull distributions. Communications in Statistics: Theory and Methods, 31(1):37–48, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsionas:2003:PRB

[Tsi03] Efthymios G. Tsionas. Pareto regression: A Bayesian analysis. Communications in Statistics: Theory and Methods, 32 (6):1213–1225, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsionas:2008:BIG

[Tsi08] Efthymios G. Tsionas. Bayesian inference in generalized error and generalized Student-t regression models. Communications in Statistics: Theory and Methods, 37(3):388–407, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsionas:2009:BIU

[Tsi09] Efthymios G. Tsionas. Bayesian inference using artificial augmenting regressions. Communications in Statistics: Theory and Methods, 38(9):1361–1370, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsou:2003:CTP

[Tso03] Tsung-Shan Tsou. Comparing two population means and variances: A parametric robust way. Communications in Statistics: Theory and Methods, 32(10):2013–2029, 2003.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsou:2005:RIC

[Tso05] Tsung-Shan Tsou. Robust inference for the correlation coefficient — a parametric method. Communications in Statistics: Theory and Methods, 34(1):147–162, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsou:2009:PLP

[Tso09a] Tsung-Shan Tsou. Performing legitimate parametric regression analysis without knowing the true underlying random mechanisms. Communications in Statistics: Theory and Methods, 38(10):1680–1689, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsou:2009:RST

[Tso09b] Tsung-Shan Tsou. A robust score test for testing several coefficients of variation with unknown underlying distributions. Communications in Statistics: Theory and Methods, 38(9):1350–1360, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taneichi:2002:FNT

[TSSI02] Nobuhiro Taneichi, Yuri Sekiya, Akio Suzukawa, and Hideyuki Imai. A formula for normalizing transformation of some statistics. Communications in Statistics: Theory and Methods, 31(2):163–179, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsay:2003:SGR

[TT03] Jyh-Jian Tsay and C. Andy Tsao. Statistical Gambler’s ruin problem. Communications in Statistics: Theory and Methods, 32(7):1337–1359, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tian:2007:SDW

[TT07] Yongge Tian and Yoshio Takane. On sum decompositions of weighted least-squares estimators for the partitioned linear model. Communications in Statistics: Theory and Methods, 37(1):55–69, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tekle:2008:HED

[TTB08] Fetene B. Tekle, Frans E. S. Tan, and Martijn P. F. Berger. Highly efficient designs for logistic models with categorical variables. Communications in Statistics: Theory and Methods, 37(5):746–759, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tan:2001:EBS

[TTX01] Ming Tan, Guo-Liang Tian, and Xiaoping Xiong. Explicit Bayesian solution for incomplete pre-post test problems using inverse Bayes formulae. Communications in Statistics: Theory and Methods, 30(6):1111–1129, 2001. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tu:2001:RST

[Tu01] Dongsheng Tu. Rao’s score tests in survival analysis: Examples and Bartlett type adjustments. Communications in Statistics: Theory and Methods, 30(8–9):1521–1532, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Torigoe:2006:RLE

[TU06] Norio Torigoe and Katsumi Ujiie. On the restricted Liu estimator in the Gauss–Markov model. Communications in Statistics: Theory and Methods, 35(9):1713–1722, 2006. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Teles:2008:TUR

[TWH08] Paulo Teles, William W. S. Wei, and Erin M. Hodgess. Testing a unit root based on aggregate time series. Communications in Statistics: Theory and Methods, 37(4):565–590, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tang:2001:LIN

[TWW01] Nian-Sheng Tang, Bo-Cheng Wei, and Xue-Ren Wang. Local influence in nonlinear reproductive dispersion models. Communications in Statistics: Theory and Methods, 30(3): 435–449, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tang:2003:SAW

[TXG03] Y. Tang, M. Xie, and T. N. Goh. Statistical analysis of a Weibull extension model. Communications in Statistics: Theory and Methods, 32(5):913–928, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tan:2000:SSS

[TY00] Wai-Yuan Tan and Zhengzheng Ye. Some state space models of HIV epidemic and its applications for the estimation of HIV infection and incubation. Communications in Statistics: Theory and Methods, 29(5–6):1059–1088, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanizaki:2001:PAM

[TZ01] Hisashi Tanizaki and Xingyun Zhang. Posterior analysis of the multiplicative heteroscedasticity model. Communications in Statistics: Theory and Methods, 30(5):855–874, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tzavelas:2008:SCE

[Tza08] George Tzavelas. Sufficient conditions for the existence of a solution for the log-likelihood equations in the threeparameter gamma distribution. Communications in Statistics: Theory and Methods, 37(9):1371–1382, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tzavelas:2009:ETP

[Tza09] George Tzavelas. Estimation in the three-parameter gamma distribution based on the profile log-likelihood function. Communications in Statistics: Theory and Methods, 38(5): 573–583, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanaka:2002:LIC

[TZY02] Y. Tanaka, F. Zhang, and W. S. Yang. On local influence in canonical correlation analysis. Communications in Statistics: Theory and Methods, 31(12):2325–2347, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Uddin:2005:RFE

[UNN05] Md. Kamal Uddin and U. V. Naik-Nimbalkar. Regression function estimation using spline wavelets. Communications in Statistics: Theory and Methods, 34(4):823–832, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Uribe-Opazo:2008:IST

[UOFC08] Miguel A. Uribe-Opazo, Silvia L. P. Ferrari, and Gauss M. Cordeiro. Improved score tests in symmetric linear regression models. Communications in Statistics: Theory and Methods, 37(2):261–276, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Upadhyay:2003:CBW

[UP03] S. K. Upadhyay and M. Peshwani. Choice between Weibull and lognormal models: A simulation based Bayesian study. Communications in Statistics: Theory and Methods, 32(2): 381–405, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ula:2003:IPM

[US03] Taylan A. Ula, Ph.D. and Abdullah A. Smadi, Ph.D. Identification of periodic moving-average models. Communications in Statistics: Theory and Methods, 32(12):2465–2475, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Usami:2002:SAP

[Usa02] Yoshihiro Usami. Some asymptotic properties of the least squares estimators of a polynomial regression with a heteroskedastic error. Communications in Statistics: Theory and Methods, 31(4):625–637, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Usami:2001:EER

[UT01] Yoshihiro Usami and Yasuyuki Toyooka. An examination of estimated residuals in a regression with an infinite order parametric model. Communications in Statistics: Theory and Methods, 30(10):1989–2004, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ula:2004:MML

[UY04] Taylan A. Ula and Ceylan Yozgatligil. Modified maximumlikelihood method for non-normal time series revisited. Communications in Statistics: Theory and Methods, 33(2):397– 417, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vasdekis:2008:CRC

[Vas08] Vassilis G. S. Vasdekis. A comparison of REML and covariance adjustment method in the estimation of growth curve models. Communications in Statistics: Theory and Methods, 37(20):3287–3297, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vaughan:2002:GSH

[Vau02] David C. Vaughan. The generalized secant hyperbolic distribution and its properties. Communications in Statistics: Theory and Methods, 31(2):219–238, 2002. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vasconcellos:2000:BCE

[VC00] Klaus L. P. Vasconcellos and Gauss M. Cordeiro. Bias corrected estimates in multivariate Student t regression models. Communications in Statistics: Theory and Methods, 29(4): 797–822, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

vanDorp:2003:GTS

[vDK03] J. René van Dorp and Samuel Kotz. Generalizations of twosided power distributions and their convolution. Communications in Statistics: Theory and Methods, 32(9):1703–1723, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

vanderMeulen:2008:NNV

[vdM08] Egbert A. van der Meulen. A nonrandomized, nonconservative version of the Fisher exact test. Communications in Statistics: Theory and Methods, 37(5):699–708, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vexler:2008:MTS

[Vex08] Albert Vexler. Martingale type statistics applied to change points detection. Communications in Statistics: Theory and Methods, 37(8):1207–1224, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vexler:2006:GLM

[VG06] A. Vexler and G. Gurevich. Guaranteed local maximum likelihood detection of a change point in nonparametric logistic regression. Communications in Statistics: Theory and Methods, 35(4):711–726, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vexler:2009:AMP

[VG09a] Albert Vexler and Gregory Gurevich. Average most powerful tests for a segmented regression. Communications in Statistics: Theory and Methods, 38(13):2214–2231, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

VillasenorAlva:2009:GSW

[VG09b] José A. Villasenor Alva and Elizabeth González Estrada. A generalization of Shapiro–Wilk’s test for multivariate normality. Communications in Statistics: Theory and Methods, 38(11):1870–1883, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vollandt:2001:SSD

[VH01] Rüdiger Vollandt and Manfred Horn. Sample size determination for multiple many-one and pairwise comparisons of proportions using the arcsin transformation. Communications in Statistics: Theory and Methods, 30(12):2481–2498, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vollandt:2004:NMS

[VH04] Rüdiger Vollandt and Manfred Horn. A nonparametric method for selecting a good population. Communications in Statistics: Theory and Methods, 33(7):1541–1555, 2004. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Volkmer:2007:CCS

[VH07] H. Volkmer and G. G. Hamedani. Characterization of continuous symmetric distributions based on families of skewed distributions. Communications in Statistics: Theory and Methods, 36(13):2329–2338, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vollandt:2000:SSD

[VHS00] Rüdiger Vollandt, Manfred Horn, and Pranab K. Sen. Sample size determination of Steel’s nonparametric many-one test. Communications in Statistics: Theory and Methods, 29 (12):2895–2914, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vinogradov:2004:PVF

[Vin04] Vladimir Vinogradov. On the power-variance family of probability distributions. Communications in Statistics: Theory and Methods, 33(5):1007–1029, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vinogradov:2005:E

[Vin05] Vladimir Vinogradov. Errata. Communications in Statistics: Theory and Methods, 33(10):2573, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vinogradov:2007:IDE

[Vin07] Vladimir Vinogradov. On infinitely divisible exponential dispersion model related to Poisson-exponential distribution. Communications in Statistics: Theory and Methods, 36(2): 253–263, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Visk:2009:PEA

[Vis09] Helle Visk. On the parameter estimation of the asymmetric multivariate Laplace distribution. Communications in Statistics: Theory and Methods, 38(4):461–470, 2009. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Verrill:2007:CBH

[VJ07] Steve Verrill and Richard A. Johnson. Confidence bounds and hypothesis tests for normal distribution coefficients of variation. Communications in Statistics: Theory and Methods, 36(12):2187–2206, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vijayakumar:2000:SCM

[VKT00] A. Vijayakumar, B. Krishna Kumar, and B. Thilaka. Stochastic compartmental models with branching particles and disasters: sojourn time and related characteristics. Communications in Statistics: Theory and Methods, 29(2):291– 318, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Villani:2006:MSN

[VL06] Mattias Villani and Rolf Larsson. The multivariate split normal distribution and asymmetric principal components analysis. Communications in Statistics: Theory and Methods, 35 (6):1123–1140, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vilca-Labra:2006:NFL

[VLLS06] Filidor Vilca-Labra and Víctor Leiva-Sánchez. A new fatigue life model based on the family of skew-elliptical distributions. Communications in Statistics: Theory and Methods, 35(2): 229–244, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Villegas:2008:BTM

[VPS08] Miguel A. Gómez Villegas, Javier Portela, and Luis Sanz. A Bayesian test for the mean of the power exponential distribution. Communications in Statistics: Theory and Methods, 37 (18):2865–2876, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vassiliou:2004:ABF

[VT04] P.-C. G. Vassiliou and Helen L. Tsakiridou. Asymptotic behavior of first passage probabilities in the perturbed nonhomogeneous semi-Markov systems. Communications in Statistics: Theory and Methods, 33(3):651–679, 2004. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vassiliou:2004:NSM

[VTKB04] A. Vassiliou, D. G. Tambouratzis, M. V. Koutras, and S. Bersimis. A new similarity measure and its use in determining the number of clusters in a multivariate data set. Communications in Statistics: Theory and Methods, 33(7):1643–1666, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vaghefi:2009:PIM

[VTN09] A. Vaghefi, Sam D. Tajbakhsh, and R. Noorossana. Phase II monitoring of nonlinear profiles. Communications in Statistics: Theory and Methods, 38(11):1834–1851, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2005:CRE

[Wan05a] Lichun Wang. Convergence rates of empirical Bayes estimation for the parameter of the uniform distribution U(0, θ) under random censorship. Communications in Statistics: Theory and Methods, 34(12):2209–2220, 2005. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2005:AEM

[Wan05b] Lihong Wang. Asymptotics of L_{1}-estimators in moving average time series models. Communications in Statistics: Theory and Methods, 33(1):107–118, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2006:MVT

[Wan06a] Hsiuying Wang. Modified p-value of two-sided test for normal distribution with restricted parameter space. Communications in Statistics: Theory and Methods, 35(8):1361–1374, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2006:APA

[Wan06b] Xiaogang Wang. Asymptotic properties of adaptive likelihood weights by cross-validation. Communications in Statistics: Theory and Methods, 35(7):1257–1270, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2007:AEP

[Wan07a] P. C. Wang. Analysis of experimental Poisson data. Communications in Statistics: Theory and Methods, 36(6):1059– 1068, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2007:EBI

[Wan07b] Qihua Wang. M-estimators based on inverse probability weighted estimating equations with response missing at random. Communications in Statistics: Theory and Methods, 36(6):1091–1103, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2008:IES

[Wan08] Bingxing Wang. Interval estimation for the scaled halflogistic distribution under progressive Type-II censoring. Communications in Statistics: Theory and Methods, 38(3): 364–371, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2009:CEW

[Wan09a] Zhu Wang. Correction to “Estimation of the Weibull Parameters with Grouped Data”. Communications in Statistics: Theory and Methods, 38(20):3755, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [CC88].

Wang:2009:FSP

[Wan09b] Zijun Wang. Finite sample performances of the model selection approach in nonparametric model specification for time series. Communications in Statistics: Theory and Methods, 38(14):2302–2320, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2000:MDM

[WC00] Dong Q. Wang and F. Critchley. Multiple deletion measures and conditional influence in regression model. Communications in Statistics: Theory and Methods, 29(11):2391–2404, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2007:EBP

[WC07] Xiaoming Wang and K. C. Carriere. Exponential-bound property of estimators and variable selection in generalized additive models. Communications in Statistics: Theory and Methods, 36(6):1105–1122, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2008:HTP

[WC08] Chien-Wei Wu and Y. C. Chang. A hypothesis testing procedure on assessing process performance for asymmetric tolerances. Communications in Statistics: Theory and Methods, 37(12):1959–1976, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2004:DAP

[WCL04] Dong Q. Wang, F. Critchley, and Ivy Liu. Diagnostics analysis and perturbations in a clustered sampling model. Communications in Statistics: Theory and Methods, 33(11):2709– 2721, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2008:IMQ

[WCL08] Xiaoxia Wu, Hanfeng Chen, and Yanyan Liu. Interval mapping for quantitative trait loci detection in finite mixture models with general kernel functions. Communications in Statistics: Theory and Methods, 37(5):803–814, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2003:MSD

[WCS03] Dong Q. Wang, F. Critchley, and Peter J. Smith. The multiple sets of deletion measures and masking in regression. Communications in Statistics: Theory and Methods, 32(2): 407–413, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2000:DPC

[WCT00] Chinying Wang, T. Timothy Chen, and Irvin Tyan. Designs for Phase I cancer clinical trials with differentiation of graded toxicity. Communications in Statistics: Theory and Methods, 29(5–6):975–987, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

William:2004:MCR

[WD04] Martin L. William and T. M. Durairajan. Minimum convex risk equivariant finite population prediction for linear functions in regression models. Communications in Statistics: Theory and Methods, 33(5):1171–1185, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Whitney:2000:PEH

[WDRB00] James E. Whitney II, Kerron Duncan, Maria Richardson, and Isaac Bankman. Parameter estimation in a highly non-linear model using simultaneous perturbation stochastic approximation. Communications in Statistics: Theory and Methods, 29(5–6):1247–1256, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weiss:2009:NCA

[Wei09] Christian H. Weiß. A new class of autoregressive models for time series of binomial counts. Communications in Statistics: Theory and Methods, 38(4):447–460, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wenzel:2002:PCM

[Wen02] Thomas Wenzel. Pitman-closeness as a measure to evaluate the quality of forecasts. Communications in Statistics: Theory and Methods, 31(4):535–550, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Watamori:2005:CRM

[WF05] Yoko Watamori and Teruo Fujioka. Confidence regions for the mean direction of the von Mises–Fisher distribution. Communications in Statistics: Theory and Methods, 34(3): 671–678, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Woldie:2001:PFI

[WFC01] Mammo Woldie, J. Leroy Folks, and John P. Chandler. Power function for inverse Gaussian regression models. Communications in Statistics: Theory and Methods, 30(5):787– 797, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2008:EML

[WFZ08] Ping Wu, Yun Fang, and Li-Xing Zhu. Estimating moments in linear mixed models. Communications in Statistics: Theory and Methods, 37(16):2582–2594, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wiper:2008:OBI

[WGP08] M. P. Wiper, F. J. Girón, and Arthur Pewsey. Objective Bayesian inference for the half-normal and half- t distributions. Communications in Statistics: Theory and Methods, 37(20):3165–3185, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wada:2000:RAT

[WH00] Cicilia Y. Wada and Luiz K. Hotta. Restricted alternatives tests in a bivariate exponential model with covariates. Communications in Statistics: Theory and Methods, 29(1): 193–210, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2008:OLA

[WH08] Molin Wang and John J. Hanfelt. Orthogonal locally ancillary estimating functions with application to stratified clustered data. Communications in Statistics: Theory and Methods, 37(18):2836–2853, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2003:MLO

[WHZ03] Yao Wang, Anwar M. Hossain, and William J. Zimmer. Monotone log-odds rate distributions in reliability analysis. Communications in Statistics: Theory and Methods, 32 (11):2227–2244, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Willink:2003:RBC

[Wil03] R. Willink. Relationships between central moments and cumulants, with formulae for the central moments of gamma distributions. Communications in Statistics: Theory and Methods, 32(4):701–704, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Willink:2005:BBN

[Wil05a] R. Willink. Bounds on the bivariate normal distribution function. Communications in Statistics: Theory and Methods, 33(10):2281–2297, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Willink:2005:CIT

[Wil05b] R. Willink. A confidence interval and test for the mean of an asymmetric distribution. Communications in Statistics: Theory and Methods, 34(4):753–766, 2005. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Willink:2007:CIM

[Wil07] R. Willink. On a confidence interval for the mean of an asymmetric distribution. Communications in Statistics: Theory and Methods, 36(12):2235–2236, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Williamson:2009:BAT

[Wil09] Patricia Pepple Williamson. Bayesian alternatives to the F test for two population variances. Communications in Statistics: Theory and Methods, 38(14):2288–2301, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2000:MBB

[WJ00] Qi-Hua Wang and Bing-Yi Jing. A martingale-based bootstrap inference with censored data. Communications in Statistics: Theory and Methods, 29(2):401–415, 2000. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2006:EEB

[WJ06] Qihua Wang and Bing-Yi Jing. Edgeworth expansion and bootstrap approximation for Studentized product-limit estimator with truncated and censored data. Communications in Statistics: Theory and Methods, 35(4):609–623, 2006. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Watnik:2001:NLM

[WJB01] Mitchell Watnik, Wesley Johnson, and Edward J. Bedrick. Nonnested linear model selection revisited. Communications in Statistics: Theory and Methods, 30(1):1–20, 2001. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2000:TSP

[WL00] Ji-Cheng Wu and Lii-Yuh Leu. A two-stage procedure with accelerated test for selecting the best exponential population. Communications in Statistics: Theory and Methods, 29(2):369–380, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wen:2001:ABE

[WL01a] Daling Wen and Martin S. Levy. Admissibility of Bayes estimates under BLINEX loss for the normal mean problem. Communications in Statistics: Theory and Methods, 30(1): 155–163, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wen:2001:BBA

[WL01b] Daling Wen and Martin S. Levy. BLINEX: a bounded asymmetric loss function with application to Bayesian estimation. Communications in Statistics: Theory and Methods, 30(1): 147–153, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2003:OPE

[WL03] Jong-Wuu Wu and Pai-Ling Li. Optimal parameter estimation of the extreme value distribution based on a Type II censored sample. Communications in Statistics: Theory and Methods, 32(3):533–554, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2005:ESF

[WL05] Jiantian Wang and Yingfu Li. Estimators for survival function when censoring times are known. Communications in Statistics: Theory and Methods, 34(2):449–459, 2005. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2008:SPD

[WL08] Kaiping Wang and Lu Lin. Semi-parametric density estimation for time-series with multiplicative adjustment. Communications in Statistics: Theory and Methods, 37(8):1274– 1283, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2009:MAM

[WLQ09] Kaiping Wang, Lu Lin, and Ruihua Qi. Multiplicative adjustment method for semiparametric regression with mixing dependent data. Communications in Statistics: Theory and Methods, 38(20):3654–3665, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wong:2009:P

[Won09] Weng Kee Wong, Guest Editor. Preface. Communications in Statistics: Theory and Methods, 38(18):3267, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wood:2002:AGF

[Woo02] G. R. Wood. Assessing goodness of fit for Poisson and negative binomial models with low mean. Communications in Statistics: Theory and Methods, 31(11):1977–2001, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wilson:2002:MCV

[WP02] Craig A. Wilson and Mark E. Payton. Modelling the coefficient of variation in factorial experiments. Communications in Statistics: Theory and Methods, 31(3):463–476, 2002. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2007:LBS

[WQC07] Zheng-Hong Wang, Hong Qin, and Kashinath Chatterjee. Lower bounds on the symmetric L_{2}-discrepancy and their application. Communications in Statistics: Theory and Methods, 36(13):2413–2423, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2000:SST

[WS00] Dai Wang and Fengzhu Sun. Sample sizes for the transmission disequilibrium tests: TDT, s-TDT and 1-TDT. Communications in Statistics: Theory and Methods, 29(5– 6):1129–1142, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Walter:2005:WLB

[WS05] Gilbert G. Walter and Xiaoping Shen. Wavelet like behavior of Slepian functions and their use in density estimation. Communications in Statistics: Theory and Methods, 34(3): 687–711, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Williams:2009:MVH

[WSB09] James D. Williams, Joe H. Sullivan, and Jeffrey B. Birch. Maximum value of Hotelling’s T^{2} statistics based on the successive differences covariance matrix estimator. Communications in Statistics: Theory and Methods, 38(4):471–483, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wesley:2008:ACI

[WSPP08] Wayne R. Wesley, James R. Simpson, Peter A. Parker, and Joseph R. Pignatiello, Jr. Analytical characterization of the information matrix for split-plot designs. Communications in Statistics: Theory and Methods, 38(1):90–102, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wesley:2009:MMP

[WSPP09] Wayne R. Wesley, James R. Simpson, Peter A. Parker, and Joseph J. Pignatiello, Jr. Maximum and minimum prediction variance for spherical and cuboidal split plot designs. Communications in Statistics: Theory and Methods, 38(13):2251– 2266, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2001:NEN

[WTZL01] Y. Wu, K. W. Tam, M. M. Zen, and F. Li. A note on estimating the number of superimposed exponential signals by the cross-validation approach. Communications in Statistics: Theory and Methods, 30(8–9):1579–1589, 2001. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2000:ESC

[Wu00] Jiang-Ming Johnny Wu. Extreme statistics from the cosine distribution. Communications in Statistics: Theory and Methods, 29(5–6):1397–1403, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2002:SST

[Wu02] Yanhong Wu. Supplementary score test in mixture model. Communications in Statistics: Theory and Methods, 31(5): 753–779, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2006:CUT

[Wu06] Samuel S. Wu. Combining univariate tests for multivariate location problem. Communications in Statistics: Theory and Methods, 35(8):1483–1494, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wulff:2005:BUE

[Wul05] Shaun S. Wulff. Best unbiased estimation in unbalanced split plot designs. Communications in Statistics: Theory and Methods, 34(3):631–649, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2006:ELS

[WV06] Lichun Wang and Noel Veraverbeke. Empirical likelihood in a semi-parametric model for missing response data. Communications in Statistics: Theory and Methods, 35(4):625–639, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2002:MKR

[WW02] Minta Wu and Robert F. Woolson. Modification of Kelly and Rice method for synergism tests. Communications in Statistics: Theory and Methods, 31(6):1019–1030, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2003:NDV

[WW03] J. Wu and A. C. M. Wong. A note on determining the pvalue of Bartlett’s test of homogeneity of variances. Communications in Statistics: Theory and Methods, 32(1):91–101, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2005:EMA

[WW05] Samuel S. Wu and Martin T. Wells. An extension of the Metropolis algorithm. Communications in Statistics: Theory and Methods, 34(3):585–596, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2007:HTR

[WW07] Shun-Fang Wang and Xue-Ren Wang. Homogeneity test of risk differences of marginal and conditional probabilities in several incomplete correlated 2 × 2 tables. Communications in Statistics: Theory and Methods, 36(16):2877–2890, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wawrzosek:2009:TEM

[WWJ09] Jacek Wawrzosek and Mirosława Wesołowska-Janczarek. Testability and estimability in multivariate linear normal model with various restrictions. Communications in Statistics: Theory and Methods, 38(6):828–841, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wei:2008:AER

[WWY08] Jiaqin Wei, Rongming Wang, and Dingjun Yao. The asymptotic estimate of ruin probability under a class of risk model in the presence of heavy tails. Communications in Statistics: Theory and Methods, 37(15):2331–2341, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2000:LSP

[WYY00] Colin O. Wu, Kai F. Yu, and Vivian W. S. Yuan. Large sample properties and confidence bands for component-wise varying-coefficient regression with longitudinal dependent variable. Communications in Statistics: Theory and Methods, 29(5–6):1017–1037, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2005:API

[WYZ05] Xianyi Wu, Jinhong You, and Xian Zhou. Asymptotic properties of the ISE in nonparametric regressions with serially correlated errors. Communications in Statistics: Theory and Methods, 34(4):943–953, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2001:EPL

[WZ01] Qihua Wang and Lixing Zhu. Estimation in partly linear error-in-covariable models with censored data. Communications in Statistics: Theory and Methods, 30(1):41–54, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wei:2007:SBL

[WZ07] Laisheng Wei and Weiping Zhang. The superiorities of Bayes linear minimum risk estimation in linear model. Communications in Statistics: Theory and Methods, 36(5):917–926, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xekalaki:2004:CBP

[XD04] Evdokia Xekalaki and Caterina Dimaki. Characterizations of bivariate Pareto and Yule distributions. Communications in Statistics: Theory and Methods, 33(12):3033–3042, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xue-Dong:2009:BAS

[XD09] Chen Xue-Dong. Bayesian analysis of semiparametric mixedeffects models for zero-inflated count data. Communications in Statistics: Theory and Methods, 38(11):1815–1833, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2009:APS

[XG09] Hongyan Xu and Varghese George. Assessment of population structure and its effects on genome-wide association studies. Communications in Statistics: Theory and Methods, 38(16– 17):2843–2855, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xia:2009:MSC

[XGZ09] Zhiming Xia, Pengjiang Guo, and Wenzhi Zhao. Monitoring structural changes in generalized linear models. Communications in Statistics: Theory and Methods, 38(11):1927–1947, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2008:TDT

[XK08] Lu Xu and John E. Kolassa. Testing the difference of two binomial proportions: Comparison of continuity corrections for saddlepoint approximation. Communications in Statistics: Theory and Methods, 37(14):2213–2218, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xia:2008:APM

[XKWW08] Tian Xia, Fanchao Kong, Shunfang Wang, and Xueren Wang. Asymptotic properties of the maximum quasilikelihood estimator in quasi-likelihood nonlinear models. Communications in Statistics: Theory and Methods, 37(15): 2358–2368, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2006:MCM

[XL06] Youzhi Xu and Jinlin Li. A Markov chain model for forecasting the management quality of Chinese a share listed companies. Communications in Statistics: Theory and Methods, 35 (10):1909–1919, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2009:GFL

[XL09] Wangli Xu and Yiqiang Lu. Goodness-of-fit for longitudinal count data with overdispersion. Communications in Statistics: Theory and Methods, 38(20):3745–3754, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2007:LRO

[XLZL07] Maochao Xu, Xiaohu Li, Peng Zhao, and Zhouping Li. Likelihood ratio order of m-spacings in multiple-outlier models. Communications in Statistics: Theory and Methods, 36 (8):1507–1525, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xiong:2000:CPS

[XM00] Chengjie Xiong and George A. Milliken. Change-points in stochastic ordering. Communications in Statistics: Theory and Methods, 29(2):381–400, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xiong:2008:NIE

[XM08] Shifeng Xiong and Weiyan Mu. Nonparametric interval estimation in one-way random-effects models. Communications in Statistics: Theory and Methods, 38(3):322–331, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xiong:2008:GIC

[XMX08] Shifeng Xiong, Weiyan Mu, and Xingzhong Xu. Generalized inference for a class of linear models under heteroscedasticity. Communications in Statistics: Theory and Methods, 37 (8):1225–1236, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xekalaki:2004:BDD

[XP04] Evdokia Xekalaki and John Panaretos. A binomial distribution with dependent trials and its use in stochastic model evaluation. Communications in Statistics: Theory and Methods, 33(12):3043–3058, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xie:2008:QLA

[XSD^{+}08] Changchun Xie, R. S. Singh, A. F. Desmond, Xuewen Lu, and Eric Ormsby. Quasi-likelihood approach to bioavailability and bioequivalence analysis. Communications in Statistics: Theory and Methods, 37(10):1641–1658, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2009:ALP

[Xu09] Liwen Xu. Admissible linear predictors in the superpopulation model with respect to inequality constraints. Communications in Statistics: Theory and Methods, 38(15):2528– 2540, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2008:MLE

[XW08a] Jing Xu and Jinde Wang. Maximum likelihood estimation of linear models for longitudinal data with inequality constraints. Communications in Statistics: Theory and Methods, 37(6):931–946, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2008:NGV

[XW08b] Liwen Xu and Songgui Wang. A new generalized p-value and its upper bound for ANOVA under unequal error variances. Communications in Statistics: Theory and Methods, 37(7):1002–1010, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xia:2007:CRS

[XWW07] Tian Xia, Shun-Fang Wang, and Xue-Ren Wang. Convergence rate of strong consistency of the maximum likelihood estimator in exponential family nonlinear models. Communications in Statistics: Theory and Methods, 36(1):103–115, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2007:ESL

[XY07a] Jianwen Xu and Hu Yang. Estimation in singular linear models with stochastic linear restrictions. Communications in Statistics: Theory and Methods, 36(10):1945–1951, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2007:CSL

[XY07b] Qinfeng Xu and Jinhong You. Covariate selection for linear errors-in-variables regression models. Communications in Statistics: Theory and Methods, 36(2):375–386, 2007. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xekalaki:2008:GWP

[XZ08] Evdokia Xekalaki and Mimoza Zografi. The generalized Waring process and its application. Communications in Statistics: Theory and Methods, 37(12):1835–1854, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yata:2009:PCN

[YA09] Kazuyoshi Yata and Makoto Aoshima. PCA consistency for non-Gaussian data in high dimension, low sample size context. Communications in Statistics: Theory and Methods, 38(16–17):2634–2652, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamada:2007:CEP

[Yam07] Takayuki Yamada. On comparisons of exact powers of bivariate GMANOVA tests. Communications in Statistics: Theory and Methods, 36(2):399–413, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yanagihara:2000:AEN

[Yan00a] Hirokazu Yanagihara. Asymptotic expansion of the null distribution of one-way ANOVA test statistic for heteroscedastic case under nonnormality. Communications in Statistics: Theory and Methods, 29(2):463–476, 2000. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2000:FNG

[Yan00b] Lijian Yang. Finite nonparametric GARCH model for foreign exchange volatility. Communications in Statistics: Theory and Methods, 29(5–6):1347–1365, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yan:2007:NCN

[Yan07] Liqing Yan. A note on comparison of (non) linear estimators using Pitman’s measure of closeness. Communications in Statistics: Theory and Methods, 36(15):2687–2694, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yao:2008:DSS

[YBCH08] Junchao Yao, Marco Burkschat, Huaihou Chen, and Taizhong Hu. Dependence structure of spacings of order statistics. Communications in Statistics: Theory and Methods, 37(15):2390–2403, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

You:2002:PEP

[YC02] Jinhong You and Gemai Chen. Parameter estimation in a partly linear regression model with random coefficient autoregressive errors. Communications in Statistics: Theory and Methods, 31(7):1137–1158, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

You:2003:JES

[YC03] Jinhong You and Gemai Chen. Jackknife estimation for smooth functions of the parametric component in partially linear regression models. Communications in Statistics: Theory and Methods, 32(9):1817–1833, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yue:2005:SMB

[YC05] Jack C. Yue and Murray K. Clayton. A similarity measure based on species proportions. Communications in Statistics: Theory and Methods, 34(11):2123–2131, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2009:FEM

[YC09] Jingyun Yang and Vernon M. Chinchilli. Fixed-effects modeling of Cohen’s kappa for bivariate multinomial data. Communications in Statistics: Theory and Methods, 38(20):3634– 3653, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2009:ILE

[YCL09] Hu Yang, Xinfeng Chang, and Deqiang Liu. Improvement of the Liu estimator in weighted mixed regression. Communications in Statistics: Theory and Methods, 38(2):285–292, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:2008:RBC

[YCW^{+}08] Qiqing Yu, Rick Chappell, G. Y. C. Wong, Yuting Hsu, and Marcin Mazur. Relationship between the Cox, Lehmann, Weibull, and accelerated lifetime models. Communications in Statistics: Theory and Methods, 37(9):1458–1470, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2000:EMM

[YD00] Yonghong (Fred) Yang and Norman R. Draper. The effects of model mis-specifications in linear measurement error models. Communications in Statistics: Theory and Methods, 29 (9–10):2123–2142, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/03610920008832599.

Yeh:2004:GMO

[Yeh04] Hsiaw-Chan Yeh. The generalized Marshall–Olkin type multivariate Pareto distributions. Communications in Statistics: Theory and Methods, 33(5):1053–1068, 2004. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yeh:2009:SRM

[Yeh09] Hsiaw-Chan Yeh. Some related minima stability and minima infinite divisibility of the general multivariate Pareto distributions. Communications in Statistics: Theory and Methods, 38(4):497–510, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Young:2009:AUS

[YGKD09] Linda J. Young, Carol A. Gotway, Greg Kearney, and Chris DuClos. Assessing uncertainty in support-adjusted spatial misalignment problems. Communications in Statistics: Theory and Methods, 38(16–17):3249–3264, 2009. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yi:2006:MBS

[YH06] Grace Y. Yi and Wenqing He. Methods for bivariate survival data with mismeasured covariates under an accelerated failure time model. Communications in Statistics: Theory and Methods, 35(8):1539–1554, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yao:2008:DSC

[YH08] Junchao Yao and Taizhong Hu. Dependence structure of a class of symmetric distributions. Communications in Statistics: Theory and Methods, 37(9):1338–1346, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yi:2001:ASO

[Yi01] Seongbaek Yi. Asymptotic stability of the OSCV smoothing parameter selection. Communications in Statistics: Theory and Methods, 30(10):2033–2044, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yoshida:2004:ULC

[YIS04] Kiyotaka Yoshida, Hideyuki Imai, and Yoshiharu Sato. Upper limits for criteria for tests of dimensionality under elliptical populations. Communications in Statistics: Theory and Methods, 33(11):2799–2816, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yun:2006:ALT

[YKC06] Won Young Yun, Eun Suk Kim, and Ji Hwan Cha. Accelerated life test data analysis for repairable systems. Communications in Statistics: Theory and Methods, 35(10):1803– 1814, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2006:NLB

[YL06a] Gui-Jun Yang and Min-Qian Liu. A note on the lower bounds on maximum number of clear two-factor interactions for (2^{m−p_{III}}) and (2^{m−p_{IV}}) designs. Communications in Statistics: Theory and Methods, 35(5):849–860, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

You:2006:PTR

[YL06b] Na You and Liping Liu. Proportional trapping-removal model with a known ratio. Communications in Statistics: Theory and Methods, 35(1):33–41, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2009:MSE

[YLX09] Hu Yang, Yalian Li, and Jianwen Xu. Mean squared error matrix comparisons of some restricted almost unbiased estimators. Communications in Statistics: Theory and Methods, 38(14):2321–2332, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2009:SRF

[YLZ09] Jianfeng Yang, Min-Qian Liu, and Runchu Zhang. Some results on fractional factorial split-plot designs with multi-level factors. Communications in Statistics: Theory and Methods, 38(20):3623–3633, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yafune:2005:NSS

[YNI05] Akifumi Yafune, Mamoru Narukawa, and Makio Ishiguro. A note on sample size determination for Akaike information criterion (AIC) approach to clinical data analysis. Communications in Statistics: Theory and Methods, 34(12):2331– 2343, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Young:2005:EEC

[YSH05] Dean M. Young, Stephen M. Scariano, and Cecil R. Hallum. Estimation-equivalent covariance structures for the least squares and MINQUE estimators of the linear model variance. Communications in Statistics: Theory and Methods, 34(3):625–629, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:2009:ELM

[YSZ09] Wen Yu, Yunting Sun, and Ming Zheng. Empirical likelihood method for censored median regression models. Communications in Statistics: Theory and Methods, 38(8):1170–1183, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamato:2007:JCC

[YTN07] Hajime Yamato, Koichiro Toda, and Toshifumi Nomachi. Jackknifing a convex combination of one-sample U-statistics. Communications in Statistics: Theory and Methods, 36(12): 2223–2234, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:2006:E

[Yu06] K. Yu. Erratum. Communications in Statistics: Theory and Methods, 35(7):1359, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:2007:MSA

[Yu07] Hong-Fwu Yu. Mis-specification analysis between normal and extreme value distributions for a linear regression model. Communications in Statistics: Theory and Methods, 36(3): 499–521, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yuan:2005:ASP

[Yua05] Ming Yuan. Automatic smoothing for Poisson regression. Communications in Statistics: Theory and Methods, 34(3):

603–617, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yue:2004:NOS

[Yue04] Jack C. Yue. A note on optimal strategies of a generalized two-stage bandit problem. Communications in Statistics: Theory and Methods, 33(7):1577–1585, 2004. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yi:2007:SIA

[YW07] Yanqing Yi and Xikui Wang. Statistical inference of adaptive designs with binary responses. Communications in Statistics: Theory and Methods, 36(8):1643–1653, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ye:2008:GIC

[YW08] Ren-Dao Ye and Song-Gui Wang. Generalized inferences on the common mean in MANOVA models. Communications in Statistics: Theory and Methods, 37(14):2291–2303, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yanez:2001:UME

[YWK01] N. David Yanez III, Gregory R. Warnes, and Richard A. Kronmal. A univariate measurement error model for longitudinal change. Communications in Statistics: Theory and Methods, 30(2):279–287, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yan:2008:EPD

[YWL08] Zaizai Yan, Jingyu Wang, and Junfeng Lai. An efficiency and protection degree-based comparison among the quantitative randomized response strategies. Communications in Statistics: Theory and Methods, 38(3):400–408, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2008:TPS

[YX08] Hu Yang and Jianwen Xu. Tuning parameter selection and various good fitting characteristics for the Liu-type estimator in linear regression. Communications in Statistics: Theory and Methods, 37(20):3204–3215, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2001:TSE

[YYAZ01] Shie-Shien Yang, Qifeng Yu, and Munther A. Al-Zaid. A two-stage estimation procedure for linear mixed-effects models. Communications in Statistics: Theory and Methods, 30 (12):2637–2653, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamashita:2007:SAM

[YYK07] Toshie Yamashita, Keizo Yamashita, and Ryotaro Kamimura. A stepwise AIC method for variable selection in linear regression. Communications in Statistics: Theory and Methods, 36 (13):2395–2403, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:2009:QLR

[YYL09] Lili Yu, Ruifeng Yu, and Liang Liu. Quasi-likelihood for right-censored data in the generalized linear model. Communications in Statistics: Theory and Methods, 38(13):2187– 2200, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yuen:2005:ERP

[YYW05] Kam C. Yuen, Hailiang Yang, and Rongming Wang. On Erlang(2) risk process perturbed by diffusion. Communications in Statistics: Theory and Methods, 34(11):2197–2208, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2006:MGP

[YYY06] Aijun Yang, Hui Yu, and Zhenhai Yang. The MLE of geometric parameter for a geometric process. Communications in Statistics: Theory and Methods, 35(10):1921–1930, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:2001:EIF

[YZ01] Chang Yu and Daniel Zelterman. Exact inference for family disease clusters. Communications in Statistics: Theory and Methods, 30(11):2293–2305, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:2005:TPA

[YZ05] Keming Yu and Jin Zhang. A three-parameter asymmetric Laplace distribution and its extension. Communications in Statistics: Theory and Methods, 34(9–10):1867–1879, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2007:RTK

[YZ07] Hu Yang and Chunmei Zhang. The research on two kinds of restricted biased estimators based on mean squared error matrix. Communications in Statistics: Theory and Methods, 37(1):70–80, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yin:2008:RSC

[YZZ08] Changming Yin, Hong Zhang, and Lincheng Zhao. Rate of strong consistency of maximum quasi-likelihood estimator in multivariate generalized linear models. Communications in Statistics: Theory and Methods, 37(19):3115–3123, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2005:SDC

[ZBLG05] Lingyun Zhang, M. S. Bebbington, C. D. Lai, and K. Govindaraju. On statistical design of the S^{2} control chart. Communications in Statistics: Theory and Methods, 34(1):229– 244, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zuoxiang:2009:TBG

[ZBN09] Peng Zuoxiang, Tong Bin, and Saralees Nadarajah. Tail behavior of the general error distribution. Communications in Statistics: Theory and Methods, 38(11):1884–1892, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2002:NEC

[ZC02] Lingyun Zhang and Gemai Chen. A note on EWMA charts for monitoring mean changes in normal processes. Communications in Statistics: Theory and Methods, 31(4):649–661, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2009:BME

[ZC09] Xiao Zhang and Gary Cutter. A Bayesian method of estimating kappa coefficient with application to a rheumatoid arthritis study. Communications in Statistics: Theory and Methods, 38(18):3432–3444, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zen:2009:NCR

[ZCL09] Mei-Mei Zen, Chia-Hao Chan, and Yi-Hsiung Lin. A note on criterion-robust optimal designs for model discrimination and parameter estimation in polynomial regression models. Communications in Statistics: Theory and Methods, 38(5): 584–593, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zewotir:2008:MCD

[Zew08] T. Zewotir. Multiple cases deletion diagnostics for linear mixed models. Communications in Statistics: Theory and Methods, 37(7):1071–1084, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2003:NAI

[ZF03] Junhua Zhang and Weiwu Fang. A new approach of information discrepancy to analysis of questionnaire data. Communications in Statistics: Theory and Methods, 32(2):435–457, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2000:MHT

[ZFY00] Zhong-Zhan Zhang, Yoshinori Fuji, and Takashi Yanagawa. On Mantel–Haenszel type estimators in simple nested casecontrol studies. Communications in Statistics: Theory and Methods, 29(11):2507–2521, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2003:OSE

[ZGS03] Bao-Xue Zhang, Wei Gao, and Ning-Zhong Shi. Onesided estimating and testing problems for scale models from grouped samples. Communications in Statistics: Theory and Methods, 32(12):2339–2352, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2001:GLR

[Zha01] Shunpu Zhang. Generalized likelihood ratio test for the shoulder condition in line transect sampling. Communications in Statistics: Theory and Methods, 30(11):2343–2354, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2006:MBC

[Zha06] Tonglin Zhang. A modification for Bayesian credible intervals. Communications in Statistics: Theory and Methods, 35 (9):1703–1711, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2007:ODT

[Zha07a] Chongqi Zhang. Optimal designs for trigonometric regression. Communications in Statistics: Theory and Methods, 36(4):755–766, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2007:INC

[Zha07b] Riquan Zhang. Inferences on nonparametric component for partially linear models. Communications in Statistics: Theory and Methods, 36(3):549–562, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2007:ETP

[Zha07c] H. B. Zhao. Exact tests in panel data using generalized p-values. Communications in Statistics: Theory and Methods, 37(1):18–36, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2008:EMC

[Zha08] Zhiwei Zhang. Estimating a marginal causal odds ratio subject to confounding. Communications in Statistics: Theory and Methods, 38(3):309–321, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2009:RAC

[Zha09] Yang Zhao. Regression analysis with covariates missing at random: A piece-wise nonparametric model for missing covariates. Communications in Statistics: Theory and Methods, 38(20):3736–3744, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zheng:2000:BEM

[Zhe00] Beiyao Zheng. Bayesian estimation of multidimensional item response theory model using Gibbs sampling. Communications in Statistics: Theory and Methods, 29(5–6):1405–1417, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zheng:2003:ILT

[Zhe03] Zukang Zheng. Improvement of life table method. Communications in Statistics: Theory and Methods, 32(4):827–839, 2003. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhou:2001:IVM

[Zho01] Julie Zhou. Integer-valued, minimax robust designs for approximately linear models with correlated errors. Communications in Statistics: Theory and Methods, 30(1):21–39, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zielinski:2006:SSQ

[Zie06] Ryszard Zieliński. Small-sample quantile estimators in a large nonparametric model. Communications in Statistics: Theory and Methods, 35(7):1223–1241, 2006. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zielinski:2009:ONQ

[Zie09] Ryszard Zieliński. Optimal nonparametric quantile estimators. Towards a general theory. A survey. Communications in Statistics: Theory and Methods, 38(7):980–992, 2009. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2008:SAD

[ZK08] Juan Zhang and John E. Kolassa. Saddlepoint approximation for the distribution of the modified signed root of likelihood ratio statistics near the mean. Communications in Statistics: Theory and Methods, 37(2):194–203, 2008. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zheng:2005:ELP

[ZL05] Ming Zheng and Sihua Li. Empirical likelihood in partial linear error-in-covariable model with censored data. Communications in Statistics: Theory and Methods, 34(2):389–404, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2006:IEF

[ZL06] Riquan Zhang and Guoying Li. Integrated estimation of functional-coefficient regression models with different smoothing variables. Communications in Statistics: Theory and Methods, 35(6):1093–1100, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2009:MTM

[ZLH^{+}09] Bin Zhang, Haiqun Lin, David J. Hunter, Tuhina Neogi, Barton Wise, Ernest Choy, Jingbo Niu, and Yuqing Zhang. A multistate transition model for osteoarthritis pain change. Communications in Statistics: Theory and Methods, 38(18): 3297–3306, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2005:EPS

[ZLY05] Likai Zhang, Liping Liu, and Na You. Estimating population size in logistic capture-recapture models with a known sex ratio. Communications in Statistics: Theory and Methods, 34(1):37–44, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2005:CRT

[ZM05] Racheal Zhang and Gary C. McDonald. Characterization of ridge trace behavior. Communications in Statistics: Theory and Methods, 34(7):1487–1501, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zacks:2009:EAP

[ZM09a] Shelemyahu Zacks and Nitis Mukhopadhyay. On exact and asymptotic properties of two-stage and sequential estimation of the normal mean under LINEX loss. Communications in Statistics: Theory and Methods, 38(16–17):2992– 3014, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2009:PLP

[ZM09b] Huixiu Zhao and Wen-Qing Ma. A pairwise likelihood procedure for analyzing exchangeable binary data with random cluster sizes. Communications in Statistics: Theory and Methods, 38(5):594–606, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zerbet:2003:NSD

[ZN03] A. Zerbet and M. Nikulin. A new statistic for detecting outliers in exponential case. Communications in Statistics: Theory and Methods, 32(3):573–583, 2003. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zheng:2004:FIM

[ZP04] Gang Zheng and Sangun Park. On the Fisher information in multiply censored and progressively censored data. Communications in Statistics: Theory and Methods, 33(8):1821– 1835, 2004. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2006:IMR

[ZQ06] Yichuan Zhao and Gengsheng Qin. Inference for the mean residual life function via empirical likelihood. Communications in Statistics: Theory and Methods, 35(6):1025–1036, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2007:ILF

[ZQ07] Yichuan Zhao and Gengsheng Qin. Inference for a linear functional of cumulative hazard function via empirical likelihood. Communications in Statistics: Theory and Methods, 36(2):313–327, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhong:2002:TAT

[ZS02] Bob Zhong and Jun Shao. Testing the agreement of two quantitative assays in individual means. Communications in Statistics: Theory and Methods, 31(8):1283–1299, 2002. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zahedi:2006:PER

[ZS06] Hassan Zahedi and M. Shakil. Properties of entropies of record values in reliability and life testing context. Communications in Statistics: Theory and Methods, 35(6):997– 1010, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zaier:2009:BAT

[ZT09] L. Hedhili Zaier and A. Trabelsi. A benchmarking approach to temporal disaggregation of economic time series by related series. Communications in Statistics: Theory and Methods, 38(19):3561–3582, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zuma:2007:BAC

[Zum07] Khangelani Zuma. A Bayesian analysis of correlated intervalcensored data. Communications in Statistics: Theory and Methods, 36(4):725–730, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhu:2000:OTA

[ZW00] Wei Zhu and Weng Kee Wong. Optimum treatment allocation for dual-objective clinical trials with binary outcomes. Communications in Statistics: Theory and Methods, 29(5–6):957–974, 2000. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhu:2008:LEA

[ZW08] Fukang Zhu and Dehui Wang. Local estimation in AR models with nonparametric ARCH errors. Communications in Statistics: Theory and Methods, 37(10):1591–1609, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2007:ESP

[ZXM07] Junlong Zhao, Xingzhong Xu, and Jianjun Ma. Extending save and PHD. Communications in Statistics: Theory and Methods, 36(8):1591–1606, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhong:2007:RER

[ZY07] Zhen Zhong and Hu Yang. Ridge estimation to the restricted linear model. Communications in Statistics: Theory and Methods, 36(11):2099–2115, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhu:2008:RPD

[ZY08] Jinxia Zhu and Hailiang Yang. Ruin probabilities of a dual Markov-modulated risk model. Communications in Statistics: Theory and Methods, 37(20):3298–3307, 2008. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2006:SAS

[ZYL06] Xian Zhao, Haisheng Yang, and Jinlin Li. A sensitivity analysis for system reliability. Communications in Statistics: Theory and Methods, 35(10):1845–1856, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhu:2008:NTI

[ZYSZ08] Chao Zhu, Kam C. Yuen, Jianguo Sun, and Xingqiu Zhao. A nonparametric test for interval-censored failure time data with unequal censoring. Communications in Statistics: Theory and Methods, 37(12):1895–1904, 2008. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2006:CDP

[ZYZ06] Yu Zhang, Musheng Yang, and Yanxin Zhang. Concurrent design for process quality, statistical tolerance, and SPC. Communications in Statistics: Theory and Methods, 35 (10):1869–1882, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2001:CLT

[ZZ01] L. C. Zhao and H. Zhang. A central limit theorem for testing exponentiality. Communications in Statistics: Theory and Methods, 30(6):1163–1170, 2001. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhu:2009:SAS

[ZZ09] Jianzhong Zhu and Wang Zhou. Saddlepoint approximation for sample quantiles with some applications. Communications in Statistics: Theory and Methods, 38(13):2241–2250, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2007:LLR

[ZZW07] Xiaobing Zhao, Xian Zhou, and Xianyi Wu. Local linear regression in proportional hazards model with censored data. Communications in Statistics: Theory and Methods, 36 (15):2761–2776, 2007. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).